A Bibliography of NHFB’s Master Collection

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 April 2024 Version 4.68

Title word cross-reference

+ [Kle06, Wir76]. 1.13198824 … [Vis00]. = [Wir76]. e [Mao94]. E = mc^{2} [Bod00, KN19a]. γ [Gra88]. π [Bec93, Bla97, EL04]. \sqrt{−1} [Nah98].

-GLA [Gra88].

.NET [AHG^{+}04, BS03].

00 [Mis90]. 000-year [Mao07].

1 [Knu11, MGH80]. 1.0 [Gra00a]. 1.1 [GHLP03]. 1.2 [QO93a]. 10 [MM07, MMG07]. 100-digit [BLWW04]. 101-1983 [Gra83a]. 1057 [IBM86]. 109 [Con05]. 11 [ACHM82, Dig82, EM79, LE80]. 1100 [FF78]. 112 [FW07]. 1164/WTL [Cor86]. 16-Bit [Dol79, MW82]. 16/32 [Mot84]. 16/32-Bit [Mot84]. 1883 [Win03a]. 1905 [Rig05]. 1914 [Mar14b]. 1918 [Ger20]. 1919 [Mac03]. 1945 [DW46, Sza84]. 1966 [Ame66]. 1974 [Erd74]. 1983 [Par84]. 1985 [Ame85]. 1986 [Ame86b]. 1989 [Ame89, AH89]. 198x [Ame87a]. 1991 [MA91]. 1992 [Ame92, Zla92]. 1994 [USE94]. 1999 [CS99]. 199x [Ame90]. 19th [BCDH09].

2 [CLK99b, Chr86, MS87, Wir83]. 2.0 [Kan96]. 20 [Gor81]. 2006 [Ber06]. 2007 [SM07]. 2018 [Joh18, Kol19, Mat19a, Mat19b, SK19]. 20th [SO11]. 21st [Pin14, Bal97, BB04c]. 24-dots [Jap83]. 25th [SBH^{+}04]. 26 [DB04]. 28 [CS99]. 2nd [IEE88].

3 [Gwe96, LS78, Nel91, Sed93, Sum09, Ruc15]. 3-D [Gwe96]. 3.5-billion-year [Shu08a]. 32-Bit [Mot84, Mot85a, Mit86, Mot85b].

4.3BSD [LMKQ89]. 4.4BSD [MBKQ96].

5 [Gen93, GBR05].

6.22 [Nor94]. 6000 [Bel90, Hos91, WR94a, WR94b]. 6234-1983 [Jap83]. 64 [Mar00]. 64-bit [ET03, Int90, Int91, Cor86]. 6626-1978 [Jap78]. 6th [Lio96].

7 [Rob99]. 7185-1982 [Int82]. 754 [Hou81]. 754-1985 [IEE85b]. 77 [Ell90, Ett93, Kat78, MS88]. 770X3.97 [Ame83b]. 770X3.97-1983 [Ame83b]. 7th [Zla92].

8086/8088 [LG84, MW82, Ski85, Yeu84]. 8086/88 [Sca84]. ’89 [HHS89]. 8X [AO90, MR87].

9 [MGH^{+}03a, MGH^{+}03b]. 90 [ABW92, BGA90, Cou91, Ker93, MR90, Tho97]. ’92 [Zla92]. 95 [ABM^{+}97].

= [Maz97, RB08].

Abecedarium [Fir93]. aboard [Bur98]. Absolute [Sha99a, Cre11]. Absorption [Dye65, Nak62]. Abstract [BC06]. Abstracting [Lor19]. Abstracts [Che94]. abuse [DeL86]. academic [BB04a]. Academy [Sok00]. Acceleration [Gle99]. Access [Nor82]. Accessory [dTL18]. Accompany [KCM^{+}76, Kyk76]. account [Nic87]. Accuracy [Hig96, Hig02, BLWW04]. Accurately [Cli90, SW90]. ACE [Cop12]. Achievements [Bel37, Woo80]. Acrobat [Mer97b, Mer97a, Ame93, GS93, Mer98]. Acrobat/PDF [Mer97a, Mer98]. ACS [DCJ^{+}97]. Action [BBC^{+}07, ST12, HPR16]. Ada [Ame83a, Ano79b, Ano79a, May82, Pyl81, Weg80, You83]. Adam [Bic09]. Adapt [New14]. Add [Gwe96]. Addressing [FA89]. Administration [BF05, Fri91, Hun92, MJW^{+}06, NSS89, NSS^{+}01, NSH02, NSHW10, AT&T86f]. Administrator [MP92]. Admit [ST12]. Adobe [Inc90a, Inc90b, Inc, Ame93, GS93]. Adolescence [GREC91b]. Advanced [The98, Bro77, CW66, Dav65, DS51, GW98, Her81, JR90, JL16, KD93, KD95, KD99, Kre93, KN94, Kun86, MGH^{+}03a, Muc97, Nor82, Roc85, Roc04, Sal95, SKM19, Ste92, SR05]. Adventure [Bro15]. Adventures [Ban99, FL06, KM84, Lor18, FL88, GM97, FLH85]. AES [Ope91a]. Affirmative [ST12]. afraid [MZP97]. Africa [Hoc98, Hoc99]. After [HH15, Gor15]. Again [Eis06]. Against [DB04, Jer02]. Age [Ess04, Pri21, Pyn10, RH97, Stu18, Tap96, Whi88, Whi90a, Acz09, Hol08, Mag09, SH16, SG17, Tho08b, Fag10, Fag19]. aid [Ame70]. Aided [Far88, Far90, Far93, Far97, LS89]. AIP [Ano90]. Aircraft [RMV42]. Alamos [Ano96a, Bet91, Con05]. Alan [Cop12, Swi19]. Albert [BB05, Sch49a, Sch49b, Woo80]. Alexander [Hel04, Shu08b]. Algebra [HM88, Joh93, WR71]. Algebraic [Ray87, Wil65, Ata40]. algebras [TW95]. Algorithm [BB91, Gri86, Ski98, Ski08, Sun90, BM77]. Algorithms [AHU74, Akl85, ANTT02, CLR90, CLR01, CLRS09, Ear85, Edm08, FBY92, GHN^{+}90, Gil78, Gon84, GK82, GK78, Gus97, Hig96, Hig02, HJ83, HJ88, HS78, JL96, KR81, Knu97, Knu98a, Knu00, Knu08, Knu68, Knu69, Knu73a, Kor02, Kro85, Luk77, Mal72, MT86, Mat19a, MB00, Mul97, Omo94, Par00, Pav82, Sch94, Sch96, Sed83, Sed88, Sed90, Sed92, Sed93, Sed98, Sed02, SW11, Ste94a, Wil86a, Wir76, Wir86, CBB^{+}99, Knu11].

Allemagne [Mic89]. Alliance [dTL18, Bla02]. Allowing [WM92]. almanac [Cal04]. Almost [BH17, Tau61b]. Alpha [Bha96, Cor92, Sit92, SW95]. Alphabet [Fir93, Zap07a, Int99, Zap07b]. Alphabetgeschichten [Zap07b]. Alsos [Gou96]. AltaVista [SRR97]. ALU [Cor86]. Amateur [LB14]. Ambiguity [SB07]. Ambition [Jon08, Kea18]. America [Bla02, DB04, Fag17, Gra11, Jac14, Kev79, Mar14a, Nic87, Tho81, Tho08b, Win05, Win14, dT08]. American [Ame86a, Ame88a, Ame88b, Org91, Ame89, Ame92, Ame90, BS05, De 19, Hel04, Mun17, Ric06, Sch90, Ted06]. Americans [Boo58]. Amsterdam [Sho14]. analyse [Cou38]. Analysis [AHU74, BC81, BC84, BN60, Cet91, Cod81, Ehr73, GK82, Kah92, Knu00, KS52, Mor83, Pen65, Ric83, SW80, Str74, TD00, Tau63b, Ueb97, WMD65, Cou38]. Analytic [FZ61, Oak49, Oak58]. Analytical [GW98, Hut52, Bab10, Lud09]. Analyzer [Fel76, LS79, Pax88]. Analyzers [Gra88]. Analyzing [Mil92, Stu84]. Anatomy [All78]. ancient [AK11, Bar80, Wei16]. Andy [Ted06]. angels [Pin11]. Anglicko [CNVH93]. Anglický [CNVH93]. Angular [Sch15b]. Animals [Hau00]. Animating [PW94]. Annals [McP98]. Annihilation [Ave76]. Annotated [AHG^{+}04, ES90, JPS07, Sch90]. Annual [SBH^{+}04]. ANS [ANS76]. Ansätze [Hen92]. ANSI [Ame89, Ame66, Ame78, Ame87a, Ame85, Ame86b, Ame83b, ABW92, ABM^{+}97, Dyb96, IEE85b, Sch90]. ANSI/IEEE [Ame83b, IEE85b]. ANSI/ISO [ABW92]. Answer [TG89]. answered [Ado97]. Answers [Kit63]. Anti [RMV42]. Anti-Aircraft [RMV42]. Anticipating [Mac11]. Antimatter [QN08]. Antoine [Hut52]. Anything [SRR97]. APA [Tri98, Tri99]. Apache [LL97]. API [SFR04, X/O93a]. APIs [Ste98]. Apology [Har40]. Appendix [Tri99, Tri98]. Application [Cou38, Ope91a, Wei80, Gut11]. Applications [AT&T86a, AT&T86b, Bic48, Cha78, Chu60, Cla90, Cot63, Den66, Dye65, EM79, Egu18, Erd74, FSK10, GB18, HJ76, Hou81, Hwa84, JR89, JR90, Joh96, Joh18, Kun83, PW35, PP86, Rob59, Sei98, Sri96, Vul92, WC00, You89, You90, YP92, Cor05, Leb72, Mer97a, Mil15, Tor15]. Applied [Art84, Art85, Ban99, Fow97, HY81, Har00, Kur94, Sch94, Sch96, War85, BR89]. Approach [AK01, Ara06, BYG92, BL63, BS60, Com84, FBD05, Fro77, Har83, Har87, HP90, HP96, HP03, HPAD^{+}06, HP12, HP19, HK05, KmWH10, MB00, PH90, PH96, PH03, Smi90, Tru03, Cow90, FGMvH14, Sin07]. approaches [Hen92]. Approximation [Gal08]. Approximations [HCL^{+}68, Has55, Luk69a, Luk69b, Luk75]. Apps [Mat19b]. aqueous [Lat52]. Arabic [AK11, Kha86]. Arc [Kea00]. Archibald [Gut12]. Architecture [AIM95, Bae80, Bha96, BB97, Com88, Cor92, Dol79, ET03, Gut04, HH07, HP90, HP96, HP03, HPAD^{+}06, HP12, HP19, HJ83, HJ88, Hsu01, Kan87, KH92, Kan96, Kog81, LE80, LG84, Mar90, MM01, MSSW94, MA86, MP00, MB88, Omo94, PH90, PH96, PH03, PW17, SPA92, Sit92, SW95, Sto07, Sun87, Tab87, TZ86, WG94, WS94, Bae10, MM07, Sha99b, Sta10]. Architectures [AK01, Com00b, HP92, MT86, PN04]. archive [KW90]. Archives [Bar06b]. Areas [Mil16b]. Arithmetic [BZ11, BCDH09, CCG^{+}84, Coo80, Coo81b, Hou81, IEE85b, Kor02, Loc17, Mal72, MBdD^{+}10, Omo94, Ove01, Par00, SO11, SR06, Ste81, Swa90a, Swa90b, Dev11]. Arithmetical [Bus40]. arms [Far13b]. Arrays [PCGK89]. Art [Bis03, BN06, Bre09, Keh93, Knu05a, Knu05b, Knu08, Knu09, Knu11, Kun85a, Mye79, PFTV86, PTVF07, Pru84, Ray04, Ruc15, TW87, Zob97, FF95, HS08].

Articles [Gri78, Gus12]. Artificial [Ung87]. artists [Mor07]. ASA [Hei66]. Ascent [HB90a]. asked [Ado97]. Aspects [Eth10, Hsu01, Thu80, WR94b, Hen92]. Aspekten [Hen92]. Assembly [Bra84, Gor81, Laf84, Mor84, Sca84, Ski85, Yeu84]. assessment [Boi97]. Associated [Bic48, Gol98]. assumptions [Smi14]. Astonishing [BH17, Liv02]. Astronomers [Fle18]. Astronomy [PF19, Cla09, Hen97]. Astrophysics [Tau63c, dTL18]. Asymptotics [Olv74]. AT&T [AT&T86a, AT&T86b]. Atanasoff [Smi10]. Athens [Wei16]. Atlantic [Win11]. Atlas [Fro04, Mic05, PW17, Tho97, Ano76, Ano99, Ran91, Ran01]. Atmospheric [Boh87]. Atom [AH13, Far09, FS65, Rod19, Sin05, Win15, Cat04, vCH18]. Atomic [Bag10, BBSR69, FR13, Gre20, HL33, Kur06, Mah09, Rho86, SM68, Bet91, Kel09, SH16, Kie13]. Atoms [Fro77]. Auditory [Ram97]. August [AFI75, SBH^{+}04, Win03a]. Author [Bry88]. Authors [DCJ^{+}97, Fly95, RV09, Smi87b, Whe21]. autobiography [Hah67]. Automata [Tau63b, Kle56]. Automatic [JL96, AOB64, Zus36, AH46]. Automating [CH87, Lib94, MJW^{+}06, RB06a]. automation [Cle88, Tor15]. automatique [Tor15]. automatisieren [RB06a]. aux [Cou38, RB05b]. available [KW90]. Avenarius [Kro90]. awakening [Mah09]. Aware [GM02]. Away [Lor19]. Awk [Col91, Dou91, AKW88, AKW23, Rob96, Rob97, Rob01, Van86]. AXIOM [JS92]. AXP [SW95].

B [BL19, Ope91a]. B5700 [Org73]. B5700/B6700 [Org73]. B6700 [Org73]. Babbage [Bab10, Swa01]. Babel [Dor18, Gor15]. background [Can01]. Bacon [GG05]. Bad [Tau93]. balance [Cre11]. Ballistic [Jul47]. Band [Fle18]. Bang [MH22, Par03, Cle09, Seg11].

Baptiste [Hut52]. Base [AHG^{+}04, Cor05]. Based [Gus12, McK79, AK01, CHT02, Lib94]. Basel [Ten09]. Basic [Ano05, RC64a, RC64b]. Basis [AKM81]. Basque [Kur99]. bastard [McW08]. Battle [Sus08]. Battles [Win11]. Bayes [McG11]. BDs [GHLP03]. Be [Far03, Liv05, GREC91a, HHPM07, Str13a, The97, The98]. Beaten [Fey05]. Beautiful [Bla15, Far03]. Beauty [FvGGM90, PR86, Hol08]. Beer [Boh87]. Before [Cle09, MH22, Gor15]. began [Cla09]. Beginner [Keh93, SL91, Sno92]. Beginners [KD93, KD95, KD99]. Beginning [DS51, Gla97, Str74, Ano96a]. Behind [Ano96a, Ber05, Col05, Rue07, CL88, Ekm16, Maz10]. Being [Bar92, Bar96]. Bell [Alt48, Geh03, Shu08b]. Belong [Sto99]. Ben [RB08]. Benchmark [GREC90, Pri89]. Benchmarks [BBB^{+}91, Cyb91]. benefits [BB04a]. Benford [BH15, Mil15]. BeOS [The97, The98]. Berkeley [Ope91e]. Bernhard [Der03]. Bernoullis [Ten09]. Beschreibung [Hen92]. Best [Pit19, Fey00]. Beta [Jen00]. Bethe [Ber81, Sch12b]. Better [BB91, HM98, Pag07, Pin11]. Betting [Ung87]. Between [Stu18, dTL18, Bla02, BB05, Sen21]. Beyond [Ame93, Cas09, Mao91, MH22, Pau91, Pav99, Rod19, Sch03b, Wat90, LM06]. Bibliography [Gut12]. Big [Hil15, Hür22, MH22, O’N16, Par03, Car94, Cle09, DeL86, Seg11]. Biggest [Rei09]. BigNum [SR06]. billion [Shu08a]. Billions [Ker18]. Binary [Gri90, IEE85b, Phi36, Ste81, Knu09]. Binding [McP74]. Biographical [Ara06, Hen92]. biographischen [Hen92]. Biography [Lor18, Mal95, Pad90, Sei96, Bod00, Kur98, Smi10]. Biology [Gus97, LP64, OALD97]. biomedical [BR89]. BIOSIS [BIO94]. Birth [Dol11, Ess04, Hil15, RH97, WB01, Win01, Fag10, Kel09, SH16]. Birthday [FvGGM90]. bis [DW46]. BISON [DS88a]. Bit [Dol79, Mot84, Mot85a, Mit86, MW82, Mot85b, Rim93, ET03, Int90, Int91, Cor86, Aug84]. Bit-Mapped [Rim93]. Bitwise [Knu09]. Black [Fle18, Imp19, Lev16, Sus08]. blank [Pin02, Pin03]. Bletchley [Pri21, Smi98b, Tur20]. blink [Par03]. Blue [CL88, DeL86]. Bluebook [Mor84]. blues [Lev16, Car94]. BMP [Mia99]. boat [DB04]. body [Mit06, Shu08a]. Bohemia [Gor20]. Bohr [AH13, CG14, FK85, Kum10, Lin08, Pai91]. bolts [Ame01a]. Bomb [Bag10, Kap20, Kur06, RS09a, Rho86, Rho95, Cas09, Far13a, Far13b, FGMvH14, Her02, Kel09, Ric06, vCH18]. Bombs [Win15]. Bonding [Com64a]. Bones [Mar15]. Book [Ber06, CG96, GRA89, Gra90, Kau78, PT97, Ros91, SL91, Sin99, TG89, BBB00, Com00a, Cox15, AAG^{+}00, ABR02]. Books [Les97]. Boolean [Knu08]. Borders [Leh98]. Born [BB05, Gre05]. boson [Kra14]. Boston [MA91]. bottle [Sei08]. Boundary [Hoh01]. Box [Bil89]. Brain [Cop12, Rue07, Sch98]. branded [X/O92]. Braun [Bid09, Gör11]. break [BL19]. Breakaway [Sun97]. Breaker [Str95]. Breakers [Mun17]. breakneck [Zac14]. Breakthroughs [Cre10, Haw05]. Brief [Cle03, Har15]. brilliant [FLS95, Mag09]. Brinch [Bri85]. bring [Jac14]. Bringing [PBMW98]. Britain [Ano76, Ano99, Far13b, Smi98a, Syk07]. Britannica [Boy16]. British [Con09, Jon78]. broad [McK79]. Broken [Tur18]. Brotherhood [Her02]. Bruno [Clo15]. Brutal [Win15]. Buch [ABR02, Mer91]. Bucks [Les97]. Bugs [Dup95, Dup98]. Build [Cha18, Cop12, Joh18, Kol19, Mat19a, Swa01, Tro18]. Building [Bai99, GB18, Cor05, Mat19b, NS21, Sri96, DB04]. built [Zim08]. Bush [Bus45]. Business [FvGGM90, Hos91, Sun97]. Bust [Lei00]. Bytes [Les97].

C

[Ame89, Jap78, Jap83, Kru94, RH11, TG89, Van99, Wal01, Ame86a, Ame88a, Ame88b, Aus99, Bel90, Blu02, Bud98, Cam87, CLG99, Dig82, ES90, FH95, HK82, HS84, HS87, HS91, HS95, HS02, Hol90a, Hun85, Hym90, Jae89, Jae91, Jae93, Jae96, Jae01, Jos99, KR78, KR88, Lap87, Lip91, Pla92, PSLM00, Plu87, RS90, SGN88, Sch90, Sch94, Sch96, Sea05, Sea20, Sed90, Sed92, Sed98, Sed02, Sri96, Str86c, Str91, Str94, Str97, Str13b, TG89, Van88, Van99, War85, dG93, dG96, dG97]. C# [HWG04, JPS07]. C-Kermit [dG93, dG96, dG97]. C/C [Blu02]. CA [USE94]. Cache [JNW08]. CAE [X/O93a]. CAGD [Far01]. Calamities [Gla99]. Calc [GHLP03]. Calculating [Ste16, AOB64, Bab37, Cou38, Mau42, Zus36, Cou33]. Calculation [Sta92, Cou38, Phi36]. Calculations [But64, DR97, RD02a, Zus36]. Calculator [DW46, ST52, AH46]. calculer [Cou33, Cou38]. calculs [Cou38]. Calculus [Bar06a, BF91, Dev93, EL89, Fat92, FZ61, Har92, HEBS93, PW94, Str19]. Calendrical [DR97, RD02a, RD02b]. California [Goo91, IEE88, Win05]. Calligraphy [Her67]. Calling [Knu89a]. Calls [AT&T86g]. Cambridge [Cat04, JL14]. came [Str13a]. Can [Gra88, War08]. Canada [Ran91]. Canadian [RS91]. Cancer [Con20]. can’t [Kle06]. Cape [Ber06]. Captain [Nic03]. Captions [Gol98]. caractères [Har04]. Carbon [Eli62]. card [ST52]. card-programmed [ST52]. Cards [Fer88, Ghi05]. care [FL88]. Carl [Ten06]. Caroline [Hos11]. Carrington [Cla09]. Cartography [Her69]. Case [OALD97]. Cassell [Sim59]. cat [MZP97]. Catalog [Kro93]. Catastrophe [Wal03]. cathedral [Cat04, Dys12]. Caught [Mur77]. cause [Chr97]. Cavendish [Cro74]. CC [Sta88]. CD [AL97]. CD-ROM [AL97]. celebrate [Woo80]. Celebrated [Boy16]. céleste [Cou38]. celestial [Cou38]. Cell [OALD97]. Celsius [Ekm16]. Celts [Syk07]. centenary [FK85]. centennial [Woo80]. Center [Mon13, Par84, PA06]. centuries [McG11]. Century [Dic03, HH99, Pin14, Ree20, Sal94, Sta98b, Dev08a, McK79, Bal97, BB04c, Har10]. CERN [Acz10, Kra14]. Certain [SB07, Gre05, TW95]. certificate [BL63]. Cesko [CNVH93]. Ceský [CNVH93]. Chairing [Pau89]. Challenge [Ber06, BLWW04]. Chambers [Pry58a, Pry58b]. Chance [Tal01]. Chances [Eth10]. Change [Rei88b]. Changed [BH17, Bro15, GP19, Hel04, Win01, Fal11, Haw05, Kur98, Mac03, Mar14b, Sta98a]. changes [Ekm16]. changing [Mil16a]. Chaos [Sch91]. Chaotic [Saa01]. Character [FLH85, Fey94, FL06, Jap78, Jap83, Mac80, Sea94, The91, Cle88, FL88, FF95, SW72]. Characters [Bat88, NF87]. Charles [GP04, Nic03, Swa01]. Chasing [Coh10, Wul12, Shu08b]. Chemical [Ben33, Ben35, Che94, Com64a, Cot63, Den66, Erd74, Lai65]. Chemistry [Bar79, BL63, CW66, DWB^{+}62, Dav65, DS51, Den66, Dor65, Gla46, GL60, Han63, Hut52, JO83, Kau57, Kit63, La 71, LÖ73, LÖ04, LP64, MM56, MM64, PW35, Pit53, Pry66, Rob59, RC64a, RC64b, San64, SG62, SP64, SLG52, Tay25b, Tay25a, WBA85, GS12]. Chemists [Str61, Tay25b, Tay25a]. Chicago [Ano93, Ano03, Ano10, Ano17, Tur07]. Chimie [Fra64]. Chimique [Eve65]. China [KIR^{+}07, Kis11, Tem91, Win08]. Chinese [Lan01, Lun99]. Chinese/English [Lan01]. Chip [Ele98, Wea08, Bae10]. Chips [Col05, Win15]. Chronicle [Zap07b, Zap07a]. Chronicles [Col05, Fal11]. Chronik [Zap07b]. Churchill [Far13a, Far13b]. Cipher [GG05]. Ciphers [AAG^{+}00, Has16]. Circles [Wir90]. Circular [For12, Thu80]. Citation [PBMW98]. Citizen [Fey98]. city [Con05, Sho14, Kie13, Ros14]. Civil [Rho15]. Civilization [Cah95, McL91]. CJKV [Lun99]. Clara [IEE88]. Clare [Gut12]. Clarification [ANS69, ANS71]. Clash [Bar06a]. Class [AHG^{+}04, CL98, CLK99a, CLK99b, Fer86, Gri86]. Classic [FL06, RB05a, Sta99, RB06a]. Classical [FS65, Jac62, Lee65]. Classified [Con20]. Classroom [PJS91, Sto99]. Claude [SG17]. Clients [Cha18]. Climate [Fag19]. Clocks [Gal03]. Clockwork [Dol11, New07]. closer [Har10]. Clouds [Boh87]. club [Ber96]. Co [Tay25b, Tay25a, Wat90]. Co-operative [Tay25b, Tay25a]. Cod [Kur98]. codages [Har04]. Code [ACHM82, FF01, Hol04, Jap78, Kea12, Lio96, Mat19a, Mun17, Pag82, Pag83, Sch94, Sch96, Sin99, SS02, Str95, McG11, Pag11, AAG^{+}00]. Codebreakers [Kah74, Kah96, Smi98b, Tur20]. Coded [Mac80]. CODEN [Ame70, Chexx]. Coder [Ben88]. Codes [DB04, Sal07b, Fag17]. Coding [Hsu01, Sal05a, Sea05]. Cold [Sha99a, Sen21, Tau93, Mac18]. Collected [Fey05, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. Collection [Ben33, Ben35, GK78, Gri78, JL96, LEMW52, Bar80]. College [Gei75]. Collider [Acz10]. Collision [Win15]. Colonial [Boo58, Hoc98, Hoc99]. Colophon [Inc]. Color [DRB89, McM88, Whi90a, Cor89]. Colored [Dys10]. Coloring [Kau78]. Colors [Wil02]. Colossus [Pri21, Ran77]. Combinations [Knu05b]. Combinatorial [Knu08, Knu11]. Comes [Smo07]. Coming [Win15, Bro10]. Command [Hol87b, Voß12]. Commands [X/O89c, AT&T86c]. Comment [Her74a, Her74b].

Commentaries [CGO07, BB05]. Commentary [Lio96]. Commercial [Nes87]. Committee [ACM77, ACM79]. Common [AIM95, BN08, BS03, X/O95a, Bro85a, Ste84, Ste90a, Tat87]. Communicating [BRR08]. Communication [RR96, RR03, Zob97, Ame70]. Communications [Cam87, Mic12, Sal05a, Ste99b]. Community [Kev79]. Compact [Han85, NO92a]. Companies [Gla99]. Companion [GMS94, GRM97, GR99, GMR^{+}08, JL14, MGB^{+}04, Ups90, Van99]. Company [Rho15, Tes07]. Compatible [Hol87a, DS88a]. Compiler [AU77, ACHM82, CT04, FH95, Hol90a, Joh79, Muc97, Pys80, SF85, SS02, Wir92, Bal19, CT11]. Compilers [ASU86, ALSU07, AK01, Bri85, Hun81]. Compiling [AU72, AU73, DM81, AU73]. Complete [ABW92, ABM^{+}97, AS90, Bha96, Leh03, Sal98, Sal00, Sal04, Sal07a, SKM19, Nor94, Sum09]. Complex [Chu60, Gal08, Hil15, Des10, GM97, WB00]. Complexity [Cet91, Kro85, Mit11, MP00, Pag88, Wil86a]. Composites [Mil02, Mil16b]. Composition [Rub88]. Compounds [Dye65, Eli62, FPZ^{+}64, SFC64]. Comprehensive [AWH13, BR89, Whe21]. Compressed [Mia99]. Compressing [WMB94, WMB99]. Compression [GBL^{+}98, PM92, RY01, Sal98, Sal00, Sal02, Sal04, Sal07a, Sal07b, Sal08, Say96, Say00, Say05]. Computation [ANTT02, Bee17, CGO07, Hey99, JS92, Jen77, Luk77, MT86, PP09, PF05, Ray87, Ueb97, ZJ96, Eck48, Par84, SM07]. Computational [FP79, GK78, Gus97, GW98, HM88, Kro85, Mor83, PS85, SM07, Van92, BBG04]. Computations [CV88, GHN^{+}90, GV83, GV89, RG88, Ric81, Ste73]. Computer [AFI75, Ame85, Ame86b, Ame88c, AHU74, Ang81, AKM81, AB87, Bae80, BB82, BBB87, BN71, BMM78, BSN82, Bis03, BB97, Bra84, BZ11, BCDH09, Cha78, Cop12, Ear85, EKP84, EKP87, Far88, Far90, Far93, Far97, Fv82, FvFH90, GL81, Gil78, GG46, Gus97, GREC91a, Har83, Har87, HH07, HCL^{+}68, HP90, HP94, HP96, HP97, HP03, HP04, HPAD^{+}06, HP12, HP19, Her81, Her72a, HS78, Hsu01, Hub84, IBM86, IEE88, IEE86, Int82, Ifr00, Jul47, KK67, Knu70, Knu96, Knu03a, Knu05a, Knu05b, Knu08, Knu09, Knu11, KD04, Kor02, KN94, Kun83, Kun85a, Kun85b, Kun86, LE80, LS89, Mac92, McC01, MS87, MP00, NS73, NS79, NS21, Omo94, Org73, Pag88, Par00, PH90, PH94, PH96, PH97, PH03, PH04, PH08, PH12, PH18, PP86, Pen65]. Computer [Pru84, Ram97, Ran77, Ran89, RS87, RA76, Rog85, RE87, Rub88, Ruc15, Sal99, Sal05a, Sal05b, Sal06a, Sal06b, Sal11, SS84, Sch83, SO11, Sco82, SL98, SBN82, Sta10, Sti40, Sto07, Sto89, Sto99, Stu84, Swa01, Swa76, Swa90a, Swa90b, Tab87, Tan88, Tan96, Tau77, Tau79, Tau76, TZ86, WH76, Wol88, Wol91, Zob97, ABR02, BH09, BD09, CL88, Com00a, HPR16, Knu99b, Smi10, Knu86d]. Computer-Aided [Far88, Far90, Far93, Far97]. Computers [Aug84, Bos99, DDSv91, Don89, Ehr73, Gri05, Has55, Her67, Hey99, HJ83, HJ88, Kog81, The92, Ran82, Sch84, Sta90, Sto99, Tau63b, SSU85, WK48]. Computing [Alt48, AB87, Ata40, CHT02, Dij82, Dow93, FJSD96, Gla99, G^{+}99, GM02, HC21, Lun99, MS88, MRS82, Nas90, NS21, Ove01, PFTV86, PTVF07, Sun97, WC00, Wal01, Wea08, Wil97, BLWW04, BGvN46, DHG08, Fro03, Tho97, Zus62]. Concepts [AvD19, Art84, Art85, BB97, EI66, HJ04, Mom01]. Concise [Sal08, Röh14]. Concisely [Joh92]. Concrete [BC06, GKP89, GKP94]. Concurrency [Bre09, Cha18, RR96, RR03, SK19, Tro18].

Concurrent [GB18]. Condensed [Ben33, Ben35]. conducted [Par84]. Conference [AFI75, AH89, Boi97, IEE88, SM07, SBH^{+}04, USE94, Zla92, Erd74, Par84]. Confessions [Ben88]. confidence [Pau89]. Confidently [Tro18]. conflict [BO76]. Confront [Ros19]. confuse [Bes04]. conjecture [O’S07, Dox00]. Connecting [Est93, Gia92, Gia90]. Connection [Hil85]. Connections [For12]. Conquest [Sha99a]. Consciousness [RK11]. Consensus [Jen00]. Conservation [Bel74]. considered [Dij68]. Constant [Hav03]. Constants [Fin03]. Constrained [Fle81]. Construction [Hun81, Pau79, Pys80, Ran89, SF85, Hen97]. Constructs [Law78, SKM19]. Contemporary [AAR90, AAR94]. Continued [BvdPSZ14, CPV^{+}08]. Continues [GREC91c]. Continuing [Ber06]. Continuity [Rei88b]. Continuous [Tau62]. Contrarian [Sto99]. Contribution [WH76]. Control [RMV42, X/O93b]. Controlled [AH46]. controversial [Fal11]. Controversy [Jen00, McG11]. conventions [DHJ^{+}12, Pos89]. Conversations [AAR90, AAR94]. Conversion [LS78]. Convolution [NF87]. Cookbook [Ado85b, CT98, Dar01, Mat19b]. Coordinates [WH76]. Copenhagen [Seg07]. Copernicus [Sob12]. Coprocessor [Mot85c, Mot85d, RJR88, JR13]. Copyright [Gol98, Kle06]. Coral [Win15]. Core [MM01, SR00]. Corfu [SM07]. Corporation [Cor19, Bla02, RH88, RH90, Sch03a]. Correctness [MP00]. correspondence [BB05]. Cosmic [Pag82, Pag83, Pag11]. Cosmology [RN68, Kea18, Seg11]. Cosmos [Gre04, Gre11, Hel04, PF19, Ste16, Bar06b, LBS17, PA06, Sob12]. Costs [ZM08, BB04a]. Could [Mit18]. Couldn’t [Liv05]. Counting [Bar92, Bar96]. Courage [Ros19]. Course [Bro77, DS51, Jae96, Kit63, MS87, Str74]. cousins [Car10]. Crack [Win05]. Cracked [AAG^{+}00, McG11]. Cracking [Ele98]. Craft [Fin91, BCW08]. Crafting [Nys21]. crapshoot [New07]. crater [Alv97]. CRC [Bey78, RG88, WBA85]. Create [Kur12, Zac14]. Created [Isa14, Mit18, Win18, Acz09, Haw05]. creates [Pag07]. Creating [WC00]. Creation [Acz10, Ave76, Boy16, Win14, Gle10]. creative [Wei16]. Creativity [CW14]. Creators [Boo92, Kel09]. Crisis [Smo14, Bur98]. Criteria [dB75]. Cro [Fag10]. Cro-Magnon [Fag10]. cross [Mer97a]. cross-platform [Mer97a]. crown [Geh03]. crucible [BD09]. Cryptographic [Gut04, SR06]. Cryptography [FS03, FSK10, Mao04, Sch94, Sch96, Sin99]. Crystal [RH97]. CSS [SR00]. Cuckoo [Sto89]. CUDA [SK10]. Cult [ZM08]. Cultural [Mao91, Smi98a]. Culture [Dyl20, Mil16a]. cures [Nah06]. Curie [Des10, Eml12]. Curious [Eis06, FLH85, FL06, FL88, FF95]. curmudgeon [And11]. Curses [Hol87a, X/O95b, Str86a]. Curve [Far91, McP74]. Curves [Far88, Far90, Far93, Far95, Far97, Far99, Far01, Gri86, Sal06a, Wil95a]. CWEB [KL93]. Cycles [Pen11]. Cyclotron [Hil15]. Czech [CNVH93]. Czechoslovakia [Zla92].

D [Ale10b, Gwe96]. Dahl [Con09]. damned [Bes04]. dangerous [Ber07]. dans [Bor90]. dansk [Lar89]. Daring [GP19]. Dark [Bid09, Gre20, Nic03, Rho95, Gat09, Ros14]. darkness [Mag09]. Darwin [Nic03]. Data [BC81, BC84, BHK20, Bud98, Ehr73, FBY92, Gil78, Gon84, GG78, HS83, KPv86, Kru84, Mat19a, PM92, RY01, Sal98, Sal00, Sal02, Sal03, Sal04, Sal05a, Sal07a, Sal07b, Sal08, Say96, Say00, Say05, TD00, TA81, Van88, Wir76, Wir86, X/O88a, O’N16, SW72, Sed98, Smi14, Whe14]. Database [BIO94, Rod90, Wie77, Wie83a, Wie83b, Wie87]. Databases [FHCD92]. datacenter [BH09]. Dateien [Mer97b]. daughters [Eml12]. Dawn [Ale10a, Pri21]. Day [Bar09, Win03a, Sza84]. DCE [X/O93b]. dead [Sch03a]. Dealers [Klo86]. Debate [Gwe96, Kum10]. debugging [Dup98]. DEC [BMM78, Sch03a]. Decay [Jen00]. Decimal [Gri90]. Decision [Knu09]. declined [Pin11]. Decoding [Mac11]. DECSYSTEM [Gor81]. DECSYSTEM-20 [Gor81]. Deep [Gre11, Ree00]. Defeat [Tur20]. Defending [Ker18]. Definite [GL81]. Definition [RS91, Tor15]. Definitions [X/O95a, X/O89b]. Definitive [Fla97, LL97, MK96, WM99]. degree [BL63]. Degrees [Ekm16, Lin04]. Delbrück [Seg11]. Deleted [Ame70]. Delight [War13, War03]. Delusions [Bor94]. Deluxe [BST^{+}96]. democracy [O’N16]. demonstration [Wor50]. Demystified [BN08, Gil03]. denial [Pin02, Pin03]. Denormalized [Coo81a]. Dense [Hen92]. Density [Erd74]. Dependence [AK01]. Dependence-based [AK01]. Deploy [Kol19, SKM19]. Descent [DM81, Han85]. Description [Ame86b, Ame87b, DRB89, LW84, Smi92, Hen92]. Descriptors [DRB89]. Design [Ano79b, AHU74, AU77, Bac86, BMM78, Blu02, Com84, Com87a, Ele98, Far88, Far90, Far93, Far97, Far91, FP79, FSK10, FH95, Gra00a, Gut04, HH07, HR75, HP94, Hol90a, Hol04, Hun81, Hwa84, Kru84, LMKQ89, LG84, LS89, MBKQ96, MNN05, MNNW15, Muc97, PH94, PH08, PH18, Pla93a, Pys80, Rei88a, Rub88, SKM19, Ski98, Ski08, SS02, Str94, Tan87, TW97, Whi88, Wie77, Wie83a, Wie83b, Wie87, Wir92, dB75, BH09, BGvN46, PH12, Swa76, CS91, Tau63b]. Designing [Cra71, Sta10]. Designs [Par00]. Desktop [Fla95, HM04, O’Q99, Rob99, RB92, X/O95a]. Destroy [Mit18]. destruction [O’N16]. Determinants [Mui60]. Deterrence [Mac92]. Deutschland [DW46]. Develop [SKM19]. Developer [The97, Nag97]. Developers [Dar01, McG98]. Development [Cow84, Hei66, JL78, JL87, Mac80, Tau77, Tau79, Tau76, Cle88, DW46, Swa76, Zus62]. Developments [Zap07a, Zap07b]. Deviations [Fey05, Smi14]. Device [SSU85]. Device-Independent [SSU85]. Devices [DS88b, KU81, Mau42]. Devil [Cor03]. Diagrams [BJ66, Knu09]. Dial [Gil93]. Dial-up [Gil93]. Diamond [Klo86]. Diatomic [Her50]. Dichroism [Thu80]. dichten [Hen92]. Dictators [Win15]. Dictionary [CNVH93, Gei75, Jae91, Jae01, KA98, Lan01, Mur77, RS91, Sim59, TM71, Wat00, WR98, Win98, Maz97, McK79, Ram69, Win03b]. died [Fin09]. Difference [Bic48, Pag07, Sen21, Swa01]. Differences [Gwe96]. different [And11]. Differential [DKM^{+}09, PP09, Str74, Tho08a]. Diffusion [Knu87]. digit [BLWW04]. Digital [AH89, BD83, BRR08, Cas79, Eks84, GBL^{+}98, GW77, GS93, Gre83, HH07, Has55, Isa14, Kar87, KK67, Knu99a, Knu87, Les97, MA91, Pri21, Ran82, Rub88, Sch00, STS03, SBH^{+}04, Tap96, Uli87, Dys12, Smi10, WK48, Gwe96, RH88, RH90, Sch03a]. Digitization [Bos99]. Dijkstra [FvGGM90]. dilemma [Chr97]. Dimension [LM66]. dimensional [DHG08]. Dinosaur [Mar15]. Dinosaurs [Mar15]. Dirac [Far09]. Direct [DER86]. Directions [Knu79]. Directory [Ano96b, Chexx, FA89]. disappearance [Mag09]. disappearing [Kea10]. disarmament [FGMvH14]. Disaster [Con20]. Disasters [Gla98]. Disc [NO92a]. Discipline [PDL04]. Discovered [Liv05, Hol08]. Discoverers [Boo83, Hos11]. Discoveries [SL98, Win11, Bar06b, Bre10, Fal11]. Discovering [BC06, Har92, PA06].

Discovery [BH17, Bro15, KD17, Pyn10, Tem91, BBG04]. Discrete [Knu03b]. discussion [BGvN46]. Disk [JNW08]. Disks [PCGK89]. Display [Inc88, KU81, Tuf83, Ado93, Hol90b]. Displaying [Gri86]. disruptor [De 19]. Dissertation [Gus12]. dissertations [Tur07]. Distributed [Wei80]. Distribution [DG87a]. Ditroff [AB89]. diversity [Pag07]. divided [Clo15]. Divider [Cor86]. Division [HF87]. Dmitrii [Gor04]. do [FL88, Maz10]. DocBook [WM99]. Doctoral [Gus12]. Doctrine [Eth10]. Document [Ado00, BC93, Dil93, Gra83a, Geh86, Geh87, GL88, KD93, KD95, KD99, Lam86, Lam94, Smi92, AT&T86d, Pos89]. Documentation [The97, CS99, KL93, Leh03, Stu84, Tri99, The98, Tri98]. Documents [AB89, GRM97, Shu94, WMB94, WMB99]. Doing [Wol88, Wol91]. Domain [Ste96]. Donald [Ruc15]. done [Dev08a, Gor15]. Don’t [Sto99]. doom [Alv97]. Dot [Jap83, Kha86, Knu87]. dots [Jap83]. Douglas [Fro03]. down [McG11, McK79]. Dr. [Goo04, Nah06]. Draft [Ame86a, Ame88a, Ame88b, ANS76, Ame87a, Ame90, Ame88c, Plu87, vN45]. Dragonfly [Bur98]. DRAM [JNW08]. Dramatic [Kea00]. Draw [GHLP03]. Drawing [Gur94a, Wir90, Ren87]. dread [Whe14]. Dream [Wal01]. Dreams [Pag88]. Driver [Bee87]. DTrace [MMG07]. Duden [SSStt97]. Duel [Ale10a]. Durchführung [Zus36]. DVI [Bee87]. Dynamic [JL96, Hen97]. Dynamics [Mac76, Tan93]. Dynasty [Ted03]. Dyson [Sch13].

Early [EI66, San07]. earth [Str13a, WB00, Ekm16, Wal03]. Earthquake [Win05]. East [Con05, Str95]. Easter [Sta92]. easy [FLS95, Fey97, Pau89]. Eats [Tru03]. Eccentric [Win08]. Eccentrics [Win14]. École [AH89].

Ecology [OALD97]. Economic [SS09]. Economy [SV99, Tap96]. Edgar [Jer02]. Edge [Win05, Gle10, Hol18]. Editing [Fin91, Plo82, SBA92]. Edition [BST^{+}96, Boy16, CLK99b, KN94, The92, NO90b, NO92b, Pog05, QO93b, QO93a, RD02a, SK19, You90, YP92, CK07, Lio96, Van99]. Editors [DCJ^{+}97, Plo82, DHJ^{+}12, RV09]. EDSAC [WR50, Wor50]. Edsger [FvGGM90]. EDVAC [Mau47, vN45]. EDVAC-type [Mau47]. Edward [BO76, Goo04, Har10, Her02]. Effective [Blo01, HM98, Mat19a, Rob96, Rob97, Rob01, Sea20, SKM19, Stu84, Zob97]. Effects [Mac76]. efficace [GHLP03]. Efficient [Ben82, KR81, Tro18]. Effiziente [Mer97b]. Effort [Tay25b, Tay25a]. EGA [Fer88]. Egg [Sto89]. Egil [MTW68a, MTW68b]. Egypt [KIR^{+}07]. Eigenfunctions [Pau79]. Eigensystem [GBDM77, SBD^{+}76]. Eigenvalue [Par80b, Par98, Wat07, Wil65]. Einbindung [Mer97b]. Einführung [Sch88, dG96, Kop88]. Einstein [Hen92, Acz99, Ari03, Ari05, BB05, Bro06, Cal04, CG14, Eis06, ECS19, FW07, Fey97, Fle18, Gal03, Gat09, GP19, Gor20, Hen97, Hol18, Imp19, Isa07, JL14, Jer02, Ken12, KN19a, Kum10, Lin08, Oha08, Pai94, Rig05, Rob05c, Sch49a, Sch49b, Sch08, Sen21, Sta98a, Ste99a, Ste01, Sto13, Wil86b, Woo80]. Einstein-Marić [ECS19]. Einstein-Turm [Hen92]. EISPACK [GBDM77, SBD^{+}76]. Elasticity [Hoh01]. Elections [HK05, Saa01]. Electricity [Duc60, Hut52]. Electrochemistry [Mac61]. Electrodynamics [Jac62]. Electromagnetic [Mit18, FM14]. Electron [WB01]. Electronic [Cop12, DG87a, FA89, Gia90, Gia92, Her66, Org91, RMV42, San64, Sla63, Whi88, Whi90a, WK48, BGvN46, ST52, GG46]. Electronic-Structure [Sla63]. electronics [Key88, Zus62]. Electrons [Ara06, Eck48].

Electrostatic [Cet91]. Elektronik [Zus62]. element [Ber07]. Elementary [Bak62, CW80, Mar00, Mor83, Mul97, Sti75, TD00]. Elements [Bro15, Dus38, Gla46, GL60, Hut52, KP74, KP78, NS21, Plo82, RA76, Rog85, She86, Sma23, SW79, SW99, SWK05, Kea10, Lat52]. Eleventh [Boy16]. Ellipses [Wir90]. Elliptic [BF71, Wil95a]. Emacs [CR91, Fin91, Gli97, RB92, SBA92]. emerged [McG11]. Empire [Bai99]. Empires [Gal03, Win15]. Enabling [Sun97, Sch12a]. Encapsulated [Ado89]. Encoding [The91, Cor89, Gil03]. Encounters [RK11]. Encryption [Ele98]. Encyclopædia [Boy16]. Encyclopedic [TM71]. end [FW07, Gra11, Gre05]. endless [Bus45]. Endurance [KD17]. Enduring [Lon16, Mor07]. enemies [Fag17]. enemy [Has05]. Energy [BO76, JT18, McP74, Rho18, Zoe09, Ber81, Cra07, Gat09, Gra11, Smi98a]. Engaging [GB18]. Engine [dB75, Bab37, Bab10, LM06, Swa01]. Engineering [ACHM82, BMM78, Bor94, Bro82, Bro95, CT04, CT11, Den66, Dev08b, FSK10, G^{+}99, GW98, Kre93, KN94, Mil84, RW04, SM07, SR66, Urr01a, Urr01b]. Engineers [Bue90, BF71, Ett93, HB91, Jen77, Ma09, MS88, Ros05, Win18, Ett87]. English [CNVH93, SSStt97, Win03b, Cry04, For12, Gor15, KA98, Lan01, Maz97, McK79, McW08, Mur77, Ram69, RR90, Sim59, TM71, Trz95, Wat00, WR98, Win98]. English/Chinese [Lan01]. English/English [SSStt97]. enhancement [Boi97]. ENIAC [GG46, HPR16, McC01]. Enigma [RK11, Sin97, Van89, DB04, McG11, Tur18]. Enrico [BB04b, SH16]. Enter [GB18]. Enterprise [Rei03, Sun97]. Entire [MW94]. Entities [Ara06]. entrepreneur [RH88, RH90]. Entries [Str86b]. Entropy [BN08]. Entwicklung [DW46, Zus62].

Entwicklungen [Zap07b]. Entwicklungslinien [Zus62]. Environment [BC84, Don89, GS73, KP84, Nes87, Ope91a, Ste92, SR05, X/O95a, Gra11]. Environments [BSS84, IEE86]. Envisioning [Tuf90]. Epic [Sin97, ET03]. Equation [Acz99, Liv05, Bod00]. Equations [Cre10, DKM^{+}09, DS83, Don89, Far03, Fle08, PP09, Str74, Ata40, Tho08a]. Equilibria [But64]. Equilibrium [And63, Den66]. Equipment [RH88, RH90, Sch03a, Zus62]. Era [BRR08, GS93, KS87]. Erdos [Sch98]. ère [Har04]. Ergodic [Tau61b]. Ernest [Her02, Hil15, Ree08]. Errata [Coo81b]. Error [Ald02, ZM08]. Errors [Knu88, Knu89b, WM92]. Erstellung [Mer97b]. Erwin [Hen92, Hen92, KN94]. Espionage [Mac18, Sto89]. Essais [Tor15]. Essays [Bro82, Bro95, Pla93a, Pla93b, Pla93c, Bro06, Pai94, Tor15]. Essential [BF05, BS03, Fri91, Gil93]. Essentials [Bal15a, Kit63, Rue07, FLS95, Tri98]. Étendue [Tor15]. Ethernet [Sei98]. Ettore [Mag09]. Etymologicon [For12]. Euclid [Mlo01]. Euler [Hav03, HHPM07, Nah06, San07, Ten09]. Eurographics [HHS89]. Europe [Ben21, Cah95, Dor14, Fro04, Mic05, Rod19]. European [Zla92]. EuroTEX [Zla92]. Eva [Gör11]. Evaluation [Gab85, Wu99]. Even [Ber06, Woi06]. events [Kle56]. Everest [Kea00]. Everyday [Mic12]. Everyone [Hah91]. Everything [Boy16, Gle99, Win03b, Blo08, Com00a]. Everywhere [Bar88]. Evidence [Rog11, Tuf97]. Evolution [BB97, EI66, Rog11, Sin99, Str94, Cou33, Fal11, Par03, Nic03, Cou33]. Evolved [Dev00]. exactly [LM66]. Examination [Org72]. Example [Bor92, Bor93, GB18, Qui02, Rob04, SK10]. Examples [BN71, BSN82, Dar01, SBN82, Weg80, CT98, SSU85]. excursions [Hol18]. execution [Zus36]. Executioners [Pad90]. Exercise [HWS^{+}88]. Exercises [Ove01, TG89]. Existence [Wil15]. Expanding [Acz99]. Expect [Lib94]. Expedition [Lev19]. Expeditions [GP19]. Experience [Boo58]. Experienced [Jae93]. experiences [Ano96a]. Experiment [BB04c]. Experimental [BBC^{+}07, DWB^{+}62, HL33, Hut52, BD09]. Experimentation [BBG04]. Experiments [ANTT02, Boh87, HM88, SG62]. Explainable [Boy16]. Explained [Boy16, Kor06, MR90, MR87, FLS95]. explains [Sen21]. Explanations [Tuf97]. Exploded [Win03a]. Exploration [Gil93]. Explore [Mat19b]. Explorers [Win14]. Explores [Dyl20]. Exploring [Cla94, Gud16, Hav03, Hey99, HEBS93, Lib94]. explosion [Sza84]. Express [Bai99, RS09a]. expression [Spe85]. Expressions [Han85]. Extended [Ame92, Ame90, AHG^{+}04]. Extending [Aus99, BC85, Mil16b]. extensible [GRA89]. Extension [GBDM77]. Extensions [Gli97, Gwe96]. extent [Tor15]. extinct [Fin09]. Extinction [New14, Kol14]. extinctions [War08]. Extraction [Wal87]. Extraordinary [Gib19, Pen11, Res18, Mag09, PA06]. Extreme [HF87]. eye [Par03]. eyewitnesses [Kel09].

F [GP04, Hen92]. Fabric [Gre04]. Fabulous [PL07, Nah06]. face [Sta98a]. Facilities [AT&T86f]. Fading [Win15]. fail [Chr97]. Failed [Gla99]. Failings [Oha08]. Failure [Ber06, Woi06]. Fall [Cah95, Smo07, dT08]. Fallacy [Ung87]. Fallen [Cok09]. Familiar [Bar80]. Family [App87, Bee87, Int91, LG84, MSSW94, Eml12]. famous [Bod00, Jer02]. Fantastic [Rei09, Win08]. FAQ [Ado97]. FAQs [CLG99]. Far [Str95]. Faraday [FM14, Hut52]. Farey [Gut11].

Farm [Ber96]. Fascicle [Knu05a, Knu05b, Knu08, Knu09]. Fast [Bla15, Cha18, KMP77, Mat19b, NF87, Sun90, WM92, BM77, Pax88, LNV89, Van92]. Faster [Gle99, WC00]. Fastest [Cop12]. Fate [Nic03]. Father [Ted03, Wat90]. Faust [Seg07]. Favorite [Ker81b, dT08]. FBI [BLM19]. Fear [Sch03b, Wea12]. Feather [Klo86]. Features [Cha18, Mat19b, Nor82, SK19]. Featuring [Sea94]. Fédérale [AH89]. feet [Mar14a]. fences [Ano96a]. Fermat [Dev08a, Sin97, Wil95a]. Fermi [BB04b, SH16]. Feynman [Ano05, BR93b, FLH85, FF95, Fey00, FBD05, Fey05, FL06, Gle92, Hey99, Kra11, Lei00, Res18, Syk94]. Fibonacci [Dev11, PL07, Vis00]. Field [Bai92, Bal62, LB14, FM14]. Fields [Ben33, Ben35, Cet91, Kho02, RH11]. Fiery [Ted03]. Fifth [SM07, Ung87]. Fight [Moo16]. Figure [Pry58a, Pry58b, Com64b]. Figures [Wel05]. File [CC88, Bor95, KL92, KL95, Man94, MW94, Nag97, Wie83b, Wie87, da 87, Jer02]. Files [BLM19, Man94, Mia99, dT08, Ado89]. Filing [Gol98]. filling [Gri86]. Filtering [KU81, NF87]. Final [Ado92, Moo16]. Find [SRR97, Roc06]. Finder [RUL78]. Finding [Man94, Fey00]. findphrases [AB89]. finely [LBS17]. finely-tuned [LBS17]. finite [Wil95b]. finte [Kle56]. Fire [RMV42, RH97]. Firewalls [CB94, CBR03]. firms [Chr97, Pag07]. First [Bai99, C^{+}89, CGG^{+}92, Grä99, GREC90, Knu70, McC01, NS21, Swa01, vN45, Eml12, Fag10, Far13b, Rud07, Bag10, Stu18]. fish [Kur98, Shu08a]. fissile [FGMvH14]. five [Sta98a]. flawed [Smi14]. flex [Pax88]. Floating [Cli90, CCG^{+}84, Cod81, Coo80, Coo81b, Hou81, IEE85b, Mot85c, Mal72, Mot85d, MBdD^{+}10, Ove01, RJR88, SW90, Ste81, Cor86]. Floating-Point [CCG^{+}84, Cod81, IEE85b, Mot85c, Mal72, Mot85d, MBdD^{+}10, RJR88, SW90, Ste81, Cor86].

Flow [BJ66]. fly [Cat04]. Fock [Fro77]. Focuses [Gwe96]. Font [Inc90a, Inc90b, LNN95, LNN00]. Fontes [Har04]. Fonts [Hoe98, Mer91, Par88, WH76, Ash90, Knu90b]. Fooled [Tal01]. Force [Mit18, Ree08]. forces [CW14, Ree00, Sch12b]. Foreign [von52]. Forgotten [Aus82]. Form [LW84]. Formal [Pri55]. Format [Inc90a, Ado00, Inc90b, CC88, BC93, CS99]. Formation [BJ66]. Formats [Bor95, Kar87, KL92, KL95]. Formatting [Geh86, Geh87, GL88]. Former [McP98]. formula [Nah06, Lor19]. Formulae [Ben33, Ben35, GO70]. Formulary [Ben33, Ben35]. Formulas [AS64, Kro65]. Fortgeschrittene [App88]. FORTLEX [Fel76]. Fortran [Ame66, Ame78, Ame87a, Ame92, Kau78, Bac78, Ett87, Her69, Her70, Her72b, HB91, PFTV86, ANS69, ANS71, ANS76, Ame90, ABW92, ABM^{+}97, AO90, BR89, Bel90, BGA90, CCN^{+}79, Cou91, Don89, Ell90, Ett93, Fel76, Hal69, HWS^{+}88, Hei66, Kat78, Ker93, Lor19, MS88, Met82, MR90, MR87, Sch89b, Tho97, TW88]. Fortran-to-Pcode [CCN^{+}79]. fortunate [LBS17]. fossil [Fal11]. Fossils [Mar15, Tho08b]. Found [Bar09, Eis06, Kra14]. Foundations [BHK20, Llo87, RS87, Sal05b, McK79]. Founding [Ted03]. Four [Com64b, Wil02, Bro06]. Four-Figure [Com64b]. Fourier [Van92, Her74a, Hut52, Tol62]. Fractal [Lap08, Man83, PS88]. Fractals [Bar88, PR86, PJS91, Sch91]. Fractions [BvdPSZ14, CPV^{+}08]. framework [X/O94a, AHG^{+}04]. Frameworks [Van92]. France [Ros19]. Francisco [USE94]. Fraud [Rei09]. Free [Pry66]. FreeBSD [JL16, Leh03, LJ15, MNN05, MNNW15]. Freeman [Sch13]. French [Cou33, Cou38, RS91, Tor15]. Frenzy [Rin02]. frequently [Ado97]. Freundlich [Hen92]. Friar [GG05]. fridge [Sen21]. Friedrich [Ten06]. Friendly [Bor94, Hen92]. Friends [Ker05]. friendship [BB05]. Frontier [Ree08, Bus45]. Frontiers [Kun85b, MSWA86]. Fuchs [Gre20]. Full [Pad90]. Full-scale [Pad90]. fun [Ash90, Knu90b]. Function [Bee17]. Functional [Ame87b, Joh18, Pey87]. Functions [AS64, AAR99, CW80, CPV^{+}08, FHCD92, Hol87a, Knu08, Luk69a, Luk69b, Luk75, Luk77, Mar00, Mos89, Mul97, Olv74, OLBC10, Tau61b, ZJ96, DHG08, Leb72, Tho97]. Fundamental [Ear85, Knu97, Knu68, Knu73a, Swa65]. Fundamentals [Fv82, HS83, HS78, JW76, Ma09, Pav99, Rob04, Sed98]. Further [Ban99, FL88]. fusion [Sei08, Tau93]. Future [Bha22, Mac11, Ree18, Smo14, Ung87, GC02, Mah09, War08].

G. [AAD16]. Gadsby [Wri39]. Gaiji

[Sea94]. Galactic [GS73]. gambit [Shu08b]. gambler [Maz10]. Gambling [EE07, Eth10]. Game [Mac92, Dev08a]. Games [EE07, Tau63c]. Gamma [Hav03]. Gamow [Seg11, Sta99]. Garbage [JL96]. Gauss [Ten06]. gave [AK11, Fag10]. GAWK [GAW87]. GCA [Gra83a]. GDB [Rob05a]. Geeks [Isa14]. Gems [Gla90, Kir92, Pae95, PF05, Arv91, Fer04, Hec94, Ngu08]. GenCode [Gra83a]. Gene [Dev00, CGO07]. General [Fel76, LEMW52, PF05, SK10, BL63, Wil86b]. General-purpose [PF05, SK10]. Generalized [Bry88, Fee88, FGP93, Gra83a, Smi87a, Smi87b]. Generals [Kap20]. Generating [AB89, Knu05b, TW88]. Generation [ACHM82, Gra00a, Hür22, Sch83, SS02, Ung87, Hol08, Kir05, Zac14]. Generator [DS88a, Gra88, LS79, Pax88]. Generic [Aus99]. Genetic [Kea12, Syk07]. Genius [Gle92, Kan91, Kea12, Lin04, Liv05, Mac92, Oha08, Ree08, Sch13, Sei96, Sob95, Syk94, Tem91, Mag09, Sch08, Wei16].

Geniuses [Isa14, Pri21, Seg11]. genomes [Pää14]. genomics [Seg11]. geography [Wei16]. Geological [Sny82]. Geology [Mac11, Win01, Gud16]. Geometric [BBB87, Cha78, Far88, Far90, Far93, Far97, LS89, Mor85, Sal99]. Geometry [Cas05, Dra12, Far95, Far99, FP79, FZ61, Man83, Mlo01, Oak49, Oak58, PP86, PS85, Tau62]. George [Can01, Car10, Seg11, Sta99]. German [ABR02, DW46, Ger20, Hen92, RB06a, Ros14, SSStt97, Ste99a, Ste01, Zap07b, Zus36, Zus62]. Germans [Ros19]. Germany [Mic89, Bla02, DW46, Ric06, SO11, Wei18]. Ghost [Hoc98, Hoc99]. Giant [Aus82]. GIF [Mia99]. Gifted [LB14]. Gigabit [Sei98]. Gigabytes [WMB94, WMB99]. GIMP [Bus99, KK99]. Girls [Moo16, Mun17, Kie13]. Gives [Coh10]. GKS [Ame85, EKP84, EKP87, HGDS83, HGDS86]. GLA [Gra88]. Glass [Boh87, Dys10]. GLIM [FGP93]. GLIMPSE [MW94]. Global [Wal03, Gor15, War08]. Glossary [Hub84]. Glue [Van86]. Glyphes [Har04]. GNU [CR91, CS99, Gli97, Mec05, Rob96, SBA92, Sta88]. Go [Dij68, Knu74, Bal18, Bal19, DK16]. God [Acz99, Ghi05, Haw05]. Gödel [FW07, Gol05, Hol18]. Godfather [Ran77]. Goldbach [Dox00]. Golden [HF98, Liv02, Tho08b]. Goldfarb [GP04]. Golub [CGO07]. Good [BR93b, Gar95, Rho15]. Goodbye [Kir05]. Google [LM06]. Gossip [Dev00]. got [Maz10]. Goudsmit [vCH18]. Gourmet [Spi86, Spi90]. GPGPU [Tho08a]. GPS [Mil16a]. GPU [Fer04, Ngu08, PF05, SK10]. GPUs [Tho08a]. GR [Wat07]. Graduate [Str74]. Graduated [Weg80]. Graham [Shu08b]. Grammar [She86]. Grammars [CU77]. Grand [Ash90, Cla94, Knu89a, Knu90b]. grap [GL88, BK86]. Graph [Sed02]. Graphic [Whi88, ACM77, ACM79]. Graphical [Ame85, Ame88c, HGDS83, HGDS86, Joh96, Pav99]. Graphics [Ame85, Ame86b, Ame88c, Ang81, Art84, Art85, Arv91, BBB87, BC84, Bee89, Cha78, Che02, Con85, Ear85, EKP84, EKP87, Fiu89, Fv82, FvFH90, GI886, Gil78, Gla90, GRM97, GMR^{+}08, Gwe96, Har83, Har87, Hec94, Hoe98, Hub84, Jap78, KL92, KL95, KPv86, Kir92, Kun83, Kun85a, Kun85b, Kun86, NS73, NS79, Pae95, Pav82, PP86, PF05, Ran89, Rim93, RA76, Rog85, RE87, Sal99, Sal06a, Sal06b, Sal11, Sco82, SSU85, Whi83, WH76, BR89, Bus99, Fer04, Ker82]. Graphs [AS64, BK86]. gravitation [Ber68]. Gravity [Eis06]. Great [Acz11, Bel37, Bus99, Con20, Cre10, Far03, Kea00, Pyn10, Rue07, SL98, Tur93, Wal03, Whi90b, Win05, Win11, Chr97, Kum10]. Greatest [Bar06a, Dev02, Hür22, Mac18, Sab03, Sin97, Sob95, Der03, dS03]. greatness [Rig05]. Greece [SM07, SBH^{+}04]. Greed [Hoc98, Hoc99]. Greek [Wat00]. Greely [Lev19]. green [War08, Can01]. Group [ACP70, Cot63, Isa14, SBH^{+}04, Tau61b, Tay25b, Tay25a, Tin64, Cat04, SW72]. Groups [BL66, Pag07]. Grove [Ted06]. Guerrillas [Has16]. Guide [Ame01a, Ame93, ABB^{+}92, ABB^{+}95, ABB^{+}99, AWH13, AS90, The97, Bee80, Bee79, Bha96, Blo01, BGA90, Bry88, Cam87, Che02, Chr86, Coo80, Coo81b, DCJ^{+}97, DBMS79, Eng87, Far93, Far97, Far01, Fer88, Fla97, Fle08, Fly95, FC76, Fly73, FF78, For96, GBDM77, Gil93, HHR95, Hol89, Int99, Int83b, Keh93, KD93, KD95, KD99, KD04, Kro93, Kyk76, Lam86, Lam94, LL97, LNN95, LNN00, Mia99, Mic12, MGH^{+}96, MGH^{+}03a, MGH^{+}03b, MGH80, MP92, MK96, Nag97, NDW93, Nor85, OQL88, Ope91b, Ope91d, Pin14, Plo82, QO90, QO93b, QO93a, Rob96, RR96, Rob05a, RB08, RB92, Rod19, Sal02, SV99, SKM19, SBD^{+}76, SS88, Spi86, Spi90, Tat87, TY82, TH01, Tri99, Trz95, WM99, Wol92, Zim95, Zlo91, X/O88c, X/O88a, X/O88d, X/O88b, X/O89c, X/O89d, X/O89b, DHJ^{+}12, Dup98]. guide [Gil03, Ken12, Lew91, Nor94, Pau89, RV09, Tho97, Tri98, X/O89a, X/O94b, AT&T86h]. Guidebook [Tro06a, Tro06b]. Guided [Mit11]. Guidelines [Gol98, Smi87b]. Gutenberg [GS93, Kir05].

H [AAD16, CGO07, Mai18, Mur77, Kle06]. Hacker [CB94, CBR03, War03, War13]. Hackers [Isa14]. Hadron [Acz10]. Hahn [Hah67, She83]. Half [Clo15]. Half-life [Clo15]. Halftones [BR93a, Knu87]. Halftoning [Uli87]. Hall [Ber96]. Hand [Ess04]. Hand-Loom [Ess04]. Handbook [AS64, ABW92, ABM^{+}97, ABO06, Bee17, Bey78, Bor95, BED96, BSMM07, BF71, CV88, Con85, CPV^{+}08, Cor92, FB94, Fly95, Gib96, Gib99, GR90, GP04, Gon84, GS93, Hig93, HH99, Int99, KK67, Mod09, MBdD^{+}10, NSS89, NSS^{+}01, NSH02, NSHW10, OLBC10, PPB98, RW04, RY01, Red93, Red96, RS09b, SS02, WBA85, Zwi92, BL19, Ker10, You97, vL92]. Handling [Fly98]. Hands [HK05, Joh18, KmWH10, Kol19, Mat19a, Tro18]. Hands-on [HK05, Joh18, KmWH10, Kol19, Mat19a, Tro18]. Hans [Ber81, Sch12b]. Hansen [Bri85]. happened [Mar14a, Rud07]. Hardcopy [DS88b]. Hardware [AIM95, BMM78, Fer86, HP94, HP97, HP04, HJ76, Hsu01, Int83a, Int87, Int90, Par00, PH94, PH97, PH04, PH08, PH18, TT00, TZ86, PH12]. Hardware/Software [HP94, HP97, HP04, PH94, PH97, PH04, PH08, TZ86, PH12]. Hardy [AAD16]. harmful [Dij68]. Hartley [LNV89]. Hartree [Fro77, Fro03]. Hash [FHCD92]. Hawking [Sus08]. Headers [X/O89d]. heartbeat [Ber11]. Heat [Hut52]. heaven [Sob12]. heavens [Wul12]. Hecke [TW95]. Hedwig [BB05].

Heisenberg [Cas09, Cre10, CG14, Lin08]. held [Erd74, MA91, SBH^{+}04]. Helgoland [Rov21]. Hell [Rho15]. hello [Kir05]. Help [ST12]. Helped [Tur20, Kie13]. her [Eml12]. here [Pai94]. Heretic [Sto99]. Herman [Aus82]. Heroes [Ale10a, Ari03, Ari05]. Heroic [Cah95, Win11]. heroine [Fag17]. Heroism [Hoc98, Hoc99]. Herschel [Hos11]. Hershey [WH76]. Hidden [Ald02, Far09, For12, Gre11, Tal01, Des10, Far13a, Roc06]. Hierarchical [Ame88c]. Higgs [Kra14]. High [Dow93, Egu18, FJSD96, HJ04, KU81, PF05, Sei98, Sou99, Sto99, Van82, WC00, BLWW04, JR13, Mau42, Tho08a]. high-accuracy [BLWW04]. High-Level [Van82]. High-Performance [FJSD96, PF05, JR13]. High-Speed [Sei98]. higher [BL63, DHG08]. higher-dimensional [DHG08]. highest [Kea18]. Hilfe [Zus36]. Himmler [Pad90]. historians [Kel09]. historic [Cre11]. Historical [Rog01, Hen92]. Histories [WB01]. History [Aug84, Bac78, Bag10, Bal15b, BG96, Bur07, CKCFR03, Cla94, Cle03, Coh10, Cok09, Dyl20, Eis06, Eve83, Fag19, HC21, Hei66, HF87, HF98, Ifr00, Kap99, Kap20, Ker19, Kev79, Leh98, Lor19, Mac80, Mao91, Mao07, Nas90, Qua18, RS09a, RH11, RS09b, WP85, Wil97, vLPO23, Bag11, Bec93, Ber07, CBB^{+}99, Des10, Far13a, GS12, Goo91, Gut11, Har15, Haw05, Kea10, Kol14, Kur99, Kur03, Kur17, Mah09, Mal11, Maz10, McW08, Mor07, New07, Ree14a, Rho18, Sch03c, Sei08, Sho14, Shu08a, Smi98a]. Hitler [Bal14, Ber96, Cor03, Gör11, Ker05, O’R08, Pad90, vCH18]. Hoax [Sok00]. Hole [Fle18, Sus08, Lev16]. Holes [Imp19]. Hollerith [Aus82]. Holocaust [Bla02]. Holub [Hol04]. Homo [BH17]. honor [Kea18]. Hoover [Jer02]. hot [Sen21]. House [BN06, AK11, Gei75]. HTML [Fly95, GR99, Gra00a, MK96]. HTTP [Ste96]. Hubble [Zim08]. Human [BH17, Gri05, Kur12, Oha08, Wil15, Fal11, Pin02, Pin03, Rho18, Shu08a]. Humanity [Ree18]. humankind [Har15]. Humans [Fin09, New14, Bic09, Fag10, Str13a]. Humboldt [Hel04]. Hundred [Ove01, Rob05c]. hunt [Gat09, vCH18]. hunted [McG11]. Hurts [ST12]. Hydrodynamics [Tau63c]. Hydrogen [Rho95, Sin05]. Hylleraas [MTW68a, MTW68b]. Hyperspace [Mlo01]. Hypothesis [Sab03, MS15, Roc06].

I/O [Har99]. i860 [Int90, Int91, Mar90]. IA [Mar00]. IA-64 [Mar00]. iAPX [Int83a, Int85a, Int85b, Int83b]. IBM [Bel90, Bla02, Bra84, Car94, CL88, Cor19, DeL86, DHJ^{+}12, Hos91, Mis90, Laf84, Mor84, Nor82, Nor85, SS84, ST52, SSU85, Ted03, Wat90, WR94a, WR94b]. ICCMSE [SM07]. Ice [Fag19, Lev19, Pol10, Fag10]. Iceland [Gud16]. Icon [GG83, GG86, GG90]. idea [Ros10]. Ideas [Ein54, Ein82, MZP97]. Identification [FPZ^{+}64, SFC64]. IEEE [BCDH09, SO11, Ame83b, Hou81, Ove01, Cor86]. IEEE-754 [Hou81]. If [GREC91a, Bor94]. IFIP [Boi97, Gri78]. ignited [Gre05]. Ihre [RB06a, Mer97b]. II [AU73, Arv91, Bac78, BG96, Con20, Erd74, FC76, Her45, Ker05, Kie13, MA91, Mun17, Pla93b, Röh14, Ser18, Sha99b, Tau61b, Tau79, WR94a]. III [Bac78, Her66, Kir92, Pla93c, Tau63a]. ills [Nah06]. Illuminating [TW87]. illusion [Maz10]. Illustrate [Gus17]. Illustrated [Aug84, Mac76, SA98, Ste94b, Ste96, Sto07, Syk94, WS95, Tho97]. Illustrating [GRM97]. Illustrations [Cas05]. Illustré [Lar06]. Image [CC88, Cas79, DG87b, Eks84, GW77, Gre83, Mia99, Pav82, PM92, Sch83]. Images [Gri81, PS88, Tuf97, WMB94, WMB99, FF95]. Imaginary [Nah98]. Imaging [AH89, MA91]. imagining [Ari03, Ari05]. Impatient [ABH90, AL92, AL97]. imperfect [Gle10]. Implementation [Blu02, CS91, Coo80, Coo81b, FH95, GG86, LNV89, LMKQ89, MBKQ96, MNN05, MNNW15, Muc97, Mul97, Omo94, Pey87, Tan87, TW97, WS94, WS95]. Implementations [Bha96, Wei80]. Implementing [SR06]. Impossibility [Bar98]. Impress [GHLP03]. Inc. [CW14]. Including [Ame70, DS51, Gol98, Ove01]. Incompleteness [Gol05]. increases [O’N16]. independence [Gra11]. Independent [GHHH82, GSS68, KM84, Ram89, SSU85, CCG^{+}84, Ker81a]. Index [Che94, CH88, CH87]. Indexes [AB89]. Indexing [Ame01a, WMB94, WMB99]. India [KIR^{+}07, Kea00]. Individual [Gil93]. Indivisible [Win14]. Industrial [Hil15, Jap78, Jap83, Osb79]. Industry [Ben33, Ben35, Ros10]. indx [AB89]. inequality [O’N16]. Inexpensive [PCGK89]. Infinite [Mao91, Sma23, Str19, Wil95b]. Infinity [Cle03, Kan91, Mao91, Rei06]. influential [Har10]. Informatik [ABR02]. Information [Ame85, Ame86b, Ame88c, Aus82, Ess04, FBY92, Ger00, Jap78, Lun93, Lun99, RH97, SV99, Tuf83, Tuf90, Ame70, SG17]. Infrared [Her45, Nak62]. Infrastructure [MJW^{+}06, X/O95a]. Ingenuity [Bro15]. inglese [Maz97]. inglese-italiano [Maz97]. Ingredients [Pin99]. Initial [ANS69]. inner [Shu08a]. Innocence [Stu18, FW07]. innovations [Tes07]. innovator [Chr97]. Innovators [Isa14]. Inorganic [BN60, BL63, CW66, Dor65, KS52]. Inquiry [Gib19]. ins [Mer97b]. Insanity [Win98]. inspiration [CW14]. Installation [Bee79]. Instinct [Pin94]. Institute [Goo91]. institutional [Hen92]. institutionellen [Hen92]. Instruction [Sta90, Tab87]. Instructor [KCM^{+}76]. instrument [BGvN46]. Instrumental [SW80, WMD65]. Instrumente [DW46]. instruments [DW46]. integers [Haw05]. Integrals [BF71]. Integrating [BT98, GR99]. Integration [Zwi92, DR75]. Integrator [GG46]. Intel [Org83, Col05, JR13]. Intelligence [Jon78, Tap96, Tur20, Ung87, Jac14, Ric06]. Intended [ST12]. Intensive [Bel90]. interacting [LM66]. Interactive [Ame88c, BSS84, BC84, Fv82, FGP93, Gil78, HR75, Lib94, NS73, NS79, Sco82]. Interchange [Jap78]. Interesting [Rei06]. Interface [GI886, HP94, HP97, HP04, PH94, PH97, PH04, PH08, PH18, X/O93c, X/O89d, Ker10, PH12, Ren87]. Interfaces [IEE85a, Pav99, Ram97]. Intermediate [TD00]. Internals [CS91, MM01, Nag97, Vah96, Wea08, MM07, Sin07]. International [Ano96b, Chexx, SM07, SBH^{+}04, Cat04, AH89, Leh98]. internationalisation [X/O94b]. Internet [CB94, Est93, Gil93, CBR03, Com00a, GS96, HL96, Keh93, Kro93, Mil92, SRR97, Sta98b]. Internetworking [Com87a, Com88, CS91, Com00b, Com13]. Interpolation [Her72c]. Interpolator [Ces46]. Interpreter [Bal18]. Interpreters [Nys21]. Interprocess [Ste99b]. intertexture [Hen97]. Interview [Ano05]. Interviews [AA85, AA08]. Intimate [Cok09]. Intranet [McC98]. Intrinsics [NO90a, NO90b, NO92b, Fla92, NO93, O’R90]. Introducing [MM83]. Introduction [Ang81, App87, Bak62, Bal62, BBB87, Bel74, BvdPSZ14, Bur07, CGG^{+}92, CLR90, CLR01, CLRS09, Ell90, Eve83, FJSD96, Gla89, Gor81, Hec93, Hec96, HGDS83, HGDS86, Int85b, Jae93, Jon89, Kau57, Kho02, Knu08, KS90, Mom01, PCGK89, PW35, PS85, Pry66, RB05b, Rog01, Ros88, Rub88, Sal08, SK10, Say96, Say00, Say05, Sch84, SF85, Sco82, Ski85, Ste73, Sto07, Weg80, You83, ABR02, BH09, BH15, BD09, C^{+}89, GRA89, Grä96, Sum09]. Introductory [AvD19, Kit63, La 71, MGH^{+}03b, Pen65, Tay63]. invented [Smi10, SG17]. Invention [Ber05, Ifr00, Lin04, Tem91, Ros10]. Inventor [Mai18]. Inventors [Win14]. investigating [Ekm16]. Investigation [Boo54]. involving [Ame70]. Ionic [But64]. IP [Com88, CS91, Com00b, HJ04, Hun92, Mil92, Ste94b, Ste96, WS95]. Ipm [DW46, DW46]. Iran [Ric06]. Ireland [Cah95, Syk07]. Irish [Cah95, Mal13]. Irishmen [O’R08]. Irregulars [Con09]. Isaac [Bre10, Dol11, Dry14, Gle03]. Islam [KIR^{+}07]. Isn’t [Knu90a]. ISO [ABW92, ABM^{+}97, Int82, JW85, SS88]. ISO/ANSI [ABM^{+}97]. Israel [SS09]. issue [X/O95b]. Issues [HJ04, Smi87a, Bes04]. Italian [Maz97]. italiano [Maz97]. italiano-inglese [Maz97]. Itanium [CHT02, ET03]. Itanium-based [CHT02]. Iterative [HY81, You71]. IV [Hal69, Hec94, Her69, Tau62].

J. [BS05, Jer02, RH11, Wal01]. Jacquard [Ess04]. jaguar [GM97]. James [Mur77]. January [USE94]. Japan [AFI75, Ung87, AFI75, Has05]. Japanese [Jap78, Jap78, Jap83, Lun93, Lun99, Rei88b, Van89]. Java [AGH00, Blo01, CL98, CLK99a, CLK99b, Dar01, Fla95, GJSB00, Har99, LY97, LY99, Sun97, WN96]. java.applet [CL98]. java.awt [CL98]. java.beans [CL98]. java.io [CLK99a]. java.lang [CLK99a]. java.math [CLK99a]. java.net [CLK99a]. java.text [CLK99a]. java.util [CLK99a]. JavaScript [Fla97]. Jean [Hut52]. jewel [Geh03]. Jiao [RB08]. JIS [Jap78, Jap83, Sea94]. JIS-Kanji [Sea94]. Jobs [ZM08, RB06a]. John [AB87, Bha22, Lor18, Mac92, Smi10, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. jointly [SBH^{+}04]. Joking [FLH85]. Jonathan [Ber06]. Joseph [Hut52]. Journal [Bal15b, Ame70, Woo]. Journalists [Plo82]. journals [BB04a]. Journey [FR13, FF01, Hel04, Lei00, Shu08a]. Journeys [Dyl20, Sch98]. Joy [Bla97, Spi86, Spi90]. JPEG [Mia99, PM92]. JSTOR [Sch03c]. Judging [Har10]. July [Boi97, SO11, Mar14b, Sza84]. June [BCDH09, Erd74]. Jung [Mil10]. Jungles [Pet98]. Just [Gle99, Ree00]. Justice [Moo16, ZM08].

Kaiser [Ben21, Röh14]. Kanji [Sea94]. keep [MZP97]. Kelvin [Lin04]. Ken [RH88, RH90]. kept [Mar14a]. Kermit [Dou90, Gia90, Gia92, da 87, dG93, dG96, dG97]. Kernel [Ame85, BCDW07, HWS^{+}88, HGDS83, HGDS86, MM01, MM07]. killed [Bro10]. Kinetics [Lai65]. King [Hoc98, Hoc99]. Kingdom [Boi97, Win08]. Kings [Cla09]. Kingston [Erd74]. Klassische [RB06a]. Klaus [Gre20]. Knew [Kan91]. Knott [Com64b]. Know [Wal03, Com00a]. knowledge [AK11]. Knuth [Ruc15]. Korea [Ric06]. Korean [Lun99]. Korn [Ros93, RR02]. Krakatoa [Win03a]. Kreyszig [KN94]. Krommes [Kro90]. Krylov [Wat07]. Kurt [Gol05].

Laboratories [BF91, Alt48]. Laboratory [PW94, SFC64, Cro74]. Labs [Fat92, Geh03]. Labyrinth [Lev19]. Lagally [Her74b]. laid [McK79]. LAMS-TEX [Spi89]. Landmark [Col05]. LANG [HR75]. LANG-PAK [HR75]. Langenscheidt [KA98]. Language [Ame83a, Ame86a, Ame88a, Ame88b, Ame89, Ame92, Ame90, Ame83b, Ano79b, Ado85a, Ado85b, Ado90, Ado99, AKW88, AGH00, Bac79, BC81, BCW88, BK86, Bla15, Blo01, BS03, Bra84, CGG^{+}91b, Cle88, DK16, Dyb96, Fee88, For12, Fox87, Gra83a, Gor81, GJSB00, Gou86, Gra00a, GG83, GG86, GG90, HR75, Int82, Jae89, JL78, JL87, KR78, Ker81b, KR88, KN17, Laf84, Law78, Liv05, Lor19, Pag82, Pag83, Pin94, Pin99, Pyl81, Ram89, Rei88a, RS91, Sca84, Sco00, Sco06, Ski85, Ste84, Ste90a, Sti75, Str86c, Str91, Str97, TM71, Van82, Van99, Wal87, Woo, Yeu84, AKW23, Ale10b, Ari03, Ari05, Bic09, Cow90, Dyb87, Dyb03, HWG04, JE06, Ker82, KN19b, McK79, Mit06, Pag11, Str13b, Sum09, BED96, Bry88, Smi87a, Smi87b, TG89]. Language-Independent [Ram89]. Languages [Ben86b, BG96, BJ66, CU77, Com87b, DRB89, Gil78, Knu03a, Pey87, Sch90, Smi92, Tuc77, von52, X/O88c, Dor14, Dor18, X/O89a, AT&T86e]. LANs [Sei98]. LAPACK [ABB^{+}92, ABB^{+}95, ABB^{+}99]. Large [BCDW07, FHCD92, GL81, HWS^{+}88, Man94, You71, Ata40, Ran01, Acz10, Par84]. Large-Scale [HWS^{+}88, BCDW07, Par84]. Laser [Mac76, Mai18]. Last [Lei00, Ser18, Wil95a]. lasting [Sch03a]. Late [HL96]. Latest [Mat19b]. LaTEX2e [Grä96, KD95]. Latin [Hel04, Sim59]. Launched [Con20]. Laurent [Hut52]. Lausanne [AH89]. Lavoisier [Hut52]. Law [Fey94, Par80a, Woi06, Roc06, BN08, BH15, Mil15]. Lawrence [Her02, Hil15]. Laws [Ber06, Boo54, Gre11, Sch91]. layman [Pai94]. LCAS [Kah92]. leading [Bro06]. Learn [Egu18, Mat19a, SK19]. Learned [Gla99]. Learning [CR91, Gno84, HHR95, Hol04, NDW93, Ros93, RR02, Smi90, TSP93]. Least [LH74]. Leaves [CGG^{+}92, Tru03, C^{+}89]. Led [Ess04]. Legacy [BB04b, Lon16, Ted03, Bro06, Mor07, Sch03a, Tho08b]. Leibniz [Bar06a]. Leland [De 19]. length [CCG^{+}84, Sal07b]. Leonhard [HHPM07, San07, Ten09]. Leopold [Hoc98, Hoc99]. Lessons [Gla99]. letter [Dev08a, Wri39]. Letters [Fey05, Fir93, BB05]. leur [Cou33]. leurs [Cou33]. Level [Van82]. Leveraging [Cha18]. LEX [Mor87a, LS79, LMB92, MB90]. Lexical [Fel76, Gra88, LS79, Pax88]. Liars [Sch12a]. Libel [Gol98]. liberal [Sho14]. Libraries [CL98, CLK99a, CLK99b, Les97]. Library [AHG^{+}04, Aus99, Bee17, CGG^{+}91a, Phi86, Pla92, SGN88, BB04a, AT&T86g, AHG^{+}04, Bud98, Jos99, PSLM00]. licenses [BB04a]. Licensing [Ma09]. Licklider [Wal01]. lie [Smi14]. Lies [MH22, Sch00, Bes04]. Life [Bha22, Coh10, Dys10, Far09, Gle92, Imp19, Kan91, Kra11, LR89, Mon13, Qua18, Res18, RH11, Sei96, Tal01, Wal03, Wat90, Ber11, BO76, Bre10, Can01, Clo15, De 19, Fro03, Geh03, Gle10, Gör11, Gre05, Isa07, LBS17, Mag09, Röh14, Tau93, Ted06, WB00]. Lifetime [KD17]. Ligand [Bal62]. Light [Ros14]. Like [Dev00, Hol87b, GREC91a]. Limits [Hey99, Bar98, Key88]. Line [Dil93, BR89, Ren87, Sta98b]. line-drawing [Ren87]. Linear [DDSv91, Don89, FGP93, Joh93, Thu80, WR71, You71, Ata40]. Linearity [Sin05]. Lines [Mlo01, Wir90, Zus62]. Lingo [Dor14]. LINPACK [DBMS79]. Linux [RB06a, Ker10, Cor05, Lip99, NSH02, NSHW10, Rob04]. Linux-Tasks [RB06a]. Lions’ [Lio96]. Lisp [Gno84, All78, Kes88, Tat87, Bro85a, Gab85, Ste84, Ste90a]. LISPer [FF87]. List [GREC91a]. Literate [AO90, Gur94a, Hym90, Knu84b, Knu92, Sew89]. Literature [AH13, Bar80]. Little [Ben86b, FF87, Fag19]. live [Sch03a]. lived [Pai94]. Lives [Acz11, Bat88, Bel37, Gre20, Mar15, Moo16, SL98, ZM08, Eml12, Her02]. Livre [Ser89]. Logic [CM85, Llo87, Pri55, Tau61a]. logical [BGvN46]. London [Ber06]. Londonderry [Ker05]. Lone [Sob95, Str13a]. long [Sch03a]. Longitude [Sob95, SA98]. Look [GP19, Ghi05, Har10, Mah09, Mil90, YP92]. Looking [Hol04]. Looks [Saa01]. Loom [Ess04]. Lord [Ker05]. Losing [Kea18]. Lost [Eis06, Kea12, Mor07, Pää14]. Love [AH13, Kea12, CG14, Fag17, Kea10]. Loved [Win08]. loyalties [Her02]. luck [Maz10]. Lucy [JE06]. Lunch [Ben21].

m1 [Ben90]. Ma [Kle06]. Mac [BF05, Pog05, Sin07]. Machine [Alt48, Blu02, Bus40, GHHH82, Hil85, LY97, LY99, Lor19, SS02, Sto07, Wal01, AOB64, Ata40, Lud09]. Machine-Independent [GHHH82]. Machines [BJ66, BCDW07, Cra05, SN05, BH09, Cou33, Cou38, Mau47, Zus36]. Macintosh [C^{+}89, App87]. Macro [Ben90, Col76, Spi86, Spi90]. Madame [Des10]. Made [Bal97, Fag19, Rho15, Wal01, Bic09, Dev08a, Nic87]. Madison [Par84]. Madman [Win98]. madness [Kea10]. magic [CL88]. Magma [BC06]. Magnet [Mit18]. Magnetic [Mac76, Rob59, Thu80]. Magnetism [Duc60]. magnificent [McW08]. Magnon [Fag10]. Mail [DG87a, FA89]. Maiman [Mai18]. Maintainable [Mat19a]. Major [Com87b, Ken12]. Majorana [Mag09]. Make [Mec05, OT91, Sus08, Tal86, Tal88, BL19]. Makes [Rei06]. Making [Bal15b, Ben33, Ben35, Dou90, Ker05, Rho86, Rho95, Rov21, Wal94, Win98, Blo08, HPR16, Sch12b]. Makropakete [App88]. Man [Ber05, Bha22, Bro82, Bro95, Far09, Kan91, Kra11, Nic03, Pau91, Res18, Win08, Bro06, Dev11, Ekm16, Fry22, HHPM07, Pää14, Smi10]. Man-Month [Bro82, Bro95]. Management [HDF^{+}86, JL96, Rod90, X/O88a, X/O88d]. managers [Tes07]. Managing [Mec05, OT88, OT91, Ste91, Tal86, Tal88, WMB94, WMB99]. Manchester [Swi19]. Manhattan [Kel09, Ree14a, Ree14b, Ree20]. Manual [Ame94, Ame01b, AT&T83a, AT&T83b, Ano79a, Ado85a, Ado90, Ano93, Ano03, Ano10, Ano17, BC93, Cas05, CGG^{+}91a, CGG^{+}91b, CW80, DCJ^{+}97, ES90, GAW87, FB93, Fla92, Gas92b, Gas92a, Gol98, HS84, HS87, HS91, HS95, HS02, Hei93, Hel90, Hel91a, Hel91b, Int83a, Int85a, Int87, Int86, Int90, Int91, JW74, JW85, KN94, KCM^{+}76, KK99, Lam86, Lam94, LF78, Mot84, Mot85a, Mot85c, Mic12, Mot85b, Mot85d, Mot89, Nye88a, Nye88b, Nye90, NO90a, NO90b, Nye92, NO92b, NO93, O’R90, Pog05, SPA92, Sal11, Sch89a, SFC64, Sit92, SW95, Ski98, Ski08, Sun87, WG94, von52, Ame10, Ano90, Gen93, RV09, Tur07, Whe21, AT&T86c, AT&T86g, AT&T86f, AT&T86d, AT&T86e]. Manufacture [FP79]. Manuscript [GG05, Org91]. Manuscripts [Dry14]. Many [Dys10, Ker18, Nah06]. Many-Colored [Dys10]. Map [Sny82, Wil02, Win01]. Maple [C^{+}89, MAP, BF91, CGG^{+}91a, CGG^{+}91b, CGG^{+}92, Dev93, EL89, Fat92, Har92, HHR95, Hec93, Hec96, HEBS93, Joh93, KN94, MGH^{+}96, MGH^{+}03a, MGH^{+}03b, Red93, Red96]. Mapped [Kea00, Rim93]. Maps [Gal03, Mic89, Tho81]. March [IEE88]. Marić [ECS19]. Marie [Eml12]. Markets [Tal01]. Marks [Gol98]. Markup [Bry88, Fee88, Gra83a, Org91, Smi87a, Smi87b]. Martyrs [Ale10a]. Mary [Sin99]. Mass [New14, War08]. Massively [KmWH10]. master [KW90]. Mastering [Jae96, SK19, Whi83]. Mastery [JL16, LJ15]. Mastodon [Tho08b]. Matching [KR81, KMP77, Spe85]. Material [DS51, FGMvH14]. Materials [Bal97]. Math [Dev00, Grä96, Grä00b, Grä07, Grä16, SR06, Wel05, O’N16]. MathCW [Bee17]. Mathematica [PW94, Wol91, Cra91, Tho97, Tro06a, Tro06b, Wol88, Wol92]. Mathematical [AS64, AA85, AAR90, AAR94, AA08, AW05, AWH13, Bar06a, Bee17, Bey78, Boa06, Bor93, CKCFR03, Cas05, Pry58a, Pry58b, Cla94, Com64b, Cow84, Dev00, Dev02, DG87a, EE07, Fin03, Fiu89, FF01, Fow97, HF87, HF98, Hig93, KLR89, Leh98, LM66, Liv05, LW84, Luk75, Luk77, LS89, Mos89, OLBC10, OALD97, Pet98, Ric81, RA76, Sch98, Sch11, Sin97, SB07, Trz95, Bab37, DW46, Haw05, Nah06, Tho97]. Mathematical-Function [Bee17]. Mathematician [Har40, Rue07, Saa01, Can01]. Mathematicians [Acz11, Bel37, Tur93, Ten09]. Mathematics [Ale10a, AvD19, BBC^{+}07, Ban99, BN06, BB04c, BRR08, BC06, Bur07, But64, Chu58, Eve83, GKP89, GKP94, GK82, GW98, HHR95, HK05, Knu03b, Kre93, KN94, Leh98, MM56, MM64, MRS82, Par84, Pit19, RW04, RH11, RS09b, Rue07, Sab03, San07, SR66, Ste16, Str74, Wol88, Wol91, Ari03, Ari05, Bel37, BBG04, BD09, BSMM07, Der03, Dev08a, Gut11, KIR^{+}07, Maz10, McL91, Rud07, dS03, PA05, Ten06]. Mathematischer [DW46]. MATLAB [Tre00, The92]. Matrices [DER86, GK78]. Matrix [CGO07, CV88, DG87b, GHN^{+}90, GBDM77, GV83, GV89, HM88, Jap83, Jen77, Kha86, Pis84, Qua89, Ric81, SBD^{+}76, Ste73, Wat07]. matter [Gat09]. Mattered [Bor94]. Matters [Mac11]. Maverick [Sch13]. Mavericks [Win14]. Max [BB05, Gre05, Seg11]. Maxwell [Fle08, FM14, Lon16]. May [Par84]. Mayfield [PPB98]. Maze [Sto89]. MC68000 [Mot84]. MC68020 [Mot85a, Mot85b]. MC68881 [Mot85c, Mot85d]. MC88100 [Mot89]. McNealy [Sou99]. MDB [MMG07]. Mean [HF87]. Meaning [EG05, Fey98, Wil15, Win03b, Sch08]. Means [Weg80]. Measure [Ald02, Bal97, Wul12]. measurement [Cre11]. mécanique [Cou38]. mechanical [Cou38]. Mechanics [And63, Dus38, Fre71, Ghi05, Han63, Har00, Lee65, Mes99, Osg65, PP09, PW35, SN17, SLT66, Sus08, Tau61a, Tay63, Tin64, Cou38, Zus62]. Mechanik [Zus62]. Mecum [Lip99]. Medal [RH11]. mediant [Gut11]. Medieval [Cah95]. Meeting [SBH^{+}04, MA91, Pau89]. megatons [GC02]. Megawatts [GC02]. Members [Gri78]. Membranes [Lap08]. Memoir [Ker19]. Memoirs [Bre10, LEMW52, Mai18]. Memorial [Nat93]. Memories [BR93b]. Memory [Cha18, DDSv91, JNW08, JL96, SK19, Tro18]. Memory-safe [Tro18]. Memory-safety [Cha18]. Men [Bel37, Win14, FM14]. Mendeleev [Gor04]. Mercer [FW07]. Merchants [Klo86]. Mesopotamia [KIR^{+}07]. message [X/O93a]. Messenger [Tan93]. Metafile [Ame86b]. book [Knu86b]. Metalanguage [Gra83a]. Metamarks [Bil89]. Meteorology [Tau63c]. Method [Fro77, DHG08, Sea94]. Methodologies [NHM97]. Methodology [Gri78, HDF^{+}86]. Methods [AvD19, AW05, AWH13, Boa06, Che00, Cla90, DBA74, DR75, DS83, DER86, Fle80, Fle81, For96, Gra90, GW98, HY81, KMN89, LS89, Mor83, Mos89, Pen65, Ric83, Sal02, Sch84, SM07, Tau76, Tre00, Ueb97, Urr01a, Wal75, Wat07, WMD65, Tho08a, Urr01b, Tau77]. metric [Mar14a]. Mexico [Ran91]. MI6 [Fry22]. Michael [Hut52]. Michelin [Mic89]. Michelle [Ano05]. Micro [Wei80, CM85]. micro-PROLOG [CM85]. Micro/Minicomputer [Wei80]. Microchip [Ber05, CBB^{+}99]. Microcomputer [Art84, Art85, Bow77, LG84]. Microcomputers [HJ76]. Microcomputers/Microprocessors [HJ76]. Microphysics [WB01]. Microprocessor [AIM95, Bae10, Cho89, Dol79, Hei93, IEE85a, Int90, Int91, Mot84, Mot85a, Mal95, Mar90, MW82, Mot85b, Mot89, MB88]. Microprocessors [HJ76, Mit86, Sto07]. Microservices [Kol19]. Microsoft [CK07, Zac14]. Microsystems [HB90a, Sou99]. Microwave [Sla42]. Middle [Win08]. Migrating [Ker93, PDL04]. Migrations [PDL04]. Milestones [CGO07]. Mileva [ECS19]. Military [Ame83a, Hil15, dTL18]. Millenium [RD02a]. Millennium [Dev02, Knu05a, PF19, Knu99b]. Millikan [Goo91]. Million [Win11]. Millions [Ker18]. Mind [Hür22, Kur12, Pin97, Mit06, SG17]. Minds [Hau00, Rue07, SL98, Mil16a]. Mini [Ben90]. Miniaturization [Key88]. Minicomputer [EM79, Wei80]. Minicourse [DKM^{+}09]. Minimal [FHCD92]. MINPACK [MGH80]. MINPACK-1 [MGH80]. MINT [GHHH82]. Minus [Nah98]. MIPS [Cho89, FB94, Gwe96, Hei93, Kan87, KH92, RJR88, Swe99, Swe07]. MIPS-X [Cho89]. MIR [Bur98]. Miracle [SS09]. miraculous [Sta98a]. mirror [Zim08]. Misleading [Bai92]. Mismatch [ST12]. Missing [Pog05, QN08]. Mistakes [Oha08]. Mixed [Hoh01]. MKS [Mor87a]. MKSAWK [Mor87b]. MLA [Gib96, Gib99, Mod09, Tri98, Tri99]. MMIX [ABR02, Knu05a, Ruc15]. MMIX-Buch [ABR02]. MMIXware [Knu99b]. Model [GSS68]. Modeling [BBB87, CCN06, Mor85, OALD97, Sal99]. Modelling [FGP93]. Models [Fow97, GSS68, LM66, X/O94a]. Modern [Ale10a, AK01, BL63, BZ11, Cor19, Dol11, Far03, Fin91, Gib19, HC21, Jae96, Kev79, Mac92, Mao04, Mit18, NS21, Plo82, SN17, SM07, SR66, Tan92, Tan08, Vul92, Win01, Win18, Bar80, Cla09, Dev08a, Fag10, HPR16, PM13, Pin02, Pin03, Knu86d]. Modula [Chr86, MS87, Nel91, Sed93, Wir83]. Modula-2 [Chr86, MS87, Wir83]. Modula-3 [Nel91, Sed93]. Modular [Joh18, Wil95a]. Molecular [Bak62, BS60, Her45, Her50, Her66, LP64, Str61]. Molecules [Dav65, GS73, Her45, Her50, Her66, Sla63, Thu80]. moment [CG14]. moments [Bag11]. Momentum [Sch15b]. Monsters [Imp19]. Month [Bro82, Bro95, Mar14b]. months [Mac03]. Monumental [Gla98]. moon [Bid09]. Most [BR93b, GG05, Liv02, Pad90, Wel05, Ado97, Ber07, Bla02, Bod00, FLS95, Har10, Jer02, Ros10, Sho14, Wei16]. Motif [NO92b, You90, Gut11, JR91, JR93, FB93, Hel91a, NO90b, Ope91b, Ope91c, Ope91d, QO93a]. Motoring [Mic05]. Mr. [FLH85, Sta99]. ms [GL88, Dou90, Gia90, Gia92, Hol87b, The92]. MS-DOS [Dou90, Gia90, Gia92, Hol87b, The92]. Much [Mac92]. Mulliken [LP64]. Multics [Org72]. Multigrid [Tho08a]. Multimedia [GBL^{+}98, Gwe96]. Multinational [Cor19]. Multiparameter [Kho02]. Multiple [SR06]. Multiplier [Cor86]. Multiplier/Divider [Cor86]. multiprocessor [HS08]. multiprocessors [Bae10]. Multithreaded [Wea08]. Multithreading [RR96]. Murder [Win98]. Murray [Mur77]. Music [Fox87, Gou86, RS87, Mit06, dS03]. Muslim [Mor07]. Must [Far03]. mv [GL88]. My [Bro06, Ker81b, Sch98, Sus08, Wat90]. MySQL [DuB99]. Mysteries [Ghi05, Win08]. Mysterious [Wel05, Mal11]. Mystery [FR13, GG05, QN08, dS03]. Mystic [Far09]. mysticism [Klo95]. Mythical [Bro82, Bro95].

NAG [Phi86]. Naked [Whe14]. Naledi [BH17]. Named [Kea00]. Nan [RB08]. Narrative [Tuf97]. NAS [BBB^{+}91]. NASA [Bur98]. Nation [SS09, Win14, Van89]. National [Ame86a, Ame88a, Ame88b, Ame89, Ame92, Ame90, Sch90, BL63]. Natural [Kap99, tVW14]. Nature [Man83, Pag82, Pag83, Ree08, Kum10, Pag11, Pin02, Pin03, Pin11, Bal15b]. Navigator [Gil93]. Navy [Bos99]. Nazi [Bla02, Jac14, Ric06]. Nazis [Ker05, Tur20]. Neanderthal [Pää14]. Neanderthals [Fin09, Mit06, PM13]. need [Com00a, MZP97]. needs [Sch12a]. Neither [GS12]. nerve [Kle56]. nets [Kle56]. Network [Cha18, Gil93, Hun92, SV99, Ste90b, Ste98, Ste99b, SFR04, Rag93, GRA89]. Network/extensible [GRA89]. Networked [Sch00, Tap96]. Networking [HJ04, KW89, Ste98, SFR04, X/O88b, Com00a]. Networks [FA89, Ogl99, Tan88, Tan96]. Neumann [AB87, Bha22, Knu70, Mac92, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. Neverending [BvdPSZ14]. Newer [Pyn10]. NeWS [GRA89, Inc]. Newsletter [MAP]. Newton [Bar06a, Bre10, Dol11, Dry14, Gle03]. Newtonian [Fre71]. Next [Gra00a, Osb79, Smo07, Zac14]. NFS [Ste91]. Nicholas [Car10, Ser18]. Niels [AH13, FK85, Pai91]. Nikola [Sei96]. Nineteenth [Sta98b]. NIS [Ste91]. NIST [OLBC10]. NNTP [Ste96]. No [Syk94, BL19]. Noah [McK79]. Nobel [Gre05, Kea18]. Noble [Bor94]. Nodes [Kro65]. Non [Ame70]. Non-Periodical [Ame70]. Noncommutative [Lap08]. Nonequilibrium [Zub74]. Nonlinear [DS83]. nonproliferation [FGMvH14]. Noon [Sou99]. nor [GS12]. North [Ric06]. Norton [Nor94]. Not-so-easy [Fey97]. notation [Sch11]. Notebooks [PW94]. Notes [Fir93, Plu87, Shu94, And11]. Nothing [Kap99]. Novel [SB07]. November [Ger20]. Noyce [Ber05]. NT [Nag97, Zac14]. NTC [RS91]. Nuclear [FKM64, GSS68, Jen00, Kap20, Mac92, RS09a, Rin02, Rob59, Sch12b, She83, Stu18, Acz09, Bet91, Cra07, Far13b, FGMvH14, GC02, Mag09, Mah09, Nic87, Ric06, Sza84, Wea12]. Nucleus [FR13]. Number [EL04, HF98, Klo95, Liv02, Mao94, Vis00]. Numbers [Cla94, Cli90, CG96, Coo81a, Dev00, Ifr00, Ker18, McL91, Pau91, PL07, Rei06, SW90, Wel05, Bes04, Dev11, MS15, Ree00, Roc06]. Numeracy [Pau91]. Numerical [DBA74, DS83, Fro77, GG46, Gra90, Hig96, Hig02, KMN89, Mil84, Mor83, NW06, Ove01, Pen65, PFTV86, PTVF07, Ric83, Sch84, Str74, Tau63b, Ueb97, Urr01a, Urr01b, Boi97, BLWW04, DR75]. Numerically [Bel90]. Numerics [AHG^{+}04, Tro06a]. numérique [Har04]. NURB [Far95, Far99]. NURBS [Far91, PT97, Rog01]. Nuts [Ame01a]. Nuts-and-bolts [Ame01a]. Nutshell [Fla95, GO90, Gt92, GO92, HM04, O’Q99, Rob99, Rob05b, TT00, Hud06].

O [Har99]. Oberon [Wir92]. Objects [Kes88]. Obsession [Mil10, Der03]. Occidental [WH76]. Occupation [Ros14]. Ocean [Win11]. Oct [MA91]. October [AH89]. Odd [Tre99]. Odyssey [Ald02, Dry14, Sch13]. Official [Ame93, The97, The98, NDW93, Zim95]. Old [Wei11]. Olsen [RH88, RH90]. On-line [Sta98b]. One [Gib19, Gri90, LM66, Nah98, Ove01, Pad90, RH77, Win14, Har10]. Only [BJ66, Str13a]. Open [PW17, Sch98, Tay92, Mil90, YP92, Tay92]. OpenBSD [PN04]. OpenGL [NDW93]. OpenOffice.org [GHLP03]. OpenSolaris [MM07, MMG07]. OpenSPARC [Wea08]. Operating [Bac86, Chr83, Chr88, Com84, Com87a, IEE85a, IEE86, Int83b, LMKQ89, MBKQ96, MNN05, MNNW15, Mil87, PDL04, Tan87, Tan92, TW97, Tan08, TSP93, Wir92]. Operation [Jac14]. Operational [Chu58]. operative [Tay25b, Tay25a]. Operators [Ave76, Bic48, Erd74, Tau63a, Tau61b]. Opinions [Ein54, Ein82]. Oppenheimer [BS05, Con05, Her02, Mon13, Sch08]. Oppermann [Kro90]. Optical [GSS68]. Optics [JW76]. Optimal [EE07]. Optimising [Sch89b]. Optimization [ACHM82, Che00, DS83, Fle80, Fle81, GMW81, Met82, NW06, WC00, Wal75]. Optimizations [SS02]. Optimized [TW88]. Optimizing [AK01]. Oracle [BF91]. Orbital [Bel74, Str61, GS72]. Orbitals [LP64, OJ71]. Order [PBMW98, Rob89]. ordered [Gor04]. Ordinary [Seg11, Str74, Syk94, Blo08]. Oregon [BCDH09]. Organic [DS51, Dye65, FPZ^{+}64, GS72, GHHH82, Rob59, RC64a, RC64b, SFC64, Str61, Thu80]. Organique [Fra64]. Organization [EM79, HP94, HP97, HP04, Org73, PH94, PH97, PH04, PH08, PH18, Wie83b, Wie87, PH12, Sta10]. Origin [MH22]. Original [LEMW52]. Origins [HL96, Mal13, Ran82, Dys12, Mit06, Seg11]. OSF [Hel91a, NO90b, Ope91b, Ope91c, Ope91d, QO93a, You90]. OSF/Motif [Hel91a, Ope91b, Ope91c, Ope91d, QO93a, You90, NO90b]. Other [Bor94, Kea12, Mil16b, Sto99, Tau62, Ber11, FL88, Kea10, Lev16, Smi14]. Otto [Bat88, Hah67, She83]. Our [BH17, Dev02, Kea12, McW08, Bar06b, Ber11, Cle09, Fal11, FvGGM90, Mil16a, Pin11, PA06, War08]. outer [Lev16]. outliers [Sch12a]. Outline [DS51]. Output [DS88b, Hol87a, TW88]. outwitted [Fag17]. overcoming [CW14]. overtook [Far13b]. Overview [Ros88]. Overviews [NHM97]. Oxford [Boi97, Maz97, Mur77, RS09b, SSStt97, Wat00, WR98, Win98, Win03b]. oxidation [Lat52].

P [FF95, Fey00, Fey05]. PA [Kan96].

PA-RISC [Kan96]. Paced [Jae96]. Pacific [Win15]. package [Spi86, Spi90, Voß12]. Packing [AW00]. pact [Cor03]. Page [DRB89]. PageRank [ANTT02, LM06, PBMW98]. Pages [Whi90b]. paging [Kur17]. Paid [Moo16]. PAK [HR75]. Palace [Con05]. Pandora [Bil89]. Panther [BF05]. Paper [Ame93, Gus17, Kur17]. Paperback [Sta99, Maz97]. Paperclip [Jac14]. Papers [AB87, Dry14, Gib96, Gib99, HB90b, Knu96, Knu00, Knu03a, Knu03b, Mod09, MTW68a, MTW68b, MA91, Ran82, Trz95, Ken12, Sta98a, Swa76, Tri98, Tur07]. paradox [Gol05]. Parallel [Akl85, BBB^{+}91, GHN^{+}90, Gre11, HJ83, HJ88, KmWH10, Kro85, MB00, Mlo01, Sch84, Tro18, Wu99]. Paramagnetic [Low60]. Paris [Mac03, Ros14]. Park [Pri21, Smi98b, Tur20]. ParseMe.1st [McG98]. Parser [DS88a]. Parsing [AU72, AU73, Han85]. Part [RH77, Sed02, Tau76, Knu11, Tau77, Tau79]. Partial [DKM^{+}09, Tho08a]. Particle [GSS68]. Particles [Swa65, LM66]. Partitions [Knu05b]. Parts [Sed98]. Pascal [Ame83b, Bri85, JW85, Bow77, Dev08a, Int82, JW74, JW85, KP81, Ker81b, LHN79, Roh84, TA81]. passages [Bar80]. Passion [Col05, Jon08]. Past [Mac11, War08]. Patent [Tes07]. Patenting [Ma09]. patents [Kle06]. paths [BBG04]. Pattern [KR81, KMP77, Spe85]. Pattern-Matching [KR81]. Patterns [Hol04, Jap83]. Paul [KCM^{+}76, Kyk76, Far09, Sch98]. Pauli [FW07, Mil10]. PC [Gia90, Gia92, Laf84, Mor84, Nor82, Nor85, TT00]. PCFORT [CCN^{+}79]. Pcode [CCN^{+}79]. PCs [Mue00]. PCYACC [Abr87]. PDF [Mer97b, Ado00, Mer97a, Mer98]. PDF-Dateien [Mer97b]. PDP [EM79]. PDP-11 [EM79]. Pearls [Ben86a, Ben88, Ben00, Ben86b]. pebble [CBB^{+}99]. Pentium [Col05, Sha99b]. People [AA85, AAR90, AAR94, Bor94, Col05, Pla93b, AA08, CL88, FL88, Van89]. Perfect [AW00, FHCD92, Sob12, Cyb91]. perfection [Bor90]. Perfectionists [Win18]. Perfectly [Fey05]. Performance [Bai92, Bel90, Cyb91, Don89, Dow93, Egu18, FJSD96, Gab85, HJ04, LNV89, WC00, Wu99, JR13, MMG07, PF05, Sta10, Tho08a]. Peril [Tap96]. perils [Kea18]. Periodic [Tau61b, Gor04, Kea10]. Periodical [Ame70, Ano96b]. Perl [CT98, Wal87, CT98]. Person [Pin14]. Personal [Bra84, Dij82, IBM86, The92, Rue07, SS84, Wal01, Bet91, CL88, Nic87, SSU85].

Personalities [Dyl20]. Perspective

[Dij82, Hos91, Rog01, Bet91]. perspectives [KN19a]. Peter [Ber06, Nor94]. Petit [Lar06, Ser89]. Petros [Dox00]. PEX [Ros88]. PEXlib [Gas92a]. PFORT [Ryd74]. PGP [Gar95, Zim95]. pH [But64]. Phenomena [tVW14]. Phi [Liv02, JR13]. PHIGS [Ame87b, Bro85b, Gas92b]. Philosopher [Sch49a, Sch49b]. Philosopher-Scientist [Sch49a, Sch49b]. philosophy [Gan95, Pai91]. Photo [Gol98]. Photograph [Kur06]. Photoshop [O’Q99]. PHP [GBR05, Hud06]. phrases [Bar80, RR90]. Physical [Bar79, Boa06, BS60, DWB^{+}62, Dav65, Erd74, Fey94, Gla46, GL60, Kit63, SG62, SLG52, Tay25b, Tay25a, Woi06]. Physicist [Dyl20, Can01, Clo15, Gre05, Sch12b]. Physicists [AvD19, AW05, AWH13, Hür22, Kev79]. Physics [Boh87, BBSR69, Ehr73, EI66, Gam85, Ger00, HL33, HH15, HG78, JT18, KCM^{+}76, Kyk76, LM66, LP64, MM56, MM64, Pag82, Pag83, Rei09, Res18, RH77, RK11, Sch15a, SH16, She83, Smo07, Smo14, SR66, Sti75, Stu18, Tip76, WBA85, vLPO23, Bal14, Blo08, BB05, Cas09, FLS95, FM14, GS12, New07, Pag11, Pai91, Ree14b, Seg07, Smi98a, Sta98a, WP85, Bag10, Ber06]. Physiology [OALD97]. Pi [AH01, Bar92, Bar96, BBB00]. PIC [Ker82]. Picture [Ame86b]. Pictures [Van82]. Pieces [Fey97, FLS95]. Pinpoint [Mil16a]. pioneer [Ekm16, Smi10]. Pioneered [Mac92]. Pioneering [Sch13]. Pioneers [Gla97, Sta98b]. Pipelined [Kog81]. pipelines [Bae10]. Pizzas [Ban99]. Place [Dys10, Kro65, PA06]. places [Wei16]. Plain [BN08, Cla92]. Plan [BL19]. Planar [Thu80]. Planck [CG14]. planet [Ber11, dT08]. Planning [ACM77, ACM79]. Plastic [Rei09]. Platform [AIM95, CLK99b, Mer97a]. Platforms [SN05]. Plausible [BB04c]. Play [EE07, SG17]. pleasure [Fey00]. Pluto [Bro10, dT08]. Plutonium [Ber07]. PNG [Mia99]. Pocket [Lan01, Rob05a]. Poincaré [Bel37, Gal03, O’S07]. Point [Cli90, CCG^{+}84, Cod81, Coo80, Coo81b, Hou81, IEE85b, Mot85c, Mal72, Mot85d, MBdD^{+}10, Ove01, RJR88, SW72, SW90, Ste81, Cor86]. Polar [Lev19]. policies [Nic87]. Political [RS09a]. Politics [Col05, Ele98, BB05, CL88, Far13a, Van89]. polity [Pai91]. Polyatomic [Her45, Her66]. PolyAWK [Cor87]. Polynomials [Gal08]. Polytechnique [AH89]. Pontecorvo [Clo15]. Pope [SH16]. Portability [Bro77, HWS^{+}88, Jae89, X/O88c, X/O88a, X/O88d, X/O88b, X/O89c, X/O89d, X/O89b]. Portable [Ado00, Bee17, BC93, Hal69, IEE86, Lap87, Law78, RS90, Sri96, Wal82, BR89, Lew91]. Porting [Sta88]. Portland [BCDH09]. Portuguese [WR98]. Positive [GL81]. POSIX [Lew91]. POSIX.1 [Lew91, Zlo91]. POSIX.4 [Gal95]. Post [GS93]. Post-Gutenberg [GS93]. PostgreSQL [Mom01]. PostScript [Ado93, Ado85a, Ado85b, Inc88, Ado90, Ado99, Bra89, Cas05, GRM97, Hol87c, Hol89, Hol90b, Rei88a, Rot88, Smi90]. potentials [Lat52]. Power [Cra07, GM02, MJW^{+}06, Sch91, Bet91, CL88, DeL86, GC02, GBR05, JR91, JR93, Mah09, Pag07, Van89, WS94]. Power2 [WR94b]. powerful [Bla02, Ros10]. PowerPC [AIM95, MSSW94, WS94, WR94b]. Powers [Str19, tVW14, Bab37, Ber11]. powłoki [RB06b]. Practical [Ang81, Ben33, Ben35, BED96, Bow01, Far93, Far95, Far97, Far99, Far01, FS03, FSK10, Fle80, Fle81, Fly98, For96, GS96, GMW81, HN10, Les97, Nak62, RR96, Shu94, Trz95, Wal87, van90, van94, ABR02, Cow90, Gil03]. Practice [FvFH90, KP99, Mao04]. practitioners [Tho97]. Pragmatic [TH01].

Pragmatics [Sco00, Sco06]. Prague [Zla92]. praxisnaher [ABR02]. Précis [Fra64]. Precision [LNN95, LNN00, Mar00, SR06, Win18]. Prediction [Sin05, Wu99]. Predictors [RMV42]. Prehistory [Ifr00, Cle09]. Preliminary [Ano79a, Bil89, BGvN46]. Preparation [CH88, CH87, Her70, KD93, KD95, KD99, Lam86, Lam94, Mau47, Org91, AT&T86d]. Prepare [Gus12]. Preparing [Shu94]. Preprocessor [Law78]. Preprocessors [TW88]. Present [Acz10, Lom11]. Presentations [Voß12]. Preserving [Gal08]. President [Bus45]. Presidents [Kap20]. Press [Gol98]. Pretty [Gar95]. Previews [BIO94]. Prime [Bro94, Der03, MS15, Wel05, Roc06]. Primer [Cla92, HK82, Laf84, Lip91, LL83, MW82, PM84, Sof91]. primes [dS03]. Prince [Ten06]. principal [Swa76]. principes [Cou33]. Principle [LEMW52]. Principles [AU77, ASU86, ALSU07, BSN82, Cha78, Com88, Com00b, Den66, ET03, FSK10, FvFH90, GBL^{+}98, HG78, Mac61, NS73, NS79, NS21, RC64a, RC64b, SM68, SBN82, SW80, WS94, Cou33]. Print [SW90]. Printer [Par88]. Printers [Kha86, von52, Jap83]. Printing [Gou86, Kha86, McM88]. Privacy [Gar95, Sal03]. private [Eml12]. Prize [Kea18]. Pro [Sha99b]. Probabilistic [Eth10]. Probability [DS02, Dev08b, Kho07]. Problem [Bee89, Bow77, May82, Par80b, Par98, Sab03, Sin97, Sob95, Wat07, Wil65, Wil02, Der03]. Problems [Dev02, Ehr73, Hoh01, JO83, Kit63, LH74, Oak58, SLG52, Mau47]. Procedural [Rog85]. procedure [Pau89, X/O93b, Zus36]. Procedures [Cha78, DHG08]. Proceedings [AH89, SM07, USE94, Zla92, AFI75, Boi97, Erd74, Par84, BCDH09, IEE88, SO11, SBH^{+}04]. Processes [Kho02, Sma23, SN05].

Processing [Aus82, Cas79, CH88, Dil93, DG87b, Eks84, GW77, Gre83, Lun93, Lun99, Pav82, Smi92]. Processor [Ben90]. Processors [Col76, ET03, KmWH10, MSSW94]. Products [Ben33, Ben35, Gla99, X/O92]. Professional [Mic12]. Professor [Win98]. Profiles [AA85, AA08]. Profound [Bet91]. Program [Knu70, Knu90a, Knu86a, Knu86c, Kru84, Rei88a, Sew89, Jac14]. programmed [ST52]. Programmer [Ame88c, AT&T83a, AT&T83b, AS90, BGA90, Cam87, FB94, Fer88, Fla91, Hol89, Int85a, Int86, Int91, LF78, Mot84, Mia99, Nor85, Ope91b, Ope91c, Org83, Tat87, TH01, Zlo91, Fos05, Gil03, Lew91, AT&T86c, AT&T86g, AT&T86f, AT&T86d, AT&T86e, AT&T86h]. Programmers [ET03, Fla95, Gra88, Jae93, CT98]. Programmiertechnik [Mer91]. Programmiertechniken [App88]. Programmierung [RB06a]. Programming [Ame83a, Ame86a, Ame88a, Ame88b, Ame89, Ame92, Ame90, Ame83b, Ano79b, Ado93, AKW88, AGH00, Aus99, AO90, Bac79, BSS84, BCW88, Bee17, Ben86a, Ben88, Ben00, Ben86b, BG96, Bla15, BO17, Blo01, Bor94, Bra84, Bro85a, Cha18, CM85, CU77, Dig82, DK16, Dyb96, EM79, Ell90, EKP84, EKP87, Gal95, Gas92b, Gas92a, GB18, Gor81, Gra90, Gri78, GG83, GG86, GG90, Gur94a, Har83, Har87, HM98, Hel90, Hel91a, Hel91b, Her69, Her72b, HJ83, HJ88, Hol87c, Hol90b, Int82, JR89, JR90, Joh18, KP74, KR78, KP78, Ker81b, KP84, KR88, KP99, Kes88, KmWH10, KN17, Knu84b, Knu92, Knu05a, Knu05b, Knu08, Knu09, Knu11, Knu74, Lap87, LE80, LG84, Llo87, Lor19, Mat19a, Mat19b, MGH^{+}96, MGH^{+}03a, MGH^{+}03b, NDW93, Nel91, Nor82, Nye88a, NO90a, NO92b]. Programming [NO93, Pav99, Pey87, PF05, Pla93a, Pla93b, Pla93c, Pyl81, Ray04, Rob96, RR96, Rob97, Rob01, RR03, Rob04, Roc85, Roc04, Ruc15, SK10, Sca84, Sch90, SBA92, Sco00, Sco06, Sew89, SKM19, Ski85, Ste90b, Ste92, Ste98, Ste99b, SFR04, SR05, Str86a, Str86c, Str91, Str97, Sum09, TH01, TG89, Tuc77, Van99, Wal82, War85, Weg80, Wir83, Yeu84, You89, You90, YP92, X/O88c, AKW23, Ale10b, Cow90, Dyb87, Dyb03, Fer04, GBR05, HWG04, HS08, JR13, JR91, JR93, Ker10, KN19b, Kur94, Rag93, Str13b, X/O89a]. Programowanie [RB06b]. Programs [Bel90, Ben82, Bor94, Fly98, HM98, LS78, Lib94, Mos89, Van86, Wir76, Lew91, Sch89b, Tho97]. Progress [ANS69]. Project [Kel09, PA05, Ree14a, Ree14b, Ree20, Wir92]. Projections [Sal06b, Sny82]. Projective [Far95, Far99, PP86]. Projects [Gla99, Gra90, Mec05, OT91, Tal86, Tal88]. Proliferation [RS09a]. PROLOG [CM85]. Prometheus [BS05]. Promise [Tap96, Wei18]. Proof [Knu90a, Gol05]. Proofreader [Gol98]. Propagators [LÖ73, LÖ04]. Properties [EK69, Mal72, TW95]. Prophet [Ber81]. proportion [Woo80]. proposal [Ame88c]. Proposals [Cod81]. Proposed [Ame86a, Ame88a, Ame88b, ANS76, Ame87a, Ame90, AOB64, CCG^{+}84, Coo80, Coo81b, Hou81, Ste81, Lud09]. prose [Dup98]. Prospects [Ree18]. protect [Tes07]. Protocol [Nye90, Nye92, SGN88, Sch89a, da 87]. Protocols [Com88, Com00b, Mil92, Sch94, Sch96, Ste94b, Ste96]. proverbs [Bar80]. Proving [GP19]. Pseudospectral [For96]. psychology [Maz10]. public [Bes04]. Publication [Ame94, Ame01b, Ame10, Ear87, IBM86, Mis90, Inc]. Publiceren [dBvLV88]. Publishers [Fly95]. Publishing [GS93, Mer98, Mer97a, Whe21]. Punctuation [Tru03]. Purpose [Fel76, Pla93a, Pla93b, Pla93c, PF05, SK10]. Pursuit [AW00, Mil10]. Purveyors [Wor]. Putting [Wil86b]. Puzzles [Dev02].

Pysanka [Del76]. Pythagoras [Cre10]. Pythagorean [Mao07]. Python [Sum09].

Quadrature [GO70, Str74, Kro65]. Quality [Boi97, KU81]. Quanta [EI66]. Quantitative [BN60, HP90, HP96, HP03, HPAD^{+}06, HP12, HP19, KS52, PH90, PH96, PH03, Tuf83]. Quantities [Tuf97]. Quantum [AH13, Dus38, FBD05, Gam85, Ghi05, Han63, Har00, Hür22, Jon08, JO83, Kau57, Kra11, Kum10, La 71, LB14, LÖ73, LÖ04, Mes99, Pag82, Pag83, PW35, Pit53, RK11, Rov21, SN17, San64, Sin99, Sla63, SLT66, Sus08, Tau61a, Tin64, Tre99, Bag11, Cas09, CG14, GS12, Gre05, Pag11, Sto13]. quark [GM97]. Quarter [Sal94]. Queen [Sin99, Erd74]. Query [BED96]. Quest [Cle03, Fle18, Sin97, Swa01, Cre11, Roc06, Sto13]. Questions [Kit63, Ado97]. Quick [Fla95, HM04, O’Q99, Rob99]. Quotations [Bar80].

R [Wal01]. R. [LP64]. R2000 [Kan87]. R3010 [RJR88]. R4000 [Hei93]. Race [GREC91c, Bid09, Cat04, Far13b, Wul12, Zac14]. Racing [Ban99, Has05]. Radical [Pry66, Qua18, Gle10]. Radioactivity [Mal11]. Radiochemistry [FKM64]. Radium [Moo16]. Radix [CCG^{+}84]. Radix- [CCG^{+}84]. Railroad [Bai99]. Raman [Her45]. Ramanujan [Kan91]. Random [Kho02, Vis00, Gei75]. Randomized [KR81]. Randomness [Pet98, Tal01]. Ranking [PBMW98]. rankings [LM06]. Rare [WB00]. Raster [AH89, Con85, Fiu89, KU81, KPv86, MA91, Wir90]. Ratio [HF87, Liv02]. Rationale [Ano79b]. Ray [Gla89]. Raymond [Gut12]. Rayner [Fro03]. Re [Ash90]. reactions [GS72]. Reactive [GB18, Joh18, Kol19]. Read [Cli90]. Readable [Mat19a]. Reader [PW17, AAD16]. readers [Kir05]. Reading [Str86b]. Readings [AT&T86a, AT&T86b, BN71]. Real [BR93a, Gal08, Gal95, Rot88, Tur18, ECS19, Fer04, Goo04]. real-time [Fer04]. Reality [Gre04, Gre11, Har88, Kum10]. Really [Hau00, Dev08a]. Reason [Pag88]. Reasonable [Fey05]. Reasoning [BB04c]. Reborn [Smo14]. Rechenautomat [DW46]. Rechengeräte [Zus62]. Rechengeräte-Entwicklung [Zus62]. Rechenmaschinen [Zus36]. Rechnungen [Zus36]. Recipes [PFTV86, PTVF07]. reckoned [HHPM07]. Recognition [Sea94]. Recollections [Gla97]. recordings [Ber96]. Rectangular [NF87]. Recursion [Roh84]. Recursive [DM81, Han85]. Recursive-descent [Han85]. rediscovered [PM13]. REDUCE [Ray87]. Reduced [BN08, Erd74, Sta90, Tab87]. Reducing [BC06]. Reduction [Cet91]. Redundant [PCGK89]. Reefs [Win15]. Reference [AHG^{+}04, Ano79a, ABW92, ABM^{+}97, Ado85a, Inc88, Ado90, Ado99, Ado00, AIM95, Bha96, BC93, CGG^{+}91a, CGG^{+}91b, Eij92, ES90, FB93, Fla92, Fla95, HS84, HS87, HS91, HS95, HS02, HM04, Hol89, Int83a, Int85a, Int87, Int86, Int90, Int91, Lam86, Lam94, LF78, LNN95, LNN00, Mot84, Nye88b, Nye90, NO90b, Nye92, O’Q99, O’R90, Ope91c, Rob99, Sal98, Sal00, Sal04, Sal07a, SGN88, Sch89a, SKM19, Sit92, SW95, Wol92, Fos05, Gen93, Jos99]. references [Ame70]. Referenz [dG96]. Reflections [Dys10, Sto99, Tho84]. Refutation [Kah92]. Regular [Spe85]. Reich [Bal14, Bid09]. Reign [Ted03]. Related [Smi87a, Smi92, SW72]. Relativitätstheorie [Hen92]. Relativity [Acz99, EI66, Fre68, Dra12, EG05, Eis06, Fey97, Hen92, Hen97, LEMW52, RN68, Rob05c, Wil86b]. Relay [Ces46]. Release [Fla91, MP92, NDW93, NO90b, Red96, Gen93]. remaking [HPR16]. Remarkable [Sta98b]. Remember [New14]. remote [X/O93b]. Renaissance [AK11].

RenderMan [Ups90]. Repairing [Mue00, Ogl99]. Repelling [CB94, CBR03]. Repertoire [WH76]. Report [ANS71, JW74, JW85, RMV42, ACM77, ACM79, Wal87, vN45, Bus45]. Reports [Her70]. Representation [Kle56]. Representing [Ara06]. Research [Ele98, Gib96, Gib99, Gus12, Mod09, Par84, BR89, BCW08, Bus45, Tri98, Tur07]. researchers [RV09, Tur07]. Researches [Hut52]. Resilient [Cha18]. Resolute [Bat88]. Resonance [Low60, Mac76, Rob59]. Retargetable [FH95]. Retrieval [FBY92, Ame70]. Reveal [Mal72]. Revealed [Kur12, Mar15]. Reveals [Str19]. Review [Ber06, Kit63]. Revision [Ado92, Ame87b, CC88, Ope91a, Ope91b, Ope91c]. Revisited [AH13]. Revolution [Ger20, Isa14, Osb79, Rov21, SRR97, Ser18, Wal01, Dev11, Gre05]. revolutionized [FM14, Sob12]. rewriting [PM13]. Rex [Mor01, Alv97]. REXX [Cow90, Fos05]. Richard [BR93b, Cla09, FF95, Fey00, Fey05, Gle92, Kra11, Lei00, Syk94]. riddle [Ber68]. RIDT [MA91]. Riemann [Der03, Lap08, MS15, Roc06, Sab03]. Right [GP19, Wil86b]. Ring [Con09, TW95]. Ring-theoretic [TW95]. Rings [Tau63a]. RISC [Bel90, Hos91, WR94a, WR94b, Cho89, HP92, Mis90, Kan87, KH92, Kan96, Knu99b, Knu05a, MSSW94, Mot89, PW17, Tab87, Tab90]. RISC-V [PW17]. Rise [Ale10a, Cah95, Pag88, She83, Smo07, Sou99, Wea12, dT08]. rivalry [Acz09]. Road [Ano76, Ano99, Fro04, Ker05, Bet91, Car10, Nic87, Ran91, Ran01]. Roald [Con09]. Robert [BS05, Ber05, Con05, Her02, Mon13, Rob89]. Robust [GB18]. Rock [Zoe09]. Roger [GG05]. Roget [RR90]. Role [Cah95, Tal01]. ROM [AL97]. romantic [Hol08]. Rome [Cah95]. Root [Nah98]. Roots [Syk07]. rose [Sza84]. Routines [GBDM77, Mor84, SBD^{+}76, AT&T86g]. royal [Car10, Dol11]. RT [IBM86]. Ruby [TH01]. Rule [Ove01, McG11]. Rules [BJ66, Pin99, Rob89, SV99, Pau89]. Ruminations [Pau91]. Run [Swe99, Swe07]. Runaways [Gla98]. Running [Osb79]. Runtime [BS03]. Russell [FW07]. Russia [Ser18]. Russian [McG11]. Rust [SKM19, Bal15a, Bla15, BO17, Cha18, Egu18, GB18, Joh18, KN17, KN19b, Kol19, Mat19a, Mat19b, SKM19, SK19, Tro18]. Rutherford [Ree08]. Ruthless [Pad90].

S [LP64, BC81, BC84, BC85, BCW88].

S.O.S [OJ71]. Sa [Tor15]. SA23 [IBM86, Mis90]. SA23-1057 [IBM86]. SA23-2619-00 [Mis90]. Safari [Pet98]. Safe [Bla15, Sus08, Tro18]. Safety [SK19, Cha18]. Saga [Rin02, Zim08]. Sailed [Nic03]. Salt [Kur03]. Salute [FvGGM90]. Sam [vCH18]. Samba [BT98, ECBK00]. San [USE94]. Santa [IEE88]. Sapiens [Har15]. save [Cra07]. Saved [Cah95, AK11, Fry22]. savvy [Tes07]. Saxons [Syk07]. Scalable [GREC90, Kol19]. Scale [HWS^{+}88, BH09, BCDW07, Pad90, Ran01, Par84]. Scan [KU81]. scandalous [De 19]. Scanning [BR93a]. Scatter [New14]. Scheme [Dyb96, Dyb87, Dyb96, Dyb03]. scholars [vL92]. school [Goo91]. schools [Pag07]. Schrödinger [MZP97]. Scicos [CCN06]. Science [AKM81, Bis03, BHK20, Bus45, Cre10, Dyl20, Far03, Gib19, Gle92, Gus97, Hil15, Jon08, Knu96, Kra11, MS87, Mil16b, PS88, RW04, Res18, SM07, Smo07, Tem91, Urr01a, Zob97, AK11, ABR02, Bar98, Bet91, Cor03, Cox15, Des10, Eml12, Far13a, FW07, Fro03, Gor15, Gre05, Hol08, Kea18, LM06, Lin08, Mal11, MZP97, PM13, Pol03, Ree14a, Sei08, Smi98a, Tau93, Urr01b]. Sciences [Bey78, Boa06, Cra91, Dev08b, Fow97, Hig93, Klo86, Pag88]. Scientific [Bal15b, CHT02, Dyl20, FJSD96, G^{+}99, Gor15, Gus12, Gus17, Her72a, JS92, Jon78, Kev79, Mac92, MT86, MTW68a, MTW68b, Mil10, Nas90, NHM97, PPB98, PFTV86, PTVF07, Rei09, Rod19, Sob95, Whe21, Acz09, BR89, Bus45, Hah67, Klo95, Par84, RV09]. Scientist [Fey98, Sch49a, Sch49b, Win08, Jer02]. Scientists [BLM19, Bue90, BF71, Ett93, HB91, Jen77, Ma09, MS88, Ros05, SL98, Cat04, Cor03, Ett87, Har10, Jac14, Mor07]. SCILAB [Urr01a, Urr01b, CCN06, G^{+}99]. Scilab/Scicos [CCN06]. Scores [Fox87]. Scots [Sin99]. Scott [Sou99]. Script [RB08]. Scripting [MJW^{+}06, RB05a, RB06a]. scripts [RB05b]. Sea [Bos99, Win11]. Search [Lap08, MW94, SRR97, Sun90, Woi06, LM06, O’S07, Pää14, Wei16]. Searching [BYG92, Knu98b, Knu73b, Ste94a, WM92, BM77, Sed98, dS03]. Second [ANS71, CLK99b, MS87, Tan93, MA91, AFI75, BN08, Stu18]. Secrecy [Sin99]. Secret [Bag10, Con20, DB04, Has16, Kah96, Kap20, Kar01, Kur12, Tur20, Ber96, Con05, Jac14, Jer02, Shu08b]. Secrets [Ele98, Sch00, Str19]. Secure [Mat19b, PN04, Rei03, Sea05]. Security [Bis03, CB94, CBR03, GS96, Gut04, McC98, Sal03, Sal05b, Sch00, Sch03b, Gra11, X/O89a]. sed [Dou91]. See [Fle18, Hel04, Swe99, Swe07]. Seeking [Pyn10]. Seismic [HWS^{+}88]. selbsttätigen [Zus36]. Selected [CGO07, Dij82, Knu96, Knu00, Knu03a, Knu03b, MTW68a, MTW68b, Ran82]. Self [Jae96]. Self-Paced [Jae96]. Seminumerical [Knu98a, Knu69]. Sense [BN08, Pin14, Rov21]. Sensibly [Sch03b]. September [CS99, SM07, SBH^{+}04, Zla92]. Sequel [Bow01]. Sequence [AH46, Gut11]. Sequences [Gus97, Vis00]. Sequential [Kro85, MB00]. Serial [BIO94, Cam87]. Series [Org73, Tol62]. Server [BF05]. Servers [Cha18]. Service [Che94]. Services [X/O88b]. Serving [Bal14]. ses [Tor15]. Set [Jap78, Sta90, Tab87, Cle88]. Sets [Mac80, Tau61a]. Seven [Ald02, Bro15, Pry58a, Pry58b, Dev02]. Seven-Figure [Pry58a, Pry58b]. Seven-Year [Ald02]. seventeenth [Dev08a]. seventeenth-century [Dev08a]. Seventh [KN94]. SFTRAN [Bee79, FC76, Fly73, FF78, Law78, LS78]. SFTRAN/1100 [FF78]. SFTRAN3 [LF78]. SGML [Bry88, Fee88, Fly98, Gra83a, GR90, McG98, Smi87b, SS88, Smi92, Sof91, van90, van94]. Shōkai [RB06c]. Shadow [Fle18, Gor04]. Sham [Sok00]. Shannon [SG17]. Shape [Gal08, LW84, O’S07, Ree00]. Shape-Preserving [Gal08]. Shaped [Cor19, Gib19, Zoe09, McL91]. Shapes [Bil89]. Shared [DDSv91]. Sharing [RT74, RT79]. Shell [GSS68, Hol87b, Rei03, RB05a, RB06a, Ros93, RR02, RB05b, RB08]. Shell-Programmierung [RB06a]. Shells [Qui02]. shieru [RB06c]. Shock [Dic03]. Shook [Gam85, Rei09, Sok00]. Shooting [Cok09]. Shoots [Tru03]. short [Fey00, Tau93]. Showstopper [Zac14]. SIAM [BLWW04]. side [Bid09]. sieve [DHG08]. Significance [ZM08]. Silicon [Win15, Ber05, Wei16]. Similar [Man94]. Simple [Boh87, Knu90a, Ros91, Bae10, GM97]. Simplest [But64]. Simulation [CCN06]. Simulations [Ehr73]. singing [Mit06]. Single [Hür22]. Sir [Bre10]. site [BB04a, Sza84]. Six [FLS95, Fey97, Mac03, Ree00]. sixteen [Kro65]. sixteen-place [Kro65]. sixth [Kol14]. Sixty [FR13, Dor14]. skryptów [RB06b]. Sky [Bar92, Bar96, Cok09, War08]. Skyrocket [Egu18]. SLALOM [BB91, GREC90, GREC91c, GREC91a, GREC91b]. slate [Pin02, Pin03]. Slicing [Ban99]. Slovník [CNVH93]. smashed [Fag17]. Smith [Win01]. Sneaking [Ghi05]. Snowball [Wal03]. societies [Pag07]. Society [Gib19, Sch12a, Dol11]. Sockets [Ste98, SFR04]. Software [Bee17, Bor94, Bro82, Bro95, Bro77, Che02, CW80, Cow84, Don89, DG87a, Fer86, Gla97, Gla98, GHHH82, HP94, HP97, HP04, HJ76, Hsu01, KMN89, KP76, KP81, McG98, Mil84, Mye79, PH94, PH97, PH04, PH08, PH18, Pla93a, Pla93b, Pla93c, Ran89, Ray87, Ric81, Ric83, SKM19, Tau77, Tau79, Tau76, TZ86, Tro18, Ueb97, WC00, Wor, Boi97, Kle06, PH12, Spe85]. Sokal [Sok00]. Solaris [MM01, MM07, MMG07, PDL04, Rob99]. Solid [Sti75]. Solids [Low60, Sla63]. Solubility [But64]. soluble [LM66]. Solution [GL81, Hoh01, Str74, You71, Ata40]. Solutions [CT98, Dar01, HJ04, HN10, Oak58, TG89, Lat52, Van99]. Solve [Sin97, dS03]. Solved [Liv05, Sob95, Wil02]. solvers [Tho08a]. Solving [Bow77, DDSv91, LH74, May82]. Some [Bic48, Fir93, Rue07, Woo80]. Son [Ted03, Wat90]. songs [Lev16]. sonic [Cox15]. Sort [Lor75]. Sorting [Akl85, Knu98b, Knu73b, Lor75, Sed98]. soul [Bal14, Lin08, Seg07]. sound [Cox15]. Source [Che94, Leh03, Lio96, Sch94, Sch96, BBB00]. Sourcebook [Gra00a, KIR^{+}07]. Sources [BIO94, Cow84, Bar80, KW90]. Space [Fey97, Gre04, Gri86, KD17, Bid09, Lev16]. Space-filling [Gri86]. Spacetimes [Lap08]. Spanish [Ram69, Rho15]. SPARC [Sun87, WG94]. sparked [Par03]. Sparse [DER86, GL81, Pis84]. Spawned [Wal03]. Speaking [Ram97]. Special [AAR99, CK07, CPV^{+}08, Fre68, Leb72, Luk69a, Luk69b, Olv74, ZJ96, Dra12, LEMW52]. Specification [Pos89, Ado89, CC88, AS90, GJSB00, Int82, LY97, LY99, MSSW94, Ope91a, X/O93a]. Specifications [IEE85a]. Specifying [Van82]. Spectra [Bak62, Dav65, Her45, Her50, Her66, OJ71, San64, SM68]. Spectral [Tre00]. Spectroscopic [Mac76]. Spectroscopy [Dye65, Nak62]. Speed [Mar00, Sei98, Mau42]. Spies [Has16, Fag17]. Spin [Pau79]. Spinning [Mit18]. Spirit [Gib19]. Splines [BBB87]. split [Cat04]. spoon [Kea10]. spot [Tes07]. Spreadsheets [CKCFR03]. Spring [Ros05]. Sputnik [Dic03]. Spy [Con09, Gre20, Mac18, Sto89, Clo15]. Spying [Ric06]. Spymaster [Fry22]. SQL [BED96, Bow01]. Square [Doo93, Nah98]. Squares [LH74]. Stability [Hig96, Hig02]. Stalin [Has05]. Stalking [Roc06]. stand [CW14]. Standard [Ame83a, Ame86a, Ame88a, Ame88b, Ame89, Ame92, AHG^{+}04, Aus99, CLK99b, CCG^{+}84, Cod81, Coo80, Coo81b, Don89, EKP84, EKP87, Fee88, Gra83a, Hou81, IEE85a, IEE86, IEE85b, Jap78, Jap83, Jae91, Jos99, Org91, PM92, PSLM00, Plu87, QO93b, Sch90, Smi14, Ste81, The91, ZM08, Zlo91, Gil03, JPS07, Lew91, Rig05, Ame90, Bry88, Bud98, Ell90, Jae96, Jae01, JW85, Pla92, Smi87a, Smi87b, The96, The00, The03, The06, Cor89]. Standardization [Hei66]. Standardized [Tau77, Tau79, Tau76]. Standards [ANS69, ANS71, GBL^{+}98, ACM77, ACM79, Smi92, Cle88, Tau79]. Stanford [De 19]. Stanford’s [KW90]. Star [Coh10, Ber11]. Stars [Cok09]. Start [SS09]. Start-up [SS09]. Starting [Doo93]. State [Bee89, Kar01, Sti75]. stateless [Van89]. statement [Dij68]. Statements [Knu74]. States [Bos99, Win14, Lat52, Bus45, Far13b, Ran91, Ran01]. Station [Tur20, Smi98b]. Statistical [And63, FGP93, Urr01a, ZM08, Zub74, Urr01b]. Statistician [Lor18]. Statistics [DS02, Dev08b, TD00, Bes04, Smi14, Whe14]. Status [ACM77, ACM79]. Stay [HL96]. steam [Ros10]. Stephen [Sus08]. Steps [Grä99]. Stereochemistry [Eli62]. Still [PM92]. STL [Aus99]. Stochastic [DKM^{+}09]. Storage [Ame86b, Ame70]. store [X/O93a]. Stories [Cry04, Win11, Zap07b, Zap07a, Ano96a, Ber11]. Storms [Win11]. Story [Acz10, BH17, Cah95, DB04, Gam85, Hoc98, Jon08, Kah96, Kar01, Liv02, Mac18, Mao94, McL91, Mlo01, Mun17, Nah98, Ree20, SS09, Sob95, Sou99, Sta98b, Tur18, Wal03, Win03b, Win08, Bag11, ECS19, Fag17, Hoc99, Kea18, Kie13, PM13, RH88, RH90, Ros10, Sza84, Wri39]. Strange [Acz11, Dry14, Sei08]. Strangelove [Goo04]. strangeness [Woo80]. Strangest [Far09]. Strategic [SV99, Bla02]. Strategies [Hoe98, Sun97]. Street [FW07]. String [Ber06, Smo07, Ste94a, Woi06, BM77]. Strings [Gus97, Lap08, KMP77]. stripping [Whe14]. Stroll [For12]. Structural [Che00]. Structure [ANTT02, BS60, Dav65, EK69, Fiu89, Her45, Her50, Her66, Org72, Sla63, Wei80]. Structured [DG87b, Ett87, Ett93, Fly98, KL93, Knu74, TW88, BED96]. Structures [BN71, BSN82, Bud98, FBY92, Gil78, Gon84, GG78, HS83, KPv86, Kru84, Mat19a, SBN82, TA81, Van88, Wir76, Wir86, Sed98]. Structuring [Pos89, dB75]. Struggle [Cop12, Bal14, Lin08, Seg07]. Struggled [Cor19]. Struve [Bat88]. Student [EL89, Fle08, The92, Ken12]. Students [ST12, BL63, Tur07]. Studies [EE07, OALD97, Tan93, Bil89]. Study [DS51, Kyk76, RB08, BLWW04]. Stuff [BR93b]. Style [Ano93, Ano03, Ano10, Ano17, DCJ^{+}97, KP74, KP78, LHN79, Mic12, Ope91d, Pin14, SW79, SW99, SWK05, Tri99, Ano90, DHJ^{+}12, RV09, Tri98, Tur07]. Stylebook [Gol98]. submarines [McG11]. Subset [Hal69]. Subspace [Wat07]. Substring [Sun90]. Success [Cor19]. Sudden [Ros19]. Suffice [Wil02]. sukuriputo [RB06c]. Sumer [CKCFR03]. Summary [Hei66]. Sun [Eng87, Rho95, Sza84, Ber11, Cla09, Coh10, HB90a, Sei08, Sou99]. Sunburst [HB90a]. SunTechnology [HB90b]. Supercomputer [Cyb91, GREC90, KS87]. Supercomputers [Fer86, Hwa84, MT86].

Supercomputing [Bai92, MSWA86]. Superpowers [Win15]. Supplement [CLK99b, Fla91, RC64b, Ruc15]. Supplementary [X/O89b]. Support [AT&T86e]. Supported [Law78]. Surely [FLH85]. Surface [Far91, Her72b, Her72c]. Surfaces [Far88, Far90, Far93, Far95, Far97, Far99, Far01, Gri81, Sal06a]. Surfboards [Win15]. surrender [Has05]. Surveillance [BLM19]. Survey [PA05, Sny82]. Survive [New14]. survived [Fin09]. Surviving [GREC91b]. survivors [Str13a]. SVR4 [Rob99]. Swabian [Sto13]. Switzerland [AH89]. Symbolic [Kes88]. Symbolics [Tro06b]. Symbols [WH76]. Symmetric [Par80b, Par98]. Symmetry [Bel74, Dor65, Fey97, Liv05, OJ71, Sch15a, Sin05, GS72]. Symposium [BCDH09, SO11, Woo80]. Synonym [RUL78]. Syntax [Bac79, Voß12]. Synthesis [Spi89]. System [Ame85, Ame88c, AT&T86a, AT&T86b, AT&T86h, AB89, Inc88, AIM95, AS90, AO90, Bac86, BC81, BC85, BMM78, Chr83, Chr88, Fla91, Fri91, Geh86, Geh87, HR75, Her72a, Her70, HGDS83, HGDS86, Mis90, IEE86, JS92, KL93, Lam86, Lam94, LMKQ89, Man94, MM83, MM90, MBKQ96, MNN05, MNNW15, Nag97, NSS89, NSS^{+}01, NSHW10, OQL88, Org72, Org73, Org83, PDL04, QO90, Rag93, RT74, RT79, Rub88, SK19, TY82, TSP93, Wir92, Wol88, Wol91, X/O89d, BR89, Cre11, FGP93, GL88, Ker10, Mar14a, Sha99b, AT&T86g, AT&T86f, Ado93, Bel90, GO90, GO92, GRA89, HWS^{+}88, Hos91, Jon89, MP92, QO93b, QO93a, SGN88, WR94a, WR94b, YP92]. System/6000 [Mis90, Bel90, Hos91, WR94a, WR94b]. Systematic [SFC64]. Systems [Ame85, Ame86b, Ame88c, Bae80, Com84, Com87a, CHT02, DDSv91, Ear87, EM79, Erd74, Fer86, Gab85, GL81, IEE85a, Int83b, JNW08, LG84, Lor75, MW94, Mil87, Nel91, NS21, Omo94, RR03, Rod90, SKM19, SN05, Tab90, Tan87, Tan92, TW97, Tan08, Wei80, Wu99, You71, You89, You90, Ata40, Sin07, Tay92, X/O92].

T. [Alv97]. T1 [Wea08]. T1/T2 [Wea08]. T2 [Wea08]. table [Gor04, Kea10]. Tables [AS64, ACP70, CKCFR03, Pry58a, Pry58b, Com64b, FPZ^{+}64, MM90, Kro65, SW72, WH76]. Tabulations [RD02b]. Tag [CC88]. Tale [BH17, Kea00, Lin04, Nah98, Win98, Cla09, Dev08a]. Tales [Kea12, Klo86, Kea10]. tall [Ano96a]. Tangled [Qua18, Her02]. Tasks [RB06a]. taught [CG14]. TC2 [Boi97]. TC2/WG2.5 [Boi97]. Tcl [HM98, Joh96, Lib94, You97]. Tcl-based [Lib94]. Tcl/Tk [HM98]. TCP [Com88, CS91, Com00b, Com13, HJ04, Hun92, Mil92, Ste94b, Ste96, WS95]. TCP/IP [Com88, CS91, Com00b, HJ04, Hun92, Mil92, Ste94b, Ste96, WS95]. teacher [FLS95]. tear [Gle10]. Tech [Sto99]. Technical [ABO06, Ame01a, App87, MAP, Mic12, PPB98, Ros05, Shu94, WR94b, Zap07a, Zap07b]. Techniques [ASU86, ALSU07, Dil93, Eks84, Joh18, Knu09, KD04, Mat19a, NHM97, PF05, RE87, War85, WH76, Fer04, MMG07]. technischer [Zap07b]. technologies [Chr97]. Technology [Ger00, IBM86, Mis90, Kun85a, NO92a, Pis84, Pla93c, Sei98, WR94a, Wil97, Mil16a, Goo91]. Telegraph [Sta98b]. Telephone [Alt48, Shu08b]. telescope [Gat09, Zim08]. tell [War08]. Teller [BO76, Goo04, Har10, Her02]. Template [Bud98, PSLM00, Aus99]. Ten [Jon08, tVW14]. Tensors [BL66]. TERMCAP [Str86b, SMO88]. Terminal [Mer91]. terminfo [SMO88]. Terminology [Mic12]. Terror [Hoc98, Hoc99, Hol08]. Tesla [Sei96]. Test [Kol19, Wil86b]. Testing [GK78, Mye79]. TEX/ [KW90]. TEXbook [Knu84a, Sal95]. TEXcetera [Bro94]. Texinfo [CS99]. TEXnician [Eij92]. Text [BYG92, Fin91, Fly98, KU81, Kha86, SBA92, WM92]. Textbook [KS52, Str74, BL63]. texte [Bor90]. Texts [Str74]. Texture [Gre04]. Textured [Cet91]. TEXtyl [Ren87]. Their [Luk69a, Luk69b, Luk75, Mar15, Moo16, SL98, tVW14, Bar80, Cou33, Lat52, Leb72, PM13]. Them [Rue07]. Theodore [Mai18, Mor01]. Theorem [Her74a, Her74b, Mao07, Ove01, Wil95a]. theoretic [TW95]. Theoretical [Ara06, AKM81, Tan93, Hen92, Tor15]. theoretical-historical [Hen92]. theoriengeschichtlichen [Hen92]. théorique [Tor15]. Theory [AU72, AU73, AB87, ACP70, Bal62, Ber06, Cot63, Eis06, FBD05, Gam85, Hut52, Kun83, LB14, LEMW52, Mac92, Mao04, Mil02, Mil16b, Mui60, Sla63, Sma23, Smo07, Str61, Tau61a, Tau61b, Tau63b, Tin64, Woi06, McG11, Mil15, Hen92, Hen97, Tau63c]. There [BL19]. Thermodynamics [LR61, Zub74]. Thermodynamique [Eve65]. thesaurus [RR90]. theses [Tur07]. Thesis [FBD05]. thing [Gor04]. Things [Ald02, Fey00]. Think [Cle03, Edm08, Hau00, FL88]. thinkers [Bro06, Mor07]. Thinking [Bar92, Bar96, Dev00, Pin14, Rei90, Sch03b, Klo95, MZP97, Sei08]. Third [Pyn10, Van99, Knu99b, Bid09]. Thirty [Gam85]. Thought [Boo54, Kur12, Hol18]. thoughtful [And11]. Thoughts [Fey98]. Thousands [Ben33, Ben35]. Threads [RR03]. threatens [O’N16]. three [Car10]. thrive [Sch12a]. Throughput [Wea08]. Thumb [Kea12, Ove01]. TIFF [Ado92, CC88]. Tiger [Pog05]. Time [Bar06a, Bro94, Dev02, Gal03, Gre04, Gri90, Pen11, RT74, RT79, Smo14, Sob95, tVW14, Fer04, Fey97, Har88]. Time-Sharing [RT74, RT79]. Timely [Ben33, Ben35]. Times [Imp19, RH11, Sei96, BO76, BB05, Pai91, Tau93, Ted06]. Timescales [tVW14]. Tipler [KCM^{+}76, Kyk76]. Tips [Dil93, Gra83b, Shu94, Fer04]. Titanic [Win11]. Titles [Ame70]. Tk [HM98, Joh96]. Today [Rei88b]. Tokyo [AFI75]. Tolerance [Tru03]. Tompkins [Sta99]. tongue [Bic09, McW08]. Too [Hür22, Ker18]. Tool [BF91, MW94]. Toolbox [Hun85]. Toolkit [Fla92, Lib94, NO90a, NO90b, NO92b, NO93, O’R90, SML98, AS90]. Tools [ASU86, ALSU07, Che02, Fly98, GHHH82, JL78, JL87, KP76, KP81, KD04, MJW^{+}06, MMG07, AT&T86e]. Topic [Eij92]. Topics [The98, Tau62]. Topography [AH89, MA91]. tortured [Smi14]. Tour [Mit11, Rue07]. Tourist [Mic05]. Tower [Hen92, Hen97]. Trace [Mar15]. traced [Bar80]. Tracing [Gla89]. Track [Fey05]. Tracking [Sto89]. Tragedies [McC01]. Tragedy [Jon08, Lin04, BS05, Cla09, Wei18]. Tragic [Lev19]. Training [Jae96]. Trains [Her72b, Her72c]. traitement [Bor90]. Traitor [Mac18]. Transactions [Ste96]. Transcontinental [Bai99]. Transfer [Ame86b, da 87]. Transform [RY01, LNV89, Van92]. Transformations [Sal06b]. Transformed [Ald02, Bar06b]. Transit [Lom11]. Transitions [Mac76]. Translation [AU72, AU73, Lor19]. Translator [CCN^{+}79]. Translators [von52]. Transmission [Sla42]. Transport [X/O93c]. Traveler [Cla94]. Travelling [Rod19]. Treatise [Mui60, Tay25b, Tay25a]. Tree [Qua18]. Trees [Gus97]. Trenches [HN10]. Trends [Cyb91]. Trial [IEE85a, IEE86]. Trial-Use [IEE85a, IEE86]. Tribute [LP64]. tributes [Nat93]. Tricks [Knu09, Fer04]. Trilogy [AH13]. Trinity [Nic87, Sza84]. triumph [BS05]. Triumphant [Lev19, McG11]. Triumphs [McC01]. troff [GL88, KM84, Ker81a]. Trouble [Smo07]. Troubled [Ted03, Mag09]. Troubleshooting [Mil92, Mer97a]. True [Dry14, Sob95, CW14, Fag17, Kea10]. Truman [Has05]. Trust [Tho84, Sch12a].

Trusting [Tho84]. truth [Cra07]. Tsars [Ser18]. tube [Mau42]. Tübingen [SO11]. TUG [SBH^{+}04]. Tukey [Lor18]. tuned [LBS17]. Tuning [Bel90]. Turbulent [RH11]. Turing [Swi19, BJ66, Cop12, Dys12]. Turkish [KA98]. Turm [Hen92]. Turtles [Ban99]. Tutorial [Ado85b, CGG^{+}92, Hwa84, Pri89, RG88, C^{+}89, Jos99]. Tuva [Lei00]. twentieth [McK79, Har10]. twenty [Bro06, Dor18]. twenty-four [Bro06]. twice [Sza84]. Two [BJ66, Fal11, FM14, McG11]. Type [Inc90a, Inc90b, Cra71, Gra83b, LNN95, LNN00, Rub88, SK19, SG74, WH76, Mau47]. Typefaces [Kar87, Knu86d]. Types [GG78]. Typesetter [KM84, Ker81a]. Typesetter-Independent [KM84, Ker81a]. Typesetting [Fox87, Geh86, Geh87, Knu79, KD04, MM90, Spi86, Spi90, WH76, BK86, GL88, Ker82]. Typographic [Her70]. Typography [BD83, Her81, Her72a, Her69, Knu99a, Rub88, SBH^{+}04, STS03].

U [DB04]. U-boat [DB04]. U.S. [Sny82]. Ulrich [Ano96b]. ultimate [RH88, RH90]. Ultra [DB04]. unabridged [McK79]. Unbound [Hoe98]. Uncertain [Pol03, Sch03b, BB05]. Uncertainty [Lin08, Cas09, CG14]. Uncle [Dox00]. uncommon [WB00]. Unconstrained [DS83, Fle80]. Uncovered [Hür22]. Underflow [Coo81a]. Understanding [Bro85b, ET03, Fly98, Hol87c, Lun93, Bar06b]. Undertaking [Bat88]. unexpected [Cla09]. unfinished [Dev08a]. Unicode [Gil03, Kor06, The96, The00, The03, The06]. Unified [MB00]. Unify [Ber06]. Union [Leh98]. United [Boi97, Win14, Bos99, Bus45, Far13b, Ran91, Ran01]. Unity [Woi06]. Universal [Ifr00, KA98]. Universe [Acz99, Bar09, Dol11, Dys10, GP19, MH22, Pen11, Smo14, Ste16, Str19, Wei11, Bar06b, Cle09, Dys12, Gat09, Gle10, Hos11, Isa07, LBS17, O’S07, PA06, Ree00, Sen21, WB00, Zim08]. Universes [Gre11]. Universities [ST12]. University [Erd74, Par84]. UNIX [Ker19, RB06a, AT&T83a, AT&T83b, AT&T86c, AT&T86g, AT&T86f, AT&T86d, AT&T86e, AT&T86h, AL92, AL97, Bac86, BT98, Bou90, Chr83, Chr88, Gan95, GS96, Geh86, Geh87, GL88, Gt92, Hol87a, KP84, Ker10, KW89, Lap87, LMKQ89, Lew91, LR89, Lio96, LL83, MM83, MM90, Mil90, NSS89, NSS^{+}01, NSHW10, PDL04, Qui02, Rag93, Ray04, RT74, RT79, RR96, Rob99, RR03, RB92, Roc85, Roc04, Rod90, Sal94, SBA92, SF85, Ste90b, Ste92, Ste96, Ste98, Ste99b, SFR04, TY82, TSP93, Vah96, Hol87b, Kur94, Rob05b, SR05, AT&T86a, AT&T86b]. UNIX-Compatible [Hol87a]. Unix-Like [Hol87b]. UNIX-to-UNIX [PDL04]. Unix/Linux [RB06a]. Unix/Linux-Tasks [RB06a]. Unleashed [AH01]. unlikely [Fag17]. Unlocked [Win08]. Unmaking [FGMvH14, Car94, Gra11]. unnatural [Kol14]. Unraveling [Ghi05]. Unravelling [FR13]. Unseeable [Fle18]. unseen [CW14]. Unsolved [Dev02, GG05, Sab03, Der03]. Unspoken [dTL18]. Unthinkable [Cle03]. Untold [Cah95, DB04, Mun17, Kie13, McW08]. Unusual [GG05]. Unveils [Ste16]. Update [GREC91c]. Upgrading [Mue00, Ogl99]. upon [McK79]. Uranium [Acz09, Rin02, Zoe09, Ber96]. USA [AFI75, BCDH09, USE94]. Usage [Mic12]. Use [BBB87, Far95, Far99, Gra88, IEE85a, IEE86, Int99, Nes87, Pau79, von52, DeL86, Kle06, Mau42]. Used [Sny82]. Usenet [Tod87, OT88]. USENIX [USE94]. User [ABB^{+}92, ABB^{+}95, ABB^{+}99, Bee80, DBMS79, Eng87, FC76, Fly73, FF78, Gil93, Hei93, JW74, JW85, KK67, Kro93, KK99, Lam86, Lam94, Mot85a, Mot85c, MGH80, Mot85b, Mot85d, Mot89, OQL88, Ope91a, Pav99, QO90, QO93b, QO93a, Ram97, Rob96, SS88, TY82, Zim95]. Users [Bou90, Fly95, KD93, KD95, KD99, SBH^{+}04]. Uses [Cla90]. Using [Aus99, Bee17, Bow77, BED96, Bud98, CK07, Don89, ECBK00, Gia90, Gia92, SKM19, Sta88, TA81, Tod87, dG93, dG97, STS03, Wri39]. Utilities [X/O89c, AT&T86c, Ope91e]. UUCP [OT88, Tod87].

V

[AT&T86h, CGG^{+}91a, CGG^{+}91b, CGG^{+}92, Dev93, Fat92, HHR95, Joh93, MGH^{+}96, Pae95, PW17, Rag93, Red96, Tau63b]. v1.2 [CLK99b]. vacuum [Mau42]. Vade [Lip99]. valiant [Sto13]. Valley [Ber05, Wei16]. Valuable [Ben33, Ben35, Tes07]. Value [Hoh01]. Variable [Sal07b]. Variable-length [Sal07b]. Variables [Chu60]. Variational [Che00]. Various [Don89]. Vast [Win11]. VAX [ACHM82, Dig82, LE80]. VAX-11 [ACHM82, Dig82, LE80]. Vector [DDSv91]. Vectors [BL66]. Vegreville [Del76]. Vendors [Gwe96]. Venus [Lom11, Wul12]. Verfahren [Zus36]. Verification [Gut04]. Verifier [Ryd74]. Versatile [SN05]. Version [AL97, Ado00, Inc90b, Pos89, Ado89, Fla91, Hel91a, KL93, Kru94, LNN95, LNN00, Nye88a, Nye88b, OQL88, PFTV86, The91, The96, The00, The03, The06, WG94, Bal18, CS99, X/O93c, X/O94b, SPA92]. Very [Sun90]. VGA [Fer88]. VI [Fer86, Tau63c]. via [DG87a, Roh84]. Victorian [Smi98a, Sta98b]. Video [Gwe96]. Vietnamese [Lun99]. View [BMM78, Org83, Pen11, Fal11, PA06]. viewpoints [Bet91]. Vikings [Syk07]. violence [Pin11]. Violinist [Kea12]. Virtual [Ash90, Blu02, Cra05, Knu90b, LY97, LY99, MJW^{+}06, SN05]. Virtualization [HN10]. Vision [BB82, Gle10, Par03]. visionaries [Zim08]. Visionary [Bha22]. Vista [CK07]. Visual [Kun85a, Smi90, Tuf83, Tuf97, Kru94, You97]. Visualization [Cet91, NHM97, SML98, Wu99]. VMS [Bou90]. VMware [MJW^{+}06]. Vol [GL88]. Volume [AHG^{+}04, BS03, CLK99b, Com00b, Fle80, Fle81, Knu05a, Knu05b, Knu08, Knu09, Knu11, Ope91a, Ste94b, Ste96, WR94a, WS95, FK85, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c]. Volumes [The91, Ruc15]. Voting [HK05, Saa01]. Voyager [Pyn10]. vs [Kro90].

W [FvGGM90, Lor18]. walked [Hol18]. War [Bag10, Ben21, Car10, Con20, DB04, Has16, HH15, Jon78, Kap20, Kea12, Ker05, Kie13, Mun17, Pri21, Sus08, Zoe09, dTL18, Cor03, Far13a, Jer02, Mac18, Rho15]. warehouse [BH09]. warehouse-scale [BH09]. warming [War08]. Wars [Bar06a, Acz09, Stu18]. Wartime [Con09]. Was [Eis06, Kea00, Moo16, Tur18, Wil86b, Wil02, Gor15]. Washington [Con09]. Water [EK69]. Watson [Ted03]. Wave [Her72b, Her72c]. Way [Hel04, CW14]. ways [Smi14]. Weapons [O’N16, Bet91, GC02]. Weaving [Ram89, Sew89]. Web [Ess04, Mer97b, Mur77, PBMW98, ANTT02, GR99, Mer97b, Mer98]. Webster [McK79, TM71]. we’d [GREC91a]. weights [Kro65]. Weimar [Wei18]. weird [Tau93]. well [Gor04]. well-ordered [Gor04]. Went [Ben21, Bos99, Fin09, Ros14]. Were [Gri05, Nic87]. Wernher [Bid09]. West [Rin02]. WG2.3 [Gri78]. WG2.5 [Boi97]. Whatever [Mar14a]. Where [HL96]. which [Voß12]. Who [Kan91, Mac92, MZP97, Nic03, Sob95, Win08, Fag17, Fin09, Fry22, Gre05, Kie13, Smi10, Win14, Zim08]. Whole [Kro93]. Whose [Knu90a]. Wide [Mer97b]. wife [ECS19]. Wild [Hau00, Osb79]. Wilderness [Acz11]. Wilhelm [Röh14, Bat88, Car10]. Will [New14]. William [Hos11, Win01]. Wily [CB94, CBR03]. win [Kie13]. Window [AS90, Fla91, GO90, GO92, GRA89, HDF^{+}86, JR89, JR90, Jon89, Mlo01, MP92, OQL88, QO90, QO93b, QO93a, SGN88, You89, You90, YP92, X/O88d]. Windowing [Hol87a]. Windows [BST^{+}96, CK07, BT98, Nag97, Zac14]. Wire [Gol98]. Wiretap [Ele98]. Wisconsin [Par84]. wisdom [AK11]. wishful [Sei08]. Without [Leh98, Mar15, Pol10, Wri39]. Wizard [Jon78, Sei96]. Wizards [Ash90, HL96, Knu89a, Knu90b]. Woit [Ber06]. woman [Fag17]. Women [FF01, HH15, Mun17, Des10, Kie13]. won [Cat04]. wonder [Hol08]. wonders [Cox15]. Won’t [ST12]. Word [CCG^{+}84]. Word-length-independent [CCG^{+}84]. Words [Mur77, Pin99, SG74, Kel09, RR90, Wri39]. Work [BB04b, Wal94, Bro06, Can01]. Working [Boi97]. Works [CGO07, Pin97, Tau61a, Tau61b, Tau63a, Tau62, Tau63b, Tau63c, Blo08, Com00a, Fey00]. Workstations [Ear87, IEE88]. World [Ald02, BR93a, Bro15, Com87b, Cop12, Cor19, Cre11, Dol11, Fin91, Gal95, Gia90, Gia92, GG05, GB18, Hel04, Hür22, Ker18, Liv02, McC01, McP98, Mer97b, Mit18, Nic03, Rei09, Rho15, Rot88, Sch00, Sch03b, Sin97, Sta99, Stu18, Sus08, Win01, Win03a, Win05, Win15, Win18, Zoe09, Ari03, Ari05, Ber07, Bod00, Bro06, Cox15, Cra07, Dev08a, Dor18, Gre05, Jer02, Kea10, Kur98, Kur99, Kur03, Mac03, Mar14b, Pol03, Pol10, Ram69, Ros10, Sho14, Ste99a, Ste01, Wei16, Car10, Con20, Ker05, Kie13, Mun17]. Worlds [Pyn10]. Worldwide [The91, Cle88]. WorldWideWeb [Fly95]. would [McG11]. Write [Gus17]. Writer [Int83b, GHLP03]. Writers [Ame01a, Gib96, Gib99, Mod09, DHJ^{+}12, Kir05, Tri98, Tur07]. Writing [ABO06, Bal18, Bal19, Ben82, Dup95, Gli97, Gur94b, HM98, Hig93, Hol87b, Kah96, KLR89, PPB98, Pin14, Pit19, Ros05, Str86b, Stu84, Trz95, Zob97, Dup98, Lew91, Whe21]. Writings [Dij82, Bre10]. Written [Kea12]. Wrong [Ber06, Woi06]. WTL [Cor86].

X [BF05, Ado93, AS90, Cho89, Fla91, Fla92, GO90, GO92, JR89, JR90, Jon89, MP92, Nye90, NO90a, NO90b, Nye92, NO92b, NO93, OQL88, O’R90, Pav99, Pog05, QO90, QO93b, QO93a, SGN88, Sch89a, Sin07, Smi98b, Tay92, Tur18, You89, You90, YP92]. X-Open [X/O93a, X/O94a, X/O94b]. X/Open [Tay92]. X11 [MP92, NO90b]. X3.122 [Ame86b]. X3.122-1986 [Ame86b]. X3.124 [Ame85]. X3.124-1985 [Ame85]. X3.144.1988 [Ame88c]. X3.159 [Ame89]. X3.159-1989 [Ame89]. X3.198 [Ame92, Ame90]. X3.9 [Ame66, Ame78, Ame87a]. X3.9-1966 [Ame66]. Xanthi [SBH^{+}04]. XBM [Mia99]. XCDE [X/O95a]. Xeon [JR13]. Xerox [Cor89]. XHTML [Gra00a]. Xi [RB08]. XINU [Com84, Com87a]. Xlib [Nye88a, Nye88b]. XML [SBH^{+}04, Fly98, GP04, GR99, HM04]. XMS [X/O93a]. XPG4 [X/O92]. XSI [X/O89c, X/O89d]. Xt [You89, You90, Laf84, Mor84, YP92]. XTI [X/O93c, Ste98]. Xue [RB08]. XView [Hel90, Hel91b].

Yacc [Joh79, DS88a, Mor87a, LMB92, MB90]. YACC-compatible [DS88a]. YACL [Sri96]. Year [Ald02, FR13, KD17, Mao07, Shu08a, Sta98a]. Years [Gam85, Tem91, BL19, Rob05c, Rud07]. Young [FF01]. You’re [FLH85]. Yourself [Ker18]. Youth [Ros19].

Z [Tur18]. Zen [Keh93]. Zeno [Bel37]. Zero [Kap99, Rei06, Sha99a, Tru03]. Zeros [Lap08]. ZFS [JL16, LJ15]. Zhi [RB08]. Zillions [Ker18]. Zugang [ABR02]. zur [ABR02, Zus36, Zus62].

References

Albers:1985:MPP

[AA85] Donald J. Albers and Gerald L. Alexanderson, editors. Mathematical People: Profiles and Interviews. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1985. ISBN 0-8176-3191-7, 3-7643-3191-7. xvi + 2 + 372 pp. LCCN QA28 .M37 1985.

Albers:2008:MPP

[AA08] Donald J. Albers and Gerald L. Alexanderson, editors. Mathematical people: profiles and interviews. A. K. Peters, Ltd., Wellesley, MA, USA, second edition, 2008. ISBN 1-56881-340-6. xxvi + 386 pp. LCCN QA28 .M37 2008.

Albers:2016:GHH

[AAD16] Donald J. Albers, Gerald L. Alexanderson, and William Dunham, editors. The G. H. Hardy reader. Spectrum. Cambridge University Press, Cambridge, UK, 2016. ISBN 1-107-13555-9 (hardcover), 1-107-59464-2 (paperback). xv + 1 + 395 pp. LCCN QA29.H23 G44 2015.

Almgren:2000:HWC

[AAG^{+}00] Fredrik Almgren, Gunnar Andersson, Torbjörn Granlund, Lars Ivansson, and Staffan Ulfberg. How we cracked the Code Book ciphers. Technical report, ????, ????, October 11, 2000. 40 pp. URL http:

//frode.home.cern.ch/frode/
crypto/codebook_solution.
pdf;   http://www.simonsingh.
com/cipher.htm. See [Sin99].

Albers:1990:MMP

[AAR90] Donald J. Albers, Gerald L.

Alexanderson, and Constance Reid, editors. More Mathematical People: Contemporary Conversations. Harcourt Brace Jovanovich, Boston, MA, USA, 1990. ISBN 0-15-158175-4. xviii + 375 pp. LCCN QA28 .M67 1990. US$29.95.

Albers:1994:MMP

[AAR94] Donald J. Albers, Gerald L.

Alexanderson, and Constance Reid, editors. More Mathematical People: Contemporary Conversations. Academic Press, New York, NY, USA, 1994. ISBN 0-12-048251-7. xviii + 375 pp. LCCN QA28 .M67 1994.

Andrews:1999:SF

[AAR99] George E. Andrews, Richard Askey, and Ranjan Roy. Special Functions, volume 71 of Encyclopedia of mathematics and its applications. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-62321-9 (hardcover), 0-521-78988-5 (paperback). xvi + 664 pp. LCCN QA351 .A74 1999. US$90.00 (hardcover), US$34.95 (paperback).

Aspray:1987:PJN

[AB87] William Aspray and Arthur Burks, editors. Papers of John von Neumann on Computing and Computer Theory, volume 12 of Charles Babbage Institute reprint series for the history of computing. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-22030-X. xviii + 624 pp. LCCN QA76.5 .P31451 1987. See [Knu70].

Abe:index

[AB89] Kris K. Abe and Daniel M.

                                        Berry. indx and findphrases, A

system for generating indexes for ditroff documents. Software— Practice and Experience, 19(1): 1–34, January 1989.

Anderson:1992:LUG

[ABB^{+}92] E. Anderson, Z. Bai, C. Bischof, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, A. McKenney, S. Ostrouchov, and D. Sorensen. LA-PACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. ISBN 0-89871-294-7. xv + 235 pp. LCCN QA76.73.F25 L36 1992.

Anderson:1995:LUG

[ABB^{+}95] E. Anderson, Z. Bai, C. Bischof, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, A. McKenney, S. Ostrouchov, and D. Sorensen. LA-PACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 1995. ISBN 0-89871-345-5 (paperback). xix + 325 pp. LCCN QA76.73.F25 L36 1995.

Anderson:1999:LUG

[ABB^{+}99] E. Anderson, Z. Bai, C. Bischof, S. Blackford, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, A. McKenney, and D. Sorensen. LAPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, third edition, 1999. ISBN 0-89871-447-8. xxi + 407 pp. LCCN QA76.73.F25 L36 1999.

Abrahams:1990:TI

[ABH90] Paul W. Abrahams, Karl Berry, and Kathryn A. Hargreaves. TEX for the Impatient. Addison-Wes-ley, Reading, MA, USA, 1990. ISBN 0-201-51375-7. xvii + 357 pp. LCCN Z253.4.T47 A27 1990.

Adams:1997:FHC

[ABM^{+}97] Jeanne C. Adams, Walter S.

Brainerd, Jeanne T. Martin, Brian T. Smith, and Jerrold L. Wagener, editors. Fortran 95 Handbook: Complete ISO/ANSI Reference. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, November 1997. ISBN 0-262-51096-0. xii + 711 pp. LCCN QA76.73.F25 F6 1997. URL http://

www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0262510960;   http://www.
mitpress.com/book-home.tcl?
isbn=0262510960.

Alred:2006:HTW

[ABO06] Gerald J. Alred, Charles T. Brusaw, and Walter E. Oliu. Handbook of Technical Writing. Bedford/St. Martins, Boston, MA, USA, eighth edition, 2006. ISBN 0-312-43613-0 (paperback), 0-312-35267-0 (hardcover). xxiv + 636 pp. LCCN T11 .B78 2006. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0708/2005921350-
                                        b.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0708/
                                        2005921350-d.html.

Abraxas:pcyacc

[Abr87] Abraxas Software, Inc. PCYACC 2.0, 1987. 7033 SW Macadam Ave., Portland, OR 97219.

Anlauff:2002:MBP

[ABR02] Heidi Anlauff, Axel Böttcher, and Martin Ruckert. Das MMIX-Buch: ein praxisnaher Zugang zur Informatik. (German) [The MMIX Book: a practical introduction to computer science]. Springer-Lehrbuch. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-42408-3 (paperback), 3-642-56233-7. xiv + 327 pp. LCCN QA76.8.M6 A55 2002. EUR 24.95. URL http://www.

                                        informatik.fh-muenchen.de/~
                                        mmix/MMIXBuch/.

Adams:1992:FHC

[ABW92] Jeanne C. Adams, Walter S. Brainerd, and Jerrold L. Wagener. Fortran 90 Handbook:

Complete ANSI/ISO Reference. McGraw-Hill, New York, NY, USA, 1992. ISBN 0-07-000406-4. xi + 740 pp. LCCN QA76.73.F28 F67 1992. US$79.50. See [Ame92].

Anklam:1982:ECV

[ACHM82] Patricia Anklam, David Cutler, Roger Heinen, Jr., and M. Donald MacLaren. Engineering a Compiler: VAX-11 Code Generation and Optimization. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1982. ISBN 0-932376-19-3. xv + 269 pp. LCCN QA76.8.V37 E53 1982.

SIGGRAPH:core77

[ACM77] ACM/SIGGRAPH. Status report of the Graphic Standards Planning Committee of ACM/ SIGGRAPH. ACM SIGGRAPH —Computer Graphics, 11(3), 1977.

SIGGRAPH:core79

[ACM79] ACM/SIGGRAPH. Status report of the Graphic Standards Planning Committee of ACM/ SIGGRAPH. ACM SIGGRAPH —Computer Graphics, 13(3), August 1979.

Atkins:1970:TGT

[ACP70] P. W. (Peter William) Atkins, M. S. Child, and C. S. G. (Courtenay Stanley Goss) Phillips. Tables for Group Theory. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1970. ISBN 0-19-855131-2. 32 pp. LCCN QA171 .A851.

Aczel:1999:GEE

[Acz99] Amir D. Aczel. God’s Equation: Einstein, Relativity, and the Expanding Universe. Four Walls Eight Windows, New York, NY, USA, 1999. ISBN 1-56858-139-4. xvii + 236 pp. LCCN QB981 .A35 1999.

Aczel:2009:UWS

[Acz09] Amir D. Aczel. Uranium wars: the scientific rivalry that created the nuclear age. Palgrave Macmillan, New York, NY, USA, 2009. ISBN 0-230-61374-8. 248 + 8 pp. LCCN QC773.A1 A28 2009.

Aczel:2010:PCS

[Acz10] Amir D. Aczel. Present at the Creation: the Story of CERN and the Large Hadron Collider. Harmony Books, New York, NY, USA, 2010. ISBN 0-307-59167-0. xvi + 271 + 8 pp. LCCN QC787.P73 A29 2010.

Aczel:2011:SWL

[Acz11] Amir D. Aczel. A Strange Wilderness: the Lives of the Great Mathematicians. Sterling, New York, NY, USA, 2011. ISBN 1-4027-8584-4 (hardback), 1-4027-9085-6 (e-book). xix + 284 pp. LCCN QA21 .A29 2011.

Adobe:1985:PLR

[Ado85a] Adobe Systems Incorporated. PostScript Language Reference Manual. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-10174-2. ix + 321 pp. LCCN QA76.73.P67 A33 1985.

Adobe:1985:PLT

[Ado85b] Adobe Systems Incorporated.

PostScript Language Tutorial and Cookbook. Addison-Wes-ley, Reading, MA, USA, 1985. ISBN 0-201-10179-3. x + 243 pp. LCCN QA76.73.P67 A34 1985.

Adobe:epsf-spec

[Ado89] Adobe PostScript Developer Support Group. Encapsulated PostScript Files specification version 2.0. Technical Report PN LPS5002, Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400, June 5 1989. Available electronically from ps-

file-server@adobe.com in re-

sponse to an e-mail request

send Documents EPSF.ps.Zba.

The request send Index will return a complete index for the server. The version 3.0 specification is published in [Ado90, Appendix H].

Adobe:1990:PLR

[Ado90] Adobe Systems Incorporated.

PostScript Language Reference Manual. Addison-Wesley, Reading, MA, USA, second edition, 1990. ISBN 0-201-18127-4. viii + 764 pp. LCCN QA76.73.P67 P67 1990.

ADA:1992:TRF

[Ado92] Adobe Developers Association. TIFF Revision 6.0: Final. Adobe Systems Incorporated, 1585 Charleston Road P.O. Box 7900 Mountain View, CA 94039-7900, June

3, 1992. URL http://home.
earthlink.net/~ritter/tiff/
 (The Unofficial TIFF Home
Page);   http://partners.
adobe.com/asn/tech/tiff/specification.

jsp. Includes TIFF Specification Supplement 1 (enhancements for Adobe PageMaker 6.0) [14-Sep-1995] and TIFF Specification Supplement 2 (enhancements for Adobe Photoshop) [22-Mar-2002]. Hypertext linked for Web access.

Adobe:1993:PDP

[Ado93] Adobe Systems Incorporated. Programming the Display Post-Script System with X. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-62203-3. LCCN QA76.73.P67 D57 1993. US$29.95.

Adobe:1997:AFA

[Ado97]  Adobe Press.  Adobe FAQ:

Adobe’s most frequently asked questions answered. Adobe Press, Mountain View, CA, USA, 1997. ISBN 1-56830-372-6. xvi + 774 pp. LCCN QA76.754 .A36 1997. Includes CD-ROM.

Adobe:1999:PLR

[Ado99] Adobe Systems Incorporated. PostScript Language Reference. Addison-Wesley, Reading, MA, USA, third edition, 1999. ISBN 0-201-37922-8. xii + 897 pp. LCCN QA76.73.P67 P67 1999. US$49.95, CDN$74.95. URL

http://partners.adobe.com/
asn/developer/PDFS/TN/PLRM.

pdf;   http://partners.adobe.
com/supportservice/devrelations/
PDFS/TN/PLRM.pdf. This new

edition defines PostScript Language Level 3. An electronic version of the book is available at the Adobe Web site, and is also included in a CD-ROM attached to the book.

Adobe:2000:PRA

[Ado00] Adobe Systems Incorporated.

PDF Reference: Adobe Portable Document Format, Version 1.3. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-61588-6 (paperback). xvi + 679 pp. LCCN QA76.76.T49 P38 2000. US$49.95. URL

http://partners.adobe.com/
asn/developer/acrosdk/DOCS/
PDFRef.pdf.

AFIPS:1975:SUJ

[AFI75] Second USA–Japan Computer Conference proceedings: August 26–28, 1975, Tokyo, Japan. AFIPS Press, Montvale, NJ, USA, 1975. LCCN QA76 .U2 1975.

Arnold:2000:JPL

[AGH00] Ken Arnold, James Gosling, and David Holmes. The Java Programming Language. Addison-Wesley, Reading, MA, USA, third edition, 2000. ISBN 0-201-70433-1 (paperback). xxiv + 595 pp. LCCN QA76.73.J38 A76 2000. US$37.95.

Aiken:1946:ASC

[AH46] H. H. Aiken and G. M. Hopper.

The Automatic Sequence Controlled Calculator. Electrical Engineering, 65(??):384–391, 449– 454, 522–528, 1946. Reprinted in [Ran82, §5.2].

Andre:ridt

[AH89] Jacques André and Roger Hersch, editors. Raster Imaging and Digital Topography: Proceedings of the International Conference, École Polytechnique Fédérale, Lausanne, Switzerland, October 1989. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-37490-1. LCCN Z253.3 .R37 1989.

Arndt:2001:PU

[AH01] Jörg Arndt and Christoph Haenel. Pi — Unleashed. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-66572-2 (paperback), 3-642-56735-5 (e-book). xii + 270 pp. LCCN QA484.A7513 2001. US$. Includes CD-ROM. Translated from the German by Catriona and David Lischka.

Aaserud:2013:LLQ

[AH13] Finn Aaserud and J. L. Heilbron. Love, Literature, and the Quantum Atom: Niels Bohr’s 1913 Trilogy Revisited. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-968028-0 (hardcover). viii + 284 pp. LCCN QC774.B64 A19 2013.

Abrams:2004:NFS

[AHG^{+}04] Brad Abrams, Anders Hejlsberg, Brian Grunkemeyer, Joel Marcey, Kit George, Krzysztof Cwalina, and Jeffrey Richter. .NET Framework Standard Library Annotated Reference. Volume 1: Base Class Library and Extended Numerics Library. Microsoft .NET development series. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-15489-4 (hardcover). xxvi + 528 pp. LCCN QA76.76.M52 A27 2004.

URL http://www.aw-bc.com/
catalog/academic/product/
0,1144,0321154894,00.html;
http://www.loc.gov/catdir/
toc/ecip0411/2003024327.html.

Foreword by Joel Marcey.

Aho:1974:DAC

[AHU74] Alfred V. Aho, John E. Hopcroft, and Jeffrey D. Ullman. The Design and Analysis of Computer Algorithms. Addison-Wes-ley, Reading, MA, USA, 1974. ISBN 0-201-00029-6. x + 470 pp. LCCN QA76.6 .A284 1974.

Apple:1995:PMC

[AIM95] Apple Computer, Inc., IBM Corporation, and Motorola, Inc. PowerPC Microprocessor Common Hardware Reference Platform: a System Architecture. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1995. ISBN 1-55860-394-8. xxiv + 309 pp. LCCN QA76.89.P67P74 1995. US$39.95.

Allen:2001:OCM

[AK01] Randy Allen and Ken Kennedy.

Optimizing Compilers for Modern Architectures: a Dependencebased Approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2001. ISBN 1-55860-286-0, 0-585-45699-2 (e-book). xxv + 790 pp. LCCN QA76.76.C65

                                        A45 2001. URL http://www.
                                        loc.gov/catdir/description/
                                        els031/2001092381.html;
                                        http://www.loc.gov/catdir/
                                        toc/els031/2001092381.html.

Al-Khalili:2011:HWH

[AK11] Jim Al-Khalili. The house of wisdom: how Arabic science saved ancient knowledge and gave us the Renaissance. Penguin, London, UK and New York, NY, USA, 2011. ISBN 1-59420-279-6 (hardcover). xxix + 302 + 8 pp. LCCN Q127.A5 A4 2011.

Akl:1985:PSA

[Akl85] Selim G. Akl. Parallel Sorting Algorithms. Academic Press, New York, NY, USA, 1985. ISBN 0-12-047680-0. xiii + 229 pp. LCCN QA76.5 .A363 1985.

Arbib:1981:BTC

[AKM81] Michael A. Arbib, A. J. Kfoury, and Robert N. Moll. A Basis for Theoretical Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 0-387-90573-1. viii + 220 pp. LCCN QA267 .A715.

Aho:1988:APL

[AKW88] Alfred V. Aho, Brian W. Kernighan, and Peter J. Weinberger. The AWK Programming Language. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-07981-X. x + 210 pp. LCCN QA76.73.A95 A35 1988. See also [GAW87, Mor87b, Cor87].

Aho:2023:APL

[AKW23] Alfred V. Aho, Brian W. Kernighan, and Peter J. Weinberger. The AWK programming language. Addison-Wesley professional computing series. Addison-Wesley, Reading, MA, USA, second edition, 2023. ISBN 0-13-826972-6. xiii + 214 pp. LCCN QA76.73.A95.

Abrahams:1992:UI

[AL92] Paul W. Abrahams and Bruce R.

Larson. UNIX for the Impatient. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-55703-7. xxvii + 559 pp. LCCN QA76.76.O63 A27 1992. Excellent, and thorough, coverage of UNIX, with chapters on the file system, utilities, shells, editors, Emacs, data manipulation, mail, network communications and resources, the X Window System, and a comparison of MS-DOS and UNIX.

Abrahams:1997:UIC

[AL97] Paul W. Abrahams and Bruce R.

Larson. UNIX for the Impatient: CD-ROM Version. Addison-Wesley, Reading, MA, USA, second edition, 1997. ISBN 0-201-41979-3. xxxvi + 824 + CD-31 pp. LCCN QA76.76.O63 A27 1992. US$49.00.

Alder:2002:MAT

[Ald02] Ken Alder. The Measure of All Things: The Seven-Year Odyssey and Hidden Error That Transformed the World. Free Press, New York, NY, USA, 2002. ISBN 0-7432-1675-X. x + 422 pp. LCCN QB291 .A43 2002. US$27.00. In June 1792, the erudite and cosmopolitan Jean-Baptiste-Joseph Delambre and the cautious and scrupulous Pierre-François-Andre Mechain set out from Paris — one north to Dunkirk, the other south to Barcelona to calculate the length of the meter. In the face of death threats from village revolutionary councils, superstitious peasants, and civil war, they had only their wits and their letters to each other for support. Their findings would be used to create what we now know as the metric system. Despite their painstaking and Herculean efforts, Mechain made a mistake in his calculations that he covered up. The guilty knowledge of his error drove him to the brink of madness, and in the end, he died in an attempt to correct himself. Only then was his mistake discovered. Delambre decided to seal all evidence of the error in a vault at the Paris Observatory. Two hundred year later, historian Ken Alder discovered the truth. With scintillating prose and wry wit, Alder uses these previously overlooked letters, diaries, and journals to bring to life a remarkable time when everything was open to question and the light of reason made every dream seem possible.

Alexander:2010:DDH

[Ale10a] Amir Alexander. Duel at Dawn:

Heroes, Martyrs, and the Rise of Modern Mathematics. Harvard University Press, Cambridge, MA, USA, 2010. ISBN 0-674-04661-7. 307 pp. LCCN QA10.7 .A44 2010.

Alexandrescu:2010:DPL

[Ale10b] Andrei Alexandrescu. The D programming language. Addison-Wesley, Reading, MA, USA, 2010. ISBN 0-321-65953-8 (hardcover), 0-321-63536-1 (paperback). xxvii + 463 pp. LCCN QA76.73.D138 A44 2010; QA76.73.D138.

Allen:1978:AL

[All78] John R. Allen. Anatomy of LISP.

McGraw-Hill, New York, NY, USA, 1978. ISBN 0-07-001115-X. xviii + 446 pp. LCCN QA76.73 .L23 A43 1978.

Aho:2007:CPT

[ALSU07] Alfred V. Aho, Monica S.

Lam, Ravi Sethi, and Jeffrey D. Ullman, editors. Compilers: Principles, Techniques, and Tools. Pearson/Addison Wesley, Boston, MA, USA, second edition, 2007. ISBN 0-321-48681-1 (hardcover), 0-321-49169-6 (paperback). xxiv + 1009 pp. LCCN QA76.76.C65 A37 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0618/2006024333.html.

Alt:1948:BTLa

[Alt48] Franz L. Alt. A Bell Telephone Laboratories’ computing machine—I. Mathematical Tables and Other Aids to Computation, 3(21):1–13, January 1948. CODEN MTTCAS. ISSN 0891-6837. Reprinted in [Ran82, §6.4].

Alvarez:1997:RCD

[Alv97] Walter Alvarez. T. rex and the crater of doom. Princeton University Press, Princeton, NJ, USA, 1997. ISBN 0-691-13103-1. xxii + 185 pp. LCCN QE721.2.E97 ALV

                                        1997. URL http://www.jstor.
                                        org/stable/j.ctt28553d.

ANSI:ftn66

[Ame66] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. ANSI Fortran X3.9-1966, 1966. Approved March 7, 1966 (also known as Fortran 66). See also subsequent clarifications [ANS69] and [ANS71], and history [Hei66].

ASTM:1970:CPT

[Ame70] American Society for Testing and Materials, 1916 Race St., Philadelphia, PA 19103, USA. CODEN for Periodical Titles (Including Non-Periodical Titles and Deleted CODEN): An aid to the storage and retrieval of information and to communication involving journal references, third edition, 1970. US$215.00. ASTM Data Series DS 23 B. Two volumes. Includes CODEN values for 102,146 periodical titles and 7,361 non-periodical publications.

ANSI:ftn77

[Ame78] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. ANSI Fortran X3.9–1978, 1978. Approved April 3, 1978 (also known as Fortran 77). See also draft [ANS76].

ANSI:ada

[Ame83a] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Military Standard Ada Programming Language, February 17 1983. Also MIL-STD-1815A. See also [Ano79b, Ano79a].

ANSI:pascal

[Ame83b] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. The Pascal Programming Language. ANSI/IEEE 770X3.97-1983, 1983. See also [JW74, JW85].

ANSI:gks

[Ame85] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Information Systems: Computer Graphics: Graphical Kernel System (GKS). ANSI X3.124-1985, 1985. Includes Fortran bindings to GKS.

ANSI:c86

[Ame86a] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed American National Standard Programming Language C, October 1 1986.

ANSI:metafile

[Ame86b] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Information Systems: Computer Graphics: Metafile for the Storage and Transfer of Picture Description Information. ANSI X3.122-1986, 1986.

ANSI:ftn8x

[Ame87a] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed ANSI Fortran X3.9– 198x, September 18 1987. See also [MR87].

ANSI:phigs+

[Ame87b] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. PHIGS+ Functional Description, Revision 2.0, July 20 1987.

ANSI:c88a

[Ame88a] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed American National Standard Programming Language C, January 11 1988.

ANSI:c88b

[Ame88b] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed American National Standard Programming Language C, October 31 1988.

ANSI:phigs

[Ame88c] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Information Systems: Computer Graphics: Programmer’s Hierarchical Interactive Graphical System. Draft proposal X3.144.1988, 1988.

ANSI:c89

[Ame89] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. American National Standard Programming Language C, ANSI X3.159-1989, December 14 1989.

ANSI:ftn9x

[Ame90] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed American National Standard Programming Language Fortran Extended X3.198– 199x, September 24 1990. See also [MR87].

ANSI:ftn92

[Ame92] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. American National Standard Programming Language Fortran Extended X3.198–1992, 1992. This US Standard is identical to the international standard, ISO 1539:1991. See also [ABW92, BGA90, Cou91, MR90].

Ames:1993:BPO

[Ame93] Patrick Ames. Beyond Paper: The Official Guide to Adobe Acrobat. Adobe Press, Mountain View, CA, USA, 1993. ISBN 1-56830-050-6. 127 pp. LCCN TK5105.9 .A48 1993. US$16.95.

APA:1994:PMA

[Ame94] American Psychological Association. Publication Manual of the American Psychological Association. American Psychological Association, Washington, DC, USA, fourth edition, 1994. ISBN 1-55798-243-0 (hardcover), 1-55798-241-4 (paperback). xxxii + 368 pp. LCCN BF76.7.P82 1994. US$23.00 (paperback).

Ament:2001:INB

[Ame01a] Kurt Ament. Indexing: a Nuts-and-bolts Guide for Technical Writers. William Andrew Publishing, Norwich, NY, USA, 2001. ISBN 0-8155-1481-6. x + 97 pp. LCCN T10.8 .A44 2001. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/wap041/2001033646.
html;   http://www.loc.gov/
catdir/toc/wap041/2001033646.
html;   http://www.williamandrew.
com/books.asp?id=1481.

APA:2001:PMA

[Ame01b] American Psychological Association. Publication Manual of the American Psychological Association. American Psychological Association, Washington, DC, USA, fifth edition, 2001. ISBN 1-55798-790-4 (hardcover), 1-55798-810-2 (coil bound), 1-55798-791-2 (paperback). xxviii + 439 pp. LCCN BF76.7.P83 2001; PE1475 .P976 2001; WZ345 P83 2001.

APA:2010:PMA

[Ame10] American Psychological Association. Publication manual of the American Psychological Association. American Psychological Association, Washington, DC, USA, sixth edition, 2010. ISBN 1-4338-0559-6 (hardcover), 1-4338-0561-8 (softcover), 1-4338-0562-6 (spiral bound). xviii + 272 pp. LCCN BF76.7 .P83 2010.

Andrews:1963:ESM

[And63] Frank C. Andrews. Equilibrium Statistical Mechanics. Wiley, New York, NY, USA, 1963. xi + 206 pp. LCCN QC174.8.Z534.

Anderson:2011:MDN

[And11] P. W. (Philip W.) Anderson.

More and different: notes from a thoughtful curmudgeon. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 981-4350-12-5 (hardcover), 981-4350-13-3 (paperback). ix + 412 pp. LCCN Q171 .A527 2011.

Angell:1981:PIC

[Ang81] Ian O. Angell. A Practical Introduction to Computer Graphics. Wiley, New York, NY, USA, 1981. ISBN 0-470-27251-1. viii + 146 pp. LCCN T385 .A53 1981.

Anonymous:1976:RAB

[Ano76] Anonymous. Road atlas Britain. John Bartholomew & Son, Edinburgh, Scotland, 1976. ISBN 0-85152-505-9. 120 pp. LCCN G1811.P2 B350 1975.

Ada79:refman

[Ano79a] Anonymous. Preliminary Ada reference manual. ACM SIG-PLAN Notices, 14(6A), June 1979. The final standard is [Ame83a].

Ada79:rationale

[Ano79b] Anonymous. Rationale for the design of the Ada programming language. ACM SIGPLAN Notices, 14(6B), June 1979. The final standard is [Ame83a].

Anonymous:1990:ASM

[Ano90] Anonymous. AIP style manual. American Institute of Physics, Woodbury, NY, USA, fourth edition, 1990. ISBN 0-88318-642-X (paperback). 64 pp. LCCN QC5.45 .A45; QC28 .A5 1990; T11 .A45 1990.

Anonymous:1993:CMS

[Ano93] Anonymous. The Chicago Manual of Style. University of Chicago Press, Chicago, IL, USA and London, UK, fourteenth edition, 1993. ISBN 0-226-10389-7. ix + 921 pp. LCCN Z253 .U69 1993. US$40.00.

Anonymous:1996:BTF

[Ano96a] Anonymous. Behind tall fences: stories and experiences about Los Alamos at its beginning. Los Alamos Historical Society, Los Alamos, NM, USA, 1996. ISBN 0-941232-91-3. xiv + 210 pp. LCCN QC773.3.U5 B43 1996.

Anonymous:1996:UIP

[Ano96b] Anonymous, editor. Ulrich’s International Periodical Directory. R. R. Bowker Company, 1180 Avenue of the Americas, New York, NY 10036, USA, 34th edition, 1996. ISBN 0-8352-3681-1 (vol. 4), 0-8352-3676-5 (set). ISSN 0000-0175. various pp. LCCN AP1.21 .P4. Five volumes. Volume 4 contains a numerical ISSN index.

Anonymous:1999:RAB

[Ano99] Anonymous. Road atlas 2000 Britain. Ordnance Survey and Bounty, Southampton, UK, 1999. ISBN 0-7537-0138-3. iv + 124 pp.

Anonymous:2003:CMS

[Ano03] Anonymous. The Chicago Manual of Style. University of Chicago Press, Chicago, IL, USA and London, UK, 15th edition, 2003. ISBN 0-226-10403-6. xvii + 956 pp. LCCN Z253 .U69 2003. US$55.00.

Anonymous:2005:BFI

[Ano05] Anonymous. Basic Feynman:

An interview with Michelle Feynman. Web site, 2005. Comments about the preparation of [Fey05].

Anonymous:2010:CMS

[Ano10] Anonymous. The Chicago Manual of Style. University of Chicago Press, Chicago, IL, USA and London, UK, 16th edition, 2010. ISBN 0-226-10420-6 (hardcover). xvi + 1026 pp. LCCN Z253 .U69 2010.

Anonymous:2017:CMS

[Ano17] Anonymous. The Chicago Manual of Style. University of Chicago Press, Chicago, IL, USA and London, UK, 17th edition, 2017. ISBN 0-226-28705-X (hardcover). xvi + 1144 pp. LCCN Z253 .U69 2017.

ANSI:ftn69

[ANS69] ANSI Subcommittee X3J3. Clarification of Fortran standards: Initial progress. Communications of the Association for Computing Machinery, 12:289–294, 1969. See also [Ame66].

ANSI:ftn71

[ANS71] ANSI Subcommittee X3J3. Clarification of Fortran standards: Second report. Communications of the Association for Computing Machinery, 14:628–642, 1971. See also [Ame66].

ANSI:ftn76

[ANS76] ANSI Subcommittee X3J3. Draft proposed ANS Fortran. ACM SIGPLAN Notices, 11(3), 1976. See also final standard [Ame78].

Arasu:2002:PCS

[ANTT02] Arvind Arasu, Jasmine Novak, Andrew Tomkins, and John Tomlin. PageRank computation and the structure of the Web: Experiments and algorithms. In Proceedings of the Eleventh International World Wide Web Conference, Honolulu, Hawaii, USA, 7–11 May 2002. ACM Press, New York, NY 10036, USA, 2002. ISBN 1-880672-20-0. LCCN TK5105.888 Wic 11

2002. URL http://www2002.
org/CDROM/poster/173.pdf.

Avenarius:fortran-web

[AO90] Adrian Avenarius and Siegfried Oppermann. FWEB: a literate programming system for Fortran 8X. ACM SIGPLAN Notices, 25 (1):52–58, January 1990.

Aiken:1964:PAC

[AOB64] H. H. Aiken, A. G. Oettinger, and T. C. Bartee. Proposed automatic calculating machine. IEEE Spectrum, 1(8):62–69, August 1964. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic). Previously unpublished memorandum. Reprinted in [Ran82, §5.1].

Apple:1987:TIM

[App87] Apple Computer, Inc. Technical Introduction to the Macintosh Family. Addison-Wes-ley, Reading, MA, USA, 1987. ISBN 0-201-17765-X. xix + 289 pp. LCCN QA76.8.M3 T43 1987. US$19.95.

Appelt:1988:TFP

[App88] Wolfgang Appelt. TEX für Fortgeschrittene: Programmiertechniken und Makropakete. Addison-Wesley, Reading, MA, USA, 1988. ISBN 3-89319-115-1. x + 179 pp. LCCN ????

Arabatzis:2006:REB

[Ara06] Theodore Arabatzis. Representing Electrons: a Biographical Approach to Theoretical Entities. University of Chicago Press, Chicago, IL, USA and London, UK, 2006. ISBN 0-226-02420-2 (hardcover), 0-226-02421-0 (paperback), 0-226-02422-9 (e-book). xiv + 295 pp. LCCN QC793.5.E62 A73 2011. URL http:

//public.eblib.com/choice/
publicfullrecord.aspx?p=709999.

Arianrhod:2003:EHI

[Ari03] Robyn Arianrhod. Einstein’s heroes: imagining the world through the language of mathematics. University of Queensland Press, St Lucia, Queensland, Australia, 2003. ISBN 0-7022-3408-7. viii + 323 pp. LCCN QC19.6 .A75 2003. US$24.00.

Arianrhod:2005:EHI

[Ari05] Robyn Arianrhod. Einstein’s heroes: imagining the world through the language of mathematics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-518370-3 (hardcover), 0-19-530890-5 (paperback). xiii + 323 pp. LCCN QC19.6 .A75 2005. URL ftp://

uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http:
//www.loc.gov/catdir/toc/
ecip053/2004026055.html.

Artwick:1984:ACM

[Art84] Bruce A. Artwick. Applied Concepts in Microcomputer Graphics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-039322-3. ix + 374 pp. LCCN T385 .A77 1984.

Artwick:1985:ACM

[Art85] Bruce A. Artwick. Applied Concepts in Microcomputer Graphics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1985. ISBN 0-13-580226-1 (paperback), 0-13-039322-3 (hardcover). ix + 374 pp. LCCN T385 .A77 1985.

Arvo:1991:GGI

[Arv91] James Arvo. Graphics gems II, volume 2 of Graphics Gems. Academic Press, New York, NY, USA, 1991. ISBN 0-12-064480-0. xxxii + 643 pp. LCCN T385 .G6972 1991. URL http://www.

sciencedirect.com/science/
book/9780080507545.

Abramowitz:1964:HMF

[AS64] Milton Abramowitz and Irene A.

Stegun, editors. Handbook of Mathematical Functions with Formulas, Graphs, and Mathematical Tables, volume 55 of Applied mathematics series. U. S. Department of Commerce, Washington, DC, USA, 1964. xiv + 1046 pp. LCCN QA47.A161 1972; QA 55 A16h 1972. Tenth printing, with corrections (December 1972). This book is also available on-

line at http://www.convertit.

                                        com/Go/ConvertIt/Reference/

AMS55.ASP in bitmap image format.

Asente:1990:XWS

[AS90] Paul J. Asente and Ralph R. Swick. X Window System Toolkit: The Complete Programmer’s Guide and Specification. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1990. ISBN 0-13-972191-6 (Prentice Hall), 0-13-973173-3, 0-13-975491-6 (Prentice Hall), 1-55558-051-3 (DP paperback). xxxv + 967 pp. LCCN QA76.76.W56 A74 1990.

Asher:virtual-fonts

[Ash90] Graham Asher. Re: Virtual fonts: More fun for grand wizards. TEXhax, 90(17), January 1990.

Aho:1986:CPC

[ASU86] Alfred V. Aho, Ravi Sethi, and Jeffrey D. Ullman. Compilers: Principles, Techniques, and Tools. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-10088-6 (hardcover), 0-201-10194-7 (paperback). x + 796 pp. LCCN QA76.76.C65 A371 1986. See [AU77].

Atanasoff:1940:CMS

[Ata40]  J. V. Atanasoff.  Comput-

ing machine for the solution of large systems of linear algebraic equations. Unpublished memorandum, Iowa State College, Ames, IA, USA, August 1940. Reprinted in [Ran82, §7.2].

ATT:UPM83-1

[AT&T83a] AT&T. UNIX Programmer’s Manual, volume 1. Holt, Rinehart, and Winston, New York, NY, USA, 1983. ISBN 0-03-061742-1. xiv + 425 pp. LCCN QA76.8.U65 B44 1983.

ATT:UPM83-2

[AT&T83b] AT&T. UNIX Programmer’s Manual, volume 2. Holt, Rinehart, and Winston, New York, NY, USA, 1983. ISBN 0-03-061743-X. vii + 616 pp. LCCN QA76.8.U65 B44 1983.

ATT:AUS86-1

[AT&T86a] AT&T. AT&T UNIX System Readings and Applications, volume I. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-938532-0. xiv + 397 pp. LCCN QA76.76.O63 U553 1986.

ATT:AUS86-2

[AT&T86b] AT&T. AT&T UNIX System Readings and Applications, volume II. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-939845-7. xii + 324 pp. LCCN QA76.76.O63 U553 1986.

ATT:UPM86-1

[AT&T86c] AT&T. UNIX Programmer’s Manual: Commands and Utilities, volume 1. Holt, Rinehart, and Winston, New York, NY, USA, 1986. ISBN 0-03-009317-1. xxix + 524 pp. LCCN QA76.76.O63 U548 1986.

ATT:UPM86-4

[AT&T86d] AT&T. UNIX Programmer’s Manual: Document Preparation, volume 4. Holt, Rinehart, and Winston, New York, NY, USA, 1986. ISBN 0-03-011207-9. xiii + 355 pp. LCCN QA76.76.O63 U548 1986.

ATT:UPM86-5

[AT&T86e] AT&T. UNIX Programmer’s Manual: Languages and Support Tools, volume 5. Holt, Rinehart, and Winston, New York, NY, USA, 1986. ISBN 0-03-011204-4. xviii + 618 pp. LCCN QA76.76.O63 U548 1986.

ATT:UPM86-3

[AT&T86f] AT&T. UNIX Programmer’s Manual: System Administration Facilities, volume 3. Holt, Rinehart, and Winston, New York, NY, USA, 1986. ISBN 0-03-009313-9. xiv + 142 pp. LCCN QA76.76.O63 U548 1986.

ATT:UPM86-2

[AT&T86g] AT&T. UNIX Programmer’s Manual: System Calls and Library Routines, volume 2. Holt, Rinehart, and Winston, New York, NY, USA, 1986. ISBN 0-03-009314-7. xxxv + 465 pp. LCCN QA76.76.O63 U548 1986.

ATT:USV86

[AT&T86h] AT&T. UNIX System V Programmer’s Guide. Holt, Rinehart, and Winston, New York, NY, USA, 1986. ISBN 0-13-940438-4. xiv + 832 pp. LCCN QA76.76.O63 U556 1987.

Aho:1972:TPT

[AU72] Alfred V. Aho and Jeffrey D. Ullman. The Theory of Parsing, Translation, and Compiling. I: Parsing, volume I. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1972. ISBN 0-13-914556-7. xviii + 542 pp. LCCN QA76.6 .A286 1972-73.

Aho:1973:TPT

[AU73] Alfred V. Aho and Jeffrey D. Ullman. The Theory of Parsing, Translation, and Compiling. II. Compiling, volume II. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1973. ISBN 0-13-914564-8. xiii + 460 pp. LCCN QA76.6 .A286 1972-73.

Aho:1977:PCD

[AU77] Alfred V. Aho and Jeffrey D.

Ullman. Principles of Compiler Design. Addison-Wesley, Reading, MA, USA, 1977. ISBN 0-201-00022-9 (hardcover), 0-201-10073-8. x + 604 pp. LCCN QA76.6 .A285 1977. See also the much expanded subsequent book [ASU86].

Augarten:1984:BBI

[Aug84] Stan Augarten. BIT by BIT:

An Illustrated History of Computers. Ticknor and Fields, New York, NY, USA, 1984. ISBN 0-89919-268-8, 0-89919-302-1 (paperback). ix + 324 pp. LCCN QA76.17.A94 1984.

Austrian:1982:HHF

[Aus82] Geoffrey D. Austrian. Herman Hollerith — Forgotten Giant of Information Processing. Columbia University Press, New York, NY, USA, 1982. ISBN 0-231-05146-8. xvi + 418 pp. LCCN QA76.2.H64 .A97.

Austern:1999:GPS

[Aus99] Matthew H. Austern. Generic Programming and the STL: Using and Extending the C++ Standard Template Library. Addison-Wesley professional computing series. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-30956-4. xx + 548 pp. LCCN QA76.73.C153A97 1999. US$49.95.

Altland:2019:MPI

[AvD19] Alexander Altland and Jan von Delft. Mathematics for Physicists: Introductory Concepts and Methods. Cambridge University Press, Cambridge, UK, 2019. ISBN 1-108-47122-6 (hardcover). xxi + 700 pp. LCCN QC20 .A4345 2019.

Avery:1976:CAO

[Ave76] John Avery. Creation and Annihilation Operators. McGraw-Hill International Book Company, New York, NY, USA, 1976. ISBN 0-07-002504-5. xii + 221 pp. LCCN QC174.52.06A9.

Aste:2000:PPP

[AW00] Tomaso Aste and Denis Weaire. The Pursuit of Perfect Packing. Institute of Physics Publishing, Bristol, UK and Philadelphia, PA, USA, 2000. ISBN 0-7503-0647-5 (hardcover), 0-7503-0648-3 (paperback). xi + 136 pp. LCCN QA166.7 .A78 2000.

Arfken:2005:MMP

[AW05] George B. Arfken and Hans-Jürgen Weber. Mathematical Methods for Physicists. Elsevier, Amsterdam, The Netherlands, sixth edition, 2005. ISBN 0-12-059876-0, 0-12-088584-0 (paperback). xii + 1182 pp. LCCN QA37.3 .A74 2005.

Arfken:2013:MMP

[AWH13] George B. (George Brown) Arfken, Hans-Jürgen Weber, and Frank E. Harris. Mathematical Methods for Physicists: a Comprehensive Guide. Elsevier Academic Press, Amsterdam, The Netherlands, seventh edition, 2013. ISBN 0-12-384654-4 (hardcover). xiii + 1205 pp. LCCN QA37.3 .A74 2013.

Babbage:1837:MPC

[Bab37] Charles Babbage. On the mathematical powers of the calculating engine. Manuscript held by Museum of the History of Science, Oxford, UK. Reprinted in [Ran82, §2.1]., December 1837.

Babbage:1910:BBA

[Bab10] H. P. Babbage. Babbage: Babbage’s analytical engine. Monthly Notices of the Royal Astronomical Society, 70(??):517–526, 645, 1910. CODEN MNRAA4. ISSN 0035-8711. Reprinted in [Ran82, §2.3].

Backus:fortran-history

[Bac78] John Backus. The history of FORTRAN I, II, and III. ACM SIGPLAN Notices, 13(8):165– 180, August 1978.

Backhouse:1979:SPL

[Bac79] Roland C. Backhouse. Syntax of Programming Language. Series in Computer Science, Editor: C. A. R. Hoare. Prentice-Hall International, Upper Saddle River, NJ 07458, USA, 1979. ISBN 0-13-879999-7. xv + 301 pp. LCCN QA76.7 .B3. US$18.95.

Bach:1986:DUO

[Bac86] Maurice J. Bach. The Design of the UNIX Operating System. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-201799-7. xiv + 471 pp. LCCN QA76.76.O63 B33 1986.

Baer:1980:CSA

[Bae80] Jean-Loup Baer. Computer Systems Architecture. Computer Science Press, 11 Taft Court, Rockville, MD 20850, USA, 1980. ISBN 0-914894-15-3. xiii + 626 pp. LCCN QA76.9.A73 B33.

Baer:2010:MAS

[Bae10]  Jean-Loup Baer.  Micropro-

cessor architecture: from simple pipelines to chip multiprocessors. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-76992-2 (hardcover). xiv + 367 pp. LCCN QA76.5 .B227 2010.

Baggott:2010:FWP

[Bag10] Jim Baggott. The First War of Physics: The Secret History of the Atomic Bomb, 1939–1949. Pegasus Books, New York, NY, USA, 2010. ISBN 1-60598-084-6. xiv + 576 pp. LCCN QC773 .B24 2010. URL http:/

/pegasusbooks.us/pdf/Spg10_
Catalog.pdf.

Baggott:2011:QSH

[Bag11] Jim Baggott. The quantum story: a history in 40 moments. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-956684-4 (hardcover), 0-19-965597-9 (paperback). xix + 469 + 16 pp. LCCN QC173.98 .B34 2011.

Bailey:1992:MPS

[Bai92] David H. Bailey. Misleading performance in the supercomputing field. In Proceedings Supercomputing ’92, pages 155–158. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2630-5, 0-89791-537-2. LCCN QA76.5 S87a 1992.

Bain:1999:EEB

[Bai99] David Haward Bain. Empire Express: Building the First Transcontinental Railroad. Penguin, London, UK and New York, NY, USA, 1999. ISBN 0-670-80889-X, 0-14-008499-1. xiii + 797 + 32 pp. LCCN HE2751 .B24 1999. URL http:

//books.google.com/books?

                                        isbn=067080889X;   http://
                                        lcweb.loc.gov/catdir/toc/
                                        99033375.html;   http://www.
                                        booknotes.org/Watch/155004-
                                        1/David+Haward+Bain.aspx.

Bak:1962:EIM

[Bak62] Børge Bak. Elementary Introduction to Molecular Spectra. North-Holland, Amsterdam, The Netherlands, second edition, 1962. xi + 144 pp. LCCN QC451 .B2 1962.

Ballhausen:1962:ILF

[Bal62] Carl J. Ballhausen. Introduction to Ligand Field Theory. Mc-Graw-Hill, New York, NY, USA, 1962. ix + 298 pp. LCCN QD475 .B3.

Ball:1997:MMN

[Bal97] Philip Ball. Made to Measure: New Materials for the 21st Century. Princeton University Press, Princeton, NJ, USA, 1997. ISBN 0-691-02733-1. vii + 458 pp. LCCN TA403.B2247 1997.

Ball:2014:SRS

[Bal14] Philip Ball. Serving the Reich: the struggle for the soul of physics under Hitler. University of Chicago Press, Chicago, IL, USA and London, UK, 2014. ISBN 0-226-20457-X (hardcover), 0-226-20460-X (ebook). ix + 303 pp. LCCN QC773.3.G3 B35 2014.

Balbaert:2015:RE

[Bal15a] Ivo Balbaert. Rust Essentials. Packt Publishing, Birmingham, UK, 2015. ISBN 1-78528-213-1. x + 161 pp. LCCN QA76.73.R87 B35 2015. URL http://

proquest.safaribooksonline.
com/9781785285769.

Baldwin:2015:MNH

[Bal15b] Melinda Baldwin. Making Nature: the History of a Scientific Journal. University of Chicago Press, Chicago, IL, USA and London, UK, 2015. ISBN 0-226-26145-X (hardcover), 0-226-26159-X (e-book). 309 pp. LCCN Q1.N23 B35 2015.

Ball:2018:WIG

[Bal18] Thorsten Ball. Writing an Interpreter in Go: version 1.6. Thorsten Ball, ????, 2018. ISBN 3-9820161-1-8 (paperback). 263 pp. LCCN QA76.76.C65 .B35 2018.

Ball:2019:WCG

[Bal19] Thorsten Ball. Writing a compiler in Go. Thorsten Ball, Aschaffenburg, Germany, version 1.1. edition, 2019. ISBN 3-9820161-0-X. 351 pp. LCCN QA76.73.G63 B348 2019.

Banks:1999:SPR

[Ban99] Robert B. Banks. Slicing Pizzas, Racing Turtles, and Further Adventures in Applied Mathematics. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-05947-0. xi + 286 pp. LCCN QA93 .B358 1999. US$24.95.

Barrow:1979:PC

[Bar79] Gordon M. Barrow. Physical Chemistry. McGraw-Hill, New York, NY, USA, fourth edition, 1979. ISBN 0-07-003825-2 (hardcover), 0-07-066170-7 (paperback). xvi + 832 pp. LCCN QD453.2.B37 1979.

Bartlett:1980:FQC

[Bar80] John Bartlett. Familiar Quotations: a collection of passages, phrases, and proverbs traced to their sources in ancient and modern literature. Little, Brown and Company, Boston, Toronto, London, fifteenth edition, 1980. ISBN 0-316-08275-9. lviii + 1540 pp. LCCN PN6081 .B3 1980.

Barnsley:1988:FE

[Bar88] Michael Barnsley. Fractals Everywhere. Academic Press, New York, NY, USA, 1988. ISBN 0-12-079062-9. xii + 394 pp. LCCN QA614.86 .B37 1988.

Barrow:1992:PSC

[Bar92] John D. Barrow. Pi in the Sky: Counting, Thinking, and Being. Clarendon Press, Oxford, UK, 1992. ISBN 0-19-853956-8. ix + 317 pp. LCCN QA36 .B37 1992. US$30.00 (Oxford Univ. Press).

Barrow:1996:PSC

[Bar96] John D. Barrow. Pi in the Sky: Counting, Thinking, and Being. Little, Brown and Company, Boston, Toronto, London, 1996. ISBN 0-316-08259-7. ix + 317 pp. LCCN QA36 .B37 1994.

Barrow:1998:ILS

[Bar98] John D. Barrow. Impossibility: the limits of science and the science of limits. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-851890-0. xiii + 279 pp. LCCN Q175 .B2245 1998.

Bardi:2006:CWN

[Bar06a] Jason Socrates Bardi. The Calculus Wars: Newton, Leibniz, and the Greatest Mathematical Clash of All Time. Thunder’s Mouth Press, New York, NY, USA, 2006. ISBN 1-56025-992-2, 1-56025-706-7. viii + 277 pp. LCCN QA303 .B2896 2006.

Bartusiak:2006:AUD

[Bar06b] Marcia Bartusiak, editor. Archives of the universe: 100 discoveries that transformed our understanding of the cosmos. Vintage Books, New York, NY, USA, 2006. ISBN 0-375-71368-9. xvii + 695 pp. LCCN QB15 .A75 2006. US$18.95.

Bartusiak:2009:DWF

[Bar09] Marcia Bartusiak. The Day We Found the Universe. Pantheon Books, New York, NY, USA, 2009. ISBN 0-375-42429-6. xviii + 337 pp. LCCN QB15 .B37 2009.

Batten:1988:RUC

[Bat88] Alan H. (Alan Henry) Batten. Resolute and Undertaking Characters: The Lives of Wilhelm and Otto Struve, volume 139 of Astrophysics and space science library. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1988. ISBN 90-277-2652-3. xxv + 259 pp. LCCN QB36.S75 B38 1988. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0902/87030386-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0902/
                                        87030386-t.html.

Ballard:1982:CV

[BB82] Dana H. Ballard and Christopher M. Brown. Computer Vision. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1982. ISBN 0-13-165316-4. xx + 523 pp. LCCN TA1632 .B34 1982.

Bjorstad:SR-4-11-57

[BB91] Petter E. Bjorstad and Erik Boman. SLALOM: a better algorithm. Supercomputing Review, 4(11):57–62, November 1991.

Blaauw:1997:CAC

[BB97] Gerrit A. Blaauw and Frederick P. Brooks, Jr. Computer Architecture: Concepts and Evolution. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-10557-8. xlviii + 1213 pp. LCCN QA76.9.A73 B57 1997. US$59.95.

Bergstrom:2004:CBL

[BB04a] Carl T. Bergstrom and Theodore C. Bergstrom. The costs and benefits of library site licenses to academic journals. Proceedings of the National Academy of Sciences of the United States of America, 101(3):897–902, January 20, 2004. CODEN PNASA6.

ISSN 0027-8424 (print), 1091-6490 (electronic). URL http:

//www.pnas.org/cgi/content/
abstract/0305628101v1.

Bernardini:2004:EFH

[BB04b] C. (Carlo) Bernardini and Luisa Bonolis, editors. Enrico Fermi: His Work and Legacy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 88-7438-015-1 (SIF, Bologna, Italy), 3-540-22141-7 (Springer-Verlag, Berlin, Heidelberg, New York). xii + 410 pp. LCCN QC774.F4 C6613 2004. URL

http://www.loc.gov/catdir/
enhancements/fy0818/2004108212-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0818/
2004108212-t.html.

Borwein:2004:MEP

[BB04c] Jonathan M. Borwein and David H. Bailey. Mathematics by Experiment: Plausible Reasoning in the 21st Century. A. K. Peters, Ltd., Wellesley, MA, USA, 2004. ISBN 1-56881-211-6. x + 288 pp. LCCN QA76.95 .B67 2003. US$45.00. URL http:/

/docserver.carma.newcastle.
edu.au/272/.

Born:2005:BEL

[BB05] Max Born and Hedwig Born, editors. The Born–Einstein letters: friendship, politics, and physics in uncertain times: correspondence between Albert Einstein and Max and Hedwig Born from 1916 to 1955 with commentaries by Max Born. Macmillan Publishing Company, New York, NY, USA, 2005. ISBN 1-4039-4496-2. xxxii + 235 pp. LCCN QC16.E5 A4 2005. URL http:

//www.loc.gov/catdir/bios/
hol059/2004061027.html;
http://www.loc.gov/catdir/
description/hol054/2004061027.
html;   http://www.loc.gov/
catdir/enhancements/fy0618/
2004061027-t.html. Note on

the new edition by Gustav Born. New preface by Diana Buchwald and Kip S. Thorne. Foreword by Bertrand Russell. Introduction by Werner Heisenberg. Translated by Irene Born.

Bartels:1987:ISU

[BBB87] Richard H. Bartels, John C. Beatty, and Brian A. Barsky. An Introduction to Splines for Use in Computer Graphics and Geometric Modeling. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1987. ISBN 0-934613-27-3. xiv + 476 pp. LCCN T385 .B3651 1987. US$38.95.

Bailey:IJSA-5-3-63

[BBB^{+}91] D. H. Bailey, E. Barszcz, J. T. Barton, D. S. Browning, R. L. Carter, L. Dagum, R. A. Fatoohi, P. O. Frederickson, T. A. Lasinski, R. S. Schreiber, H. D. Simon, V. Venkatakrishnan, and S. K. Weeratunga. The NAS parallel benchmarks. International Journal of Supercomputer Applications and High Performance Computing, 5(3):63–73, Fall 1991.

Berggren:2000:PSB

[BBB00] Lennart Berggren, Jonathan Borwein, and Peter Borwein, editors. Pi: a source book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 0-387-98946-3 (hardcover). xx + 736 pp. LCCN QA484 .P5 2000.

Bailey:2007:EMA

[BBC^{+}07] David H. Bailey, Jonathan M.

Borwein, Neil J. Calkin, Roland Girgensohn, D. Russell Luke, and Victor Moll. Experimental Mathematics in Action. A. K. Peters, Ltd., Wellesley, MA, USA, 2007. ISBN 1-56881-271-X. xii + 322 pp. LCCN QA8.7 .E97 2007. URL http:/

/docserver.carma.newcastle.
edu.au/1733/.

Borwein:2004:EMCa

[BBG04] Jonathan M. Borwein, David H.

Bailey, and Roland Girgensohn. Experimentation in mathematics: computational paths to discovery. A. K. Peters, Ltd., Wellesley, MA, USA, 2004. ISBN 1-56881-136-5. x + 357 pp. LCCN QA12 .B67 2004.

Born:1969:AP

[BBSR69] Max Born, R. J. (Roger John)

Blin-Stoyle, and J. M. Radcliffe. Atomic Physics. Dover books on physics and chemistry. Dover, New York, NY, USA, eighth edition, 1969. ISBN 0-486-65984-4. xiv + 495 + 11 pp. LCCN QC776 .B5713 1989.

US$11.95. URL http://www.

                                        loc.gov/catdir/description/
                                        dover031/89012033.html.

Becker:1981:LSD

[BC81] Richard A. Becker and John M. Chambers. S: a Language and System for Data Analysis. AT&T Bell Laboratories, Murray Hill, NJ, USA, 1981.

Becker:1984:IED

[BC84] Richard A. Becker and John M. Chambers. S: An Interactive Environment for Data Analysis and Graphics. Wadsworth, Pacific Grove, CA, USA, 1984. ISBN 0-534-03313-X. xii + 550 pp. LCCN QA76.73.S15 B44 1984.

Becker:1985:ES

[BC85] Richard A. Becker and John M. Chambers. Extending the S System. Wadsworth, Pacific Grove, CA, USA, 1985. ISBN 0-534-05016-6. viii + 166 pp. LCCN QA276.4 .B4241 1985.

Bienz:1993:PDF

[BC93] Tim Bienz and Richard Cohn. Portable Document Format Reference Manual. Addison-Wes-ley, Reading, MA, USA, 1993. ISBN 0-201-62628-4. xii + 214 pp. LCCN QA76.9.F5P67 1993. US$24.95.

Bosma:2006:DMM

[BC06] Wieb Bosma and John Cannon, editors. Discovering Mathematics with Magma: Reducing the Abstract to the Concrete, volume 19 of Algorithms and computation in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-37632-1 (hardcover), 3-540-37634-8 (e-book). ISSN 1431-1550. xxiv + 374 pp. LCCN QA155.7.E4 D57 2006. URL

http://www.loc.gov/catdir/
enhancements/fy0824/2006931477-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0824/
2006931477-d.html;   http:
//www.loc.gov/catdir/toc/
fy0707/2006931477.html;
http://www.springerlink.com/
content/978-3-642-01960-9;
 http://www.springerlink.
com/openurl.asp?genre=book&
isbn=978-3-540-37632-3.

Bruguera:2009:PIS

[BCDH09] Javier D. Bruguera, Marius Cornea, Debjit DasSarma, and John Harrison, editors. Proceedings of the 19th IEEE Symposium on Computer Arithmetic, June 8–10, 2009, Portland, Oregon, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 0-7695-3670-0, 1-4244-4329-6. ISSN 1063-6889. LCCN QA76.6 .S887

2009. URL http://www.ac.
usc.es/arith19/.

Bottou:2007:LSK

[BCDW07] Léon Bottou, Olivier Chapelle, Dennis DeCost, and Jason Weston, editors. Large-scale Kernel Machines. Neural information processing series. MIT Press, Cambridge, MA, USA, 2007. ISBN 0-262-02625-2 (hardcover), 0-262-25579-0 (e-book). xii + 396 pp. LCCN QA76.9.D35

                                        L38 2007. URL http://www.
                                        loc.gov/catdir/toc/ecip078/
                                        2007000980.html.

Becker:1988:NPL

[BCW88] Richard A. Becker, John M. Chambers, and Allan R. Wilks. The New S Programming Language. Wadsworth and Brooks/ Cole, Pacific Grove, CA, USA and Pacific Grove, CA, USA, 1988. ISBN 0-534-09192-X, 0-534-09193-8. xvii + 702 pp. LCCN QA76.73.S15 B43 1988.

Booth:2008:CR

[BCW08] Wayne C. Booth, Gregory G. Colomb, and Joseph M. Williams. The craft of research. Chicago guides to writing, editing, and publishing. University of Chicago Press, Chicago, IL, USA and London, UK, third edition, 2008. ISBN 0-226-06565-0 (cloth), 0-226-06566-9 (paperback). xvii + 317 pp. LCCN Q180.55.M4 B66 2008. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0828/2007042761-
                                        b.htm;   http://www.loc.gov/
                                        catdir/enhancements/fy0828/
                                        2007042761-d.htm;   http:
                                        //www.loc.gov/catdir/toc/
                                        ecip082/2007042761.htm.

Bigelow:digital-typography

[BD83] Charles Bigelow and Donald Day. Digital typography. Scientific American, 249(2):106–119, August 1983.

Borwein:2009:CCI

[BD09] Jonathan M. Borwein and Keith J. Devlin. The computer as crucible: an introduction to experimental mathematics. A. K. Peters, Ltd., Wellesley, MA, USA, 2009. ISBN 1-56881-343-0. xi + 158 pp. LCCN QA8.7 .B67 2009. URL http:/

/docserver.carma.newcastle.
edu.au/1730/;   http://www.
loc.gov/catdir/toc/fy0904/
2008022180.html.

Beckmann:1993:HP

[Bec93] Petr Beckmann. A history of π. Barnes and Noble, New York, NY, USA, 1993. ISBN 0-88029-418-3. 200 pp. LCCN QA484 .B4 1971. US$6.98. Reprint of the third edition of 1971.

Bowman:1996:PSH

[BED96] Judith S. Bowman, Sandra L.

Emerson, and Marcy Darnovsky. The Practical SQL Handbook: Using Structured Query Language. Addison-Wesley, Reading, MA, USA, third edition, 1996. ISBN 0-201-44787-8. xxvi + 454 pp. LCCN QA76.73.S67 B69 1996. US$39.95. Includes CD-ROM.

Beebe:sf3-installation

[Bee79] Nelson H. F. Beebe. SFTRAN 3 Installation Guide. College of Science Computer, Department of Physics, University of Utah, Salt Lake City, UT 84112, USA, July 1979.

Beebe:plot79

[Bee80] Nelson H. F. Beebe. A user’s guide to <PLOT79>. Technical report, University of Utah, 1980.

Beebe:dvi-drivers

[Bee87] Nelson H. F. Beebe. A TEXDVI driver family. TEXniques, 5:71– 114, August 1987. Proceedings of the Eighth Annual Meeting of the TEX Users Group.

Beebe:tex-graphics

[Bee89] Nelson H. F. Beebe. TEX and graphics: The state of the problem. Cahiers GUTenberg, 2:13– 53, 1989. Presented to: Congrès GUTenberg, Paris, France, 16–17 May 1989.

Beebe:2017:MFC

[Bee17]  Nelson H. F. Beebe.  The

Mathematical-Function Computation Handbook: Programming Using the MathCW Portable Software Library. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-64109-3 (hardcover), 3-319-64110-7 (ebook). xxxvi + 1114 pp. LCCN QA75.5-76.95. URL http:/

                                        /www.springer.com/us/book/
                                        9783319641096.

Bell:1937:MML

[Bel37] Eric Temple Bell. Men of mathematics: The Lives and Achievements of the Great Mathematicians from Zeno to Poincaré. Simon and Schuster, New York, NY, USA, 1937. ISBN 0-671-62818-6. xxi + 592 pp. LCCN QA28 .B4. US$17.00.

Bellamy:1974:ICO

[Bel74] Anthony J. Bellamy. An Introduction to Conservation of Orbital Symmetry. Longman Group Limited, London, England, 1974. ISBN 0-582-44089-0. 77 pp. LCCN QD476.B363.

Bell:rs6000-tuning

[Bel90] Ron Bell. IBM RISC System/ 6000 performance tuning for numerically intensive Fortran and C programs. Technical Report GG24-3611-00, IBM Corporation, August 1990.

Bennett:1933:CFC

[Ben33] H. Bennett, editor. The Chemical Formulary: a Condensed Collection of Valuable, Timely, Practical Formulae for Making Thousands of Products in All Fields of Industry, volume 1. Chemical Publishing Co., Inc., New York, NY, USA, 1933. x + 604 pp. LCCN TP151 .B439.

Bennett:1935:CFC

[Ben35] H. Bennett, editor. The Chemical Formulary: a Condensed Collection of Valuable, Timely, Practical Formulae for Making Thousands of Products in All Fields of Industry, volume 2. Chemical Publishing Co., Inc., New York, NY, USA, 1935. ix + 570 pp. LCCN TP151 .B439.

Bentley:1982:WEP

[Ben82] Jon Louis Bentley. Writing Efficient Programs. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1982. ISBN 0-13-970251-2 (hardcover), 0-13-970244-X (paperback). xvi + 170 pp. LCCN QA76.6 .B455 1982. US$22.95 (hardcover), US$14.95 (paperback).

Bentley:1986:PP

[Ben86a] Jon Louis Bentley. Programming Pearls. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-10331-1 (paperback). viii + 195 pp. LCCN QA76.6.B453 1986. Reprinted with corrections.

Bentley:pic

[Ben86b] Jon Louis Bentley. Programming pearls: Little languages. Communications of the Association for Computing Machinery, 29(8): 711–721, August 1986. Description of the pic language.

Bentley:1988:MPP

[Ben88] Jon Louis Bentley. More Programming Pearls: Confessions of a Coder. Addison-Wes-ley, Reading, MA, USA, 1988. ISBN 0-201-11889-0. viii + 207 pp. LCCN QA76.6 .B452 1988. US$18.75.

Bentley:mini-macro-processor

[Ben90] Jon Bentley. m1: a mini macro processor. Computer Language Magazine, 7(6):47–61, June 1990.

Bentley:2000:PP

[Ben00] Jon Louis Bentley. Programming Pearls. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-65788-0 (paperback), 0-13-449802-X (ebook), 0-13-449805-4. xi + 239 pp. LCCN QA76.6 .B454 2000. US$24.95.

Bennett:2021:LKW

[Ben21] Michael Bennett. Lunch with the Kaiser: Why Europe Went to War in 1914. ????, Las Vegas, NV, USA, 2021. ISBN 1-67100-082-X. 197 pp. LCCN ????

Bergmann:1968:RG

[Ber68] Peter Gabriel Bergmann. The riddle of gravitation. Scribner, New York, NY, USA, 1968. xvi + 270 pp. LCCN QC6 .B454.

Bernstein:1981:PEH

[Ber81] Jeremy Bernstein. Prophet of energy, Hans Bethe. Dutton, New York, NY, USA, 1981. ISBN 0-525-47677-6 (paperback). xii + 212 pp. LCCN QC16.B4 B47. US$7.25.

Bernstein:1996:HUC

[Ber96] Jeremy Bernstein. Hitler’s uranium club: the secret recordings at Farm Hall. American Institute of Physics, Woodbury, NY, USA, 1996. ISBN 1-56396-258-6. xxx + 427 + 4 pp. LCCN QC773.3.G3 B47 1995. US$34.95. Introduction by David Cassidy.

Berlin:2005:MBM

[Ber05] Leslie Berlin. The Man Behind the Microchip: Robert Noyce and the Invention of Silicon Valley. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-516343-5. xi + 402 pp. LCCN TK7807.N69 B47 2005.

Bergman:2006:BRE

[Ber06] Aaron Bergman. Book review: Not Even Wrong: The Failure of String Theory and the Continuing Challenge to Unify the Laws of Physics, Peter Woit, Jonathan Cape, London 2006. World Wide Web document, August 18, 2006. URL

http://zippy.ph.utexas.edu/
~abergman/Review.pdf.

Bernstein:2007:PHW

[Ber07] Jeremy Bernstein. Plutonium: a history of the world’s most dangerous element. Joseph Henry Press, Washington, DC, USA, 2007. ISBN 0-309-10296-0 (hardcover), 1-280-84457-4, 0-309-10773-3 (e-book). x + 194 + 8 pp. LCCN QD181.P9

B47 2007. URL http://www.
loc.gov/catdir/toc/ecip075/
2006038466.html.

Berman:2011:SHO

[Ber11] Bob Berman. The Sun’s heartbeat: and other stories from the life of the star that powers our planet. Little, Brown and Company, New York, NY, USA, 2011. ISBN 0-316-09101-4. viii + 290 pp. LCCN QB521.4 .B47 2011.

Best:2004:MDL

[Bes04] Joel Best. More damned lies and statistics: how numbers confuse public issues. University of California Press, Berkeley, CA, USA, 2004. ISBN 0-520-23830-3 (cloth). xvii + 200 pp. LCCN HM535 .B474 2004. URL ftp://

uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http:/
/www.loc.gov/catdir/bios/
ucal052/2003028076.html;
http://www.loc.gov/catdir/
description/ucal051/2003028076.
html;   http://www.loc.gov/
catdir/toc/ecip0412/2003028076.
html.

Bethe:1991:RAP

[Bet91]  Hans Albrecht Bethe.  The

road from Los Alamos: Profound perspective and personal viewpoints on atomic weapons, nuclear power, and science. Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1991. ISBN 0-88318-707-8, 0-671-74012-1. xvii + 286 pp. LCCN U264 .B455 1991. US$24.95.

Beyer:1978:CHM

[Bey78] William H. Beyer, editor. CRC Handbook of Mathematical Sciences. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, fifth edition, 1978. ISBN 0-8493-0655-8. xii + 982 pp. LCCN QA47.H324 1978.

Byrd:1971:HEI

[BF71] Paul F. Byrd and Morris D.

Friedman. Handbook of Elliptic Integrals for Engineers and Scientists, volume 67 of Die Grundlehren der mathematischen Wissenschaften in Einzeldarstellungen. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1971. ISBN 0-387-05318-2 (New York). xvi + 358 pp. LCCN QA343 .B95 1971.

Bauldry:1991:CLM

[BF91] William C. Bauldry and Joseph R. Fiedler. Calculus Laboratories with Maple: a Tool, Not an Oracle. Brooks/Cole, Pacific Grove, CA, USA, 1991. ISBN 0-534-13788-1. xii + 144 pp. LCCN MLCM 90/01887 (Q).

Bartosh:2005:EMX

[BF05] Michael Bartosh and Ryan Faas. Essential Mac OS X Panther Server Administration. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00635-7 (paperback). xxii + 822 pp. LCCN QA76.76.O63 B38 2005.

Bergin:1996:HPL

[BG96] Thomas J. Bergin, Jr. and Richard G. Gibson, Jr., editors. History of Programming Languages II. ACM Press and Addison-Wes-ley, New York, NY 10036, USA and Reading, MA, USA, 1996. ISBN 0-201-89502-1. LCCN QA76.7 .H558 1996. Drawn from the Second ACM SIGPLAN History of Programming Languages Conference.

Brainerd:1990:PGF

[BGA90] Walter S. Brainerd, Charles H.

Goldberg, and Jeanne C. Adams. Programmer’s Guide to Fortran 90. McGraw-Hill, New York, NY, USA, 1990. ISBN 0-07-000248-7. vii + 410 pp. LCCN QA76.73.F25 B735 1990. US$37.95. See [Ame92].

Burks:1946:PDL

[BGvN46] Arthur W. Burks, Herman H.

Goldstine, and John von Neumann. Preliminary discussion of the logical design of an electronic computing instrument. Technical report, Institute for Advanced Study, Princeton, NJ, USA, June 28, 1946. 42 pp. Report to the U.S. Army Ordnance Department under contract W-36-034-OKD-7481. Reprinted in [Tau63b, Paper 2], [BN71], [Swa76, pp. 221–259], [Ran82, §8.3], and [AB87, pp. 97–146]. Second edition dated 2 September 1947.

Barroso:2009:DCI

[BH09] Luiz André Barroso and Urs Hölzle. The datacenter as a computer: an introduction to the design of warehouse-scale machines, volume 6 of Synthesis lectures on computer architecture. Morgan and Claypool, San Rafael, CA, USA, 2009. ISBN 1-59829-556-X (paperback), 1-59829-557-8 (ebook). ISSN 1932-3243 (??invalid checksum??). xi + 107 pp. LCCN TK5105.5 .B273 2009.

Berger:2015:IBL

[BH15] Arno Berger and Theodore P. Hill. An introduction to Benford’s Law. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-16306-5, 1-4008-6658-8 (e-book). viii + 248 pp. LCCN QA273.6 .B474 2015. URL http://site.

                                        ebrary.com/id/11040167;
                                        http://www.degruyter.com/
                                        viewbooktoc/product/465875.

Berger:2017:AHA

[BH17] Lee R. Berger and John (John David) Hawks. Almost Human: the Astonishing Tale of Homo Naledi and the Discovery That Changed Our Human Story. National Geographic, Washington, DC, USA, 2017. ISBN 1-4262-1811-7 (hardcover). 239 pp. LCCN GN284.5 .B47 2017.

Bhandarkar:1996:AIA

[Bha96] Dileep P. Bhandarkar. Alpha Implementations and Architecture: Complete Reference and Guide. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1996. ISBN 1-55558-130-7 (paperback). xviii + 328 pp. LCCN QA76.8.A176B47 1996. US$41.95.

Bhattacharya:2022:MFV

[Bha22] Ananyo Bhattacharya. Man from the Future: the Visionary Life of John von Neumann. W. W. Norton & Co., New York, NY, USA, 2022. ISBN 1-324-00399-5. xiv + 353 pp. LCCN QA29.V66 B43

Blum:2020:FDS

[BHK20] Avrim Blum, John Hopcroft, and Ravi Kannan. Foundations of Data Science. Cambridge University Press, Cambridge, UK, 2020. ISBN 1-108-48506-5 (hardcover), 1-108-75552-6 (e-book). viii + 424 pp. LCCN QA76 .B5675 2020.

Bickley:1948:DAO

[Bic48]  W. G. Bickley.  Difference

and associated operators, with some applications. Journal of mathematics and physics / Massachusetts Institute of Technology, 27(3):183–192, October 1948.

Bickerton:2009:ATH

[Bic09] Derek Bickerton. Adam’s tongue:

how humans made language, how language made humans. Hill and Wang, New York, NY, USA, 2009. ISBN 0-8090-2281-8. 286 pp. LCCN P106 .B4667 2009.

Biddle:2009:DSM

[Bid09] Wayne Biddle. Dark side of the moon: Wernher von Braun, the Third Reich, and the space race. W. W. Norton & Co., New York, NY, USA, 2009. ISBN 0-393-05910-3 (hardcover). xiv + 220 + 8 pp. LCCN TL781.85.V6 B53 2009.

Billawala:metamarks

[Bil89] Neenie Billawala. Metamarks:

Preliminary studies for a Pandora’s Box of shapes. Technical Report STAN-CS-89-1256, Stanford University, Department of Computer Science, Stanford, CA, USA, May 1989.

BIOSIS:1994:SSB

[BIO94] BIOSIS, 210 Arch Street, Philadelphia, PA 19103-1399, USA. Serial Sources for the BIO-SIS Previews Database, December 1994. vii + 450 pp. US$65.00. URL http://

www.biosis.org/htmls/press/

960126.html. Contains journal full names and abbreviations used in the BIOSIS and Biological Abstracts databases, plus ISSN and CODEN values, for 6,395 active and 10,935 inactive serials in 41 languages from 94 countries. Also available on CD ROM.

Bishop:2003:CSA

[Bis03] Matt Bishop. Computer Security: Art and Science. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-44099-7 (hardcover). xli + 1084 pp. LCCN QA76.9.A25 B56 2002. US$79.99, CAN$120.99.

Boehm:1966:FDT

[BJ66] C. Boehm and G. Jacopini. Flow diagrams, Turing machines, and languages with only two formation rules. Communications of the Association for Computing Machinery, 9(5):366–371, May 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bentley:grap

[BK86] Jon Louis Bentley and Brian W. Kernighan. GRAP: A language for typesetting graphs. Communications of the Association for Computing Machinery, 29(8): 782–792, August 1986.

Bell:1963:MAI

[BL63] C. F. (Colin Frank) Bell and K. A. K. (Kenneth Alan Keeler) Lott. Modern Approach to Inorganic Chemistry: a textbook for higher national certificate and general degree students. Butterworths, London, England, 1963. xi + 295 pp. LCCN QD33 .B38 1963.

Bak:1966:VTG

[BL66] Thor A. Bak and Jonas Lichtenberg. Vectors, Tensors, and Groups, volume 1 of Mathematics for Scientists. W. A. Benjamin, Inc., New York, NY, USA, 1966. LCCN QA37 .B35 1967.

Berners-Lee:2019:TNP

[BL19] Mike Berners-Lee. There is no Plan(et) B: a handbook for the make or break years. Cambridge University Press, Cambridge, UK, 2019. ISBN 1-108-42424-4 (hardcover), 1-108-43958-6 (paperback). xii + 288 pp. LCCN TJ163.2 .B4745 2019.

Blatner:1997:JP

[Bla97] David Blatner. The Joy of π.

Walker and Co., New York, NY, USA, 1997. ISBN 0-8027-1332-7 (hardcover), 0-8027-7562-4 (paperback). xiii + 129 pp. LCCN QA484 .B55 1997. URL http:/

/www.walkerbooks.com/books/
catalog.php?key=4.

Black:2002:IHS

[Bla02] Edwin Black. IBM and the Holocaust: the strategic alliance between Nazi Germany and America’s most powerful corporation. Three Rivers Press, New York, NY, USA, 2002. ISBN 0-609-80899-0 (paperback). 551 pp. LCCN HD9696.2.U64 I253 2002.

Blandy:2015:RPL

[Bla15] Jim Blandy. The Rust Programming Language: Fast, Safe, and Beautiful. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2015. ISBN 1-4919-2544-2. ???? pp. LCCN ???? Video file (1h10m).

Brown:2019:SUS

[BLM19] J. Patrick Brown, Beryl Lipton, and Michael Morisy. Scientists under Surveillance: the FBI Files. MIT Press, Cambridge, MA, USA, 2019. ISBN 0-262-53688-9 (paperback), 0-262-35302-4 (e-book). xviii + 413 pp. LCCN Q125 .S4366

2019. URL http://mitpress.
mit.edu/9780262536882. Fore-

word by Steven Aftergood and introduction by Walter V. Robinson.

Bloch:2001:EJP

[Blo01] Joshua Bloch. Effective Java: Programming Language Guide. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-31005-8, 3-642-56735-5 (e-book). xvi + 252 pp. LCCN QA76.73.J38 B57 2001.

Bloomfield:2008:HEW

[Blo08] Louis Bloomfield. How everything works: making physics out of the ordinary. Wiley, New York, NY, USA, 2008. ISBN 0-471-74817-X (hardcover). xv + 720 pp. LCCN QC24.5 .B56 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0649/2006296744-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0708/
2006296744-b.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0728/2006296744-t.html.

Blunden:2002:VMD

[Blu02] Bill Blunden. Virtual Machine Design and Implementation in C/C++. Wordware Publishing, Plano, TX, USA, 2002. ISBN 1-55622-903-8 (paperback), 0-585-40313-9 (e-book). xvii + 668 pp. LCCN QA76.9.V5 B59 2002.

Bornemann:2004:SDC

[BLWW04] Folkma Bornemann, Dirk Laurie, Stan Wagon, and Jörg Waldvogel. The SIAM 100-digit challenge: a study in high-accuracy numerical computing. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2004. ISBN 0-89871-561-X. xii + 306 pp. LCCN QA297 .S4782 2004. URL

http://www.ams.org/msnmain?
fmt=doc&fn=105&id=2076374&
l=20&pg3=IID&r=1&s3=254928&
v3=Bornemann%2C%20Folkmar.

Boyer:string-search

[BM77] R. S. Boyer and J. S. Moore. A fast string searching algorithm. Communications of the Association for Computing Machinery, 20(10):762–772, October 1977. See also [KMP77, Sun90, BYG92].

Bell:1978:CED

[BMM78] C. Gordon Bell, J. Craig Mudge, and John E. McNamara. Computer Engineering: a DEC View of Hardware System Design. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1978. ISBN 0-932376-00-2. xxii + 585 pp. LCCN TK7885B36 1978.

Belcher:1960:QIA

[BN60] Ronald Belcher and A. J. Nutten. Quantitative Inorganic Analysis. Butterworths Scientific Publications, London, England, second edition, 1960. x + 390 pp. LCCN QD45 .B45 1960.

Bell:1971:CSR

[BN71] C. Gordon Bell and Allen Newell, editors. Computer Structures: Readings and Examples. Mc-Graw-Hill, New York, NY, USA, 1971. ISBN 0-07-004357-4. xix + 668 pp. LCCN TK7888.3 .B4.

Bliss:2006:AHM

[BN06] Anna Campbell Bliss and Skylar Nielsen. Art for a House of Mathematics. American Mathematical Society, Providence, RI, USA, 2006. ISBN 0-9754915-1-2. 53 pp. LCCN N72.M3 B55 2006.

Ben-Naim:2008:EDS

[BN08] Arieh Ben-Naim. Entropy Demystified: The Second Law Reduced to Plain Common Sense. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, revised edition, 2008. ISBN 981-283-225-4. xxxi + 225 pp. LCCN QC318.E57 B46 2007.

Blumberg:1976:ECL

[BO76] Stanley A. Blumberg and Gwinn Owens. Energy and conflict: the life and times of Edward Teller. Putnam, New York, NY, USA, 1976. ISBN 0-399-11551-X. xvii + 492 + 4 pp. LCCN QC16.T37 B58 1976. US$12.95.

URL http://alsos.wlu.edu/
information.aspx?id=367.

Blandy:2017:PR

[BO17] Jim Blandy and Jason Orendorff.

Programming Rust. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2017. ISBN 1-4919-2728-3 (paperback), 1-4919-2727-5, 1-4919-2723-2 (e-book), 1-4919-2725-9 (e-book). xx + 598 pp. LCCN QA76.73.R88 B53 2017. URL http://

proquest.safaribooksonline.
com/9781491927274.

Boas:2006:MMP

[Boa06]  Mary L. Boas.  Mathemat-

ical Methods in the Physical Sciences. Wiley, New York, NY, USA, third edition, 2006. ISBN 0-471-19826-9, 0-471-36580-7. xviii + 839 pp.

LCCN QA37.3 .B63 2006. URL

http://www.loc.gov/catdir/
enhancements/fy0626/2005279918-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0626/
2005279918-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0626/2005279918-t.html.

Bodanis:2000:BWM

[Bod00] David Bodanis. E = mc^{2}: a biography of the world’s most famous equation. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 2000. ISBN 0-8027-1352-1 (hardcover). ix + 337 pp. LCCN QC73.8.C6 B63 2000.

Bohren:1987:CGB

[Boh87] Craig F. Bohren. Clouds in a Glass of Beer — Simple Experiments in Atmospheric Physics. Wiley, New York, NY, USA, 1987. ISBN 0-471-62482-9. xv + 195 pp. LCCN QC861.2.B64 1987.

Boisvert:1997:QNS

[Boi97] Ronald F. Boisvert, editor. Quality of numerical software: assessment and enhancement / proceedings of the IFIP TC2/WG2.5 Working Conference on the Quality of Numerical Software, Assessment and Enhancement, Oxford, United Kingdom, 8–12 July 1996. Chapman and Hall, Ltd., London, UK, 1997. ISBN 0-412-80530-

8. LCCN QA297.I35 1996. US$146.50.

Boole:1854:ILT

[Boo54] George Boole. An Investigation of the Laws of Thought. Dover, New York, NY, USA, 1854. 424 pp. LCCN BC135 .B7 1951.

Boorstin:1958:ACE

[Boo58] Daniel J. Boorstin. The Americans: The Colonial Experience. Vintage Books, New York, NY, USA, 1958. ISBN 0-394-70513-0. 434 pp. LCCN E162.B68 1964. US$12.00.

Boorstin:1983:D

[Boo83] Daniel J. Boorstin. The Discoverers. Vintage Books, New York, NY, USA, 1983. ISBN 0-394-72625-1. xvi + 745 pp. LCCN CB69 .B66 1985.

Boorstin:1992:C

[Boo92] Daniel J. Boorstin. The Creators.

Vintage Books, New York, NY, USA, 1992. ISBN 0-394-54395-5 (hardcover), 0-679-74375-8 (paperback). xiv + 811 pp. LCCN CB69.B65 1993.

Borceux:1990:LPT

[Bor90] Francis Borceux. LaTEX: la perfection dans le traitement du texte. Éditions Ciaoco, Artel, Bruxelles, Belgium, 1990. ISBN 2-87085-194-4. LCCN ????

Borde:1992:TE

[Bor92] Arvind Borde. TEX by Example.

Academic Press, New York, NY, USA, 1992. ISBN 0-12-117650-9. xiv + 169 pp. LCCN Z253.4.T47 B67 1992.

Borde:1993:MTE

[Bor93] Arvind Borde. Mathematical TEX by Example. Academic Press, New York, NY, USA, 1993. ISBN 0-12-117645-2, 0-12-155940-8. xii + 352 pp. LCCN Z253.4.T47 B67 1993.

Borenstein:1994:PIP

[Bor94] Nathaniel S. Borenstein. Programming as if People Mattered: Friendly Programs, Software Engineering, and Other Noble Delusions. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03763-9 (paperback), 0-691-08752-0 (hardcover), 0-691-60788-5, 0-691-63640-0 (hardcover), 1-4008-6253-1 (e-book), 1-68015-900-3 (e-book). xiv + 186 pp. LCCN QA76.758.

Born:1995:FFH

[Bor95] Günter Born. The File Formats Handbook. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1995. ISBN 1-85032-117-5, 1-85032-128-0. xvii + 1274 pp. LCCN QA76.9.F5 B67 1995. US$59.95.

Boslaugh:1999:WCW

[Bos99] David L. Boslaugh. When Computers Went to Sea: The Digitization of the United States Navy. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN 0-7695-0024-2. xxiv + 467 pp. LCCN Q175.N438 1997; VB212 .B67 1999. US$35.00. URL

http://www.loc.gov/catdir/
enhancements/fy0706/99024731-
d.html;   http://www.loc.
gov/catdir/toc/wiley031/99024731.

html. IEEE catalog number BP000024.

Bourne:1990:UVU

[Bou90] Philip E. Bourne. UNIX for VMS Users. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1990. ISBN 1-55558-034-3. xvi + 368 pp. LCCN QA76.76.O63 B67 1989.

Bowles:1977:MPS

[Bow77] Kenneth L. Bowles. Microcomputer Problem Solving Using Pascal. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 3-540-90286-4, 1-4615-9998-9 (e-book), 3-662-38578-3 (e-book). 563 pp. LCCN QA76.73.P2 .B68. URL http:/

/link.springer.com/10.1007/
978-3-662-38578-4.

Bowman:2001:PSS

[Bow01] Judith S. Bowman. Practical SQL: the Sequel. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-61638-6 (paperback). xv + 329 pp. LCCN QA76.73.S67 B695 2001. Includes CD-ROM.

Boyles:2016:EEE

[Boy16] Denis Boyles. Everything Explained that is Explainable: On the Creation of the Encyclopædia Britannica’s Celebrated Eleventh Edition, 1910–1911. Alfred A. Knopf, New York, NY, USA, 2016. ISBN 0-307-26917-5 (hardcover), 0-307-38978-2 (paperback), 1-101-94777-2 (ebook). xviii + 442 pp. LCCN AE5.E3633 B69 2016.

Beebe:plot79-biomed

[BR89] Nelson H. F. Beebe and R. P. C. Rodgers. <PLOT79>: a comprehensive portable Fortran scientific line graphics system, as applied to biomedical research. Computers in Biology and Medicine, 19(6):385–402, 1989.

Blatner:1993:RWS

[BR93a] David Blatner and Steve Roth. Real World Scanning and Halftones. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1993. ISBN 1-56609-093-8. xx + 275 pp. LCCN T384 .B52 1993. US$24.95.

Brown:1993:MGS

[BR93b] Laurie M. Brown and John S. Rigden, editors. Most of the Good Stuff: Memories of Richard Feynman. American Institute of Physics, Woodbury, NY, USA, 1993. ISBN 0-88318-870-8. 181 + 16 pp. LCCN QC16.F49 A3 1993. Commentary by Joan Feynman, John Wheeler, Hans Bethe, Julian Schwinger, Murray Gell-Mann, Daniel Hillis, David Goodstein, Freeman Dyson, and Laurie Brown.

Bradley:1984:ALP

[Bra84] David J. Bradley. Assembly Language Programming for the IBM Personal Computer. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-049189-6. xii + 340 pp. LCCN QA76.8.I2594 B7 1984. The author is one of the designers of the IBM PC. The book covers the 8088 and 8087 instruction sets, DOS and MASM, the IBM PC hardware, and the ROM BIOS. Somewhat more technical than [Laf84]. See also [Sca84].

Braswell:1989:IP

[Bra89] Frank Merritt Braswell. Inside PostScript. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1989. ISBN 0-938151-10-X. LCCN QA76.73.P67 B73 1989. US$37.50.

Breshears:2009:AC

[Bre09] Clay Breshears. The Art of Concurrency. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2009. ISBN 0-596-80242-0. xiii + 285 pp. LCCN QA76.642 .B74 2009; Internet.

Brewster:2010:MLW

[Bre10] David Brewster. Memoirs of the life, writings, and discoveries of Sir Isaac Newton. Cambridge library collection. Cambridge University Press, Cambridge, UK, 2010. ISBN 1-108-02556-0 (vol. 1), 1-108-02557-9 (vol. 2), 1-108-02558-7 (set). xxi + 430 (vol. 1), xi + 434 (vol. 2) pp. LCCN QC16.N7 B8 1855.

BrinchHansen:1985:BHP

[Bri85] Per Brinch Hansen. Brinch Hansen on Pascal Compilers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-083098-4. x + 310 pp. LCCN QA76.73.P2 B75 1985. US$25.00.

Brown:1977:SPA

[Bro77] P. J. Brown, editor. Software Portability: An Advanced Course. Cambridge University Press, Cambridge, UK, 1977. ISBN 0-521-21485-8. xiv + 328 pp. LCCN QA76.6 .S6351. US$14.00.

Brooks:1982:MMM

[Bro82] Frederick P. Brooks, Jr. The Mythical Man-Month: Essays on Software Engineering. Addison-Wesley, Reading, MA, USA, 1982. ISBN 0-201-00650-2. xi + 195 pp. LCCN QA 76.6 B75 1982.

Brooks:1985:PCL

[Bro85a] Rodney A. Brooks. Programming in Common Lisp. Wiley, New York, NY, USA, 1985. ISBN 0-471-81888-7. xv + 303 pp. LCCN QA76.73.L23 B76 1985. US$18.95.

Brown:1985:UP

[Bro85b] Maxine Brown. Understanding PHIGS. Megatek Corporation, 9645 Scranton Rd, San Diego, CA 92121, USA, 1985. LCCN T385 .B761 1985.

Brown:1994:PTT

[Bro94] Vicki Brown, editor. Prime Time TEXcetera. Prime Time Freeware, 370 Altair Way, Suite 150, Sunnyvale, CA, USA, 1994. ISBN 1-881957-10-1. 96 pp. US$60.00. Includes CD-ROM.

Brooks:1995:MMM

[Bro95] Frederick P. Brooks, Jr. The Mythical Man-Month — Essays on Software Engineering. Addison-Wesley, Reading, MA, USA, anniversary edition, 1995. ISBN 0-201-83595-9. xiii + 322 pp. LCCN QA76.758 .B75 1995. US$26.76.

Brockman:2006:MEE

[Bro06] John Brockman, editor. My Einstein: essays by twenty-four of the world’s leading thinkers on the man, his work, and his legacy. Pantheon Books, New York, NY, USA, 2006. ISBN 0-375-42345-1. xvi + 261 pp. LCCN QC16.E5 M9 2006. URL

http://www.loc.gov/catdir/
enhancements/fy0643/2005048286-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0643/
2005048286-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0643/2005048286-s.html;
http://www.loc.gov/catdir/
enhancements/fy0661/2005048286-
t.html.

Brown:2010:HKP

[Bro10] Mike Brown. How I killed Pluto and why it had it coming. Spiegel and Grau, New York, NY, USA, 2010. ISBN 0-385-53108-7. xiii + 267 pp. LCCN QB701 .B77 2010.

Browne:2015:SEC

[Bro15] John Browne. Seven Elements That Changed the World: An Adventure of Ingenuity and Discovery. Pegasus Books, New York, NY, USA, 2015. ISBN 1-60598-691-7. xix + 279 pp. LCCN ????

Borwein:2008:CMD

[BRR08] Jonathan M. Borwein, E. M. (Eugenio M.) Rocha, and José-Francisco Rodrigues, editors. Communicating Mathematics in the Digital Era. A. K. Peters, Ltd., Wellesley, MA, USA, 2008. ISBN 1-56881-410-0. xii + 325 pp. LCCN QA76.95 .C59 2008. URL http://www.

                                        loc.gov/catdir/toc/fy0903/
                                        2008022183.html.

Bryan:1988:SAG

[Bry88]  Martin Bryan.  SGML: An

Author’s Guide to the Standard Generalized Markup Language. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-17535-5. xvii + 364 pp. LCCN QA76.73.S44 B79 1988. UK£16.95.

Brand:1960:MSP

[BS60] J. C. D. (John Charles Drury) Brand and J. C. (James Clare) Speakman. Molecular Structure: The Physical Approach. Edward Arnold (Publishers) Ltd., London, England, 1960. viii + 312 pp. LCCN QD501.B786 1964.

Box:2003:ENC

[BS03] Don Box and Chris Sells. Essential .NET: Volume 1: The Common Language Runtime. Microsoft .NET development series. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-73411-7 (paperback). xxiii + 405 pp. LCCN QA76.76.M52 B69 2003. Foreword by James S. Miller.

Bird:2005:APT

[BS05] Kai Bird and Martin J. Sherwin. American Prometheus: the triumph and tragedy of J. Robert Oppenheimer. Alfred A. Knopf, New York, NY, USA, 2005. ISBN 0-375-72626-8, 0-375-41202-6. xiii + 721 + 32 pp. LCCN QC16.O62 B57 2005. URL ftp://

uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http:/
/site.ebrary.com/lib/ucsc/
Doc?id=10078784.

Bronshtein:2007:HM

[BSMM07] I. N. (Ilıa Nikolaevich) Bronshteĭn, A. Semendyayev, Gerhard Musiol, and Heiner Mühlig, editors. Handbook of mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fifth edition, 2007. ISBN 3-540-72121-5. xliii + 1159 pp. LCCN QA40 HAN 2007; QA40 .S6813 2007.

Bell:1982:CSP

[BSN82] C. Gordon Bell, Daniel P.

Siewiorek, and Allen Newell. Computer Structures: Principles and Examples. McGraw-Hill computer science series. Mc-Graw-Hill, New York, NY, USA, second. edition, 1982. ISBN 0-07-057302-6. xvi + 926 pp. LCCN QA76.9.A73 C65; TK7888.3 .B37 1982.

Barstow:1984:IPE

[BSS84] David R. Barstow, Howard E. Shrobe, and Erik Sandewall. Interactive Programming Environments. McGraw-Hill, New York, NY, USA, 1984. ISBN 0-07-003885-6. xii + 609 pp. LCCN QA76.6 .I5251 1984. US$34.95.

Boyce:1996:IW

[BST^{+}96] Jim Boyce, Paul J. Sanna, Rob Tidrow, William Steen, Jonathan J. Chau, Scott Fuller, Kevin Pagan, Russell Jacobs, and R. James Ruehlin. Inside Windows 95: Deluxe Edition. New Riders Publishing, Carmel, IN, USA, 1996. ISBN 1-56205-695-6. xxx + 1228 pp. LCCN QA76.76.O63 B685 1996. US$59.99, CAN$84.95. Includes CD-ROM.

Blair:1998:SIU

[BT98] John D. Blair and The Samba Team. Samba: Integrating UNIX and Windows. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, June 1998. ISBN 1-57831-006-7. xviii + 298 pp. LCCN QA76.76.O63 B55 1998. US$29.95. URL http:/

/www.clbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
1578310067;   http://www.
ssc.com/ssc/samba/. From the

publisher: “Includes CD-ROM containing version 1.9.18 of the Samba server, a library of useful tools and scripts, the Samba mailing list archives, and all examples discussed in the book.”.

Budd:1998:DSC

[Bud98] Timothy Budd. Data Structures in C++ Using the Standard Template Library. Addison-Wes-ley, Reading, MA, USA, 1998. ISBN 0-201-30879-7. xxiv + 544 pp. LCCN QA76.73.C15B8 1998.

Buerger:1990:LES

[Bue90] David J. Buerger. LaTEX for Engineers and Scientists. McGraw-Hill, New York, NY, USA, 1990. ISBN 0-07-008845-4. xvi + 198 pp. LCCN Z253.4.L38 B84 1990.

Burrough:1998:DNC

[Bur98] Bryan Burrough. Dragonfly:

NASA and the crisis aboard the MIR. HarperCollins College Publishers, New York, NY, USA, 1998. ISBN 0-88730-783-3. 528 pp. LCCN TL867 .B87 1998.

Burton:2007:HMI

[Bur07] David M. Burton. The History of Mathematics: an Introduction. McGraw-Hill, New York, NY, USA, sixth edition, 2007. ISBN 0-07-305189-6. xii + 788 pp. LCCN QA21 .B96 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0625/2005051123-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0625/
2005051123-t.html.

Bush:1940:AM

[Bus40] V. Bush. Arithmetical machine. Vannevar Bush Papers, Container 18, Folder: Caldwell, Samuel, 1939–1940, 1940. Reprinted in [Ran82, §7.3]. Copyright interest in the unpublished writings of Vannevar Bush has been dedicated to the public.

Bush:1945:SEF

[Bus45] Vannevar Bush. Science, the endless frontier: a report to the President: Bush report on scientific research in the United States. United States Government Printing Office, Washington, DC, USA, 1945. xxvi + 220 pp. LCCN Q127.U6 A53 1945. URL https://

                                        ia600207.us.archive.org/12/
                                        items/scienceendlessfr00unit/
                                        scienceendlessfr00unit.pdf.

Busch:1999:GGG

[Bus99] David D. Busch. Great graphics with GIMP. Prima Publishing, Rocklin, CA, USA, 1999. ISBN 0-7615-2407-X. xxx + 370 pp. LCCN T385 .B8664 2000. US$40.00. Edited by Kevin Harreld and Kim Spilker.

Butler:1964:SPC

[But64] James Newton Butler. Solubility and pH Calculations: The Mathematics of the Simplest Ionic Equilibria. Addison-Wes-ley, Reading, MA, USA, 1964. viii + 104 pp. LCCN QD42 .B87.

Borwein:2014:NFI

[BvdPSZ14] Jonathan M. Borwein, Alfred Jacobus van der Poorten, Jeffrey Outlaw Shallit, and Wadim Zudilin. Neverending Fractions: an Introduction to Continued Fractions, volume 23 of Australian Mathematical Society lecture series. Cambridge University Press, Cambridge, UK, 2014. ISBN 0-521-18649-8. x + 212 pp. LCCN QA295 .B667 2014. URL http://

docserver.carma.newcastle.
edu.au/1722/;   http://ebooks.
cambridge.org/ebook.jsf?bid=
CBO9780511902659.

Baeza-Yates:j-CACM-35-10-74

[BYG92] Ricardo Baeza-Yates and Gaston H. Gonnet. A new approach to text searching. Communications of the Association for Computing Machinery, 35(10): 74–82, October 1992. This paper describes a new linear-time string search algorithm that can handle limited regular-expression pattern matching without backtracking. See also [KMP77], [BM77], [KR81], [Sun90], and [WM92].

Brent:2011:MCA

[BZ11] Richard P. Brent and Paul Zimmermann. Modern Computer Arithmetic, volume 18 of Cambridge monographs on applied and computational mathematics. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-19469-5 (hardcover). xvi + 221 pp. LCCN QA76.9.C62 BRE 2011. URL http://

                                        www.loria.fr/~zimmerma/mca/
                                        pub226.html.

Char:1989:FLM

[C^{+}89] Bruce W. Char et al. First leaves for the Macintosh: a tutorial introduction to Maple. Brooks/ Cole, Pacific Grove, CA, USA, 1989. ISBN 0-534-10222-0. xv + 140 pp. LCCN QA155.7.E4 F57 1989.

Cahill:1995:HIS

[Cah95] Thomas Cahill. How the Irish Saved Civilization: the Untold Story of Ireland’s Heroic Role from the Fall of Rome to the Rise of Medieval Europe. Nan A. Talese, Doubleday, New York, NY, USA, 1995. ISBN 0-385-41848-5. x + 246 pp. LCCN DA930.5 .C34 1995. US$20.00. URL http:

                                        //www.loc.gov/catdir/bios/
                                        random057/94028130.html;
                                        http://www.loc.gov/catdir/
                                        description/random043/94028130.
                                        html;   http://www.loc.gov/
                                        catdir/samples/random041/
                                        94028130.html.

Calaprice:2004:EA

[Cal04] Alice Calaprice. The Einstein almanac. The Johns Hopkins University Press, Baltimore, MD, USA, 2004. ISBN 0-8018-8021-

1. xvii + 176 pp. LCCN QC16.E5 C35 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/pub/
                                        etext/gutenberg/;   http:/
                                        /www.loc.gov/catdir/bios/
                                        jhu051/2004009048.html;

http://www.loc.gov/catdir/
description/jhu051/2004009048.
html;   http://www.loc.gov/
catdir/toc/ecip0417/2004009048.
html.

Campbell:1987:CPG

[Cam87] Joe Campbell. C Programmer’s Guide to Serial Communications. Howard W. Sams, Indianapolis, IN 46268, USA, 1987. ISBN 0-672-22584-0. xii + 655 pp. LCCN QA76.73.C15 C36 1987. US$22.95.

Cannell:2001:GGM

[Can01] D. M. (Doris Mary) Cannell.

George Green: mathematician and physicist, 1793–1841: the background to his life and work. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 2001. ISBN 0-89871-463-X. xxxiv + 316 + 10 pp. LCCN QC16.G64 C36 2001.

URL http://epubs.siam.org/
ebooks/siam/other_titles_
in_applied_mathematics/ot73;
http://www.loc.gov/catdir/
enhancements/fy0726/00041938-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0726/
00041938-t.html.

Carroll:1994:BBU

[Car94] Paul Carroll. Big Blues: the unmaking of IBM. Crown Trade Paperbacks, New York, NY, USA, 1994. ISBN 0-517-88221-3. v + 377 pp. LCCN HD9696.C64 I48317 1994. URL http:

//www.loc.gov/catdir/bios/
random057/95145899.html.

Carter:2010:GNW

[Car10]  Miranda Carter.  George,

Nicholas and Wilhelm: three royal cousins and the road to World War I. Alfred A. Knopf, New York, NY, 2010. ISBN 1-4000-4363-8 (hardcover), 1-4000-7912-8 (paperback). xxv + 498 + 32 pp. LCCN D517 .C34 2010.

Castleman:1979:DIP

[Cas79] Kenneth R. Castleman. Digital Image Processing. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1979. ISBN 0-13-212365-7. xvi + 429 pp. LCCN TA1632 .C3 1979.

Casselman:2005:MIM

[Cas05] Bill Casselman. Mathematical Illustrations: a Manual of Geometry and PostScript. Cambridge University Press, Cambridge, UK, 2005. ISBN 0-521-83921-1 (hardcover), 0-521-54788-1 (paperback). xviii + 318 pp. LCCN QA76.73.P67 C37 2004. US$39.99 (paperback), US$90.00 (hardcover). URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                        http://www.loc.gov/catdir/
                                        description/cam041/2004045886.
                                        html;   http://www.loc.gov/
                                        catdir/samples/cam051/2004045886.
                                        html;   http://www.loc.gov/
                                        catdir/toc/cam041/2004045886.
                                        html.

Cassidy:2009:BUH

[Cas09] David C. Cassidy. Beyond uncertainty: Heisenberg, quantum physics, and the bomb. Bellevue Literary Press, New York, NY, USA, 2009. ISBN 1-934137-13-8. 480 pp. LCCN QC16.W518 C37 2008.

Cathcart:2004:FCH

[Cat04] Brian Cathcart. The fly in the cathedral: how a group of Cambridge scientists won the international race to split the atom. Farrar, Strauss, and Giroux, New York, NY, USA, 2004. ISBN 0-374-15716-2 (hardcover). xii + 308 + 4 pp. LCCN Q141 .C2515 2004.

Cheswick:1994:FIS

[CB94] William R. Cheswick and Steven M. Bellovin. Firewalls and Internet Security: Repelling the Wily Hacker. Addison-Wes-ley, Reading, MA, USA, 1994. ISBN 0-201-63357-4. xiv + 306 pp. LCCN TK5105.875.I57C44 1994. US$24.95.

Chabert:1999:HAH

[CBB^{+}99] Jean-Luc Chabert, Évelyne Barbin, Jacques Borowczyk, Michel Guillemot, Anne Michel-Pajus, Ahmed Djebbar, and Jean-Claude Martzloff, editors. A history of algorithms: from the pebble to the microchip. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-63369-3 (softcover). ix + 524 pp. LCCN QA58 .H5813 1998. URL http://

www.springer.com/west/home/
math/cse?SGWID=4-10045-22-
1455224-0.

Cheswick:2003:FIS

[CBR03] William R. Cheswick, Steven M. Bellovin, and Aviel D. Rubin. Firewalls and Internet Security: Repelling the Wily Hacker. Addison-Wesley, Reading, MA, USA, second edition, 2003. ISBN 0-201-63466-X. xiv + 433 pp. LCCN TK5105.875.I57C44 2003. US$49.99, CAN$77.99.

Aldus:tiff

[CC88] Aldus Corporation and Microsoft Corporation. Tag image file format (TIFF) specification revision 5.0. Technical report, Aldus Corporation, 411 First Avenue South, Suite 200, Seattle, WA 98104, Tel: (206) 622-5500, and Microsoft Corporation, 16011 NE 36th Way, Box 97017, Redmond, WA 98073-9717, Tel: (206) 882-8080, August 8 1988.

Cody:1984:PRW

[CCG^{+}84] William J. Cody, Jr., Jerome T. Coonen, David M. Gay, K. Hanson, David G. Hough, William Kahan, Richard Karpinski, John F. Palmer, Frederic N. Ris, and David Stevenson. A proposed radix- and wordlength-independent standard for floating-point arithmetic. IEEE Micro, 4(4):86–100, July/August 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Castaneda:1979:PTP

[CCN^{+}79] Fernando Castaneda, Frederick Chow, Peter Nye, Dan Sleator, and Gio Wiederhold. PCFORT

— a Fortran-to-Pcode translator. Technical Report STAN-CS-79-714, Stanford University, Department of Computer Science, Stanford, CA, USA, January 1979. vi + 85 pp. URL http:

//i.stanford.edu/pub/cstr/
reports/cs/tr/79/714/CS-TR-
79-714.pdf.

Campbell:2006:MSS

[CCN06] S. L. (Stephen La Vern) Campbell, Jean-Philippe Chancelier, and Ramine Nikoukhah. Modeling and Simulation in Scilab/Scicos. Springer Science+Business Media, New York, NY, USA, 2006. ISBN 0-387-27802-8 (hardcover). xi + 313 pp. LCCN Q183.9 .C36

2006. URL http://www.loc.
gov/catdir/enhancements/fy0663/
2005930797-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0814/2005930797-b.html;
http://www.loc.gov/catdir/
toc/fy0705/2005930797.html.

Cesareo:1946:RI

[Ces46] O. Cesareo. The relay interpolator. Bell Laboratories Record, 23(??):457–460, 1946. CODEN BLRCAB. ISSN 0005-8564. Reprinted in [Ran82, §6.2].

Cetron:1991:CRA

[Cet91] Edward J. Cetron. Complexity Reduction for Analysis and Visualization of Textured Electrostatic Fields. PhD thesis, Department of Bioengineering, University of Utah, Salt Lake City, UT, USA, June 1991. xiii + 124 pp.

Conway:1996:BN

[CG96] John H. Conway and Richard K. Guy. The Book of Numbers. Copernicus (a division of Springer-Verlag New York, Inc.), 175 Fifth Avenue, New York, NY 10010, USA, 1996. ISBN 0-387-97993-X. ix + 310 pp. LCCN QA241 .C6897 1996. US$35.00.

Crease:2014:QMH

[CG14] Robert P. Crease and Alfred S. Goldhaber. The quantum moment: how Planck, Bohr, Einstein, and Heisenberg taught us to love uncertainty. W. W. Norton & Co., New York, NY, USA, 2014. ISBN 0-393-06792-0 (hardcover). vii + 332 pp. LCCN QC174.123 .C74 2014.

Char:1991:MLVb

[CGG^{+}91a] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton Leong, Michael B. Monagan, and Stephen M. Watt. Maple Library V Reference Manual. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97592-6, 3-540-97592-6. xxv + 698 pp. LCCN QA155.7.E4 M353 1991.

Char:1991:MVLa

[CGG^{+}91b] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton Leong, Michael B. Monagan, and Stephen M. Watt. Maple V: Language Reference Manual. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97622-1 (New York), 3-540-97622-1 (Berlin). xv + 267 pp. LCCN QA155.7.E4 M36 1991. US$24.95, FF 193,00.

Char:1992:FLT

[CGG^{+}92] Bruce W. Char, Keith O. Geddes, Gaston H. Gonnet, Benton Leong, Michael B. Monagan, and Stephen M. Watt. First Leaves: a Tutorial Introduction to Maple V. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97621-3, 3-540-97621-3. xvii + 253 pp. LCCN QA155.7.E4 F56 1992. Also available in Japanese, ISBN 4-431-70651-8.

Chan:2007:MMC

[CGO07] Raymond H. Chan, Chen Greif, and Dianne P. O’Leary, editors. Milestones in Matrix Computation: the Selected Works of Gene H. Golub with Commentaries. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-920681-3. xi + 565 + 3 pp. LCCN QA188 .G67 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0737/2007276086-
d.html.

Chen:UCB-TR-87-347

[CH87] Pehong Chen and Michael A.

Harrison. Automating index preparation. Technical Report 87/347, Computer Science Division, University of California, Berkeley, CA, USA, March 1987.

This is an expanded version of

[CH88].

Chen:SPE-19-9-897

[CH88] Pehong Chen and Michael A. Harrison. Index preparation and processing. Software—Practice and Experience, 19(9):897–915, September 1988. The LaTEX text of this paper is included in the makeindex software distribution. See [CH87].

Chasen:1978:GPP

[Cha78] Sylvan H. Chasen. Geometric Principles and Procedures for Computer Graphics Applications. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1978. ISBN 0-13-352559-7. xiv + 241 pp. LCCN T385 .C46.

Chanda:2018:NPR

[Cha18]  Abhishek Chanda.  Network

Programming Rust: Build Fast and Resilient Network Servers and Clients by Leveraging Rust’s Memory-safety and Concurrency Features. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78862-171-9 (e-book), 1-78862-489-0 (paperback). iii + 265 pp. LCCN QA76.73.R88. URL

                                        http://proquest.safaribooksonline.
                                        com/?fpi=9781788624893.

CASSI:1994:CAS

[Che94] Chemical Abstracts Service, American Chemical Society, Columbus, OH, USA. Chemical Abstracts Service Source Index, 1994. Contains journal full names and abbreviations used in the Chemical Abstracts database, plus ISSN and CO-DEN values, for 18,656 active serials, 12,230 changed-title serials, 7,486 inactive serials, 16,373 conference proceedings, and 14,239 monographs containing collections of papers.

CAS:19xx:ICD

[Chexx] Chemical Abstracts Service. International CODEN Directory. Chemical Abstracts Service, 2540 Olentangy River Road, Columbus, OH, USA, 19xx. ISSN 0364-3670. URL http:/

/www.cas.org/PRINTED/coden.

html. From the publisher’s Web site: “CODEN are unique, six-character codes that identify serial and nonserial publications produced worldwide. CO-DEN are assigned not only to chemistry-related literature but to publications in all subject areas.”.

Cherkaev:2000:VMS

[Che00] Andrej Cherkaev. Variational Methods for Structural Optimization, volume 140 of Applied mathematical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-98462-3 (hardcover). ISSN 0066-5452 (print), 2196-968X (electronic). xxvi + 545 pp. LCCN QA1 .A647 v.140; TA658.8 .C53

2000. URL https://www.math.
utah.edu/book/vmso. NHFB

was the chief technical editor for this book, and wrote the author/editor indexing software, the TEX DVI-to-PostScript driver and the initial Web site for the book, at https:

                                        //www.math.utah.edu/books/
                                        vmso/index.html.

Chen:2002:GGS

[Che02] Jim X. Chen. Guide to Graphics Software Tools. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95049-4 (hardcover), 0-585-47254-8, 0-387-22430-0 (e-book). xxiv + 503 pp. LCCN T385 .C473 2003. US$59.95. Includes CD-ROM.

Chow:1989:MXR

[Cho89] Paul Chow, editor. The MIPS-X RISC Microprocessor. The Kluwer international series in engineering and computer science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1989. ISBN 0-7923-9045-8. xxiv + 231 pp. LCCN QA76.8.M524 M57 1989.

Christian:1983:UOS

[Chr83] Kaare Christian. The UNIX Operating System. Wiley-Inter-science, New York, NY, USA, 1983. ISBN 0-471-87542-2 (hardcover) and 0-471-89052-9 (paperback). xviii + 318 pp. LCCN QA76.8.U65 C45 1983.

Christian:1986:GM

[Chr86]  Kaare Christian.  A Guide

to Modula-2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-96242-5. xix + 436 pp. LCCN QA76.73.M63 C494 1986. US$28.00.

Christian:1988:UOS

[Chr88] Kaare Christian. The UNIX Operating System. Wiley-Inter-science, New York, NY, USA, second edition, 1988. ISBN 0-471-84782-8 (hardcover), 0-471-84781-X (paperback). xxii + 455 pp. LCCN QA76.8.U65 C45 1988.

Christensen:1997:IDW

[Chr97] Clayton M. Christensen. The innovator’s dilemma: when new technologies cause great firms to fail. The Management of Innovation and Change Series. Harvard Business School, Boston, MA, USA, 1997. ISBN 0-87584-585-1. xxiv + 225 pp. LCCN HD53 .C49 1997.

Cornea:2002:SCI

[CHT02] Marius Cornea, John Harrison, and Ping Tak Peter Tang. Scientific Computing on Itaniumbased Systems. Intel Corporation, Santa Clara, CA, USA, 2002. ISBN 0-9712887-7-1. xvii + 406 pp. LCCN QA76.8.I83 C67 2002. US$69.95. URL http:

//www.intel.com/intelpress/
sum_scientific.htm.

Churchill:1958:OM

[Chu58] Ruel V. Churchill. Operational Mathematics. McGraw-Hill, New York, NY, USA, second edition, 1958. 337 pp. LCCN QA432 .C45

Churchill:1960:CVA

[Chu60] Ruel V. Churchill. Complex Variables and Applications. Mc-Graw-Hill, New York, NY, USA, second edition, 1960. 297 pp. LCCN QA331 .C45 1960.

Cowart:2007:SEU

[CK07] Robert Cowart and Brian Knittel. Special edition: Using Microsoft Windows Vista. Que Corporation, Indianapolis, IN, USA, 2007. ISBN 0-7897-3472-9 (paperback). xlv + 1458 pp. LCCN QA76.76.O63 C683376

2007.  URL http://www.
loc.gov/catdir/toc/ecip074/
2006037794.html.

Campbell-Kelly:2003:HMT

[CKCFR03] Martin Campbell-Kelly, Mary Croarken, Raymond Flood, and Eleanor Robson, editors. The History of Mathematical Tables: From Sumer to Spreadsheets. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2003. ISBN 0-19-850841-7. viii + 361 pp. LCCN QA47 .H57 2003. URL

http://www.loc.gov/catdir/
enhancements/fy0620/2004298837-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0620/
2004298837-t.html.

Chposky:1988:BMP

[CL88] James Chposky and Ted Leonsis. Blue magic: the people, power, and politics behind the IBM personal computer. Facts on File, New York, NY, USA, 1988. ISBN 0-8160-1391-8. xi + 228 pp. LCCN HD9696.C64 I4832 1988.

Chan:1998:JCLb

[CL98] Patrick Chan and Rosanna Lee. The Java Class Libraries:

java.applet, java.awt, java.beans,

volume 2. Addison-Wes-ley, Reading, MA, USA, second edition, 1998. ISBN 0-201-31003-1. xxix + 1682 pp. LCCN QA76.73.J38C47 1998. US$54.95. URL http:

//cseng.aw.com/bookdetail.
qry?ISBN=0-201-31003-1;
http://www2.awl.com/cgi-bin/
htsearch?restrict=&exclude=
&config=htdig&method=boolean&
format=builtin%2Dlong&words=
Chan%20AND%20Lee&page=2.

Clark:TAU90

[Cla90] Malcolm Clark, editor. TEX:

Applications, Uses, Methods. Ellis Horwood, New York, NY, USA, 1990. ISBN 0-13-912296-

6. LCCN Z253.4.T47 C45 1990. US$69.95.

Clark:1992:PTP

[Cla92] Malcolm Clark. A Plain TEX Primer. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1992. ISBN 0-19-853784-0 (hardcover), 0-19-853724-7 (paperback). 481 pp. LCCN Z253.4.T47 C46 1992.

Clawson:1994:MTE

[Cla94] Calvin C. Clawson. The Mathematical Traveler: Exploring the Grand History of Numbers. Plenum Press, New York, NY, USA; London, UK, 1994.

ISBN 0-306-44645-6. x + 307 pp. LCCN QA141 .C52 1994. US$25.95.

Clark:2009:SKU

[Cla09] Stuart (Stuart G.) Clark. The Sun Kings: the unexpected tragedy of Richard Carrington and the tale of how modern astronomy began. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-14126-6 (paperback). xii + 211 pp. LCCN M09.E06300. URL

                                        http://press.princeton.edu/
                                        titles/8370.html.

Clews:1988:LAW

[Cle88] John Clews. Language automation worldwide: the development of character set standards. Number 5962 in British Library R & D reports. SESAME Computer Projects, Harrogate, North Yorkshire, UK, 1988. ISBN 1-870095-01-4. 104 pp. LCCN QA76.5 .C54 1988.

Clegg:2003:BHI

[Cle03] Brian Clegg. A Brief History of Infinity: The Quest to Think the Unthinkable. Robinson, London, UK, 2003. ISBN 1-84119-650-9. 255 pp. LCCN BD411 .C57 2003. UK£8.99.

Clegg:2009:BBB

[Cle09] Brian Clegg. Before the Big Bang: the prehistory of our universe. St. Martin’s Press, New York, NY, USA, 2009. ISBN 0-312-38547-1. 306 pp. LCCN QB981 .C627 2009. URL

http://www.loc.gov/catdir/
enhancements/fy0908/2008046035-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0908/
2008046035-d.html.

Cline:1999:CF

[CLG99] Marshall Cline, Greg Lomow, and Mike Girou. C++ FAQs. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-30983-1. xxxiii + 587 pp. LCCN QA76.73.C153C55 1999. US$39.95.

Clinger:floating-point-input

[Cli90] William D. Clinger. How to read floating point numbers accurately. ACM SIGPLAN Notices, 25(6):92–101, June 1990. See also output algorithm in

[SW90] and [Knu90a].

Chan:1999:JCLb

[CLK99a] Patrick Chan, Rosanna Lee, and Doug Kramer. The Java Class

Libraries: java.io, java.lang,
java.math, java.net, java.text,

java.util, volume 1. Addison-Wesley, Reading, MA, USA, second edition, 1999. ISBN 0-201-31002-3. xxvi + 2050 pp. LCCN QA76.73.J38 C47 1998. US$59.99.

Chan:1999:JCLc

[CLK99b] Patrick Chan, Rosanna Lee, and Douglas Kramer. The Java Class Libraries, Second Edition, Volume 1: Supplement for Java 2 Platform, Standard Edition, v1.2. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-48552-4. xxix + 1157 pp.

LCCN QA76.73.J38 C47 1998. US$34.95.

Close:2015:HLD

[Clo15] Frank E. Close. Half-life: the divided life of Bruno Pontecorvo, physicist or spy. Basic Books, New York, NY, USA, 2015. ISBN 0-465-06998-3 (hardcover), 0-465-04487-5 (e-book), 1-78074-582-6 (e-book). xix + 378 pp. LCCN QC774.P66 C56 2014. US$29.99.

Cormen:1990:IA

[CLR90] Thomas H. Cormen, Charles E. (Eric) Leiserson, and Ronald L. Rivest. Introduction to Algorithms. MIT Press, Cambridge, MA, USA, 1990. ISBN 0-262-03141-8, 0-07-013143-0 (McGraw-Hill). xvii + 1028 pp. LCCN QA76.6 .C662 1990.

Cormen:2001:IA

[CLR01] Thomas H. Cormen, Charles E. (Eric) Leiserson, and Ronald L. Rivest. Introduction to Algorithms. MIT Press, Cambridge, MA, USA, second edition, 2001. ISBN 0-262-53196-8 (paperback), 0-262-03293-7 (hardcover), 0-07-013151-1 (McGraw-Hill), 0-07-297054-5 (McGraw-Hill with CD-ROM). xxi + 1180 pp. LCCN QA76.6 .I5858 2001.

Cormen:2009:IA

[CLRS09] Thomas H. Cormen, Charles Eric Leiserson, Ronald L. Rivest, and Clifford Stein, editors. Introduction to Algorithms. MIT Press, Cambridge, MA, USA, third edition, 2009. ISBN 0-262-03384-4 (hardcover), 0-262-53305-7 (paperback). xix + 1292 pp. LCCN QA76.6 .C662 2009.

Clark:1985:MPP

[CM85] K. L. Clark and F. G. Mc-Cabe. micro-PROLOG: Programming in Logic. Series in Computer Science, Editor: C. A. R. Hoare. Prentice-Hall International, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-581264-X (paperback). xi + 401 pp. LCCN QA76.73.M5 C55 1984. US$17.95.

Chroskova:1993:ACC

[CNVH93] Blank Chrosková, Bohumila Navrátilová, Miloš Vantuch, and Eva Hrabcová. Anglicko– Český a Česko–Anglický Slovník — English–Czech and Czech– English Dictionary. Vydavatelství Montaněx, Ostrava, Czechoslovakia, 1993. ISBN 80-85300-57-5. 176 pp. LCCN PG4640 .A485 1993x.

Cody:fps-analysis

[Cod81] William J. Cody, Jr. Analysis of proposals for the floating-point standard. Computer, 14(3):63– 69, March 1981. See [IEE85b].

Cohen:2010:CSH

[Coh10] Richard Cohen. Chasing the Sun:

a History of the Star That Gives Us Life. Random House, New York, NY, USA, 2010. ISBN 1-4000-6875-4. xxxi + 574 pp. LCCN QB521 .C625 2010.

Cokinos:2009:FSI

[Cok09] Christopher Cokinos. The Fallen Sky: an Intimate History of Shooting Stars. Jeremy P. Tarcher/Penguin, New York, NY, USA, 2009. ISBN 1-58542-720-9. 517 pp. LCCN QB755 .C65 2009. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy1001/2009017493-
                                        b.html;   http://www.loc.
                                        gov/catdir/enhancements/fy1001/
                                        2009017493-d.html.

Cole:1976:MP

[Col76] A. J. Cole. Macro Processors, volume 4 of Cambridge Computer Science Texts, Editors: E. S. Page and C. M. Reeves and D. E. Conway. Cambridge University Press, Cambridge, UK, 1976. ISBN 0-521-29024-4. viii + 230 pp. LCCN QA76.C358. US$7.95.

Collinson:awk

[Col91] Peter Collinson. Awk. SunExpert, 2(1):33–36, January 1991.

Colwell:2005:PCP

[Col05] Robert P. Colwell. The Pentium Chronicles: The People, Passion, and Politics Behind Intel’s Landmark Chips. Wiley, New York, NY, USA, 2005. ISBN 0-471-73617-1. xix + 187 pp. LCCN HD9696.S44 I563 2006. US$24.95.

Companion:1964:CB

[Com64a] Audrey L. Companion. Chemical Bonding. McGraw-Hill, New York, NY, USA, 1964. vi + 155 pp. LCCN QD461 .C63.

Comrie:1964:KFF

[Com64b] L. J. Comrie. Knott’s Four-Figure Mathematical Tables. W. and R. Chambers Ltd., London, UK, 1964.

Comer:1984:OSD

[Com84] Douglas E. Comer. Operating Systems Design. The XINU Approach, volume 1. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-637539-1. LCCN QA76.6 .C6275 1984.

Comer:1987:OSD

[Com87a] Douglas E. Comer. Operating Systems Design. Internetworking with XINU, volume 2. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-637414-X. LCCN QA76.6 .C6275 1984.

Comrie:1987:WML

[Com87b] Bernard Comrie, editor. The World’s Major Languages. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1987. ISBN 0-19-520521-9 (hardcover), 0-19-506511-5 (paperback). xiii + 1025 pp. LCCN P371.W6 1987. US$35.00. Reprinted in paperback in 1990.

Comer:1988:ITI

[Com88] Douglas E. Comer. Internetworking with TCP/IP. Principles, Protocols, and Architecture. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1988. ISBN 0-13-470154-2. LCCN TK5105.5 .C59 1988. US$27.00.

Comer:2000:IBE

[Com00a] Douglas Comer. The Internet book: everything you need to know about computer networking and how the Internet works. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 2000. ISBN 0-13-030852-8. xxvi + 351 pp. LCCN TK5105.875.I57 C65 2000.

Comer:2000:ITI

[Com00b] Douglas E. Comer. Internetworking with TCP/IP: Volume 1. Principles, Protocols, and Architectures. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2000. ISBN 0-13-018380-6 (hardcover). xxix + 750 pp. LCCN TK5105.585 .C66 2000 v.1, 3 (2000-c2001).

Comer:2013:IT

[Com13] Douglas E. Comer. Internetworking with TCP. Pearson, New York, NY, USA, sixth edition, 2013. ISBN 0-13-608530-X (hardcover). xxviii + 689 pp. LCCN TK5105.585 .C66 2014.

Conrac:1985:RGH

[Con85] Conrac Division. Raster Graphics Handbook. Van Nostrand Reinhold, New York, NY, USA, 1985. ISBN 0-442-21608-4. v + 345 pp. LCCN T385 .C64 1985.

Conant:2005:EPR

[Con05] Jennet Conant. 109 East Palace: Robert Oppenheimer and the secret city of Los Alamos. Simon and Schuster, New York, NY, USA, 2005. ISBN 0-7432-5007-9. xviii + 425 + 16 pp. LCCN QC773.A1 C66 2005.

Conant:2009:IRD

[Con09] Jennet Conant. The Irregulars:

Roald Dahl and the British Spy Ring in Wartime Washington. Simon and Schuster, New York, NY, USA, 2009. ISBN 0-7432-9458-0 (hardcover), 0-7432-9459-9 (paperback). xxii + 391 pp. LCCN M10.F03629.

Conant:2020:GSC

[Con20] Jennet Conant. The Great Secret: the Classified World War II Disaster That Launched the War on Cancer. W. W. Norton and Company, Inc., New York, NY, 2020. ISBN 0-393-86843-5 (paperback), 1-324-00250-6 (hardcover), 1-324-00251-4 (e-book). xviii + 380 + 8 pp. LCCN RM260 .C63 2020; RM261 .C66 2020.

Coonen:fps-guide

[Coo80] Jerome T. Coonen. An implementation guide to a proposed standard for floating point arithmetic. Computer, 13(1):68–79, January 1980. See errata in [Coo81b]. See [IEE85b].

Coonen:ufl-denorm

[Coo81a] Jerome T. Coonan. Underflow and the denormalized numbers. Computer, 14(3):75–87, March 1981. See [IEE85b].

Coonen:fps-guide-errata

[Coo81b] Jerome T. Coonen. Errata: An implementation guide to a proposed standard for floating point arithmetic. Computer, 14(3):62, March 1981. See also [Coo80, IEE85b].

Copeland:2012:ATE

[Cop12] B. Jack Copeland, editor. Alan Turing’s Electronic Brain: the Struggle to Build the ACE, the World’s Fastest Computer. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-960915-2 (paperback). xxi + 553 pp. LCCN ???? UK £14.99.

Weitek:fp-chip

[Cor86]  Weitek Corporation.  WTL

1164/WTL 1165 64-bit IEEE floating-point multiplier/divider and ALU. Technical report, Weitek Corporation, 1060 E. Arques Ave., Sunnyvale, CA 94086-BRM-9759, USA, July 1986.

Polytron:polyawk

[Cor87] Polytron Corporation. PolyAWK, 1987. 170 NW 167th Place, Beaverton, OR 97006. See also [AKW88].

Xerox:color

[Cor89] Xerox Corporation. The Xerox Color Encoding Standard. Technical Report XNSS 288811, Xerox Systems Institute, March 1989.

DEC:AAH92

[Cor92] Digital Equipment Corporation. Alpha Architecture Handbook. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1992.

Cornwell:2003:HSS

[Cor03] John Cornwell. Hitler’s scientists: science, war, and the Devil’s pact. Viking, New York, NY, USA, 2003. ISBN 0-670-03075-9. xvi + 535 pp. LCCN Q127.G3 C67 2003.

LSBT:2005:BAL

[Cor05] Core Members of the Linux Standard Base Team. Building applications with the Linux Standard Base. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-145695-4 (hardcover). xxvi + 246 pp. LCCN QA76.76.O63 B8375 2004. URL http://www.

freestandards.org/;   http:
//www.lanana.org/; http://
www.linuxbase.org/;   http:
//www.linuxbase.org/test/
registered.html;   http://
www.phptr.com/title/0131456954.

Foreword by Theodore Ts’o. Includes CD-ROM.

Cortada:2019:IHM

[Cor19]  James W. Cortada.  IBM:

How a Multinational Corporation Struggled for Success and Shaped the Modern World. History of computing. MIT Press, Cambridge, MA, USA, 2019. ISBN 0-262-03944-3 (hardcover), 0-262-35148-X (e-book). xxii + 723 pp. LCCN HD9696.2.U6.

Cotton:1963:CAG

[Cot63] F. Albert (Frank Albert) Cotton.

Chemical Applications of Group Theory. Interscience Publishers, New York, NY, USA, 1963. ix + 295 pp. LCCN QD461 .C65.

Couffignal:1933:MCL

[Cou33] L. Couffignal. Les machines à calculer, leurs principes, leur évolution. (French) [Calculating machines, their principles, their evolution]. Gauthier-Villars, Paris, France, 1933. Extracts reprinted in [Ran82, §3.2]. Translated by Mr. R. Basu.

Couffignal:1938:AMA

[Cou38] L. Couffignal. Sur l’analyse mécanique. Application aux machines à calculer et aux calculs de la mécanique céleste. (French) [On mechanical analysis. Application to calculating machines and to calculation in celestial mechanics]. PhD thesis, Faculté des Sciences de Paris, Paris, France, 1938. Extracts reprinted in [Ran82, §2.7]. Translated by Mr. R. Basu.

Counihan:1991:F

[Cou91] Martin Counihan. Fortran 90. Pitman Publishing Ltd., London, UK, 1991. ISBN 0-273-03073-6. 309 pp. LCCN QA76.73.F25 C68 1991. See [Ame92].

Cowell:1984:SDM

[Cow84] Wayne R. Cowell, editor. Sources and Development of Mathematical Software. Prentice-Hall Series in Computational Mathematics, Cleve Moler, Advisor. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-823501-5. xii + 404 pp. LCCN QA76.95 .S68 1984.

Cowlishaw:1990:RLP

[Cow90] Michael F. Cowlishaw. The REXX language: a practical approach to programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1990. ISBN 0-13-780651-5. xii + 203 pp. LCCN QA76.73.R24 C69 1990. URL http:/

/vig.prenhall.com/catalog/
academic/product/0,1144,0137806515,
00.html.

Cox:2015:SBS

[Cox15] Trevor J. Cox. The sound book:

the science of the sonic wonders of the world. W. W. Norton & Co., New York, NY, USA, 2015. ISBN 0-393-23979-9 (hardcover), 0-393-35058-4 (paperback). 331 pp. LCCN QC225.3 .C69 2015.

Cuyt:2008:HCF

[CPV^{+}08] Annie Cuyt, Vigdis B. Petersen, Brigitte Verdonk, Haakon Waadeland, and William B. Jones. Handbook of Continued Fractions for Special Functions. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-4020-6948-0. xx + 440 pp. LCCN QA295 .H275 2008.

Cameron:1991:LGE

[CR91] Debra Cameron and Bill Rosenblatt. Learning GNU Emacs. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-84-6. xxvii + 411 pp. LCCN QA76.76.T49 C35 1991. URL http:/

/www.oreilly.com/catalog/
9780937175842.

Craig:1971:DT

[Cra71] James Craig. Designing with Type. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 1971. ISBN 0-8230-1320-0. 175 pp. LCCN 70-159564.

Crandall:1991:MSa

[Cra91] Richard E. Crandall. Mathematica for the Sciences. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-51001-4. xiii + 300 pp. LCCN Q172 .C73 1991.

Craig:2005:VM

[Cra05] Iain D. Craig. Virtual Machines. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-85233-969-1, 1-84628-246-2 (e-book). xv + 269 pp. LCCN QA76.9 CRA.

Cravens:2007:PSW

[Cra07] Gwyneth Cravens. Power to save the world: the truth about nuclear energy. Alfred A. Knopf, New York, NY, USA, 2007. ISBN 0-307-26656-7. xv + 439 pp. LCCN TK9146 .C65 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0804/2007017611-
b.ht;   http://www.loc.gov/
catdir/enhancements/fy0804/
2007017611-d.ht;   http://
www.loc.gov/catdir/enhancements/
fy0804/2007017611-s.ht;
http://www.loc.gov/catdir/
toc/ecip0716/2007017611.html.

Crease:2010:GEB

[Cre10] Robert P. Crease. The Great Equations: Breakthroughs in Science from Pythagoras to Heisenberg. W. W. Norton & Co., New York, NY, USA, 2010. ISBN 0-393-33793-6. 315 pp. LCCN Q175 .C884 2010.

Crease:2011:WBH

[Cre11] Robert P. Crease. World in the balance: the historic quest for an absolute system of measurement. W. W. Norton & Co., New York, NY, USA, 2011. ISBN 0-393-07298-3. 317 pp. LCCN QC83 .C74 2011.

Crowther:1974:CLb

[Cro74] J. G. (James Gerald) Crowther.

The Cavendish Laboratory, 1874– 1974. Science History Publications, New York, NY, USA, 1974. ISBN 0-88202-029-3 (hardcover), 1-349-01670-5 (e-book). xvi + 464 pp. LCCN QC51.G72 C353 1974.

Crystal:2004:SE

[Cry04] David Crystal. The Stories of English. The Overlook Press, New York, NY, USA, 2004. ISBN 1-58567-601-2, 1-58567-719-1 (paperback), 1-4683-0617-0 (e-book). vi + 584 pp. LCCN PE1074.7 .C79 2004.

Comer:1991:ITI

[CS91] Douglas E. Comer and David L.

Stevens. Internetworking with TCP/IP. Design, Implementation, and Internals, volume 2. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-465378-5, 0-13-472242-6. LCCN TK5105.5 .C59 1991-92.

Chassell:1999:TGD

[CS99] Robert J. Chassell and Richard M. Stallman. Texinfo: The GNU Documentation Format (for Texinfo version 4.0, 28 September 1999). Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1999. ISBN 1-882114-67-1. x + 244 pp. LCCN QA76.76.T49 C53 1999. US$25.00.

Christiansen:1998:PC

[CT98] Tohm Christiansen and Nathan Torkington. Perl Cookbook: Solutions and examples for Perl programmers. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 1-56592-243-3. xxxiv + 757 pp. LCCN QA76.73.P22 C57 1998. US$39.95. URL http:

//www.oreilly.com/catalog/
cookbook/.

Cooper:2004:EC

[CT04] Keith D. (Keith Daniel) Cooper and Linda Torczon. Engineering a Compiler. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2004. ISBN 1-55860-698-X (paperback). xxx + 801 pp. LCCN QA76.76.C65 C675 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
http://www.loc.gov/catdir/
description/els051/2004268209.

html;   http://www.mkp.com/
engineeringacompiler/.

Cooper:2011:EC

[CT11] Keith D. (Keith Daniel) Cooper and Linda Torczon. Engineering a compiler. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 2011. ISBN 0-12-088478-X. xxiii + 800 pp. LCCN QA76.76.C65 C675 2011. URL http://

mkp.com/news/engineering-a-
compiler-2nd-edition.

Cleaveland:1977:GPL

[CU77] J. Craig Cleaveland and Robert C. Uzgalis. Grammars for Programming Languages, volume 4 of Programming Languages Series, Editor: Thomas E. Cheatham. Elsevier, Amsterdam, The Netherlands, 1977. ISBN 0-444-00187-5 (hardcover), 0-444-00199-9 (paperback). xiii + 154 pp. LCCN QA76.7.C571.

Coleman:1988:HMC

[CV88] Thomas F. Coleman and Charles F.

Van Loan. Handbook for Matrix Computations, volume 4 of Frontiers in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1988. ISBN 0-89871-227-0 (paperback), 1-61197-104-7 (ebook). vii + 264 pp. LCCN QA188 .C651 1988. US$34.00.

Cotton:1966:AIC

[CW66] F. Albert (Frank Albert) Cotton and Geoffrey Wilkinson. Advanced Inorganic Chemistry. Interscience Publishers, New York, NY, USA, 1966. xii + 1136 pp. LCCN QD151 .C64 1966.

Cody:1980:SME

[CW80] William J. Cody, Jr. and William Waite. Software Manual for the Elementary Functions. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1980. ISBN 0-13-822064-6. x + 269 pp. LCCN QA331 .C635 1980.

Catmull:2014:CIO

[CW14] Edwin E. Catmull and Amy Wallace. Creativity, Inc.: overcoming the unseen forces that stand in the way of true inspiration. Random House, New York, NY, USA, 2014. ISBN 0-8129-9301-2 (hardcover), 0-679-64450-4 (ebook). xvi + 340 + 8 pp. LCCN HD53 .C394 2014.

Cybenko:perfect

[Cyb91] George Cybenko. Supercomputer performance trends and the Perfect Benchmarks. Supercomputing Review, 4(4):53–60, April 1991.

daCruz:1987:KFT

[da 87] Frank da Cruz. Kermit: a File Transfer Protocol. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1987. ISBN 0-932376-88-6. xvii + 379 pp. LCCN TK5105.5 .D23 1987.

Darwin:2001:JCS

[Dar01] Ian Darwin. Java Cookbook: Solutions and Examples for Java Developers. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, June 2001. ISBN 0-596-00170-3 (paperback). xxix + 850 pp. LCCN QA76.73.J38 D348 2001. US$44.95. URL http:

//www.oreilly.com/catalog/
9780596001704;   http://www.
oreilly.com/catalog/javacook.

Davis:1965:APC

[Dav65] Jeff C. Davis, Jr. Advanced Physical Chemistry: Molecules, Structure, and Spectra. Ronald Press Company, New York, NY, USA, 1965. x + 632 pp. LCCN QD453.D36.

deBalbine:design

[dB75] Guy de Balbine. Design criteria for the structuring engine. In AFIPS:1975:SUJ [AFI75], page ?? LCCN QA76 .U2 1975.

DeBrosse:2004:SBU

[DB04] Jim DeBrosse and Colin B.

Burke. The Secret in Building 26: the Untold Story of America’s Ultra War Against the U-boat Enigma Codes. Random House, New York, NY, USA, 2004. ISBN 0-375-50807-4, 1-58836-353-8 (ebook), 0-375-75995-6. xxix + 272 pp. LCCN D810.C88 D43 2004. URL http://

www.loc.gov/catdir/samples/
random045/2003058494.html;
http://www.randomhouse.com/
catalog/display.pperl?isbn=
9781588363534.

Dahlquist:1974:NM

[DBA74] Germund Dahlquist, Åke Björck, and Ned Anderson. Numerical Methods. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1974. ISBN 0-13-627315-7. xviii + 573 pp. LCCN QA297 .D131 1969. Translated by Ned Anderson.

Dongarra:1979:LUG

[DBMS79] J. J. Dongarra, J. R. Bunch, C. B. Moler, and G. W. Stewart. LINPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1979. ISBN 0-89871-172-X (paperback). 320 pp. LCCN QA76.73 .L22 L5 1979; QA184 .L56 1982; QA214 .L56 1979.

deBruin:1988:PML

[dBvLV88] R. de Bruin, C. G. van der Laan, J. R. Luyten, and H. F. Vogt. Publiceren met LaTEX, volume 19. Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1988. ISBN 90-6196-357-5. 196 pp. LCCN Z253.4.L38 .P97 1988.

Dodd:1997:ASG

[DCJ^{+}97] Janet S. Dodd, Marvin Coyner, Madeleine Jacobs, Barbara Friedman Polansky, K. Barbara Schowen, David Weisgerber, and Larry James Winn, editors. The ACS Style Guide: a Manual for Authors and Editors. American Chemical Society, Washington, DC, USA, second edition, 1997. ISBN 0-8412-3461-2 (hardcover), 0-8412-3462-0 (paperback). xi + 460 pp. LCCN QD8.5.A25 1997. US$27.50.

Dongarra:1991:SLS

[DDSv91] Jack J. Dongarra, Iain S. Duff, Danny C. Sorensen, and Henk A. van der Vorst. Solving Linear Systems on Vector and Shared Memory Computers. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1991. ISBN 0-89871-270-X. x + 256 pp. LCCN QA184 .S65 1991.

DeWolk:2019:ADS

[De 19] Roland De Wolk. American disruptor: the scandalous life of Leland Stanford. University of California Press, Berkeley, CA, USA, 2019. ISBN 0-520-30547-7 (hardcover), 0-520-97356-9 (e-book). xvii + 299 + 23 pp. LCCN E664.S78 D4 2019.

DelZoppo:1976:VP

[Del76] Annette Del Zoppo. The Vegreville Pysanka. Ronald D. Resch, Salt Lake City, UT, USA, 1976. 16 pp.

DeLamarter:1986:BBI

[DeL86] Richard Thomas DeLamarter.

Big Blue: IBM’s use and abuse of power. Dodd, Mead, New York, NY, USA, 1986. ISBN 0-396-08515-6. xviii + 393 pp. LCCN HD9696.C64 I4835 1986.

Denbigh:1966:PCE

[Den66] Kenneth Denbigh. The Principles of Chemical Equilibrium: With Applications in Chemistry and Chemical Engineering. Cambridge University Press, Cambridge, UK, 1966. xxi + 494 pp. LCCN QD501 .D365 1966.

Duff:1986:DMS

[DER86] I. S. Duff, A. M. Erisman, and J. K. Reid. Direct Methods for Sparse Matrices. Clarendon Press, Oxford, UK, 1986. ISBN 0-19-853408-6 (hardcover), 0-19-853421-3 (paperback). xiii + 341 pp. LCCN QA188 .D841 1986. US$37.50.

Derbyshire:2003:POH

[Der03] John Derbyshire. Prime obsession: Bernhard Riemann and the greatest unsolved problem in mathematics. Joseph Henry Press, Washington, DC, USA, 2003. ISBN 0-309-08549-7. xv + 422 pp. LCCN QA246 .D47 2003. URL http://www.

loc.gov/catdir/toc/fy038/
2002156310.html;   https:/
/www.nap.edu/catalog/10532/
prime-obsession-bernhard-
riemann-and-the-greatest-
unsolved-problem-in.

DesJardins:2010:MCC

[Des10] Julie Des Jardins. The Madame Curie complex: the hidden history of women in science. Women writing science. Feminist Press at the City University of New York, New York, NY, USA, 2010. ISBN 1-55861-613-6 (paperback), 1-55861-655-1 (e-book). 312 + 8 pp. LCCN Q141 .D44 2010.

URL http://catdir.loc.gov/
catdir/enhancements/fy1008/
2009037794-b.html;   http:/
/catdir.loc.gov/catdir/enhancements/
fy1008/2009037794-d.html.

Devitt:1993:CMV

[Dev93] John S. Devitt. Calculus with Maple V. Brooks/Cole, Pacific Grove, CA, USA, 1993. ISBN 0-534-16362-9. xxii + 502 pp. LCCN QA303.5.D37 D48 1993.

Devlin:2000:MGH

[Dev00] Keith Devlin. The Math Gene:

How Mathematical Thinking Evolved and Why Numbers Are Like Gossip. Basic Books, New York, NY, USA, 2000. ISBN 0-465-01618-9. xvii + 328 pp. LCCN QA141.5 .D49 2000.

Devlin:2002:MPS

[Dev02] Keith Devlin. The Millennium Problems: the Seven Greatest Unsolved Mathematical Puzzles of Our Time. Basic Books, New York, NY, USA, 2002. ISBN 0-465-01729-0 (hardcover), 0-465-01730-4 (paperback), 2-00-353546-8. x + 237 pp. LCCN QA93 .D485 2002. US$26.00.

Devlin:2008:UGP

[Dev08a] Keith J. Devlin. The unfinished game: Pascal, Fermat, and the seventeenth-century letter that made the world modern: a tale of how mathematics is really done. Basic ideas. Basic Books, New York, NY, USA, 2008. ISBN 0-465-00910-7. x + 191 pp. LCCN QA273 .D455 2008. US$24.95.

Devore:2008:PSE

[Dev08b] Jay L. Devore. Probability and Statistics for Engineering and the Sciences. Thomson/Brooks/ Cole, Belmont, CA, USA, seventh edition, 2008. ISBN 0-495-38217-5. xvi + 720 pp. LCCN QA273 .D46 2008. URL

http://www.loc.gov/catdir/
enhancements/fy1104/2006932557-
b.html;   http://www.loc.
gov/catdir/enhancements/fy1104/
2006932557-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy1212/2006932557-t.html.

Devlin:2011:MNF

[Dev11] Keith J. Devlin. The man of numbers: Fibonacci’s arithmetic revolution. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 2011. ISBN 0-8027-7812-7 (hardcover). viii + 183 + 8 pp. LCCN QA29.F5 D48 2011.

Dongarra:netlib

[DG87a] Jack Dongarra and Eric Grosse. Distribution of mathematical software via electronic mail. Communications of the Association for Computing Machinery, 30(5):403–407, May 1987.

Dougherty:1987:MSI

[DG87b] Edward R. Dougherty and Charles R. Giardina. Matrix Structured Image Processing. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-565623-0. x + 258 pp. LCCN TA1632.D68 1987. US$44.95.

daCruz:1993:UCK

[dG93] Frank da Cruz and Christine M.

Gianone. Using C-Kermit. Digital Press and Prentice-Hall, 12 Crosby Drive, Bedford, MA 01730, USA and Upper Saddle River, NJ 07458, USA, 1993. ISBN 1-55558-108-0 (DP), 0-13-037490-3 (PH). xxi + 514 pp. LCCN TK5105.9.D33 1993. Available in German translation [dG96].

daCruz:1996:CKE

[dG96] Frank da Cruz and Christine M. Gianone. C-Kermit: Einführung und Referenz. Verlag Heinz Heise, Helsdorfer Straße 7, D-30625, Hannover, Germany, 1996. ISBN 3-88229-023-4. 557 pp. LCCN ???? 88,00 DM.

URL http://www.emedia.de/
bin/bookshop?show=3401&id=.

German translation of [dG93].

daCruz:1997:UCK

[dG97] Frank da Cruz and Christine M.

Gianone. Using C-Kermit. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, second edition, 1997. ISBN 1-55558-164-1. xxii + 662 pp. LCCN TK5105.9.D33 1997.

Diamond:2008:HDS

[DHG08] Harold G. Diamond, H. (Heini)

Halberstam, and William F. Galway. A higher-dimensional sieve method: with procedures for computing sieve functions, volume 177 of Cambridge tracts in mathematics. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-89487-5 (hardcover). xxi + 266 pp. LCCN QA246 .D53 2008.

                                        URL http://www.math.uiuc.
                                        edu/SieveTheoryBook.  The

Mathematica package of sieverelated functions described in Appendix 1, as well as a list of comments and corrigenda, will be maintained

                                        at http://www.math.uiuc.edu/
                                        SieveTheoryBook.

DeRespinis:2012:ISG

[DHJ^{+}12] Francis DeRespinis, Peter Hayward, Jana Jenkins, Amy Laird, Leslie McDonald, and Eric Radzinski. The IBM style guide: conventions for writers and editors. IBM Press/Pearson, Upper Saddle River, NJ, USA, 2012. ISBN 0-13-210130-0 (paperback). xxvi + 389 pp. LCCN T11 .I15 2012.

Dickson:2003:SSC

[Dic03] Paul Dickson. Sputnik: the Shock of the Century. Berkley Books, New York, NY, USA, 2003. ISBN 0-425-18843-4. 310 pp. LCCN TL796.5.S652.S664; TL796.5.S652.S664 2003.

DEC:1982:PVC

[Dig82] Digital Equipment Corporation. Programming in VAX-11 C. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, May 1982.

Dijkstra:goto-harmful

[Dij68] Edsger Wybe Dijkstra. Go to statement considered harmful. Communications of the Association for Computing Machinery, 11(3):147–148, March 1968. This paper inspired scores of others, published mainly in SIGPLAN Notices up to the mid-1980s. The best-known is [Knu74].

Dijkstra:1982:SWC

[Dij82] Edsger Wybe Dijkstra. Selected Writings on Computing: a Personal Perspective. Texts and Monographs in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). xvii + 362 pp. LCCN QA76.24 .D54 1982. Including a paper coauthored by C. S. Schölten.

Diller:1993:LLL

[Dil93] Antoni Diller. LaTEX Line by Line: Tips and Techniques for Document Processing. Wiley Professional Computing. Wiley, New York, NY, USA, 1993. ISBN 0-471-93471-2 (softcover). xiii + 291 pp. LCCN Z253.4.L38 D55 1993. US$29.95. Optionally accompanied by disk with examples, ISBN 0-471-93797-5.

Donovan:2016:GPL

[DK16] Alan A. A. Donovan and Brian W. Kernighan. The Go Programming Language. Addison-Wesley professional computing series. Addison-Wesley, Reading, MA, USA, 2016. ISBN 0-13-419044-0 (paperback). xvii + 380 pp. LCCN QA76.73.G63 D66 2016. URL https://

www.pearson.com/us/higher-

                                        education/program/Donovan-
                                        Go-Programming-Language-The/
                                        PGM234922.html.

Dalang:2009:MSP

[DKM^{+}09] Robert Dalang, Davar Khoshnevisan, Carl Mueller, David Nualart, and Yimin Xiao. A Minicourse on Stochastic Partial Differential Equations, volume 1962 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-540-85993-4 (softcover). xi + 216 pp. LCCN QA3 .L28 no.1962; QA3.L471.

Davie:RDC81

[DM81] A. J. T. Davie and R. Morrison. Recursive Descent Compiling. Ellis Horwood Series in Computers and their Applications, Editor: Brian Meek. Ellis Horwood, New York, NY, USA, 1981. ISBN 0-470-27270-8 (Halstead Press), 0-85312-386-1 (Ellis Horwood). 195 pp. LCCN QA76.6 .D373.

Dollhoff:16B79

[Dol79] Terry Dollhoff. 16-Bit Microprocessor Architecture. Reston Publishing Co. Inc., Reston, VA, USA, 1979. ISBN 0-8359-7001-9. xii + 471 pp. LCCN QA76.5 .D65 1979.

Dolnick:2011:CUI

[Dol11] Edward Dolnick. The Clockwork Universe: Isaac Newton, the Royal Society, and the Birth of the Modern World. Harper-Collins College Publishers, New York, NY, USA, 2011. ISBN 0-06-171951-X. xviii + 378 pp. LCCN Q127.E8 D65 2011.

Dongarra:ANL-23-89

[Don89] J. J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. Technical Memorandum 23, Argonne National Laboratory, Argonne, IL, USA, June 4 1989.

Doob:1993:TSS

[Doo93] Michael Doob. TEX: Starting from Square 1. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-56441-1, 0-387-56441-1 (US). x + 114 pp. LCCN Z253.4.T47 D66 1993. US$25.00. URL

http://www.loc.gov/catdir/
enhancements/fy0902/93026025-
d.html.

Dorain:1965:SIC

[Dor65] Paul B. Dorain. Symmetry in Inorganic Chemistry. Addison-Wesley, Reading, MA, USA, 1965. vi + 122 pp. LCCN QD475.D6.

Dorren:2014:LAE

[Dor14] Gaston Dorren. Lingo: around Europe in sixty languages. Grove Press, New York, NY, USA, 2014. ISBN 0-8021-2571-9 (softcover). 303 pp. LCCN P380 .D6713 2016. With contributions by Jenny Audring, Frauke Watson, and Alison Edwards.

Dorren:2018:BAW

[Dor18] Gaston Dorren. Babel: around the world in twenty languages. Atlantic Monthly Press, New York, NY, USA, 2018. ISBN 0-8021-2879-3 (hardcover). 361 pp. LCCN P375 .D67 2018.

Doupnik:1990:MMK

[Dou90] Joe R. Doupnik. The Making of MS-DOS Kermit. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1990. Not yet published.

Dougherty:1991:SA

[Dou91] Dale Dougherty. sed & awk. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-59-5. xxii + 394 pp. LCCN QA76.76.U84 D69 1991. URL http:

                                        //www.oreilly.com/catalog/
                                        9780937175590.

Dowd:1993:HPC

[Dow93] Kevin Dowd. High Performance Computing. RISC architectures, optimization and benchmarks; A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-032-5. xxv + 371 pp. LCCN QA76.88 .D6 1993; QA76.9.A73 D68 1993. US$25.95. URL http:

                                        //www.oreilly.com/catalog/
                                        9781565920323.

Doxiadis:2000:UPG

[Dox00] Apostolos K. Doxiadis. Uncle Petros and Goldbach’s Conjecture. Faber and Faber, London, UK, 2000. ISBN 0-571-20203-9 (paperback), 0-571-29569-X (e-book). 209 pp. LCCN PA5615.O87 T4813 2000b.

Davis:1975:MNI

[DR75] Philip J. Davis and Philip Rabinowitz. Methods of numerical integration. Computer science and applied mathematics. Academic Press, New York, NY, USA, 1975. ISBN 0-12-206350-3. xii + 459 pp. LCCN QA299.3 .D28.

Dershowitz:1997:CC

[DR97] Nachum Dershowitz and Edward M. Reingold. Calendrical Calculations. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-56474-3 (paperback), 0-521-56413-1 (hardcover). xxi + 307 pp. LCCN CE12.D47 1997.

Dray:2012:GSR

[Dra12] Tevian Dray. The Geometry of Special Relativity. Taylor and Francis, London, UK and Boca Raton, FL, USA, 2012. ISBN 1-4665-1047-1. xvii + 131 pp. LCCN QC173.65 .D73 2012.

Dvonch:color-pdl

[DRB89] Mary Ann Dvonch, Paul Roetling, and Robert Buckley. Color descriptors in page description languages. Proceedings of the Society for Information Display, 30 (2):177–181, 1989.

Dry:2014:NPS

[Dry14] Sarah Dry. The Newton Papers: the Strange and True Odyssey of Isaac Newton’s Manuscripts. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-995104-7 (hardcover). xi + 238 pp. LCCN QC16.N7 D79 2014.

Degering:1951:OCO

[DS51] Ed F. Degering and Seventy Six Assistant Editors, editors. Organic Chemistry: An Outline of the Beginning Course Including Material for Advanced Study, volume 6 of College Outline Series. Barnes and Noble, New York, NY, USA, sixth edition, 1951. xxv + 422 pp. LCCN QD256 .D4 1968. Revised, 1957.

Dennis:1983:NMU

[DS83] J. E. Dennis, Jr. and Robert B. Schnabel. Numerical Methods for Unconstrained Optimization and Nonlinear Equations. Prentice-Hall Series in Computational Mathematics, Cleve Moler, Advisor. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1983. ISBN 0-13-627216-9. xiii + 378 pp. LCCN QA402.5 .D44 1983.

Donnelly:bison

[DS88a] Charles Donnelly and Richard M. Stallman. BISON: The YACC-compatible parser generator. Technical report, Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1988. Bison was largely written by Robert Corbett, and made yacccompatible by Richard Stallman. Electronic mail: rms@prep.ai. mit.edu. Software also available via ANONYMOUS FTP

to prep.ai.mit.edu. See also

[Pax88].

Durbeck:1988:OHD

[DS88b] Robert C. Durbeck and Sol Sherr, editors. Output Hardcopy Devices. Academic Press, New York, NY, USA, 1988. ISBN 0-12-225040-0. xii + 526 pp. LCCN TK7887.7 .O981 1988. URL http://www.

sciencedirect.com/science/
book/9780122250408. See ar-

ticle by Mike Parker on printer fonts, and undesirability of white-writing engines. I have not bought this book.

DeGroot:2002:PS

[DS02] Morris H. DeGroot and Mark J.

Schervish. Probability and Statistics. Addison-Wesley, Reading, MA, USA, third edition, 2002. ISBN 0-201-52488-0, 0-321-20473-5 (paperback), 0-201-71129-X (Solutions manual). xv + 816 pp. LCCN QA273 .D35 2002.

duSautoy:2003:MPS

[dS03] Marcus du Sautoy. The music of the primes: searching to solve the greatest mystery in mathematics. HarperCollins College Publishers, New York, NY, USA, 2003. ISBN 0-06-621070-4. 335 pp. LCCN QA246 .D8 2003. US$24.95, CAN$38.95.

deGrasseTyson:2008:PFR

[dT08] Neil deGrasse Tyson. The Pluto files: the rise and fall of America’s favorite planet. W. W. Norton & Co., New York, NY, USA, 2008. ISBN 0-393-06520-0 (hardcover). xii + 194 pp. LCCN QB701 .T97 2009.

deGrasseTyson:2018:AWU

[dTL18] Neil deGrasse Tyson and Avis Lang. Accessory to War: the Unspoken Alliance Between Astrophysics and the Military. W. W. Norton & Co., New York, NY, USA, 2018. ISBN 0-393-06444-1 (hardcover). xiv + 576 pp. LCCN UG1523 .T97 2018.

DuBois:1999:M

[DuB99] Paul DuBois. MySQL. New Riders Publishing, Carmel, IN, USA, 1999. ISBN 0-7357-0921-1. xxix + 756 pp. LCCN QA76.73.S67 D83 2000. US$49.99. URL http:

                                        //www.mysql.com/Contrib/;
                                        http://www2.newriders.com/
                                        cfm/prod_book.cfm?RecordID=
                                        584.

Duckworth:1960:EM

[Duc60] Henry E. Duckworth. Electricity and Magnetism. Holt, Rinehart, and Winston, New York, NY, USA, 1960. xv + 424 pp. LCCN QC519 .D8.

Dupre:1995:BW

[Dup95] Lyn Dupré. Bugs in Writing. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-60019-6 (paperback). xxxiii + 649 pp. LCCN PEI 408.D85 1995. US$19.95, CDN$25.95.

Dupre:1998:BWG

[Dup98] Lyn Dupré. BUGS in writing: a guide to debugging your prose. Addison-Wesley, Reading, MA, USA, revised. edition, 1998. ISBN 0-201-37921-X. xxxv + 660 pp. LCCN PE1408 .D85 1998.

Dushman:1938:EQM

[Dus38] Saul Dushman. The Elements of Quantum Mechanics. Wiley, New York, NY, USA, 1938. xiii + 452 pp. LCCN QC174.1 .D8.

Dreyer:1946:REM

[DW46] H.-J. Dreyer and A. Walther. Der Rechenautomat Ipm. Entwicklung Mathematischer Instrumente in Deutschland 1939 bis 1945. (German) [The Ipm calculator. The development of mathematical instruments in Germany 1939–1945]. Bericht A3, Institut für Praktische Mathematik, Technische Hochschule, Darmstadt, West Germany, August 19, 1946. Reprinted in [Ran82, §3.3]. Translated by Mr. and Mrs. P. Jones.

Daniels:1962:EPC

[DWB^{+}62] Farrington Daniels, J. W.

Williams, Paul Bender, Robert A. Alberty, and C. D. Cornwell. Experimental Physical Chemistry. McGraw-Hill, New York, NY, USA, sixth edition, 1962. xv + 625 pp. LCCN QD457 .D2 1962.

Dybvig:1987:SPL

[Dyb87] R. Kent Dybvig. The SCHEME programming language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-791864-X (paperback). 242 pp. LCCN QA76.73.S34 D93 1987.

Dybvig:1996:SPL

[Dyb96] R. Kent Dybvig. The SCHEME Programming Language: ANSI Scheme. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1996. ISBN 0-13-454646-6. xii + 248 pp. LCCN QA76.73.S34 D93 1996. US$64.00. URL http://www.

                                        scheme.com/tspl2d/.

Dybvig:2003:SPL

[Dyb03] R. Kent Dybvig. The Scheme programming language. MIT Press, Cambridge, MA, USA, third edition, 2003. ISBN 0-262-54148-3 (paperback). xi + 295 pp. LCCN QA76.73.S34.D93; QA76.73.S34.D93 2003.

Dyer:1965:AAS

[Dye65] John R. (John Robert) Dyer. Applications of Absorption Spectroscopy of Organic Compounds. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1965. x + 147 pp. LCCN QD476 .D9.

Dylla:2020:SJP

[Dyl20] H. Frederick Dylla. Scientific Journeys: a Physicist Explores the Culture, History and Personalities of Science. Springer, Cham, Switzerland, 2020. ISBN 3-030-55799-5 (paperback), 3-030-55800-2 (e-book). xxv + 222 pp. LCCN Q125 .D955 2020.

Dyson:2010:MCG

[Dys10]  Freeman J. Dyson.  Many-

Colored Glass: Reflections on the Place of Life in the Universe. University of Virginia Press, Charlottesville, VA, USA, 2010. ISBN 0-8139-2973-3. xi + 162 pp. LCCN QH501 .D97 2010. Page–Barbour Lectures for 2004.

Dyson:2012:TCO

[Dys12] George Dyson. Turing’s cathedral: the origins of the digital universe. Pantheon Books, New York, NY, USA, 2012. ISBN 0-375-42277-3 (hardcover). xxii + 401 pp. LCCN QA76.17 .D97 2012.

Earnshaw:1985:FAC

[Ear85] Rae A. Earnshaw. Fundamental Algorithms for Computer Graphics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-13920-6 (New York), 3-540-13920-6 (Berlin). xvi + 1042 pp. LCCN QA76.6 .N379 1985.

Earnshaw:1987:WPS

[Ear87] Rae A. Earnshaw, editor. Workstations and Publication Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96527-0, 3-540-96527-0. viii + 229 pp. LCCN Z286.E43 W67 1987. US$29.50.

Eckstein:2000:US

[ECBK00] Robert Eckstein, David Collier-Brown, and Peter Kelly. Using Samba. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-449-5. xi + 398 pp. LCCN QA76.9.C55 E267 2000; QA76.9.C55 E267 2000; TK5105.5 .E35 2000. US$34.95. URL http://www.

oreilly.com/catalog/9781565924499.

Includes CD-ROM with complete mirror of Samba site with Samba 2.0.5a, including sources, documentation, binaries, and supplementary tools.

Eckert:1948:EC

[Eck48] W. J. Eckert. Electrons and computation. The Scientific Monthly, 67(5):315–323, November 1948. CODEN SCMOAA. ISSN 0096-3771. Reprinted in [Ran82, §5.3].

Esterson:2019:EWR

[ECS19] Allen Esterson, David C. Cassidy, and Ruth Lewin Sime. Einstein’s wife: the real story of Mileva Einstein-Marić. MIT Press, Cambridge, MA, USA, 2019. ISBN 0-262-03961-3 (hardcover), 0-262-53897-0 (paperback). xxi + 313 pp. LCCN QC16.E52 E88 2019.

Edmonds:2008:HTA

[Edm08] Jeff Edmonds. How to Think About Algorithms. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-84931-4 (hardcover), 0-521-61410-4 (paperback), 0-511-41278-9 (e-book), 0-511-41370-X (ebook). xiii + 448 pp. LCCN QA9.58 .E36 2008. URL

http://www.loc.gov/catdir/
enhancements/fy0808/2008001238-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0808/
2008001238-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0808/2008001238-t.html.

Ethier:2007:OPM

[EE07] Stewart N. Ethier and William R.

Eadington, editors. Optimal Play: Mathematical Studies of Games and Gambling. Institute for the Study of Gambling and Commercial Gaming, Reno, NV, USA, 2007. ISBN 0-9796873-0-6. xxvi + 550 pp. LCCN QA273 .O685 2007.

Einstein:2005:MRb

[EG05] Albert Einstein and Brian Greene. The Meaning of Relativity, volume 1921 of The Stafford Little lectures. Princeton University Press, Princeton, NJ, USA, expanded Princeton science library edition, 2005. ISBN 0-691-12027-7. xxiv + 166 pp. LCCN QC6. URL http://

press.princeton.edu/titles/
484.html.

EguiaMoraza:2018:RHP

[Egu18] Iban Eguia Moraza. Rust High Performance: Learn to Skyrocket the Performance of Your Rust Applications. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78847-823-1. 265 pp. LCCN QA76.7.

Ehrlich:1973:PCP

[Ehr73] Robert Ehrlich. Physics and Computers; Problems, Simulations, and Data Analysis. Houghton-Mifflin, Boston, MA, USA, 1973. ISBN 0-395-59472-1 (??invalid checksum??). vi + 279 pp. LCCN QC20.2.E35.

Einstein:1966:EPE

[EI66] Albert Einstein and Leopold Infeld. The Evolution of Physics from Early Concepts to Relativity and Quanta. A Touchstone book. Simon and Schuster, New York, NY, USA, 1966. ISBN 0-671-20156-5. xvi + 302 pp. LCCN QC7 .E5 1961.

Eijkhout:1992:TTT

[Eij92] Victor Eijkhout. TEX by Topic, a TEXnician’s Reference. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56882-9. vii + 307 pp. LCCN Z253.4.T47 E38

                                        1992. URL http://eijkhout.
                                        net/texbytopic/.

Einstein:1954:IO

[Ein54] Albert Einstein. Ideas and Opinions. Crown Publishers, New York, NY, USA, 1954. vi + 377 pp. LCCN AC35 .E526 1954.

Einstein:1982:IO

[Ein82] Albert Einstein. Ideas and Opinions. Three Rivers Press, New York, NY, USA, 1982. ISBN 0-517-88440-2. 377 pp. LCCN AC35 .E526 1982x.

Eisenstaedt:2006:CHR

[Eis06] Jean Eisenstaedt. The Curious History of Relativity: How Einstein’s Theory of Gravity Was Lost and Found Again, volume 15 of CNRS editions. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-11865-5. ix + 363 pp. LCCN QC173.6 EIS; 06.E06047.

Eisenberg:1969:SPW

[EK69] D. (David) Eisenberg and W. (Walter) Kauzmann. The Structure and Properties of Water. Clarendon Press, Oxford, UK, 1969. ISBN 0-19-855339-0. ix + 296 pp. LCCN QD169.W3 E35.

Ekman:2016:MBD

[Ekm16] Martin Ekman. The man behind “Degrees Celsius”: a pioneer in investigating the Earth and its changes. Summer Institute for Historical Geophysics, Godby, Åland Islands, Finland, 2016. ISBN 952-93-7732-0. 159 pp. LCCN ????

Enderle:1984:CGP

[EKP84] G. Enderle, K. Kansy, and G. Pfaff. Computer Graphics Programming. GKS: The Graphics Standard. Symbolic Computation, Editor: J. Encarnação and P. Hayes. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 0-387-11525-0 (New York), 3-540-11525-0 (Berlin). xvi + 542 pp. LCCN T385 .E531

Enderle:1987:CGP

[EKP87] G. Enderle, K. Kansy, and G. Pfaff. Computer Graphics Programming. GKS: The Graphics Standard. Symbolic Computation, Editor: J. Encarnação and P. Hayes. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1987. ISBN 3-540-16317-4 (Berlin), 0-387-16317-4 (New York). xxiii + 651 pp. LCCN T385 .E531 1987.

Ekstrom:1984:DIP

[Eks84] Michael P. Ekstrom. Digital Image Processing Techniques, volume 2 of Computational Techniques, Editors: Berni J. Alder and Sidney Fernbach. Academic Press, New York, NY, USA, 1984. ISBN 0-12-236760-X. xiii + 372 pp. LCCN TA1632 .D496 1984.

Ellis:1989:MCS

[EL89] Wade Ellis, Jr. and Ed Lodi. Maple for the Calculus Student. Brooks/Cole, Pacific Grove, CA, USA, 1989. ISBN 0-534-11874-7. vii + 67 pp. LCCN QA155.7.E4 E44 1989.

Eymard:2004:N

[EL04] Pierre Eymard and Jean-Pierre Lafon. The Number π. American Mathematical Society, Providence, RI, USA, 2004. ISBN 0-8218-3246-8. x + 322 pp. LCCN QA484 .E9613 2004. US$36.00. URL http://

                                        www.ams.org/bookpages/tnp/.

Translated by Stephen S. Wilson from the French Autour du nombre π (1999).

EFF:1998:CSE

[Ele98] Electronic Frontier Foundation.

Cracking DES: Secrets of Encryption Research, Wiretap Politics & Chip Design. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, July 1998. ISBN 1-56592-520-3. 272 pp. LCCN QA76.9.A25 C783. US$29.95. URL http:

//www.eff.org/descracker/;
 http://www.sunworld.com/
swol-07-1998/swol-07-if.html?
072098a#2.

Eliel:1962:SCC

[Eli62] Ernest L. (Ernest Ludwig) Eliel.

Stereochemistry of Carbon Compounds. McGraw-Hill, New York, NY, USA, 1962. xv + 486 pp. LCCN QD481 .E45.

Ellis:1990:FPI

[Ell90] T. M. R. Ellis. Fortran 77 Programming: With an Introduction to Fortran 90 Standard. International Computer Science Series. Addison-Wesley, Reading, MA, USA, second edition, 1990. ISBN 0-201-41638-7 (paperback). xxi + 641 pp. LCCN QA76.73.F25 E43 1990. US$38.75. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=
0-201-41638-7. Rev. ed. of:

A structured approach to FOR-TRAN 77 programming.

Eckhouse:1979:MSO

[EM79] Richard H. Eckhouse, Jr. and L. Robert Morris. Minicomputer Systems: Organization, Programming, and Applications (PDP-11). Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1979. ISBN 0-13-583914-9. xix + 491 pp. LCCN QA76.8.P2E26 1979.

Emling:2012:MCH

[Eml12] Shelley Emling. Marie Curie and her daughters: the private lives of science’s first family. Palgrave Macmillan, New York, NY, USA, 2012. ISBN 0-230-11571-3 (hardback). xx + 219 + 8 pp. LCCN QD22.C8 E46 2012.

England:1987:SUG

[Eng87] D. England. A Sun User’s Guide. Macmillan Computer Science Series, Editor: F. H. Sumner. Macmillan Education Ltd., Houndmills, Basingstoke, Hampshire RG21 2XS, UK, 1987. ISBN 0-333-44849-9. viii + 220 pp. LCCN QA76.8.S9 S85 1987.

Erdahl:1974:RDO

[Erd74] R. M. Erdahl, editor. Reduced Density Operators With Applications To Physical and Chemical Systems, II: proceedings of a conference held at Queen’s University, Kingston, June 20–22, 1974, volume 40 of Queen’s Papers in Pure and Applied Mathematics. Queen’s University, Kingston, ON, Canada, 1974. LCCN QA3.Q38 no.40.

Ellis:1990:ACR

[ES90] Margaret A. Ellis and Bjarne Stroustrup. The Annotated C++ Reference Manual. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-51459-1 (hardcover). x + 447 pp. LCCN QA76.73.C153 E35 1990.

Essinger:2004:JWH

[Ess04]  James Essinger.  Jacquard’s

Web: How a Hand-Loom Led to the Birth of the Information Age. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-280577-0. xi + 302 pp. LCCN QA76.17 .E88 2004. URL

http://www.loc.gov/catdir/
enhancements/fy0620/2004276200-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0620/
2004276200-t.html.

Estrada:1993:CI

[Est93] Susan Estrada. Connecting to the Internet. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, August 1993. ISBN 1-56592-061-9. xv + 170 pp. LCCN TK 5105.875 I57 E82c 1993. US$15.95.

Evans:2003:IAP

[ET03] James S. Evans and Gregory L.

Trimper. Itanium Architecture for Programmers: Understanding 64-bit Processors and EPIC Principles. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-101372-6. xxxiv + 529 pp. LCCN QA76.8.I83 E83 2003.

Ethier:2010:DCP

[Eth10] Stewart N. Ethier. The Doctrine of Chances: Probabilistic Aspects of Gambling. Probability and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-540-78782-8 (hardcover), 3-540-78783-6. xiv + 816 pp. LCCN QA271 .E84 2010.

Etter:1987:SFE

[Ett87] D. M. Etter. Structured FOR-TRAN 77 for engineers and scientists. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, second edition, 1987. ISBN 0-8053-2495-X (paperback). xxi + 519 pp. LCCN QA76.73.F25 E85 1987.

Etter:1993:SFE

[Ett93] Delores M. Etter. Structured Fortran 77 for Engineers and Scientists. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, fourth edition, 1993. ISBN 0-8053-1775-9. xxiii + 616 pp. LCCN QA76.73.F25 E85 1993.

Everett:1965:TC

[Eve65]  D. H. Everett.  Thermody-

namique Chimique. Dunod, Paris, France, 1965. xix + 252 pp. LCCN ???? French translation by R. Pâris of original English edition.

Eves:1983:IHM

[Eve83] Howard Eves. An Introduction to the History of Mathematics. Saunders College Publishing, Philadelphia, PA, USA, 1983. ISBN 0-03-062064-3. xviii + 593 pp. LCCN QA21.E8.

Frey:1989:VDE

[FA89] Donnalyn Frey and Rick Adams.

!%@:: a Directory of Electronic Mail Addressing and Networks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1989. ISBN 0-937175-39-0. xv + 284 pp. LCCN HE6239.E54 F73 1989.

Fagan:2010:CMH

[Fag10] Brian M. Fagan. Cro-Magnon:

how the Ice Age gave birth to the first modern humans. Bloomsbury Press, New York, NY, USA, 2010. ISBN 1-59691-582-X. xviii + 295 pp. LCCN GN286.3 .F34 2010.

Fagone:2017:WWS

[Fag17] Jason Fagone. The woman who smashed codes: a true story of love, spies, and the unlikely heroine who outwitted America’s enemies. Dey Street Books, New York, New York, 2017. ISBN 0-06-243048-3 (hardcover). xvi + 444 pp. LCCN Z103.4.U6.

Fagan:2019:LIA

[Fag19] Brian M. Fagan. The Little Ice Age: How Climate Made History 1300–1850. Basic Books, New York, NY, USA, 2019. ISBN 1-5416-1859-9 (paperback). xxi + 258 pp. LCCN QC989.A1 F34 2019.

Falk:2011:FCH

[Fal11] Dean Falk. The fossil chronicles:

how two controversial discoveries changed our view of human evolution. University of California Press, Berkeley, CA, USA, 2011. ISBN 0-520-26670-6 (hardcover). xiv + 259 pp. LCCN GN282.5 .F35 2011.

Farin:1988:CSC

[Far88] Gerald Farin. Curves and Surfaces for Computer-Aided Geometric Design. Academic Press, New York, NY, USA, 1988. ISBN 0-12-249050-9. xv + 334 + 4 pp. LCCN T385 .F371 1988.

Farin:1990:CSC

[Far90] Gerald Farin. Curves and Surfaces for Computer-Aided Geometric Design. Academic Press, New York, NY, USA, second edition, 1990. ISBN 0-12-249051-7. xvii + 444 pp. LCCN T385 .F371 1990.

Farin:NCS91

[Far91] Gerald Farin, editor. NURBS for Curve and Surface Design. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1991. ISBN 0-89871-286-6. LCCN QA224 .N85 1991.

Farin:1993:CSC

[Far93] Gerald Farin. Curves and Surfaces for Computer-Aided Geometric Design: a Practical Guide. Academic Press, New York, NY, USA, third edition, 1993. ISBN 0-12-249052-5, 0-12-249054-1, 1-4832-9699-7 (ebook). xvii + 473 pp. LCCN T385 .F37 1993.

Farin:1995:NCS

[Far95] Gerald E. Farin. NURB Curves and Surfaces: from Projective Geometry to Practical Use. A. K. Peters, Ltd., Wellesley, MA, USA, 1995. ISBN 1-56881-038-5. xii + 229 pp. LCCN QA224 .F37 1995. US$39.95.

Farin:1997:CSC

[Far97] Gerald Farin. Curves and Surfaces for Computer-Aided Geometric Design: a Practical Guide. Academic Press, New York, NY, USA, fourth edition, 1997. ISBN 0-12-249054-1. xvii + 429 pp. LCCN T385 .F37 1997.

Farin:1999:NCS

[Far99] Gerald E. Farin. NURB Curves and Surfaces: from Projective Geometry to Practical Use. A. K. Peters, Ltd., Wellesley, MA, USA, second edition, 1999. ISBN 1-56881-084-9. xv + 267 pp. LCCN QA224 .F37 1999.

Farin:2001:CSC

[Far01] Gerald E. Farin. Curves and Surfaces for CAGD: a Practical Guide. The Morgan Kaufmann series in computer graphics and geometric modeling. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, fifth edition, 2001. ISBN 1-55860-737-4, 0-08-050354-3 (e-book). xvii + 497 pp. LCCN T385 .F37 2001. US$69.95.

Farmelo:2003:IMB

[Far03] Graham Farmelo, editor. It Must Be Beautiful: Great Equations of Modern Science. Granta, London, UK, 2003. ISBN 1-86207-555-7. xviii + 284 pp. LCCN Q125 .I88 2003.

Farmelo:2009:SMH

[Far09] Graham Farmelo. The Strangest Man: The Hidden Life of Paul Dirac, Mystic of the Atom. Basic Books, New York, NY, USA, 2009. ISBN 0-465-01827-0. 539 + 8 pp. LCCN QC16.D57; QC16.D57 F37 2009. US$29.95. URL http:

                                        //bnreview.barnesandnoble.
                                        com/t5/The-Thinking-Read/
                                        The-Strangest-Man-The-Hidden-
                                        Life-of-Paul-Dirac-Mystic-
                                        of-the/ba-p/1243.

Farmelo:2013:CBHa

[Far13a] Graham Farmelo. Churchill’s bomb: a hidden history of science, war and politics. Faber and Faber, London, UK, 2013. ISBN 0-571-24978-7 (hardcover). x + 554 + 12 pp. LCCN ????

Farmelo:2013:CBHb

[Far13b] Graham Farmelo. Churchill’s bomb: how the United States overtook Britain in the first nuclear arms race. Basic Books, New York, NY, USA, 2013. ISBN 0-465-02195-6 (hardcover), 0-465-06989-4 (e-book). vi + 554 pp. LCCN UA647 .F28 2013.

Fattahi:1992:MVC

[Fat92] Abi Fattahi. Maple V Calculus Labs. Brooks/Cole, Pacific Grove, CA, USA, 1992. ISBN 0-534-19272-6. 92 + 3 pp. LCCN QA155.7.E4 F37 1992.

Ferguson:1993:MRM

[FB93] Paula Ferguson and David Brennan. Motif Reference Manual, volume 6B of The definitive guides to the X Window System. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, June 1993. ISBN 1-56592-038-4. xii + 908 pp. LCCN QA76.76.W56F47

1993.  US$34.95.  URL

http://www.loc.gov/catdir/
enhancements/fy0915/94185493-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0915/
94185493-d.html.

Farquhar:1994:MPH

[FB94] Erin Farquhar and Philip Bruce.

The MIPS Programmer’s Handbook. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1994. ISBN 1-55860-297-

6. viii + 408 pp. LCCN QA76.6 .F375 1994. US$36.95.

Feynman:2005:FTN

[FBD05] Richard P. (Richard Phillips)

Feynman, Laurie M. Brown, and P. A. M. (Paul Adrien Maurice) Dirac. Feynman’s Thesis: a New Approach to Quantum Theory. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2005. ISBN 981-256-366-0, 981-256-380-6 (paperback). xxii + 119 pp. LCCN QC174.12 .F48 2005.

Frakes:1992:IRD

[FBY92] William B. Frakes and Ricardo Baeza-Yates, editors. Information Retrieval: Data Structures and Algorithms. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-463837-9. viii + 504 pp. LCCN QA76.9.D351543 1992. US$56.00.

Flynn:sf2

[FC76] John A. Flynn and Gary P. Carr. User’s guide to SF-TRAN II. Technical Report 1846-79, Jet Propulsion Laboratory, Pasadena, CA, USA, October 1976.

Feeney:1988:SGM

[Fee88] Mary Feeney. The Standard Generalized Markup Language (SGML), volume 9 of Library & information briefing. British Library Research and Development Dept. and Library & Information Technology Centre, London, UK, 1988. ISSN 0954-1829. 15 pp. LCCN Z286.E43 F44 1988.

Feldman:fortlex

[Fel76] Stuart I. Feldman. FORTLEX: A general purpose lexical analyzer for Fortran. Technical Report 51, AT&T Bell Laboratories, Murray Hill, NJ 07974, USA, October 1976.

Fernbach:1986:SCV

[Fer86] Sidney Fernbach. Supercomputers: Class VI Systems, Hardware and Software. North-Holland, Amsterdam, The Netherlands, 1986. ISBN 0-444-87981-1. vii + 251 pp. LCCN QA76.5.S896 1986.

Ferraro:1988:PGE

[Fer88] Richard F. Ferraro. Programmer’s Guide to the EGA and VGA Cards. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-12692-3. xxvii + 607 pp. LCCN QA 76.8 I2594 .F48 1988. US$29.95.

Fernando:2004:GGP

[Fer04] Randima Fernando, editor. GPU gems: programming techniques, tips, and tricks for real-time graphics. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-22832-4. xvv + 765 pp. LCCN T385 .G6879 2004. US$45.99.

Feynman:1994:CPL

[Fey94] Richard Phillips Feynman. The Character of Physical Law. Modern Library, New York, NY, USA, modern library edition, 1994. ISBN 0-679-60127-9. xx + 167 pp. LCCN QC71 .F44 1994.

Feynman:1997:SEP

[Fey97] Richard Phillips Feynman. Six Not-so-easy Pieces: Einstein’s Relativity, Symmetry, and Space– time. Helix books. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-32841-0 (set), 0-201-15025-5 (hardcover), 0-201-32842-9 (paperback), 0-201-31151-8 (disc 1), 0-201-31152-6 (disc 2), 0-201-31153-4 (disc 3), 0-201-31154-2 (disc 4), 0-201-31155-0 (disc 5), 0-201-31156-9 (disc 6). xxvii + 152 pp. LCCN QC793.3.S9 F49 199. Originally prepared for publication by Robert B. Leighton and Matthew L. (Matthew Linzee) Sands. New introduction by Roger Penrose. See also [FLS95].

Feynman:1998:MIA

[Fey98] Richard Phillips Feynman. The Meaning of It All: Thoughts of a Citizen Scientist. Perseus Publishers, Cambridge, MA, USA, 1998. ISBN 0-7382-0166-9 (paperback). 133 pp. LCCN Q175.55 .F49 1998.

Feynman:2000:PFT

[Fey00]  Richard P. Feynman.  The

pleasure of finding things out: the best short works of Richard P. Feynman. Perseus Publishers, Cambridge, MA, USA, 2000. ISBN 0-7382-0349-1 (paperback). xvi + 270 pp. LCCN Q171 .F385 1999. Edited by Jeffrey Robbins, and foreword by Freeman Dyson.

Feynman:2005:PRD

[Fey05] Michelle Feynman, editor. Perfectly Reasonable Deviations (from the Beaten Track): the Collected Letters of Richard P. Feynman. Basic Books, New York, NY, USA, 2005. ISBN 0-7382-0636-9. xxiii + 486 pp. LCCN QC16 .F49 A4 2005. URL ftp://

uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http:
//www.loc.gov/catdir/toc/
ecip055/2005000049.html. Fore-

word by Timothy Ferris. See also interview with the editor [Ano05].

Ford:sftran

[FF78] William F. Ford and Theodore E.

Fessler. User’s guide for SF-TRAN/1100. Technical Report NASA Technical Paper 1200, NASA Lewis Research Center, April 1978.

Friedman:1987:LL

[FF87] Daniel Friedman and Matthias Felleisen. The Little LISPer. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-56038-0. xiv + 186 pp. LCCN QA76.73.L23 F74 1987.

Feynman:1995:ARP

[FF95] Richard P. (Richard Phillips)

Feynman and Michelle Feynman. The art of Richard P. Feynman: images by a curious character. GB Science Publishers SA, Basel, Switzerland, 1995. ISBN 2-88449-047-7. 173 pp. LCCN NC1075.F44 A4 1995.

Flannery:2001:CYW

[FF01] Sarah Flannery and David Flannery. In Code: a [Young Women’s] Mathematical Journey. Algonquin Books of Chapel Hill, Chapel Hill, NC, USA, 2001. ISBN 1-56512-377-8. ix + 341 pp. LCCN QA29.F6 A3 2003. US$13.95.

Feiveson:2014:UBF

[FGMvH14] Harold A. Feiveson, Alexander Glaser, Zia Mian, and Frank N. von Hippel. Unmaking the bomb: a fissile material approach to nuclear disarmament and nonproliferation. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-02774-7 (hardcover), 0-262-31918-7 (e-book). x + 4 + 277 pp. LCCN JZ5675 .F45 2014.

Francis:1993:GSS

[FGP93] Brian Francis, Mick Green, and Clive Payne, editors. The GLIM system: The Statistical System for Generalized Linear Interactive Modelling. Clarendon Press, Oxford, UK, 1993. ISBN 0-19-852231-2. xiv + 821 pp. LCCN QA279.G64 1993. Release 4 manual. Corrected reprint in 1994.

Fraser:1995:RCC

[FH95] Chris W. Fraser and David R. Hanson. A Retargetable C Compiler: Design and Implementation. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, 1995. ISBN 0-8053-1670-1. xv + 564 pp. LCCN QA76.73.C15F75 1995. US$55.95. URL https:/

/www.pearson.com/us/higher-
education/program/Hanson-
Retargetable-C-Compiler-A-
Design-and-Implementation/
PGM166351.html. From the au-

thors’ announcement: “… this book is an example of a ‘literate program.’ Like TEX: The Program by D. E. Knuth (Addison-Wesley, 1986), this book is lcc’s source code and the prose that describes it. The code is presented in the order that best suits understanding, not in the order dictated by the C programming language. The source code that appears on the diskette is extracted automatically from the book’s text files. … The distribution is available via ‘anonymous’ ftp from

ftp.cs.princeton.edu  (128.

112.152.13) in the directory pub/lcc. … Additional information about lcc and about our book is available on the

WWW at URL http://www.
cs.princeton.edu/software/
lcc.”.

Fox:1992:MPH

[FHCD92] Edward A. Fox, Lenwood S.

Heath, Qi Fan Chen, and Amjad M. Daoud. Minimal perfect hash functions for large databases. Communications of the Association for Computing Machinery, 35(1):105–121, January 1992. This is the first published algorithm for computing minimal perfect hash functions for lists of millions of words; previous algorithms were computationally infeasible for more than a few hundred words.

Finseth:1991:CTE

[Fin91] Craig A. Finseth. The Craft of Text Editing: Emacs for the Modern World. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97616-7 (New York), 3-540-97616-7 (Berlin). xii + 220 pp. LCCN QA76.76.T49 F56 1991. Contains extensive discussion of design issues for text editors, with examples from Emacs. Appendix B gives sources of numerous Emacs implementations. Appendix D summarizes the TECO command set.

Finch:2003:MC

[Fin03] Steven R. Finch. Mathematical Constants, volume 94 of Encyclopedia of mathematics and its applications. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-81805-2 (hardcover), 1-107-26335-2 (e-book), 1-107-26691-2 (e-book). xix + 602 pp. LCCN QA41 .F54 2003.

                                        URL http://algo.inria.fr/
                                        bsolve/constant/table.html;
                                         http://numbers.computation.
                                        free.fr/Constants/constants.
                                        html;   http://www.cambridge.
                                        org/us/catalogue/catalogue.
                                        asp?isbn=0521818052;   http:
                                        //www.loc.gov/catdir/description/
                                        cam031/2002074058.html;
                                        http://www.loc.gov/catdir/
                                        samples/cam034/2002074058.
                                        html;   http://www.loc.gov/
                                        catdir/toc/cam031/2002074058.
                                        html.

Finlayson:2011:HWW

[Fin09] Clive Finlayson. Humans who went extinct: why Neanderthals died out and we survived. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-923918-5 (hardcover), 0-19-923919-3 (paperback). xii + 273 pp. LCCN GN285 .F54 2009.

Firmage:1993:AAS

[Fir93] Richard A. Firmage. The Alphabet Abecedarium: Some Notes on Letters. David R. Godine, Publisher, Boston, MA, USA, 1993. ISBN 0-87923-987-5 (hardcover), 0-87923-998-0 (paperback). xi + 307 pp. LCCN P211 .F6 1993.

Fiume:1989:MSR

[Fiu89] Eugene L. Fiume. The Mathematical Structure of Raster Graphics. Academic Press, New York, NY, USA, 1989. ISBN 0-12-257960-7. xvi + 221 pp. LCCN T385 .F581 1989.

Fosdick:1996:IHP

[FJSD96] Lloyd D. Fosdick, Elizabeth R.

Jessup, Carolyn J. C. Schauble, and Gitta Domik. An Introduction to High-Performance Scientific Computing. MIT Press, Cambridge, MA, USA, 1996. ISBN 0-262-06181-3 (hardcover). xxiii + 760 pp. LCCN QA76 .A594 1996. US$70.00.

French:1985:NBC

[FK85] A. P. (Anthony Philip) French and P. J. Kennedy, editors. Niels Bohr: a centenary volume. Harvard University Press, Cambridge, MA, USA, 1985. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). xiv + 403 pp. LCCN QC16.B63 N49 1985. US$27.50.

Friedlander:1964:NR

[FKM64] Gerhart Friedlander, Joseph W.

Kennedy, and Julian Malcolm Miller. Nuclear and Radiochemistry. Wiley, New York, NY, USA, second edition, 1964. xi + 585 pp. LCCN QD601 .F71 1964.

Feynman:1988:WDY

[FL88] Richard Phillips Feynman and Ralph Leighton. What do YOU care what other people think?: Further adventures of a curious character. W. W. Norton & Co., New York, NY, USA, 1988. ISBN 0-393-02659-0. 255 pp. LCCN QC16.F49 A3 1988. US$17.95.

Feynman:2006:CFA

[FL06] Richard Phillips Feynman and Ralph Leighton. Classic Feynman: All the Adventures of a Curious Character. W. W. Norton & Co., New York, NY, USA, 2006. ISBN 0-393-06132-9. x + 511 pp. LCCN QC16.F49 A3 2006. URL ftp://

uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http:
//www.loc.gov/catdir/toc/
ecip0515/2005018928.html.

Flanagan:1991:PSR

[Fla91] David Flanagan. Programmer’s Supplement for Release 5 of the X Window System, Version 11. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-86-2. xx + 367 pp. LCCN QA76.76.W56 F5 1991.

Flanagan:1992:XTI

[Fla92] David Flanagan. X Toolkit Intrinsics Reference Manual, volume 5 of The Definitive guides to the X Window System. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 1992. ISBN 1-56592-007-4. xiii + 899 pp. LCCN QA76.76.W56 N94 1992. URL http:

//www.oreilly.com/catalog/
9781565920071;   http://www.
oreilly.com/catalog/v5.

Flanagan:1995:JND

[Fla95] David Flanagan. Java in a Nutshell: a Desktop Quick Reference for Java Programmers. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1995. ISBN 1-56592-183-6. xix + 438 pp. LCCN QA76.73.J38F553 1996. US$19.95. URL http://

www.ora.com/info/java.

Flanagan:1997:JDG

[Fla97] David Flanagan. JavaScript:

The Definitive Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1997. ISBN 1-56592-234-4. xiv + 647 pp. LCCN QA76.73.J39 F53 1997; QA76.73.V38 F53

1997.  URL http://www.
oreilly.com/catalog/9781565922341;
 http://www.oreilly.com/catalog/
jscript2.

Fletcher:1980:PMO

[Fle80] R. Fletcher. Practical Methods of Optimization. Volume 1: Unconstrained Optimization. Wiley, New York, NY, USA, 1980.

ISBN 0-471-27711-8. viii + 120

pp. LCCN QA402.5 .F43.

Fletcher:1981:PMO

[Fle81] R. Fletcher. Practical Methods of Optimization. Volume 2: Constrained Optimization. Wiley, New York, NY, USA, 1981. ISBN 0-471-27828-9. ix + 224 pp. LCCN QA402.5 .F43.

Fleisch:2008:SGM

[Fle08]  Daniel A. Fleisch.  A Stu-

dent’s Guide to Maxwell’s Equations. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-87761-X (hardcover), 0-521-70147-3 (paperback). ix + 134 pp. LCCN QC670 .F56 2008. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0803/2007037901-
                                        b.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0803/
                                        2007037901-d.html;   http:/
                                        /www.loc.gov/catdir/enhancements/
                                        fy0803/2007037901-t.html.

Fletcher:2018:ESB

[Fle18]  Seth Fletcher.  Einstein’s

Shadow: a Black Hole, a Band of Astronomers, and the Quest to See the Unseeable. Ecco, New York, NY, 2018. ISBN 0-06-231204-9 (paperback), 0-06-231202-2 (hardcover), 0-06-231203-0 (e-book). xxvii + 255 pp. LCCN QB843.B55 F595 2018.

Feynman:1985:SYJ

[FLH85] Richard P. Feynman, Ralph Leighton, and Edward Hutchings. “Surely You’re Joking, Mr.

Feynman!”: Adventures of a Curious Character. Bantam Books, New York, NY, USA, 1985. ISBN 0-553-25649-1. xi + 322 pp. LCCN QC16.F49A37 1986.

Feynman:1995:SEP

[FLS95] Richard Phillips Feynman, Robert B. Leighton, and Matthew L. (Matthew Linzee) Sands. Six easy pieces: essentials of physics, explained by its most brilliant teacher. Helix books. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-40955-0, 0-201-40956-9 (set), 0-201-48308-4 (cassettes), 0-201-40825-2 (paperback). xxix + 145 pp. LCCN QC21.2 .F52 1995b; QC21.2 .F52 1995. US$22.00.

Flynn:sf3

[Fly73] John A. Flynn. SFTRAN user guide. Technical Report Section 914, Internal Computing Memorandum 337, Jet Propulsion Laboratory, Pasadena, CA, USA, July 1973. Also catalogued as JPL Document No.1846-7.

Flynn:1995:WHH

[Fly95]  Peter Flynn.  The World-

WideWeb Handbook: An HTML Guide for Users, Authors and Publishers. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1995. ISBN 1-85032-205-8. xix + 351 pp. LCCN TK5105.888 .F56 1995. Foreword by Tim Berners-Lee. Three sections deal with (1) Getting connected to the Internet and using Internet software; (2) Writing HTML (2.0) files for the World-WideWeb; (3) Running a HTTP server and providing a Web service. Author is a member of the IETF Working Group on HTML. Text includes additional material on SGML; choice of editors, browsers and servers; copyright and intellectual property; and advance details of HTML3.

Flynn:1998:USX

[Fly98] Peter Flynn. Understanding SGML and XML Tools: Practical Programs for Handling Structured Text. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 0-7923-8169-6. xxvi + 432 pp. LCCN QA76.76.H94F59 1998.

Forbes:2014:FME

[FM14] Nancy Forbes and Basil Mahon. Faraday, Maxwell, and the electromagnetic field: how two men revolutionized physics. Prometheus Books, Amherst, NY, USA, 2014. ISBN 1-61614-942-6 (hardcover), 1-61614-943-4 (e-book). 320 pp. LCCN QC16.F2 F67 2014.

Fornberg:1996:PGP

[For96] Bengt Fornberg. A Practical Guide to Pseudospectral Methods. Cambridge Monographs on Applied and Computational Mathematics. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-49582-2. x + 231 pp. LCCN QA320

                                        F65 1996. URL http://www.
                                        loc.gov/catdir/description/

cam027/95031986.html;   http:
//www.loc.gov/catdir/toc/
cam024/95031986.html.

Forsyth:2012:ECS

[For12] Mark Forsyth. The Etymologicon: a Circular Stroll Through the Hidden Connections of the English Language. Berkley Books, New York, NY, USA, 2012. ISBN 0-425-26079-8 (paperback). xvii + 279 pp. LCCN PE1574 .F67 2012.

Fosdick:2005:RPR

[Fos05] Howard Fosdick. Rexx programmer’s reference. Wiley, New York, NY, USA, 2005. ISBN 0-7645-7996-7 (paperback), 0-7645-9695-0 (e-book). xxxiv + 686 pp. LCCN QA76.73.R24 F67 2005. URL ftp://

uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http:/
/media.wiley.com/product_
ancillary/67/07645799/DOWNLOAD/
579967_appm.pdf; http://media.
wiley.com/product_ancillary/
67/07645799/DOWNLOAD/Rexx_
Prgrs_Reference_scripts.zip;
http://www.loc.gov/catdir/
toc/ecip055/2004030560.html;
 http://www.wrox.com/WileyCDA/
WroxTitle/productCd-0764579967,
descCd-download_code.html.

Fowler:1997:MMA

[Fow97] A. C. Fowler. Mathematical Models in the Applied Sciences. Cambridge Texts in Applied Mathematics. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-46140-5 (hardcover), 0-521-46703-9 (paperback). xiv + 402 pp. LCCN QA401 .F685 1997. URL http://www.

                                        loc.gov/catdir/description/
                                        cam028/97010390.html;   http:
                                        //www.loc.gov/catdir/toc/
                                        cam024/97010390.html.

Foxley:music

[Fox87] Eric Foxley. Music: A language for typesetting music scores. Software—Practice and Experience, 17(8):485–502, August 1987.

Faux:1979:CGD

[FP79] I. D. Faux and M. J. Pratt. Computational Geometry for Design and Manufacture. Ellis Horwood Series in Mathematics and its Applications, Editor: G. M. Bell. Ellis Horwood, New York, NY, USA, 1979. ISBN 0-470-26473-X. 329 pp. LCCN QA447 .F33. UK£12.00.

Frankel:1964:TIO

[FPZ^{+}64] Max Frankel, Saul Patai, Albert Zhilkha, Zvi Rappoport, and Robert Farkas-Kadmon. Tables for Identification of Organic Compounds. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1964. x + 299 pp. LCCN QD291 .T3 1964.

Fernandez:2013:UMA

[FR13] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896– 1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). xviii + 522 pp. LCCN QC773 .F47 2013.

Francois:1964:PCO

[Fra64] E. François. Précis de Chimie Organique. Éditions de l’Enseignement, Éditions Universitaires, Bruxelles, Belgique, 1964. 94 pp. LCCN ????

French:1968:SR

[Fre68] A. P. French. Special Relativity.

The M.I.T. Introductory Physics Series. W. W. Norton & Co., New York, NY, USA, 1968. x + 286 pp. LCCN QC6 .F68.

French:1971:NM

[Fre71] A. P. French. Newtonian Mechanics. The M.I.T. Introductory Physics Series. W. W. Norton & Co., New York, NY, USA, 1971. ISBN 0-393-09958-X (hardcover), 0-393-09970-9 (paperback). xiii + 743 pp. LCCN QC125.2 .F74.

Frisch:1991:ESA

[Fri91] Æleen Frisch. Essential System Administration. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-74-9 (paperback), 0-937175-80-3 (paperback). xxiii + 440 pp. LCCN QA76.76.O63 F78 1992. URL http:/

/www.oreilly.com/catalog/
9780937175804.

FroeseFischer:1977:HFM

[Fro77] Charlotte Froese Fischer. The Hartree–Fock Method for Atoms:

A Numerical Approach. Wiley, New York, NY, USA, 1977. ISBN 0-471-25990-X. xi + 308 pp. LCCN QC173.F527.

FroeseFischer:2003:DRH

[Fro03] Charlotte Froese Fischer. Douglas Rayner Hartree: his life in science and computing. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2003. ISBN 981-238-577-0. xviii + 250 pp. LCCN Q143.H37. URL http:

//ebooks.worldscinet.com/
ISBN/9789812795014/9789812795014.
html.

Frommer:2004:FRA

[Fro04] Frommer. Frommer’s Road Atlas Europe. Frommer’s Road Atlas Series. Frommer’s, Windsor, Berkshire, UK, 2004. ISBN 0-7645-5782-3. 240 pp. LCCN G1797.21.P2 A87 2003.

Fry:2022:SMW

[Fry22] Helen Fry. Spymaster: the man who saved MI6. Yale University Press, New Haven, CT, USA, 2022. ISBN 0-300-26697-9. LCCN ????

Friedman:1965:CA

[FS65] Francis L. Friedman and Leo Sartori. The Classical Atom, volume 1 of Origins of Quantum Physics. Addison-Wesley, Reading, MA, USA, 1965. ix + 118 pp. LCCN QC 173 F914c.

Ferguson:2003:PC

[FS03] Niels Ferguson and Bruce Schneier. Practical Cryptography. Wiley, New York, NY, USA, 2003. ISBN 0-471-22894-X (hardcover), 0-471-22357-3 (paperback). xx + 410 pp. LCCN QA76.9.A25 F466 2003. URL http:/

/www.counterpane.com/book-
practical.html;   http://
www.loc.gov/catdir/bios/wiley044/
2003276249.html;   http://
www.loc.gov/catdir/description/
wiley036/2003276249.html;
http://www.loc.gov/catdir/
toc/wiley032/2003276249.html.

Ferguson:2010:CED

[FSK10] Niels Ferguson, Bruce Schneier, and Tadayoshi Kohno. Cryptography Engineering: Design Principles and Practical Applications. Wiley, New York, NY, USA, 2010. ISBN 0-470-47424-6 (paperback). xxix + 353 pp. LCCN QA76.9.A25 F466 2010.

Foley:1982:FIC

[Fv82] James D. Foley and Andries van Dam. Fundamentals of Interactive Computer Graphics. The Systems Programming Series. Addison-Wesley, Reading, MA, USA, 1982. ISBN 0-201-14468-9. xx + 664 pp. LCCN T385 .F63 1982.

Foley:1990:CGP

[FvFH90] James D. Foley, Andries van Dam, Steven K. Feiner, and John F. Hughes. Computer Graphics: Principles and Practice. The Systems Programming Series. Addison-Wesley, Reading, MA, USA, second edition, 1990.

ISBN 0-201-12110-7 (hardcover). xxiii + 1174 pp. LCCN T385 .C568 1990.

Feijen:1990:BOB

[FvGGM90] W. H. J. Feijen, A. J. M. van Gasteren, David Gries, and J. Misra, editors. Beauty is our Business: a Birthday Salute to Edsger W. Dijkstra. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. xix + 453 pp. LCCN QA76 .B326 1990. URL http:/

                                        /www.zentralblatt-math.org/
                                        zmath/en/search/?an=0718.

68004. Contains important treatment of accurate binaryto-decimal conversion [Gri90, Knu90a].

Feldman:2007:MSE

[FW07] Burton Feldman and Katherine Williams. 112 Mercer Street: Einstein, Russell, Gödel, Pauli, and the end of innocence in science. Arcade Publishing, New York, NY, USA, 2007. ISBN 1-55970-704-6. xx + 243 pp. LCCN Q141 .F345 2007. URL http://www.

                                        loc.gov/catdir/toc/ecip078/
                                        2007001194.html.

Fisher:1961:CAG

[FZ61] Robert C. Fisher and Allen D. Ziebur. Calculus and Analytic Geometry. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1961. 766 pp. LCCN QA 303 F53c.

Gomez:1999:ESC

[G^{+}99] Claude Gomez et al., editors.

Engineering and Scientific Computing with Scilab. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1999. ISBN 0-8176-4009-6, 3-7643-4009-6. xxv + 491 pp. LCCN TA345 .E5334 1999.

Gabriel:1985:PEL

[Gab85] Richard P. Gabriel. Performance and Evaluation of Lisp Systems. MIT Press Series in Computer Systems, Editor: Herb Schwetman. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-07093-6, 0-262-25619-3 (ebook). xiv + 285 pp. LCCN QA76.73.L23 G321 1985.

Gallmeister:1995:PPR

[Gal95] Bill O. Gallmeister. POSIX.4:

Programming for the Real World. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, January 1995. ISBN 1-56592-074-0. xviii + 548 pp. LCCN QA76.54 .G34 1995; QA76.76.O63 G34 1995. US$29.95. URL http:

//www.oreilly.com/catalog/
9781565920743;   http://www.
oreilly.com/catalog/posix4.

Galison:2003:ECP

[Gal03] Peter Galison. Einstein’s Clocks and Poincaré’s Maps: Empires of Time. W. W. Norton & Co., New York, NY, USA, 2003.

ISBN 0-393-02001-0 (hardcover), 0-393-32604-7 (paperback). 389 pp. LCCN QB209 .G35 2003. US$23.95.

Gal:2008:SPA

[Gal08] Sorin G. Gal. Shape-Preserving Approximation by Real and Complex Polynomials. Birkhäuser Boston Inc., Cambridge, MA, USA, 2008. ISBN 0-8176-4702-3, 0-8176-4703-1. xiii + 352 pp. LCCN QA221 .G34 2008.

                                        URL http://www.springer.
                                        com/birkhauser/mathematics/
                                        book/978-0-8176-4702-5.

Gamow:1985:TYS

[Gam85] George Gamow. Thirty Years That Shook Physics: the Story of Quantum Theory. Dover, New York, NY, USA, 1985. ISBN 0-486-24895-X (paperback). xiv + 224 + 9 pp. LCCN QC174.12 .G35 1985.

                                        US$4.95. URL http://www.
                                        loc.gov/catdir/description/
                                        dover032/85006797.html.

Gancarz:1995:UP

[Gan95]  Mike Gancarz.  The UNIX

philosophy. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1995. ISBN 1-55558-123-4. xix + 151 pp. LCCN QA76.76.O63G365 1995.

Garfinkel:1995:PPG

[Gar95] Simson Garfinkel. PGP: Pretty Good Privacy. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1995. ISBN 1-56592-098-8. xxxiii + 393 pp. LCCN QA76.9.A25G36 1995. US$24.95.

Gaskins:1992:PPMb

[Gas92a] Tom Gaskins. PEXlib Programming Manual. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 1-56592-028-7. xlv + 1105 pp. LCCN QA76.76.W56 G37 1992. US$44.95.

Gaskins:1992:PPMa

[Gas92b] Tom Gaskins. PHIGS Programming Manual. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 0-937175-85-4 (paperback), 0-937175-92-7 (hardcover). xxxix + 908 pp. LCCN QA76.76.W56 G37 1992. US$42.95 (softcover), US$52.95 (hardcover). URL http:

//www.oreilly.com/catalog/
9780937175859.

Gates:2009:ETH

[Gat09] Evalyn Gates. Einstein’s telescope: the hunt for dark matter and dark energy in the universe. W. W. Norton & Co., New York, NY, USA, 2009. ISBN 0-393-06238-4. xiv + 305 pp. LCCN QB791.3 .G38 2009.

FSF:gawk

[GAW87] The GAWK manual. Free Software Foundation, Inc. 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1987. Also available via ANONYMOUS

FTP to prep.ai.mit.edu. See

also [AKW88].

Gomez:2018:RPE

[GB18] Guillaume Gomez and Antoni Boucher. Rust Programming by Example: Enter the World of Rust by Building Engaging, Concurrent, Reactive, and Robust Applications. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78839-063-6, 1-78847-030-3 (e-book). viii + 437 pp. LCCN QA76.73.R88.

URL http://proquestcombo.
safaribooksonline.com/9781788390637.

Garbow:1977:MER

[GBDM77] B. S. Garbow, J. M. Boyle, J. J. Dongarra, and C. B. Moler. Matrix Eigensystem Routines: EIS-PACK Guide Extension, volume 51 of Lecture Notes in Computer Science, Editors: G. Goos and J. Hartmanis. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 0-387-08254-9 (New York), 3-540-08254-9 (Berlin). viii + 343 pp. LCCN QA193 .M381, QA267.A1,L43 no. 51.

Gibson:1998:DCM

[GBL^{+}98] Jerry D. Gibson, Toby Berger, Tom Lookabaugh, David Lindbergh, and Richard L. Baker. Digital Compression for Multimedia: Principles and Standards. Morgan Kaufmann series in multimedia information and systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1998. ISBN 1-55860-369-

7. xvii + 476 pp. LCCN QA76.575 .D535 1998. US$74.95, CAN$104.95, UK£49.95.

Gutmans:2005:PPP

[GBR05] Andi Gutmans, Stig Sæther Bakken, and Derick Rethans. PHP 5 power programming. Bruce Perens’ Open Source series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-147149-X (paperback). xxviii + 689 pp. LCCN QA76.73.P224 G88 2005.

Garwin:2002:MMF

[GC02] Richard L. Garwin and Georges Charpak. Megawatts and megatons: the future of nuclear power and nuclear weapons. University of Chicago Press, Chicago, IL, USA and London, UK, 2002. ISBN 0-226-28427-1 (paperback). xvii + 412 pp. LCCN QC792 .C4713 2002. URL http://

catdir.loc.gov/catdir/bios/
uchi051/2002027143.html;
http://catdir.loc.gov/catdir/
description/uchi051/2002027143.
html;   http://catdir.loc.
gov/catdir/toc/uchi051/2002027143.
html.

Gehani:1986:DFT

[Geh86] Narain Gehani. Document Formatting and Typesetting on the UNIX System. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, 1986. ISBN 0-9615336-0-9. xv + 364 pp. LCCN Z52.5.U54 G43 198. US$32.00.

Gehani:1987:DFT

[Geh87] Narain Gehani. Document Formatting and Typesetting on the UNIX System. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, second edition, 1987. ISBN 0-13-938325-5, 0-9615336-2-5. xv + 377 pp. LCCN Z52.5.U54G43 1987. US$40.90.

Gehani:2003:BLL

[Geh03] Narain Gehani. Bell Labs: life in the crown jewel. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, 2003. ISBN 0-929306-27-9. xi + 258 pp. LCCN TK415.B45 G44 2003. URL

                                        http://www.loc.gov/catdir/
                                        toc/fy043/2002012730.html.

Geiss:1975:RHC

[Gei75] Tony Geiss. Random House College Dictionary. Random House Reference, New York, NY, USA, 1975. ISBN 0-394-43500-1 (hardcover). xxxii + 1568 pp. LCCN PE1625 .R34 1975.

Genstat:1993:GRR

[Gen93] Genstat 5 Committee. Genstat 5 release 3 reference manual. Clarendon Press, Oxford, UK, 1993. ISBN 0-19-852312-2. xvi + 796 pp. LCCN QA276.4 .G443 1993. Reprinted with corrections in 1994 and 1997.

Gershenfeld:2000:PIT

[Ger00] Neil Gershenfeld. The Physics of Information Technology. Cambridge series on information and the natural sciences. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-58044-7. xiv + 370 pp. LCCN TK5103 .G45 2000.

Gerwarth:2020:NGR

[Ger20] Robert Gerwarth. November 1918: The German Revolution. Making of the Modern World. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2020. ISBN 0-19-260632-8 (electronic book), 0-19-260633-6 (electronic book), 0-19-954647-9 (hardcover). 329 pp. LCCN DD248 .G49 2020.

Goldstine:1946:ENI

[GG46] H. H. Goldstine and Adele Goldstine. The Electronic Numerical Integrator and Computer (ENIAC). Mathematical Tables and Other Aids to Computation, 2(15):97–110, July 1946. CO-DEN MTTCAS. ISSN 0891-6837. Reprinted in [Ran82, §7.7].

Gotlieb:1978:DTS

[GG78] C. C. Gotlieb and Leo R.

Gotlieb. Data Types and Structures. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1978. ISBN 0-13-197095-X. xii + 444 pp. LCCN QA76.9.D35 .G67. US$18.95.

Griswold:1983:IPL

[GG83] Ralph E. Griswold and Madge T.

Griswold. The Icon Programming Language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1983. ISBN 0-13-449777-5 (paperback). xviii + 313 pp. LCCN QA76.73.I19 G74 1983.

Griswold:1986:IIP

[GG86] Ralph E. Griswold and Madge T. Griswold. The Implementation of the Icon Programming Language. Princeton University Press, Princeton, NJ, USA, 1986. ISBN 0-691-08431-9. x + 336 pp. LCCN QA76.73.I19 G76 1986. US$39.50. URL http://www.

                                        cs.arizona.edu/icon/ibsale.
                                        htm.

Griswold:1990:IPL

[GG90] Ralph E. Griswold and Madge T. Griswold. The Icon Programming Language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1990. ISBN 0-13-447889-4. xv + 367 pp. LCCN QA76.73.I19 G74 1990. US$47.00.

Goldstone:2005:FCR

[GG05] Lawrence Goldstone and Nancy Bazelon Goldstone. The Friar and the Cipher: Roger Bacon and the Unsolved Mystery of the Most Unusual Manuscript in the World. Doubleday, New York, NY, USA, 2005. ISBN 0-7679-1473-2. xi + 320 pp. LCCN Z105.5.V65 G65 2005. URL http:

                                        //www.loc.gov/catdir/bios/
                                        random056/2004050164.html;
                                         http://www.loc.gov/catdir/
                                        description/random051/2004050164.
                                        html;   http://www.loc.gov/
                                        catdir/enhancements/fy0618/
                                        2004050164-s.html.

Godfrey:1982:MIO

[GHHH82] M. D. Godfrey, D. F. Hendry, H. J. Hermans, and R. K. Hessenberg. Machine-Independent Organic Software Tools (MINT). Academic Press, New York, NY, USA, second edition, 1982. ISBN 0-12-286982-6. xv + 369 pp. LCCN QA76.6.M319 1982.

Ghirardi:2005:SLG

[Ghi05] G. C. Ghirardi. Sneaking a Look at God’s Cards: Unraveling the Mysteries of Quantum Mechanics. Princeton University Press, Princeton, NJ, USA, revised edition, 2005. ISBN 0-691-04934-3, 0-691-12139-7, 0-691-13037-X. xix + 488 pp. LCCN Q173 .G4813 2004.

Gautier:2003:OOE

[GHLP03] Sophie Gautier, Christian Hardy, Frédéric Labbé, and Michel Pinquier. OpenOffice.org 1.1: efficace: Writer, Calc, Impress, Draw, BDs. Eyrolles, Paris, France, 2003. ISBN 2-212-11348-X (paperback). xvi + 319 pp. LCCN HF5548.4 O57O63 2004.

Gallivan:1990:PAM

[GHN^{+}90] K. A. Gallivan, Michael T.

Heath, Esmond Ng, James M. Ortega, Barry W. Peyton, R. J. Plemmons, Charles H. Romine, A. H. Sameh, and Robert G. Voigt. Parallel Algorithms for Matrix Computations. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1990. ISBN 0-89871-260-2 (paperback), 1-61197-170-5 (ebook). ix + 197 pp. LCCN QA188 .P367 1990.

GI:86

[GI886] Graphics Interface 86. Canadian Information Processing Society, 243 College St, 5th Floor, Toronto, Ontario M5T 2Y1, Canada, 1986. Vancouver, B. C. 26–30 May, 1986.

Gianone:1990:UMK

[Gia90] Christine M. Gianone. Using MS-DOS KERMIT: connecting your PC to the Electronic World. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1990. ISBN 1-55558-048-3. xxv + 244 pp. LCCN TK5105.9 .G5 1990.

Gianone:1992:UMK

[Gia92] Christine M. Gianone. Using MS-DOS KERMIT: Connecting Your PC to the Electronic World. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, second edition, 1992. ISBN 1-55558-082-3 (book + diskette), 1-55558-048-3. xxviii + 354 pp. LCCN TK5105.9 .G5 1992. US$34.95.

Gibaldi:1996:MHW

[Gib96] Joseph Gibaldi. MLA Handbook for Writers of Research Papers. Modern Language Association of America, New York, NY, USA, fourth edition, 1996. ISBN 0-87352-565-5. xviii + 293 pp. LCCN LB2369.G53 1995. US$16.75.

Gibaldi:1999:MHW

[Gib99] Joseph Gibaldi. MLA Handbook for Writers of Research Papers. Modern Language Association of America, New York, NY, USA, fifth edition, 1999. ISBN 0-87352-975-8. xviii + 332 pp. LCCN K94 .G53 1999; LB2369 .G53 1999.

Gibson:2019:SIH

[Gib19] Susannah Gibson. The Spirit of Inquiry: How One Extraordinary Society Shaped Modern Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2019. ISBN 0-19-883337-7 (hardcover). xxi + 377 pp. LCCN Q41.C194 G537 2019.

Giloi:1978:ICG

[Gil78] Wolfgang K. Giloi. Interactive Computer Graphics: Data Structures, Algorithms, Languages. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1978. ISBN 0-13-469189-X. xiii + 354 pp. LCCN T385 .G54 1978. US$18.50.

Gilster:1993:INE

[Gil93]  Paul Gilster.  The Inter-

net Navigator — The Essential Guide to Network Exploration for the Individual Dialup User. Wiley-Interscience, New York, NY, USA, 1993. ISBN 0-471-59782-1. xxiv + 470 pp. LCCN TK5105.875.I57 G55 1993. US$24.95.

Gillam:2003:UDP

[Gil03] Richard Gillam. Unicode demystified: a practical programmer’s guide to the encoding standard. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-70052-2 (paperback). xxxiii + 853 pp. LCCN QA76.6 .G5535 2002. UK£37.99.

Gosling:2000:JLS

[GJSB00] James Gosling, Bill Joy, Guy L. Steele, and Gilad Bracha. The Java Language Specification. Java series. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-31008-2. xxv + 505 pp. LCCN QA76.73.J38 G68 2000. URL

                                        http://java.sun.com/people/
                                        jag/.

Gregory:1978:CMT

[GK78] Robert T. Gregory and David L. Karney. A Collection of Matrices for Testing Computational Algorithms. Robert E. Krieger Publishing Company, Huntington, NY, USA, 1978. ISBN 0-88275-649-4. ix + 154 pp. LCCN QA188 .G72 1978.

Greene:1982:MAA

[GK82] Daniel H. Greene and Donald E. Knuth. Mathematics for the Analysis of Algorithms. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, second edition, 1982. ISBN 3-7643-3102-X. 123 pp. LCCN QA76.6 .G7423 1982.

Graham:1989:CM

[GKP89] Ronald L. Graham, Donald E. Knuth, and Oren Patashnik. Concrete Mathematics. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-14236-8. xiii + 625 pp. LCCN QA39.2 .G7331

Graham:1994:CM

[GKP94] Ronald L. Graham, Donald E.

Knuth, and Oren Patashnik. Concrete Mathematics. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-55802-5. xiii + 657 pp. LCCN QA39.2 .G733 1994.

Glasstone:1960:EPC

[GL60] Samuel Glasstone and David Lewis. Elements of Physical Chemistry. D. Van Nostrand, Princeton, NJ, USA, second edition, 1960. v + 758 pp. LCCN QD453 .G53 1960.

George:1981:CSL

[GL81] Alan George and Joseph W.

Liu. Computer Solution of Large Sparse Positive Definite Systems. Prentice-Hall Series in Computational Mathematics, Cleve Moler, Advisor. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1981. ISBN 0-13-165274-5. xii + 324 pp. LCCN QA188 .G46 1980.

Gehani:1988:DFT

[GL88] Narain Gehani and Steven Lally.

Document formatting and typesetting on the UNIX system. Vol. 2: grap, mv, ms, and troff. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, 1988. ISBN 0-9615336-3-3. xiii + 304 pp. LCCN Z52.5.U54 G431 1988. US$30.95.

Glasstone:1946:EPC

[Gla46] Samuel Glasstone. Elements of Physical Chemistry. Van Nostrand Reinhold, New York, NY, USA, 1946. vii + 695 pp. LCCN QD453.G53.

Glassner:1989:IRT

[Gla89] Andrew S. Glassner, editor. An Introduction to Ray Tracing. Academic Press, New York, NY, USA, 1989. ISBN 0-12-286160-4. xiii + 329 pp. LCCN T385 .I58 1989.

Glassner:1990:GG

[Gla90] Andrew S. Glassner, editor. Graphics Gems, volume 1 of Graphics Gems. Academic Press, New York, NY, USA, 1990. ISBN 0-12-286165-5 (paperback), 0-12-286166-3. xxix + 833 pp. LCCN T385 .G75 1990.

Glass:1997:BRS

[Gla97] Robert L. Glass. In the Beginning: Recollections of Software Pioneers. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7999-9. vi + 318 pp. LCCN QA76.76.D47I517 1998. US$29.00. IEEE catalog number BP07999.

Glass:1998:SRM

[Gla98]  Robert L. Glass.  Software

Runaways: Monumental Software Disasters. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-673443-X. xvi + 259 pp. LCCN QA76.76.F34G54 1998. US$29.99. URL http://

www.phptr.com/ptrbooks/ptr_
013673443X.html.

Glass:1999:CCL

[Gla99]  Robert L. Glass.  Comput-

ing Calamities: Lessons Learned from Products, Projects, and Companies That Failed. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-082862-9. xiv + 302 pp. LCCN HD9696.2.U62G55 199. US$29.99.

Gleick:1992:GLS

[Gle92] James Gleick. Genius: The Life and Science of Richard Feynman. Pantheon Books, New York, NY, USA, 1992. ISBN 0-679-40836-3. x + 531 pp. LCCN QC16.F49G54 1992. US$27.50.

Gleick:1999:FAJ

[Gle99] James Gleick. Faster: The Acceleration of Just About Everything. Pantheon Books, New York, NY, USA, 1999. ISBN 0-679-40837-1. x + 324 pp. LCCN QB209.G48 1999. US$24.00.

Gleick:2003:IN

[Gle03] James Gleick. Isaac Newton.

Vintage Books, New York, NY, USA, 2003. ISBN 1-4000-3295-4. xii + 272 pp. LCCN QC16.N7 .G55 2003. US$13.95; CDN$21.00.

Gleiser:2010:TEC

[Gle10] Marcelo Gleiser. A tear at the edge of creation: a radical new vision for life in an imperfect universe. Free Press, New York, NY, USA, 2010. ISBN 1-4391-0832-3. xvii + 285 pp. LCCN QB981 .G575 2010. US$25.00. URL

http://www.loc.gov/catdir/
enhancements/fy1009/2009046247-
s.html;   http://www.loc.
gov/catdir/enhancements/fy1011/
2009046247-b.html;   http:/
/www.loc.gov/catdir/enhancements/
fy1011/2009046247-d.html.

Glickstein:1997:WGE

[Gli97] Bob Glickstein. Writing GNU Emacs Extensions. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-261-

1. xviii + 215 pp. LCCN QA76.76.T49G56 1997.

Gell-Mann:1997:QJA

[GM97] Murray Gell-Mann. The quark and the jaguar: adventures in the simple and the complex. W. H. Freeman and Company, New York, NY, USA, 1997. ISBN 0-7167-2725-0 (paperback). xviii + 392 pp. LCCN QC774.G45 A3 1994. URL http://

www.gbv.de/dms/bowker/toc/
9780716725817.pdf;   http:
//www.zentralblattmath.org/
zmath/en/search/?an=0833.
00011.

Graybill:2002:PAC

[GM02] Robert Graybill and Rami Melhem, editors. Power Aware Computing. Kluwer Academic/ Plenum Publishers, New York, Boston, Dordrecht, London, Moscow, 2002. ISBN 0-306-46786-0. xxiv + 376 pp. LCCN TK7895.P68 P69 2002.

Goossens:2008:LGC

[GMR^{+}08] Michel Goossens, Frank Mittelbach, Sebastian Rahtz, Denis Roegel, and Herbert Voß, editors. The LaTEX Graphics Companion. Addison-Wesley series on tools and techniques for computer typesetting. Addison-Wes-ley, Reading, MA, USA, second edition, 2008. ISBN 0-321-50892-0. LCCN Z253.4.L38G663 2008.

Goossens:1994:LC

[GMS94] Michel Goossens, Frank Mittelbach, and Alexander Samarin. The LaTEX Companion. Tools and Techniques for Computer Typesetting. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-54199-8 (paperback), 0-201-77591-3 (hardcover). xxi + 528 + 2 pp. LCCN Z253.4.L38 G66 1994.

Gill:1981:PO

[GMW81] Philip E. Gill, Walter Murray, and Margaret H. Wright. Practical Optimization. Academic Press, New York, NY, USA, 1981. ISBN 0-12-283950-1 (hardcover), 0-12-283952-8 (paperback). xvi + 401 pp. LCCN QA402.5 .G5 1981.

Gnosis:1984:LL

[Gno84] Gnosis, Inc. Learning Lisp. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-527813-9. 200 pp. LCCN QA76.73.L23 L43 1984.

Ghizzetti:1970:QF

[GO70] Aldo Ghizzetti and Alessandro Ossicini. Quadrature formulae. Academic Press, New York, NY, USA, 1970. ISBN 0-12-281750-8. 192 pp. LCCN QA299.3 .G4813.

Gilly:1990:XWS

[GO90] Daniel Gilly and Tim O’Reilly. The X Window System in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1990. ISBN 0-937175-24-2. xi + 367 pp. LCCN QA76.76.W56 X84x.

Gilly:1992:XWS

[GO92] Daniel Gilly and Tim O’Reilly. The X Window System in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1992. ISBN 1-56592-017-1. xi + 409 pp. LCCN QA76.76.W56 X24 1992.

Goldstein:1998:APS

[Gol98] Norm Goldstein, editor. The Associated Press Stylebook and Libel Manual: Including Guidelines on Photo Captions, Filing the Wire, Proofreaders’ Marks, Copyright. Perseus Publishers, Cambridge, MA, USA, 1998. ISBN 0-201-33985-4 (paperback). 334 pp. LCCN PN4783 .A83 1998.

Goldstein:2005:IPP

[Gol05]  Rebecca Goldstein.  Incom-

pleteness: the proof and paradox of Kurt Gödel. Great discoveries. W. W. Norton & Co., New York, NY, USA, 2005. ISBN 0-393-05169-2. 296 pp. LCCN QA29.G58

G65 2005. URL http://www.
loc.gov/catdir/toc/ecip051/
2004023052.html.

Gonnet:1984:HAD

[Gon84] Gaston H. Gonnet. Handbook of Algorithms and Data Structures. International Computer Science Series. Editors: A. D. McGettrick and J. van Leeuwen. Addison-Wesley, Reading, MA, USA, 1984. ISBN 0-201-14218-X. xi + 286 pp. LCCN QA76.6 .G636 1984. US$12.95.

Goodstein:1991:MSH

[Goo91] Judith R. Goodstein. Millikan’s school: a history of the California Institute of Technology. W. W. Norton & Co., New York, NY, USA, 1991. ISBN 0-393-03017-2, 0-393-32998-4. 317 pp. LCCN T171.C219 G67 1991.

Goodchild:2004:ETRb

[Goo04] Peter Goodchild. Edward Teller, the real Dr. Strangelove. Harvard University Press, Cambridge, MA, USA, 2004. ISBN 0-674-01669-6. xxv + 469 + 16 pp. LCCN QC16.T37 G66 2004.

URL http://alsos.wlu.edu/
information.aspx?id=2237.

Gorin:1981:IDA

[Gor81] Ralph E. Gorin. Introduction to DECSYSTEM-20 Assembly Language Programming. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1981. ISBN 0-932376-12-6. xxx + 545 pp. LCCN QA76.8.D17 .G67. US$40.00.

Gordin:2004:WOT

[Gor04] Michael D. Gordin. A wellordered thing: Dmitrii Mendeleev and the shadow of the periodic table. Basic Books, New York, NY, USA, 2004. ISBN 0-465-02775-X. xx + 364 pp. LCCN QD22.M43 G67 2004. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0832/2003025533-
                                        b.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0832/
                                        2003025533-d.html;   http:
                                        //www.loc.gov/catdir/toc/
                                        ecip0411/2003025533.html.

Gortemaker:2011:EBL

[Gör11]  Heike B. Görtemaker.  Eva

Braun: life with Hitler. Alfred A. Knopf, New York, NY, USA, 2011. ISBN 0-307-59582-X. 324 pp. LCCN DD247.B66 G67 2011.

Gordin:2015:SBH

[Gor15]  Michael D. Gordin.  Scien-

tific Babel: how science was done before and after global English. University of Chicago Press, Chicago, IL, USA and London, UK, 2015. ISBN 0-226-00029-X (hardcover), 0-226-00032-X (e-book). 415 pp. LCCN Q223 .G67 2015.

Gordin:2020:EB

[Gor20] Michael D. Gordin. Einstein in Bohemia. Princeton University Press, Princeton, NJ, USA, 2020. ISBN 0-691-17737-6 (hardcover), 0-691-19984-1 (e-book). xi + 343 pp. LCCN QC16.E5 G6794 2020.

Gourlay:music-printing

[Gou86] John S. Gourlay. A language for music printing. Communications of the Association for Computing Machinery, 29(3):388–401, 1986.

Goudsmit:1996:A

[Gou96] Samuel Abraham Goudsmit. Alsos. History of modern physics and astronomy. American Institute of Physics, Woodbury, NY, USA, 1996. ISBN 1-56396-415-5. xxxviii + 259 pp. LCCN D810.S2 G6 1996.

Goldfarb:2004:CFG

[GP04] Charles F. Goldfarb and Paul Prescod. Charles F. Goldfarb’s XML Handbook. Charles F. Goldfarb definitive XML series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fifth edition, 2004. ISBN 0-13-049765-7. lv + 1222 pp. LCCN QA76.76.H94 G65 2004.

Gates:2019:PER

[GP19] S. James Gates, Jr. and Cathie Pelletier. Proving Einstein Right: the Daring Expeditions That Changed How We Look at the Universe. PublicAffairs, New York, NY, USA, 2019. ISBN 1-5417-6223-1 (e-book), 1-5417-6225-8 (hardcover). viii + 356 + 16 pp. LCCN QB544.19 .G38 2019. URL

https://historynewsnetwork.
org/article/173408.

Goldfarb:1990:SH

[GR90] Charles F. Goldfarb and Yuri Rubinsky. The SGML Handbook.

Clarendon Press, Oxford, UK, 1990. ISBN 0-19-853737-9 (hardcover). xxiv + 663 pp. LCCN Z286.E43 G64 1990. US$75.00.

Goossens:1999:LWC

[GR99] Michel Goossens and Sebastian Rahtz. The LaTEX Web Companion: Integrating TEX, HTML, and XML. Tools and Techniques for Computer Typesetting. Addison-Wesley Longman, Harlow, Essex CM20 2JE, England, 1999. ISBN 0-201-43311-7. xxii + 522 pp. LCCN QA76.76.H94G66 1999. US$36.95. With Eitan M. Gurari and Ross Moore and Robert S. Sutor.

GCA:1983:GSD

[Gra83a] Graphic Communications Association. GCA Standard 101-1983, Document Markup Metalanguage: GenCode and the Standard Generalized Markup Language (SGML). Graphic Communications Association, 100 Daingerfield Road, Alexandria, VA 22314-2804, USA, 1983. ISBN 0-89740-244-8. xxiv + 264 pp. LCCN QA76.9.T48 G33 1983. US$48.00. Adopted by Department of Defense, 10 August 1983.

Gray:1983:TT

[Gra83b] Bill Gray. Tips on Type. Van Nostrand Reinhold, New York, NY, USA, 1983. ISBN 0-442-22888-0. 128 pp. LCCN Z250.G78 1983. Little coverage of computer-based typesetting, but an excellent coverage of principles of typography, with lots of examples of fonts.

Gray:lex

[Gra88] Robert W. Gray. γ-GLA: A generator for lexical analyzers that programmers can use. In Summer USENIX ’88, pages 147– 160. USENIX, San Francisco, CA, USA, June 20–24 1988.

Gosling:1989:NBI

[GRA89] James Gosling, David S. H.

Rosenthal, and Michelle Arden. The NeWS Book: an introduction to the Network/extensible Window System. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 0-387-96915-2. vi + 235 pp. LCCN QA76.76.W56 A731 1989.

Grandine:1990:NMP

[Gra90] Thomas A. Grandine. The Numerical Methods Programming Projects Book. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-853385-3 (hardcover), 0-19-853387-X (paperback). viii + 146 pp. LCCN QA76.6 .G718 1990.

Gratzer:1996:MLI

[Grä96] George Grätzer. Math into LaTEX: an introduction to LaTEX2e and AMS-LaTEX 1.2. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1996. ISBN 0-8176-3805-9, 3-7643-3805-9. xxvii + 451 pp. LCCN Z253.4.A65G69 199. US$49.50.

Gratzer:1999:FSL

[Grä99] George Grätzer. First Steps in LaTEX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-8176-4132-7, 3-7643-4132-7. xx + 131 pp. LCCN Z253.4.L38 G74 1999. US$19.95.

Graham:2000:XLD

[Gra00a] Ian S. Graham. XHTML 1.0 Language and Design Sourcebook: the Next Generation of HTML. Wiley, New York, NY, USA, 2000. ISBN 0-471-37485-7 (paperback). xxv + 692 pp. LCCN QA76.76.H94 G7334

2000. URL http://www.loc.
gov/catdir/enhancements/fy0706/
99089119-b.html;   http://
www.loc.gov/catdir/toc/onix04/
99089119.html.

Gratzer:2000:ML

[Grä00b] George Grätzer. Math into LaTEX. Birkhäuser and Springer-Verlag, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland and Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2000. ISBN 0-8176-4131-9, 3-7643-4131-9. xxv + 584 pp. LCCN Z253.4.L38 G745 2000. US$49.95.

Gratzer:2007:ML

[Grä07] George Andrew Grätzer. Math into LaTEX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2007. ISBN 0-387-32289-2. xxxiv + 619 pp. LCCN Z253.4.L38 G745 2007. Forward by Rainer Schöpf.

Graetz:2011:EEU

[Gra11] Michael J. Graetz. The end of energy: the unmaking of America’s environment, security, and independence. MIT Press, Cambridge, MA, USA, 2011. ISBN 0-262-01567-6 (hardcover). x + 369 pp. LCCN HD9502.U52 G685 2011.

Gratzer:2016:MML

[Grä16] George Grätzer. More Math Into LaTEX. Springer International Publishing, Cham, Switzerland, fifth edition, 2016. ISBN 3-319-23795-0, 3-319-23796-9 (e-book). xxx + 609 pp. LCCN Z253.4.L38 G745 2016. URL http://www.

springerlink.com/content/
978-3-319-23796-1.

Green:1983:DIP

[Gre83] William B. Green. Digital Image Processing. Computer Science and Engineering Series. Van Nostrand Reinhold, New York, NY, USA, 1983. ISBN 0-442-28801-8. xi + 192 pp. LCCN TA1632 .G727 1983.

Greene:2004:FCS

[Gre04] Brian Greene. The Fabric of the Cosmos: Space, Time, and the Texture of Reality. Vintage Books, New York, NY, USA, 2004. ISBN 0-375-41288-3, 0-375-72720-5 (paperback). xii + 569 pp. LCCN QB982 .G74 2004.

Greenspan:2005:ECW

[Gre05] Nancy Thorndike Greenspan. The end of the certain world: the life and science of Max Born: the Nobel physicist who ignited the quantum revolution. Basic Books, New York, NY, USA, 2005. ISBN 0-7382-0693-8 (hardcover). x + 374 + 16 pp. LCCN QC16.B643 G74 2005.

Greene:2011:HRP

[Gre11] B. (Brian) Greene. The Hidden Reality: Parallel Universes and the Deep Laws of the Cosmos. Alfred A. Knopf, New York, NY, USA, 2011. ISBN 0-307-26563-3. xi + 370 pp. LCCN QC6 .G6885 2011.

Greenspan:2020:ASD

[Gre20] Nancy Thorndike Greenspan. Atomic Spy: The Dark Lives of Klaus Fuchs. Viking, New York, NY, USA, 2020. ISBN 0-593-08339-3 (hardcover), 0-593-08340-7 (paperback), 0-593-08341-5 (e-book), 2-01-905313-8. xii + 401 pp. LCCN UB271.R9 F8355 2020.

Gustafson:slalom

[GREC90] John Gustafson, Diane Rover, Stephen Elbert, and Michael Carter. SLALOM: The first scalable supercomputer benchmark. Supercomputing Review, 3(11): 56–61, November 1990.

Gustafson:slalom-3

[GREC91a] John Gustafson, Diane Rover, Stephen Elbert, and Michael Carter. SLALOM: Is your computer on the list? if not, we’d like it to be. Supercomputing Review, 4(7):52–59, July 1991.

Gustafson:slalom-4

[GREC91b] John Gustafson, Diane Rover, Stephen Elbert, and Michael Carter. SLALOM: Surviving adolescence. Supercomputing Review, 4(12):54–57, December 1991.

Gustafson:slalom-2

[GREC91c] John Gustafson, Diane Rover, Stephen Elbert, and Michael Carter. SLALOM update: The race continues. Supercomputing Review, 4(3):56–61, March 1991.

Gries:1978:PMC

[Gri78] David Gries, editor. Programming Methodology: a Collection of Articles by Members of IFIP WG2.3. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. ISBN 0-387-90329-1. xiv + 437 pp. LCCN QA76.6.P7516.

Grimson:1981:IS

[Gri81] William Eric Leifur Grimson.

From Images to Surfaces. MIT Press, Cambridge, MA, USA, 1981. ISBN 0-262-07083-9 (hardcover), 0-262-57185-4 (paperback), 0-262-25627-4 (e-book). 274 pp. LCCN QP487 .G74.

URL http://mitpress.mit.
edu/9780262070836.

Griffiths:space-fill

[Gri86] J. G. Griffiths. An algorithm for displaying a class of space-filling curves. Software—Practice and Experience, 16(5):403–411, May 1986.

Gries:1990:BDO

[Gri90] David Gries. Binary to decimal, one more time. In Feijen et al. [FvGGM90], chapter 16, pages 141–148. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL http:/

/www.zentralblatt-math.org/
zmath/en/search/?an=0718.

68004. This paper presents an alternate proof of Knuth’s algorithm [Knu90a] for conversion between decimal and fixed-point binary numbers.

Grier:2005:WCW

[Gri05] David Alan Grier. When Computers Were Human. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-09157-9 (hardcover), 0-691-13382-4 (paperback), 1-4008-4936-5 (e-book). viii + 411 pp. LCCN QA303.2 .G75 2005. URL

http://www.loc.gov/catdir/
enhancements/fy0654/2004022631-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0734/
2004022631-b.html;   http:
//www.loc.gov/catdir/toc/
ecip051/2004022631.html.

Goossens:1997:LGC

[GRM97] Michel Goossens, Sebastian Rahtz, and Frank Mittelbach. The LaTEX Graphics Companion: Illustrating Documents with TEX and PostScript. Tools and Techniques for Computer Typesetting. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-85469-4. xxi + 554 pp. LCCN Z253.4.L38G663 1997. US$39.75.

Gilchrist:1972:ORO

[GS72] T. L. (Thomas Lonsdale) Gilchrist and R. C. Storr. Organic reactions and orbital symmetry. Cambridge University Press, Cambridge, UK, 1972. ISBN 0-521-08249-8 (cloth), 0-521-09658-8 (paper). viii + 271 pp. LCCN QD476.G54.

Gordon:1973:MGE

[GS73] Mark A. Gordon and Lewis E.

Snyder. Molecules in the Galactic Environment. Wiley, New York, NY, USA, 1973. ISBN 0-471-31608-3. xv + 475 pp. LCCN QB790.M64.

Grant:1993:AAH

[GS93] Kenneth Grant and W. David Schwaderer. Adobe Acrobat Handbook: Digital Publishing in the Post-Gutenberg Era. Howard W. Sams, Indianapolis, IN 46268, USA, 1993. ISBN 0-672-30393-0 (includes diskette). xxvii + 254 pp. LCCN QA76.76.W56 G833 1993. US$19.95.

Garfinkel:1996:PUI

[GS96] Simson Garfinkel and Gene Spafford. Practical UNIX & Internet Security. Computer security. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1996. ISBN 1-56592-148-8. xxix + 971 pp. LCCN QA76.76.O63 G38 1996. US$39.95. URL

http://www.ora.com/gnn/bus/
ora/item/pus2.html;   http:
//www.oreilly.com/catalog/
puis.

Gavroglou:2012:NPN

[GS12] Kostas Gavroglou and Ana Simões. Neither physics nor chemistry: a history of quantum chemistry. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01618-4 (hardcover). xiv + 351 pp. LCCN QD462 .G38 2012.

Green:1968:NIP

[GSS68] A. E. S. Green, T. Sawada, and D. S. Saxon. The Nuclear Independent Particle Model: The Shell and Optical Models. Academic Press, New York, NY, USA, 1968. x + 363 pp. LCCN QC721 .G678 1968.

Gilly:1992:UN

[Gt92] Daniel Gilly and the staff of O’Reilly & Associates. UNIX in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1992. ISBN 1-56592-001-5. LCCN QA76.76.O63 G55 1992.

Gudmundsson:2016:EIG

[Gud16] Snæbjörn Gudmundsson. Exploring Iceland’s geology. Mál og menning, Reykjavík, Iceland, 2016. ISBN 9979-3-3625-0. 168 pp. LCCN QE279 .S63 2016. English translation by Katrina Downs-Rose and Örn Sigurdsson of Vegvísir um jardfrædi Íslands (2015).

Gurari:1994:TLD

[Gur94a] Eitan M. Gurari. TEX and LaTEX: Drawing and Literate Programming. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-025208-4 (book), 0-07-911616-7 (book + diskette). xiv + 310 pp. LCCN Z253.4.T47G87 1994. US$34.95.

Gurari:1994:WT

[Gur94b] Eitan M. Gurari. Writing with TEX. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-025207-6. xiv + 249 pp. LCCN Z253.4.T47 G87 1994.

Gusfield:1997:AST

[Gus97] Dan Gusfield. Algorithms on Strings, Trees, and Sequences: Computer Science and Computational Biology. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-58519-8 (hardcover). xviii + 534 pp. LCCN QA76.9.A43 G87 1997. US$64.95.

Gustavii:2012:HPS

[Gus12] Björn Gustavii. How to Prepare a Scientific Doctoral Dissertation Based on Research Articles. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-25510-4, 1-107-30709-0, 1-107-30929-8, 1-107-31484-4, 1-107-66904-9, 1-139-15125-8. ix + 93 pp. LCCN LB2369 .G87 2012.

Gustavii:2017:HWI

[Gus17] Björn Gustavii. How to Write and Illustrate a Scientific Paper. Cambridge University Press, Cambridge, UK, 2017. ISBN 1-316-60791-7. xiii + 225 pp. LCCN T11 .G86 2017.

Gutmann:2004:CSA

[Gut04] Peter Gutmann. Cryptographic Security Architecture: Design and Verification. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 0-387-95387-6. xviii + 320 pp. LCCN QA76.9.A25 G88 2002.

Guthery:2011:MMH

[Gut11] Scott B. Guthery. A motif of mathematics: history and application of the mediant and the Farey sequence. Docent Press, Boston, MA, USA, 2011. ISBN 1-4538-1057-9 (paperback). xx + 243 pp. LCCN QA241 .G864 2011.

Guthery:2012:BRC

[Gut12] Scott B. Guthery. Bibliography of Raymond Clare Archibald. Docent Press, Boston, MA, USA, 2012. ISBN 0-9837004-2-7. xv + 52 pp. LCCN QA36 .G88 2012.

Golub:1983:MC

[GV83] Gene H. Golub and Charles F. Van Loan. Matrix Computations. Johns Hopkins Series in the Mathematical Sciences. The Johns Hopkins University Press and North Oxford Academic, Baltimore, MD, USA and Oxford, England, 1983. ISBN 0-8018-3010-9 (hardcover), 0-8018-3011-7 (paperback), 0-946536-00-7, 0-946536-05-8 (paperback). xvi + 476 pp. LCCN QA188 .G65 1983.

URL http://www.jstor.org/
stable/2008107;   http://
www.jstor.org/stable/2030489;
 http://www.jstor.org/stable/
3616959.

Golub:1989:MC

[GV89] Gene H. Golub and Charles F.

Van Loan. Matrix Computations, volume 3 of Johns Hopkins Series in the Mathematical Sciences. The Johns Hopkins University Press, Baltimore, MD, USA, second edition, 1989. ISBN 0-8018-3772-3 (hardcover), 0-8018-3739-1 (paperback). xix + 642 pp. LCCN QA188 .G65 1989. US$14.50.

Gonzalez:1977:DIP

[GW77] Rafael C. Gonzalez and Paul Wintz. Digital Image Processing. Addison-Wesley, Reading, MA, USA, 1977. ISBN 0-201-02596-5 (hardcover), 0-201-02597-3 (paperback). xvi + 431 pp. LCCN TA1632 .G66 1977.

Gustafson:1998:ACM

[GW98] Grant B. Gustafson and Calvin H.

Wilcox. Analytical and Computational Methods of Advanced Engineering Mathematics, volume 28 of Texts in applied mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98265-5. xxii + 729 pp. LCCN TA330 .G87 1998. URL http://www.

                                        springeronline.com/sgw/cda/
                                        frontpage/0,10735,4-40109-
                                        22-1522238-0,00.html.

Gwennup:1996:DMA

[Gwe96]  Linley Gwennup.  Digital,

MIPS add multimedia extensions: Digital focuses on video, MIPS on 3-D graphics; vendors debate differences. Microprocessor report, ??(??): 24–28, November 18, 1996. ISSN 0899-9341. URL http:

                                        //www.chipanalyst.com/q/@
                                        12083418zrryvm/mpr/abstracts/
                                        abs10_15.html;   http://www.
                                        chipanalyst.com/q/@12083418zrryvm/
                                        mpr/mpr/ARTICLES/101505.pdf.

Hahn:1967:OHS

[Hah67] Otto Hahn. Otto Hahn: a scientific autobiography. MacGibbon and Kee, London, UK, 1967. xxiv + 296 + 16 pp. LCCN QD22.H2 1967. Translated and edited by Willy Ley. With an introduction by Glenn T. Seaborg.

Hahn:1991:LE

[Hah91] Jane Hahn. LaTEX for Everyone. Personal TEX Inc., 12 Madrona Street, Mill Valley, CA 94941, USA, 1991. ISBN 0-9631044-0-3. xi + 346 pp. LCCN Z253.4.L38 H34 1991. US$19.95.

Hall:pfort

[Hal69] A. D. Hall. A portable Fortran IV subset. Technical report, AT&T Bell Laboratories, Murray Hill, NJ 07974, USA, 1969.

Hanna:1963:QMC

[Han63] Melvin W. Hanna. Quantum Mechanics in Chemistry. W. A. Benjamin, Inc., New York, NY, USA, 1963. xiii + 253 pp. LCCN QD462 .H3.

Hanson:parsing

[Han85] David R. Hanson. Compact recursive-descent parsing of expressions. Software—Practice and Experience, 15(12):1205– 1212, December 1985.

Hardy:1940:MA

[Har40] G. H. (Godfrey Harold) Hardy. A Mathematician’s Apology. Cambridge University Press, Cambridge, UK, 1940. ISBN 0-521-42706-1. 153 pp. LCCN QA7.H3 1992. With a long foreword by C. P. Snow. Reprinted 1994.

Harrington:1983:CGP

[Har83] Steven Harrington. Computer Graphics: a Programming Approach. McGraw-Hill, New York, NY, USA, 1983. ISBN 0-07-026751-0. xv + 448 pp. LCCN T385 .H34 1983. US$21.95.

Harrington:1987:CGP

[Har87] Steven Harrington. Computer Graphics: A Programming Approach. McGraw-Hill, New York, NY, USA, second edition, 1987. ISBN 0-07-026753-7. xiv + 466 pp. LCCN T385 .H34 1987. US$38.95.

Harris:1988:RT

[Har88] Errol E. Harris. The reality of time. SUNY series in philosophy. State University of New York Press, Albany, NY, USA, 1988. ISBN 0-88706-860-X, 0-88706-861-8 (paperback). xii + 204 pp. LCCN BD638 .H277 1988.

Harris:1992:DCM

[Har92] Kent Harris. Discovering Calculus with Maple. Wiley, New York, NY, USA, 1992. ISBN 0-471-55156-2. 192 pp. LCCN QA303.5.D37 H37 1992.

Harold:1999:J

[Har99] Elliotte Rusty Harold. Java I/O. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-485-1. xxvi + 568 pp. LCCN QA76.73.J38 H372 1999. US$32.95. URL http:

//metalab.unc.edu/javafaq/
books/javaio/;   http://www.
oreilly.com/catalog/javaio/

.

Harrison:2000:AQM

[Har00] Walter A. (Walter Ashley) Harrison. Applied Quantum Mechanics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2000. ISBN 981-02-4375-8 (hardcover), 981-02-4394-4 (paperback). xvi + 353 pp. LCCN QC174.12 .H377 2000. URL

http://www.worldscientific.
com/worldscibooks/10.1142/
4485.

Haralambous:2004:FCG

[Har04] Yannis Haralambous. Fontes & codages: Glyphes et caractères à l’ère du numérique. O’Reilly, Paris, France, 2004. ISBN 2-84177-273-X. xx + 990 pp. LCCN Z699.35C48 H37.2004.

Hargittai:2010:JET

[Har10] István Hargittai. Judging Edward Teller: a closer look at one of the most influential scientists of the Twentieth Century. Prometheus Books, Amherst, NY, USA, 2010. ISBN 1-61614-221-9 (hardcover). 575 pp. LCCN QC16.T37 H37 2010. Afterword by Richard Garwin.

Harari:2015:SBH

[Har15] Yuval N. Harari. Sapiens: a brief history of humankind. Harper, New York, NY, USA, 2015. ISBN 0-06-231609-5 (hardcover), 0-06-231611-7 (e-book). ix + 443 pp. LCCN CB113.H4 H3713 2015. Translation from the Hebrew original by Yuval N. Harari and John Purcell and Haim Watzman.

Hastings:1955:ADC

[Has55] Cecil Hastings, Jr. Approximations for Digital Computers. The Rand series. Princeton University Press, Princeton, NJ, USA, 1955. ISBN 0-691-07914-5. viii + 201 pp. LCCN QA76 .H37. Assisted by Jeanne T. Hayward and James P. Wong, Jr.

Hasegawa:2005:RES

[Has05] Tsuyoshi Hasegawa. Racing the enemy: Stalin, Truman, and the surrender of Japan. Belknap Press of Harvard University Press, Cambridge, MA and London, UK, 2005. ISBN 0-674-01693-9 (hardcover), 0-674-03840-1 (e-book). ix + 382 pp. LCCN D813.J3 H37 2005.

Hastings:2016:SWS

[Has16] Max Hastings. The Secret War: Spies, Ciphers, and Guerrillas 1939–1945. Harper, New York, NY, 2016. ISBN 0-06-225927-X (hardcover), 0-06-225928-8 (paperback), 0-06-244156-6. xxvii + 610 + 32 pp. LCCN D810.S7 H365 2017.

Hauser:2000:WMW

[Hau00] Marc D. Hauser. Wild Minds: What Animals Really Think. Henry Holt and Company, New York, NY, USA, 2000. ISBN 0-8050-5669-6 (hardcover), 0-8050-5670-X. xx + 315 pp. LCCN QL785 .H38 2000. Illustrations by Ted Dewan.

Havil:2003:GEE

[Hav03] Julian Havil. Gamma: Exploring Euler’s Constant. Princeton University Press, Princeton, NJ, USA, 2003. ISBN 0-691-09983-9. xxiii + 266 pp. LCCN QA41 .H23 2003. US$29.95. URL http://www.

                                        loc.gov/catdir/description/
                                        prin031/2002192453.html.

Hawking:2005:GCI

[Haw05] Stephen Hawking. God created the integers: the mathematical breakthroughs that changed history. Running Press Book Publishers, Philadelphia, PA; London, UK, 2005. ISBN 0-7624-1922-9 (hardcover). xiii + 1160 pp. LCCN QA21 .G63 2005. US$29.95. URL http:/

/www.perseusbooksgroup.com/
runningpress/book_detail.
jsp?isbn=0762419229.

Hall:1990:SAS

[HB90a] Mark Hall and John Barry.

Sunburst: The Ascent of Sun Microsystems. Contemporary Books, Inc., 180 North Michigan Avenue, Chicago, IL 60601, USA, 1990. ISBN 0-8092-3989-2, 0-8092-4368-7. xvii + 297 pp. LCCN HD9696.C64 S794 1990.

Hall:1990:SP

[HB90b] Mark Hall and John Barry. The SunTechnology Papers. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-97145-9, 3-540-97145-9. x + 250 pp. LCCN HD9696.C64 S796 1990; QA76.754 .S86 1990.

Holoien:1991:FES

[HB91] Martin O. Holoien and Ali Behforooz. FORTRAN 77 for Engineers and Scientists. Brooks/ Cole, Pacific Grove, CA, USA, second edition, 1991. ISBN 0-534-14166-8. x + 463 pp. LCCN Q183.9 .H64 1991. US$28.00.

Haigh:2021:NHMe

[HC21] Thomas Haigh and Paul E.

Ceruzzi. A New History of Modern Computing. History of computing. MIT Press, Cambridge, MA, USA, 2021. ISBN 0-262-36647-9 (e-book), 0-262-54290-0 (paperback). x + 528 pp. LCCN QA76.17 .H34 2021.

Hart:1968:CAa

[HCL^{+}68] John F. Hart, E. W. Cheney, Charles L. Lawson, Hans J. Maehly, Charles K. Mesztenyi, John R. Rice, Henry G. Thatcher, Jr., and Christoph Witzgall. Computer Approximations. Robert E. Krieger Publishing Company, Huntington, NY, USA, 1968. ISBN 0-88275-642-7. x + 343 pp. LCCN QA 297 C64 1978. Reprinted 1978 with corrections.

Hopgood:1986:MWM

[HDF^{+}86] F. Robert A. Hopgood, David A. Duce, Elizabeth V. C. Fielding, Ken Robinson, and Antony S. Williams. Methodology of Window Management. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-16116-3. xv + 250 pp. LCCN T385 .C671 1985.

Holmes:1993:ECM

[HEBS93] Mark H. Holmes, Joseph G. Ecker, William E. Boyce, and William L. Siegmann. Exploring Calculus with Maple. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-52616-6. vii + 258 pp. LCCN QA303.5.D37 E97 1993.

Heck:1993:IM

[Hec93] André Heck. Introduction to Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-97662-0 (New York), 3-540-97662-0 (Berlin). xiii + 497 pp. LCCN QA155.7.E4H43 1993.

Heckbert:1994:GGI

[Hec94] Paul S. Heckbert, editor. Graphics gems IV, volume 4 of Graphics Gems. AP Professional, Boston, MA, USA, 1994. ISBN 0-12-336156-7 (with Macintosh disk), 0-12-336155-9 (with IBM disk). xvii + 575 pp. LCCN T385 .G6974 1994.

Heck:1996:IM

[Hec96] André Heck. Introduction to Maple. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1996. ISBN 0-387-94535-0 (hardcover), 1-4684-0484-9, 1-4684-0486-5. xx + 699 pp. LCCN QA155.7.E4H43 1993. US$39.00.

Heising:ftn

[Hei66] W. P. Heising. History and summary of Fortran standardization development for the ASA. Communications of the Association for Computing Machinery, 7:590–625, 1966. See also final standard [Ame66].

Heinrich:1993:MRM

[Hei93] Joe Heinrich. MIPS R4000 Microprocessor User’s Manual. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-105925-4. xxvi + 438 + A182 + B62 + C6 + D4 + E4 pp. LCCN QA76.8.M523H45 1993. US$34.00.

Heller:1990:XPM

[Hel90] Dan Heller. XView Programming Manual, volume 7. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1990. ISBN 0-937175-38-2. xxviii + 557 pp. LCCN QA76.76.W56 D44 v.7 1990.

Heller:1991:MPM

[Hel91a] Dan Heller. Motif Programming Manual for OSF/Motif Version 1.1, volume 6. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-70-6. xl + 990 pp. LCCN QA76.76.W56 H446 1991.

Heller:1991:XPM

[Hel91b] Dan Heller. XView Programming Manual, volume 7A of The Definitive guides to the X Window System. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, September 1991. ISBN 0-937175-87-0. xxxvii + 729 pp. LCCN QA76.76.W56 H447 1990.

Helferich:2004:HCA

[Hel04] Gerard Helferich. Humboldt’s Cosmos: Alexander von Humboldt and the Latin American Journey That Changed the Way We See the World. Gotham Books, New York, NY, USA, 2004. ISBN 1-59240-106-6 (paperback), 1-59240-052-3 (paperback). xxii + 360 pp. LCCN Q143.H9 A6 2005.

Hentschel:1992:ETE

[Hen92] Klaus Hentschel. Der Einstein-Turm: Erwin F. Freundlich und die Relativitätstheorie; Ansätze zu einer “dichten Beschreibung” von institutionellen, biographischen und theoriengeschichtlichen Aspekten. (German) [The Einstein Tower: Erwin F. friendly and the Theory of Relativity; Dense approaches to a description of institutional, biographical and theoretical-historical aspects]. Spektrum Akademischer Verlag, Heidelberg, Germany, 1992. ISBN 3-86025-025-6. 192 pp. LCCN QB36.F73 H46 1992. See also English translation [Hen97].

Hentschel:1997:ETI

[Hen97]  Klaus Hentschel.  The Ein-

stein Tower: an intertexture of dynamic construction, Relativity Theory, and astronomy. Writing science. Stanford University Press, Stanford, CA, USA, 1997. ISBN 0-8047-2824-0. xiv + 226 pp. LCCN QB462.65. URL http:/

/www.gbv.de/dms/goettingen/

220404755.pdf. English translation by Ann M. Hentschel of the German original [Hen92].

Herzberg:1945:MSM

[Her45] Gerhard Herzberg. Molecular Spectra and Molecular Structure: II. Infrared and Raman Spectra of Polyatomic Molecules. Van Nostrand Reinhold, New York, NY, USA, 1945. ISBN 0-442-03386-9. xiii + 632 pp. LCCN QC451 .H464.

Herzberg:1950:MSM

[Her50] Gerhard Herzberg. Molecular Spectra and Molecular Structure: I. Spectra of Diatomic Molecules. Van Nostrand Reinhold, New York, NY, USA, second edition, 1950. ISBN 0-442-03385-0. xiii + 658 pp. LCCN QC451 .H455 1950.

Herzberg:1966:MSM

[Her66] Gerhard Herzberg. Molecular Spectra and Molecular Structure: III. Electronic Spectra and Electronic Structure of Polyatomic Molecules. Van Nostrand Reinhold, New York, NY, USA, 1966. ISBN 0-442-03387-7. xviii + 745 pp. LCCN QC451 .H455 v.3.

Hershey:calligraphy

[Her67] Allen V. Hershey. Calligraphy for computers. Technical Report TR-2101, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, August 1967.

Hershey:fortran-cartography

[Her69] Allen V. Hershey. FORTRAN IV programming for cartography and typography. Technical Report TR-2339, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1969.

Hershey:fortran-typography

[Her70] Allen V. Hershey. Preparation of reports with the FORTRAN typographic system. Technical Report TN-K/27-70, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1970.

Hershey:computer-typography

[Her72a] Allen V. Hershey. A computer system for scientific typography. Computer Graphics and Image Processing, 1:373–385, 1972.

Hershey:wavetrains

[Her72b] Allen V. Hershey. FORTRAN programming for surface wave trains. Technical Report TR-2714, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1972.

Hershey:wavetrains-2

[Her72c] Allen V. Hershey. Interpolation of surface wave trains. Technical Report TR-3064, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, September 1972.

Hershey:fourier

[Her74a] Allen V. Hershey. A comment on the Fourier theorem. Technical Report TR-3145, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, June 1974.

Hershey:lagally

[Her74b] Allen V. Hershey. A comment on the lagally theorem. Technical Report TR-3133, U. S. Naval Weapons Laboratory, Dahlgren, VA 22448, USA, May 1974.

Hershey:advanced-typography

[Her81] Allen V. Hershey. Advanced computer typography. Technical Report NPS012-81-005, U. S. Naval Postgraduate School, Monterey, CA 93940, USA, December 1981.

Herken:2002:BBT

[Her02] Gregg Herken. Brotherhood of the bomb: the tangled lives and loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller. Henry Holt and Company, New York, NY, USA, 2002. ISBN 0-8050-6588-1 (hardcover). xiv + 448 pp. LCCN QC16.O62 H47 2002. URL ftp://

uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http:/
/alsos.wlu.edu/information.
aspx?id=1264;   http://www.
loc.gov/catdir/bios/hol051/
2002017219.html;   http://
www.loc.gov/catdir/description/
hol021/2002017219.html.

Hey:1999:FCE

[Hey99] Anthony J. G. Hey, editor. Feynman and Computation: Exploring the Limits of Computers. Perseus Publishers, Cambridge, MA, USA, 1999. ISBN 0-7382-0057-3. xxiii + 438 pp. LCCN QA76 .F46 1999, QC52 .F49 199. US$50.00.

Herz-Fischler:1987:MHD

[HF87] Roger Herz-Fischler. A Mathematical History of Division in Extreme and Mean Ratio. Wilfrid Laurier University Press, Waterloo, ON, Canada, 1987. ISBN 0-88920-152-8. xvi + 191 pp. LCCN A481.H47 1987. US$65.00.

Herz-Fischler:1998:MHG

[HF98] Roger Herz-Fischler. A Mathematical History of Golden Number. Dover, New York, NY, USA, 1998. ISBN 0-486-40007-7. xxii + 195 pp. LCCN QA481.H47 1998.

Hurley:1978:PP

[HG78] James P. Hurley and Claude Garrod. Principles of Physics. Houghton-Mifflin, Boston, MA, USA, 1978. ISBN 0-395-25036-6. xvii + 773 pp. LCCN QC23.H89.

Hopgood:1983:IGK

[HGDS83] F. Robert A. Hopgood, Julian R.

Gallop, David A. Duce, and Dale C. Sutcliffe. Introduction to the Graphical Kernel System (GKS). Academic Press, New York, NY, USA, 1983. ISBN 0-12-355570-1. xi + 200 pp. LCCN T385 .I57 1983. A. P. I. C. Studies in Data Processing No. 19.

Hopgood:1986:IGK

[HGDS86] F. Robert A. Hopgood, Julian R.

Gallop, David A. Duce, and Dale C. Sutcliffe. Introduction to the Graphical Kernel System (GKS). Academic Press, New York, NY, USA, second edition, 1986. ISBN 0-12-355571-X (paperback). xii + 250 pp. LCCN T385 .I571 1986. UK£12.50. Revised for the International Standard. A. P. I. C. Studies in Data Processing No. 28.

Hult:1999:NCH

[HH99] Christine Hult and Thomas N.

Huckin. New Century Handbook.

Allyn and Bacon, Boston, MA, USA, 1999. ISBN 0-205-27352-1 (hardcover), 0-205-32970-5. xxxvii + 870 pp. LCCN PE1408 .H688 1999. URL http://

occawlonline.pearsoned.com/
bookbind/pubbooks/hult2_awl/
;   http://www.ablongman.
com/hult.

Harris:2007:DDC

[HH07] David Money Harris and Sarah L. Harris. Digital Design and Computer Architecture. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2007. ISBN 0-12-370497-9. xxii + 569 pp. LCCN TK7868.D5 H34

2007. URL http://www.loc.
gov/catdir/enhancements/fy0703/
2006030554-d.html;   http:
//www.loc.gov/catdir/toc/
ecip071/2006030554.html.

Howes:2015:AWU

[HH15] Ruth H. Howes and Caroline L. Herzenberg. After the War: US Women in Physics. IOP Concise Physics Series. Morgan and Claypool Publishers, San Rafael, CA 94903, USA, 2015. ISBN 1-68174-030-3 (print), 1-68174-094-X (PDF), 1-68174-222-5 (mobi). 123 pp. LCCN QC15 .H683 2015. URL https:

//iopscience.iop.org/book/
mono/978-1-6817-4094-2.pdf.

Heyne:2007:LEM

[HHPM07] Andreas K. Heyne, Alice K. Heyne, Elena S. Pini, and Tahu Matheson. Leonhard Euler: a man to be reckoned with.

Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2007. ISBN 3-7643-8332-1. 45 pp. LCCN QA29.E8 H49 2007.

Heal:1995:MVM

[HHR95] K. M. Heal, M. L. Hansen, and K. M. Rickard. Maple V Mathematics Learning Guide. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94536-9. ix + 269 pp. LCCN QA155.7.E4 H42 1996. US$24.00. With the editorial assistance of J. S. Devitt; based in part on the work of B. W. Char.

Hansmann:eurographics89

[HHS89] W. Hansmann, F. R. A. Hopgood, and W. Strasser, editors. Eurographics ’89. North-Hol-land, Amsterdam, The Netherlands, 1989. ISBN 0-444-88013-5. LCCN T 385 E97 1989.

Higham:1993:HWM

[Hig93] Nicholas J. Higham. Handbook of Writing for the Mathematical Sciences. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, June 1993. ISBN 0-89871-314-5. xii + 241 pp. LCCN QA42.H54 1993. US$21.50.

Higham:1996:ASN

[Hig96] Nicholas J. Higham. Accuracy and Stability of Numerical Algorithms. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1996. ISBN 0-89871-355-2 (paperback). xxviii + 688 pp. LCCN QA297.H53 1996. US$39.00. URL http:

                                        //www.ma.man.ac.uk/~higham/
                                        asna.html.

Higham:2002:ASN

[Hig02] Nicholas J. Higham. Accuracy and Stability of Numerical Algorithms. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 2002. ISBN 0-89871-521-0 (hardcover), 0-89871-802-3 (e-book). xxx + 680 pp. LCCN QA297 .H53 2002.

Hillis:1985:CM

[Hil85] W. Daniel Hillis. The Connection Machine. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-08157-1, 0-262-58097-7 (paperback). xiii + 190 pp. LCCN QA267 .H4871 1985.

Hiltzik:2015:BSE

[Hil15] Michael A. Hiltzik. Big Science: Ernest Lawrence, the Cyclotron, and the Birth of the Military–Industrial Complex. Simon and Schuster, New York, NY, USA, 2015. ISBN 1-4516-7575-5 (hardcover), 1-4516-7576-3 (paperback), 1-4516-7603-4 (ebook). x + 512 pp. LCCN QC787.C8 H55 2015.

Hilburn:1976:MCC

[HJ76] John L. Hilburn and Paul M. Julich. Microcomputers/Microprocessors: Hardware, Software, and Applications. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

ISBN 0-13-580969-X. xii + 372 pp. LCCN TK7888.3 .H48. US$16.50.

Hockney:1981:PCA

[HJ83] R. W. Hockney and C. R.

Jesshope. Parallel Computers: Architecture, Programming, and Algorithms. Adam Hilger Ltd., Bristol, UK, 1981, 1983. ISBN 0-85274-422-6 (hardcover), 0-85274-752-7 (paperback). xiii + 423 pp. LCCN QA76.6 .H58 1981. UK£22.50.

Hockney:1988:PCA

[HJ88] Roger W. Hockney and Chris R.

Jesshope. Parallel Computers: Architecture, Programming, and Algorithms. Adam Hilger Ltd., Bristol, UK, second edition, 1988. ISBN 0-85274-811-6 (hardcover), 0-85274-812-4 (paperback). xv + 625 pp. LCCN QA76.5 .H57 1988.

Hassan:2004:HPT

[HJ04] Mahbub Hassan and Raj Jain, editors. High Performance TCP/ IP Networking: Concepts, Issues, and Solutions. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-064634-2, 0-13-127257-8 (paperback). xx + 383 pp. LCCN TK5105.585 .H54 2003. US$59.00.

Hancock:1982:CP

[HK82] Les Hancock and Morris Krieger.

The C Primer. McGraw-Hill, New York, NY, USA, 1982. ISBN 0-07-025981-X. ix + 235 pp. LCCN QA76.73.C15 H36 1982.

Hodge:2005:MVE

[HK05] Jonathan K. Hodge and Richard E. Klima. The Mathematics of Voting and Elections: a Hands-on Approach, volume 22 of Mathematical world. American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3798-2 (paperback), 1-4704-1194-6 (e-book). ISSN 1055-9426. xiv + 226 pp. LCCN JF1001 .H63 2005.

Harnwell:1933:EAP

[HL33] G. P. (Gaylord Probasco) Harnwell and J. J. (John Jacob) Livingood. Experimental Atomic Physics. McGraw-Hill, New York, NY, USA, 1933. xiii + 472 pp. LCCN QC173 .H38.

Hafner:1996:WWS

[HL96] Katie Hafner and Matthew Lyon. Where Wizards Stay Up Late: the Origins of the Internet. Simon and Schuster, New York, NY, USA, 1996. ISBN 0-684-81201-0 (hardcover), 0-684-83267-4 (paperback), 0-684-87216-1 (e-book), 0-7567-9221-5 (paperback). 304 pp. LCCN TK5105.875.I57H338 1996.

Hill:1988:ECM

[HM88] David R. Hill and Cleve B. Moler (consulting editor). Experiments in Computational Matrix Algebra. Random House, New York, NY, USA, 1988. ISBN 0-394-35678-0 (paperback). xii + 446 pp. LCCN QA188 .H55 1987.

Harrison:1998:ETT

[HM98] Mark Harrison and Michael McLennan. Effective Tcl/Tk Programming: Writing Better Programs with Tcl and Tk. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-63474-0. xv + 405 pp. LCCN QA76.73.T44H37 1998. US$38.50.

Harold:2004:XND

[HM04] Elliotte Rusty Harold and W. Scott Means. XML in a Nutshell: [a Desktop Quick Reference]. In a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2004. ISBN 0-596-00764-7. xix + 689 pp. LCCN QA76.76.H94; Internet; QA76.76.H94 H36 2004eb; QA76.76.H94 H36 2004.

Hess:2010:PVS

[HN10] Kenneth Hess and Amy Newman. Practical Virtualization Solutions: Virtualization from the Trenches. Negus software solutions series. Prentice Hall/ Pearson Education, Upper Saddle River, NJ, USA, 2010. ISBN 0-13-714297-8 (paperback). xxiii + 304 pp. LCCN QA76.9.V5 H47 2010.

Hochschild:1998:KLG

[Hoc98]  Adam Hochschild.  King

Leopold’s Ghost: a Story of Greed, Terror, and Heroism in Colonial Africa. Macmillan, London, UK, 1998. ISBN 0-333-66126-5 (hardcover), 0-618-00190-5 (paperback). 366 + 16 pp. LCCN DT655.H63 1999.

URL http://catdir.loc.gov/
catdir/description/hm051/
2004271848.html;   http://
catdir.loc.gov/catdir/enhancements/
fy0737/2004271848-b.html.

Hochschild:1999:KLG

[Hoc99]  Adam Hochschild.  King

Leopold’s ghost: a story of greed, terror, and heroism in Colonial Africa. Houghton Mifflin, Boston, MA, USA, 1999. ISBN 0-395-75924-2 (hardcover), 0-618-00190-5 (paperback), 0-547-52573-7 (ebook). 366 + 16 pp. LCCN DT655 .H63 1998. URL http:

//books.google.com/books?
id=VLuKAAAAMAAJ;   http://
catdir.loc.gov/catdir/description/
hm022/98016813.html;   http:
//catdir.loc.gov/catdir/enhancements/
fy0736/98016813-b.html;
http://www.houghtonmifflinbooks.
com/readers_guides/hochschild_
king_leo.shtml.

Hoenig:1998:TUL

[Hoe98] Alan Hoenig. TEX Unbound: LaTEX and TEX Strategies for Fonts, Graphics, & More. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1998. ISBN 0-19-509686-X (paperback), 0-19-509685-1 (hardcover). ix + 580 pp. LCCN Z253.4.L38H64 1997. US$60.00 (hardcover), US$35.00 (paperback). URL http://

www.oup-usa.org/gcdocs/gc_
0195096851.html.

Hohn:2001:SMB

[Hoh01] Michael Hohn. On the Solution of Mixed Boundary Value Problems in Elasticity. Ph.D. thesis, Department of Mathematics, University of Utah, Salt Lake City, UT, USA, December 2001. x + 221 pp.

Holub:curses

[Hol87a] Allen I. Holub. Curses: UNIXcompatible windowing output functions. Dr. Dobb’s Journal, 12 (7):94–104, 74–93, July 1987.

Holub:1987:CWU

[Hol87b] Allen I. Holub. On Command:

Writing a Unix-Like Shell for MS-DOS. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1987. ISBN 0-934375-29-1. 319 pp. LCCN QA76.76.O63H65 1987.

Holzgang:1987:UPP

[Hol87c] David A. Holzgang. Understanding PostScript Programming. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1987. ISBN 0-89588-396-1. xxxii + 459 pp. LCCN QA76.73.P67 H65 1987.

Holzgang:1989:PPR

[Hol89] David A. Holzgang. PostScript Programmer’s Reference Guide. Scott, Foresman and Company, Glenview, IL, USA, 1989. ISBN 0-673-38574-4. x + 486 pp. LCCN QA76.73.P67 H64 1989.

Holub:1990:CDC

[Hol90a] Allen I. Holub. Compiler Design in C. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-155045-4, 0-13-155151-5 (paperback). xviii + 924 pp. LCCN QA76.76.C65 H65 1990.

URL http://www.holub.com/
software/compiler.design.
in.c.html.

Holzgang:1990:DPP

[Hol90b] David A. Holzgang. Display PostScript Programming. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-51814-7. x + 406 pp. LCCN QA76.73.P67 H63 1990.

Holub:2004:HPL

[Hol04] Allen I. Holub. Holub on Patterns: Learning Design Patterns by Looking at Code. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-388-X. xv + 414 pp. LCCN QA76.76.P37 H65 2004. US$49.95. URL http:

//www.holub.com/software/
holubSQL/.

Holmes:2008:AWHc

[Hol08] Richard Holmes. The age of wonder: how the romantic generation discovered the beauty and terror of science. Vintage Books, New York, NY, USA, 2008. ISBN 1-4000-3187-7 (paperback), 0-307-37832-2 (e-book). xxi + 552 + 24 pp. LCCN Q127.G4 H65 2010.

Holt:2018:WEW

[Hol18] Jim Holt. When Einstein walked with Gödel: excursions to the edge of thought. Farrar, Strauss, and Giroux, New York, NY, USA, 2018. ISBN 0-374-14670-5 (hardcover), 0-374-71784-2 (ebook). xi + 368 pp. LCCN PS3608.O4943595 A6 2018.

Hoskins:1991:IRS

[Hos91] Jim Hoskins. IBM RISC System/6000: a Business Perspective. Wiley, New York, NY, USA, 1991. ISBN 0-471-53294-0. xx + 295 pp. LCCN QA76.8.I25975 H67 1991.

Hoskin:2011:DUW

[Hos11] Michael A. Hoskin. Discoverers of the universe: William and Caroline Herschel. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14833-3 (hardcover). xvi + 237 + 16 pp. LCCN QB35 .H75 2011.

Hough:fps-applications

[Hou81] David Hough. Applications of the proposed IEEE-754 standard for floating point arithmetic. Computer, 14(3):70–74, March 1981. See [IEE85b].

Hennessy:1990:CAQ

[HP90] John L. Hennessy and David A.

Patterson. Computer Architecture: a Quantitative Approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1990. ISBN 1-55860-069-8, 1-55880-169-8. xxviii + 594 pp. LCCN QA76.9.A73 P377 1990.

Heudin:1992:RA

[HP92] Jean-Claude Heudin and Christian Panetto. RISC Architectures. Chapman and Hall, Ltd., London, UK, 1992. ISBN 0-412-45340-1. ix + 261 pp. LCCN QA76.9.A93 H48 1992. US$43.95.

Hennessy:1994:COD

[HP94] John L. Hennessy and David A. Patterson. Computer Organization and Design: The Hardware/Software Interface. Morgan Kaufmann Publishers, 2929 Campus Drive, Suite 260, San Mateo, CA 94403, USA, 1994. ISBN 1-55860-281-X (paperback), 1-4832-2118-0 (e-book). xxiv + 648 pp. LCCN QA76.9 .C643 P37 1994. US$74.75.

Hennessy:1996:CAQ

[HP96] John L. Hennessy and David A. Patterson. Computer Architecture — a Quantitative Approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 1996. ISBN 1-55860-329-8. xxiii + 760 + A-77 + B-47 + C-26 + D-26 + E-13 + R-16 + I-14 pp. LCCN QA76.9.A73P377 1995. US$69.95.

Hennessy:1997:COH

[HP97] John L. Hennessy and David A. Patterson. Computer Organization: The Hardware/Software Interface. Morgan Kaufmann Publishers, 2929 Campus Drive, Suite 260, San Mateo, CA 94403, USA, second edition, 1997. ISBN 1-55860-428-6 (hardcover), 1-55860-491-X (softcover). xxix + 759 + 205 pp. LCCN QA76.9.C643H46 1997. US$78.95.

Hennessy:2003:CAQ

[HP03] John L. Hennessy and David A.

Patterson. Computer Architecture — a Quantitative Approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, third edition, 2003. ISBN 1-55860-596-7. xxi + 883 + A-87 + B-42 + C-1 + D-1 + E-1 + F-1 + G-1 + H-1 + I-1 + R-22 + I-44 pp. LCCN QA76.9.A73 P377 2003. US$89.95. URL

http://www.mkp.com/books_
catalog/catalog.asp?ISBN=1-
55860-596-7;   http://www.
mkp.com/CA3.

Hennessy:2004:COH

[HP04] John L. Hennessy and David A.

Patterson. Computer Organization: The Hardware/Software Interface. Morgan Kaufmann Publishers, 2929 Campus Drive, Suite 260, San Mateo, CA 94403, USA, third edition, 2004. ISBN 1-55860-604-1 (paperback), 0-12-088433-X (CD-ROM). xvii + 621 pp. LCCN QA76.9.C643 H46 2004. US$64.95.

Hennessy:2012:CAQ

[HP12] John L. Hennessy and David A.

Patterson. Computer Architecture: a Quantitative Approach. Morgan Kaufmann/Elsevier, Waltham, MA, USA, fifth edition, 2012. ISBN 0-12-383872-X (paperback). xxvii + 493 + 325 pp. LCCN QA76.9.A73 P377

2012. URL http://booksite.
mkp.com/9780123838728. With

contributions by Krste Asanović, Jason D. Kabos, Robert P. Colwell, Thomas M. Conte, José Duato, Diana Franklin, David Goldberg, Norman P. Jouppi, Sheng Li, Naveen Muralimanohar, Gregory D. Peterson, Timothy M. Pinkston, Parthasarathy Ranganthan, David A. Wood, and Amr Zaky.

Hennessy:2019:CAQ

[HP19] John L. Hennessy and David A. Patterson. Computer Architecture: a Quantitative Approach. Morgan Kaufmann/Elsevier, Waltham, MA, USA, sixth edition, 2019. ISBN 0-12-811905-5 (paperback). xxix + 617 + 284 pp. LCCN QA76.9.A73 H46 2019; QA76.9.A73 P377 2019.

URL http://booksite.mkp.
com/9780128119051;   https:
//www.elsevier.com/books-
and-journals/book-companion/
9780128119051.

Hennessy:2006:CAQ

[HPAD^{+}06] John L. Hennessy, David A. Patterson, Andrea C. Arpaci-Dusseau, et al. Computer Architecture: a Quantitative Approach. Morgan Kaufmann Publishers, Boston, MA, USA, fourth edition, 2006. ISBN 0-12-370490-1 (paperback), 0-08-047502-7 (e-book), 0-12-373590-

4. xxvii + 423 + 349 pp. LCCN QA76.9.A73 P377 2006. URL

http://www.loc.gov/catdir/
enhancements/fy0665/2006024358-

d.html;   http://www.loc.
gov/catdir/toc/ecip0618/2006024358.
html.

Haigh:2016:EAM

[HPR16] Thomas Haigh, Mark (Peter Mark) Priestley, and Crispin Rope. ENIAC in action: making and remaking the modern computer. History of computing. MIT Press, Cambridge, MA, USA, 2016. ISBN 0-262-03398-4 (hardcover). xvii + 341 pp. LCCN QA76.8.E53 H34 2016.

Heindel:1975:LPI

[HR75] Lee E. Heindel and Jerry T.

Roberto. LANG-PAK: An Interactive Language Design System, volume 1 of Programming Languages Series, Editor: Thomas E. Cheatham. Elsevier, Amsterdam, The Netherlands, 1975. ISBN 0-444-00154-9 (hardcover), 0-444-00162-X (paperback). xi + 184 pp. LCCN QA76.6 .H468 1975.

Horowitz:1978:FCA

[HS78] Ellis Horowitz and Sartaj Sahni.

Fundamentals of Computer Algorithms. Computer Science Press, 11 Taft Court, Rockville, MD 20850, USA, 1978. ISBN 0-914894-22-6 (hardcover). xiv + 626 pp. LCCN QA76.6 .H67 1978.

Horowitz:1976:FDS

[HS83] Ellis Horowitz and Sartaj Sahni.

Fundamentals of Data Structures. Computer Science Press, 11 Taft Court, Rockville, MD 20850, USA, 1976, 1982, 1983.

ISBN 0-914894-20-X. xii + 564 pp. LCCN QA76.9.D35 H67 1977. US$15.95.

Harbison:1984:CRM

[HS84] Samuel P. Harbison and Guy L. Steele Jr. C: a Reference Manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-110008-4 (paperback), 0-13-110016-5 (hardcover). x + 352 pp. LCCN QA76.73.C15 H38 1984.

Harbison:1987:CRM

[HS87] Samuel P. Harbison and Guy L. Steele Jr. C: A Reference Manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1987. ISBN 0-13-109802-0. xii + 404 pp. LCCN QA76.73.C15 H38 1987.

Harbison:1991:CRM

[HS91] Samuel P. Harbison and Guy L. Steele Jr. C: a Reference Manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1991. ISBN 0-13-110933-2 (paperback), 0-13-110941-3 (hardcover), 0-13-116807-X (international). viii + 392 pp. LCCN QA76.73.C15 H38 1991.

Harbison:1995:CRM

[HS95] Samuel P. Harbison and Guy L. Steele Jr. C — a Reference Manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 1995. ISBN 0-13-326232-4 (hardcover), 0-13-326224-3 (paperback). xx + 455 pp. LCCN QA76.73.C15 H38 1995. US$33.95.

Harbison:2002:CRM

[HS02] Samuel P. Harbison III and Guy L. Steele Jr. C — a Reference Manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fifth edition, 2002. ISBN 0-13-089592-X (paperback), 7-115-16188-7 (paperback). xviii + 533 pp. LCCN QA76.73.C15 H38 2002. US$45.00. URL http:/

/www.CAReferenceManual.com/
;   http://www.phptr.com/
ptrbooks/ptr_013089592X.html.

Herlihy:2008:AMP

[HS08] Maurice Herlihy and Nir Shavit.

The art of multiprocessor programming. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2008. ISBN 0-12-370591-6 (paperback), 0-08-056958-7 (ebook). xx + 508 pp. LCCN QA76.6 .H475 2008.

Hsu:2001:CAS

[Hsu01] John Y. Hsu. Computer Architecture: Software Aspects, Coding, Hardware. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001. ISBN 0-8493-1026-1, 1-351-83604-8, 1-4200-4110-X (e-book). 427 pp. LCCN A76.9.A73 H758 2001. US$89.95, UK£59.99.

Hubbard:1984:CGG

[Hub84] Stuart W. Hubbard. The Computer Graphics Glossary. Van Nostrand Reinhold, New York, NY, USA, 1984. ISBN 0-89774-072-6. vi + 94 pp. LCCN T385 .H78 1983.

Hudson:2006:PN

[Hud06] Paul Hudson. PHP in a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-10067-1 (paperback). xv + 352 pp. LCCN QA76.73.P224 H83 2006.

Hunter:1981:DCC

[Hun81] Robin Hunter. The Design and Construction of Compilers. Wiley, New York, NY, USA, 1981. ISBN 0-471-28054-2. xii + 272 pp. LCCN QA76.6.H86 1981.

Hunt:1985:CT

[Hun85] William James Hunt. The C Toolbox. Addison-Wesley, Reading, MA, USA, 1985. ISBN 0-201-11111-X. xviii + 413 pp. LCCN QA76.73.C15 H85 1985. US$19.95. An excellent introduction to the use of C for the writing of software tools, with particular emphasis (but not limited to) the IBM PC. Tools developed include a screen-based file viewing utility, sorting programs, a Btree module, assembly language primitives, and a terminal emulator.

Hunt:1992:TIN

[Hun92]  Craig Hunt.  TCP/IP Net-

work Administration. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 0-937175-82-X (paperback). xxii + 471 pp. LCCN TK5105.9 .H86.

Hurter:2022:TBS

[Hür22] Tobias Hürter. Too Big for a Single Mind: How the Greatest Generation of Physicists Uncovered the Quantum World. The Experiment, New York, NY, USA, 2022. ISBN 1-61519-920-9 (hardcover), 1-61519-921-7 (e-book). 357 pp. LCCN QC7 .H88813 2022. Translation to English by David Shaw of Das Zeitalter der Unschärfe, published by Klett-Cotta (2021).

Hutchins:1952:ECA

[Hut52] Robert Maynard Hutchins, editor. Elements of Chemistry, by Antoine Laurent Lavoisier; Analytical Theory of Heat, by Jean Baptiste Joseph Fourier; Experimental Researches in Electricity, by Michael Faraday, volume 45 of Great Books of the Western World. Encyclopaedia Britannica, Chicago, IL, USA, 1952. xii + 898 pp. LCCN ????

Hwang:1984:TSP

[Hwa84] Kai Hwang, editor. Tutorial: Supercomputers: Design and Applications. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. ISBN 0-8186-0581-2. viii + 640 pp. LCCN TK 7888.3 H82 1984.

Hejlsberg:2004:CPL

[HWG04] Anders Hejlsberg, Scott Wiltamuth, and Peter Golde. The C# programming language. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-15491-6 (hardcover). xiv + 644 pp. LCCN QA76.76.C154 H45 2004.

Hatton:seismic-kernel

[HWS^{+}88] Les Hatton, Andy Wright, Stuart Smith, Gregg Parkes, Paddy Bennett, and Robert Laws. The Seismic Kernel System: A largescale exercise in Fortran 77 portability. Software—Practice and Experience, 18(4):301–329, April 1988. Describes portability issues in a 500,000+ line system.

Hageman:1981:AIM

[HY81] Louis A. Hageman and David M. Young. Applied Iterative Methods. Computer Science and Applied Mathematics, Editor: Werner Rheinboldt. Academic Press, New York, NY, USA, 1981. ISBN 0-12-313340-8. xvii + 386 pp. LCCN QA297.8 .H34 1981.

Hyman:1990:LC

[Hym90] Marco S. Hyman. Literate C++. Computer Language Magazine, 7 (7):67–68, 70, 72, 74–77, 79, July 1990. CODEN CSAEEU. ISSN 0749-2839.

IBM:IRP86

[IBM86] IBM. IBM RT Personal Computer Technology, publication SA23-1057. IBM Corporation, San Jose, CA, USA, 1986.

IEEE:ITU85

[IEE85a] IEEE. IEEE Trial-Use Standard Specifications for Microprocessor Operating Systems Interfaces. Wiley-Interscience, New York, NY, USA, 1985. ISBN 0-471-01073-1. 173 pp. LCCN QA76.5 .I23 1985. IEEE Std. 855.

IEEE:p754

[IEE85b] IEEE Task P754. ANSI/ IEEE 754-1985, Standard for Binary Floating-Point Arithmetic. IEEE, New York, NY, USA, August 12 1985. A preliminary draft was published in the January 1980 issue of IEEE Computer, together with several companion articles [Cod81, Coo81a, Coo80, Coo81b, Hou81, Ste81]. Available from the IEEE Service Center, Piscataway, NJ, USA.

IEEE:ITU86

[IEE86] IEEE. IEEE Trial-Use Standard Portable Operating System for Computer Environments. Wiley-Interscience, New York, NY, USA, IEEE Std 1003.1 edition, April 1986. ISBN 0-471-85027-6. 207 pp. LCCN TK 275 I5 Std 1003.1.

IEEE:1988:PIC

[IEE88] IEEE, editor. Proceedings of the 2nd IEEE Conference on Computer Workstations, March 7– 10, 1988, Santa Clara, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0810-2 (paperback), 0-8186-4810-4 (microfiche), 0-8186-8810-6 (case). LCCN QA76.5 .I33 1988.

Ifrah:2000:UHN

[Ifr00] Georges Ifrah. The Universal History of Numbers from Prehistory to the Invention of the Computer. Wiley, New York, NY, USA, 2000. ISBN 0-471-37568-3, 0-471-39340-1, 0-471-39671-0, 0-9650455-0-1. xxii + 633 pp. LCCN QA141.I3713 2000. US$39.95. Translated from the French edition, Histoire universelle des chiffres, by David Bellos, E. F. Harding, Sophie Wood, and Ian Monk.

Impey:2019:EML

[Imp19] Chris Impey. Einstein’s Monsters: the Life and Times of Black Holes. W. W. Norton & Co., New York, NY, USA, 2019. ISBN 1-324-00093-7 (hardcover), 1-324-00094-5 (e-book). xix + 295 pp. LCCN QB843.B55 I47 2019.

Adobe:colophon

[Inc] Adobe Systems Incorporated. Colophon: Adobe Systems News Publication.

Adobe:1988:DPS

[Inc88] Adobe Systems Incorporated. The Display PostScript System Reference. Adobe Systems Incorporated, October 10 1988.

Adobe:1990:ATF

[Inc90a] Adobe Systems Incorporated. Adobe Type 1 Font Format. Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400, March 1990.

Adobe:AT1.190

[Inc90b] Adobe Systems Incorporated.

Adobe Type 1 Font Format— Version 1.1. Addison-Wesley, Reading, MA, USA, August 1990. ISBN 0-201-57044-0. iii + 103 pp. LCCN QA76.73.P67 A36 1990. US$14.95. URL http:

//partners.adobe.com/asn/
developer/pdfs/tn/T1_SPEC.
PDF.

ISO:pascal

[Int82] International Organization for Standardization, Geneva, Switzerland. Specification for Computer Programming Language Pascal, ISO 7185-1982, 1982.

Intel:286-hrm

[Int83a] Intel. The iAPX 286 Hardware Reference Manual. Intel Corporation, Santa Clara, CA, USA, 1983. LCCN QA76.8.I264 I14 1983. The definitive statement of the 80286 and 80287 hardware at a strongly technical level. Not an instruction set reference, but does contain instruction timing tables. See also [Int85a].

Intel:OSW83

[Int83b] Intel. iAPX 286 Operating Systems Writer’s Guide. Intel Corporation, Santa Clara, CA, USA, 1983. LCCN TK7874 .I56 198.

Intel:286-prm

[Int85a] Intel. The iAPX 286 Programmer’s Reference Manual. Intel Corporation, Santa Clara, CA, USA, 1985. The definitive statement of what the 80286 and 80287 are. A valuable reference for instruction definitions. See also [Int83a, MA86, PM84].

Intel:II85

[Int85b] Intel. Introduction to the iAPX 286. Intel Corporation, Santa Clara, CA, USA, 1985.

Intel:386-prm

[Int86] Intel. 80386 Programmer’s Reference Manual. Intel Corporation, Santa Clara, CA, USA, 1986. ISBN 1-55512-022-9. LCCN QA76.8.I2928 E5 1986. The definitive statement of what the 80386 and 80387 are. A valuable reference for instruction definitions. See also [Int87].

Intel:386-hrm

[Int87] Intel. 80386 Hardware Reference Manual. Intel Corporation, Santa Clara, CA, USA, 1987. ISBN 1-55512-069-5. LCCN TK7895.M5 E33 1986. The definitive statement of the 80386 and 80387 hardware at a strongly technical level. Not an instruction set reference, but does contain instruction timing tables. See also [Int86].

Intel:860-hrm

[Int90] Intel. i860 64-bit Microprocessor Hardware Reference Manual. Intel Corporation, Santa Clara, CA, USA, 1990. ISBN 1-55512-106-3. LCCN TK7895.M5 I57662 1990.

Intel:860-prm

[Int91] Intel. i860 64-bit Microprocessor Family Programmer’s Reference Manual. Intel Corporation, Santa Clara, CA, USA, 1991. ISBN 1-55512-135-7. LCCN QA76.8.I57 I44 1991.

IPA:1999:HIP

[Int99] International Phonetic Association. Handbook of the International Phonetic Association: A Guide to the Use of the International Phonetic Alphabet. Cambridge University Press, Cambridge, UK, August 1999. ISBN 0-521-63751-1 (paperback), 0-521-65236-7 (hardcover). viii + 204 pp. LCCN P227 .I52 1999. US$17.95.

Isaacson:2007:EHL

[Isa07]  Walter Isaacson.  Einstein:

his life and universe. Simon and Schuster, New York, NY, USA, 2007. ISBN 0-7432-6473-

8. xxii + 675 + 16 pp. LCCN QC16.E5 I76 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0708/2006051264-
d.html.

Isaacson:2014:IHG

[Isa14] Walter Isaacson. The Innovators: How a Group of Hackers, Geniuses and Geeks Created the Digital Revolution. Simon and Schuster, New York, NY, USA, 2014. ISBN 1-4711-3879-8 (hardcover), 1-4767-0869-X (cloth), 1-4711-3897-6 (paperback), 1-4104-7497-6 (cloth), 1-4767-0870-3. viii + 542 pp. LCCN QA76.2.A2.

Jackson:1962:CE

[Jac62] John David Jackson. Classical Electrodynamics. Wiley, New York, NY, USA, 1962. xvii + 641 pp. LCCN QC670 .J2.

Jacobsen:2014:OPS

[Jac14] Annie Jacobsen. Operation Paperclip: the secret intelligence program to bring Nazi scientists to America. Little, Brown and Company, Boston, Toronto, London, 2014. ISBN 0-316-22104-X, 0-316-23982-8, 0-316-25169-0 (ebook). xii + 575 + 16 pp. LCCN D810.S2.

Jaeschke:1989:PCL

[Jae89]  Rex Jaeschke.  Portability

and the C Language. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1989. ISBN 0-672-48428-5. xvi + 382 pp. LCCN QA76.76.C64 J34 1988. URL

                                        https://en.wikibooks.org/
                                        wiki/Portability_and_the_C_
                                        Language.

Jaeschke:1991:DSC

[Jae91] Rex Jaeschke. The Dictionary of Standard C. Professional Press Books, 101 Witmer Road, Horsham, PA 19044, USA, 1991. ISBN 1-878956-07-8. x + 165 pp. LCCN QA76.73.C15 J335 1991.

Jaeschke:1993:CIE

[Jae93] Rex Jaeschke. C++: An Introduction for Experienced C Programmers. CBM Books, A Division of Cardinal Business Media Inc., 101 Witmer Road, Horsham, PA 19044, USA, 1993. ISBN 1-878956-27-2. xii + 236 pp. LCCN QA76.73.C15J3354 1993.

Jaeschke:1996:MSC

[Jae96] Rex Jaeschke. Mastering Standard C: a Self-Paced Training Course in Modern C. CBM Books, A Division of Cardinal Business Media Inc., 101 Witmer Road, Horsham, PA 19044, USA, second edition, 1996. ISBN 1-878956-55-8. xvi + 472 pp. LCCN QA76.73.C15J337 1996. US$40.00.

Jaeschke:2001:DSC

[Jae01] Rex Jaeschke. The Dictionary of Standard C. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-13-090620-4. viii + 263 pp. LCCN QA76.73.C15 J335 2001.

JISCII:charset

[Jap78] Japanese Standards Association, 1-24, Akasaka 4 Chome, Minatoku, Tokyo, 107 Japan. Japanese Industrial Standard JIS C 6626-1978 Code of the Japanese Graphics Character Set for Information Interchange, 1978.

JISCII:dot-printers

[Jap83] Japanese Standards Association, 1-24, Akasaka 4 Chome, Minatoku, Tokyo, 107 Japan. Japanese Industrial Standard JIS C 6234-1983 24-dots Matrix Character Patterns for Dot Printers, 1983.

Johanson:2006:LL

[JE06] Donald C. Johanson and Blake Edgar. From Lucy to language. Simon and Schuster, New York, NY, USA, revised, updated, and expanded edition, 2006. ISBN 0-7432-8064-4. 288 pp. LCCN GN281 .J57 2006. URL

http://www.loc.gov/catdir/
enhancements/fy0705/2007270098-
t.html;   http://www.loc.
gov/catdir/enhancements/fy0713/
2007270098-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0735/2007270098-b.html;
http://www.loc.gov/catdir/
enhancements/fy1310/2007270098-

s.html. Principal photography by David L. Brill.

Jennings:1977:MCE

[Jen77] Alan Jennings. Matrix Computation for Engineers and Scientists. Wiley, New York, NY, USA, 1977. ISBN 0-471-99421-9 (hardcover), 0-471-27832-7 (paperback). xv + 330 pp. LCCN TA347 .D4J461.

Jensen:2000:CCN

[Jen00] Carsten Jensen. Controversy and Consensus: Nuclear Beta Decay 1911–1934, volume 24 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2000. ISBN 3-0348-9569-0 (paperback), 3-7643-5313-9 (hardcover), 3-0348-8444-3 (e-book). xv + 217 pp. LCCN QC793.5.B425

J46 2000. URL http://www.
springerlink.com/content/

978-3-0348-8444-0. Carsten

Jensen died of cancer a few months after presenting his doctoral dissertation in 1990 at the University of Copenhagen. Finn Aaserud, Helge Kragh, Erik Rüdinger, and Roger H. Stuewer produced this book as a slightly edited version of that work, supplying additional figures, but leaving the prose largely untouched.

Jerome:2002:EFJ

[Jer02] Fred Jerome. The Einstein file:

J. Edgar Hoover’s secret war against the world’s most famous scientist. St. Martin’s Press, New York, NY, USA, 2002. ISBN 0-312-28856-5. xxii + 358 pp. LCCN QC16.E5 J46 2002. URL

http://www.loc.gov/catdir/
toc/fy022/2001058850.html.

Johnson:1978:LDT

[JL78] Steven C. Johnson and Michael E.

Lesk. Language development tools. The Bell System Technical Journal, 57(6):2155–2176, July/ August 1978.

Johnson:1987:LDT

[JL87] Steven C. Johnson and Michael E.

Lesk. Language development tools. In UNIX System Readings and Applications, pages 245– 265. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. Reprinted from The Bell System Technical Journal, 1978.

Jones:1996:GCA

[JL96] Richard Jones and Rafael Lins.

Garbage Collection: Algorithms for Automatic Dynamic Memory Management. Wiley, New York, NY, USA, 1996. ISBN 0-471-94148-4. xxvi + 377 pp. LCCN QA76.9.G37J66 1996. US$60.00. URL http:

                                        //www.ukc.ac.uk/computer_
                                        science/Html/Jones/gc.html.

Reprinted in 1999 with improved index, and corrected errata.

Janssen:2014:CCE

[JL14] Michel Janssen and Christopher Lehner, editors. The Cambridge Companion to Einstein. Cambridge University Press, Cambridge, UK, 2014. ISBN 0-521-82834-1 (hardcover), 0-521-53542-5 (paperback). xvi + 562 pp. LCCN QC16.E5 C36 2014.

Jude:2016:FMA

[JL16] Allan Jude and Michael W. Lucas. FreeBSD Mastery: Advanced ZFS. Tilted Windmill Press, ????, 2016. ISBN 0-692-68868-4 (paperback). xx + 222 pp. LCCN ????

Jacob:2008:MSC

[JNW08] Bruce Jacob, Spencer W. Ng, and David T. Wang. Memory Systems: Cache, DRAM, Disk. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2008. ISBN 0-12-379751-9 (paperback). xxxiv + 982 pp. LCCN TK7895.M4 J336 2007. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0805/2007282120-
                                        d.html.

Jorgensen:1983:PQC

[JO83] Poul Jørgensen and Jens Oddershede. Problems in Quantum Chemistry. Addison-Wes-ley, Reading, MA, USA, 1983. ISBN 0-201-05486-8. x + 286 pp. LCCN QD462.7.J67 1982.

Johnson:yacc

[Joh79]  Steven C. Johnson.  Yacc:

Yet another compiler compiler. In UNIX Programmer’s Manual, volume 2, pages 353–387. Holt, Rinehart, and Winston, New York, NY, USA, 1979. AT&T Bell Laboratories Technical Report July 31, 1978.

Johnstone:1992:LC

[Joh92] Adrian Johnstone. LaTEX, Concisely. Ellis Horwood, New York, NY, USA, 1992. ISBN 0-13-524539-7 (paperback). xviii + 170 pp. LCCN Z253.4.L38 J64 1992. US$30.00.

Johnson:1993:LAM

[Joh93] Eugene W. Johnson. Linear Algebra with Maple V. Brooks/ Cole, Pacific Grove, CA, USA, 1993. ISBN 0-534-13069-0. xii + 142 pp. LCCN QA185.D37 J64 1993.

Johnson:1996:GAT

[Joh96] Eric F. Johnson. Graphical Applications with Tcl and Tk. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1996. ISBN 1-55851-471-6. x + 374 pp. LCCN T385.J618 1996. US$39.95.

Johnson:2018:HFP

[Joh18]  Andrew Johnson.  Hands-

on Functional Programming in Rust: Build Modular and Reactive Applications with Functional Programming Techniques in Rust 2018. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78883-935-8 (paperback). v + 278 pp. LCCN QA76.62.

Jones:1978:WWB

[Jon78] R. V. (Reginald Victor) Jones. The Wizard War: British Scientific Intelligence, 1939–1945. Coward, McCann and Geoghegan, New York, NY, USA, 1978. ISBN 0-698-10896-5. xx + 556 + 16 pp. LCCN D810.C88 J66 1978. URL

                                        https://en.wikipedia.org/
                                        wiki/Reginald_Victor_Jones.

Jones:1989:IXW

[Jon89] Oliver Jones. Introduction to the X Window System. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-499997-5. xii + 511 pp. LCCN QA76.76.W56 J66 1989.

Jones:2008:QTS

[Jon08]  Sheilla Jones.  The Quan-

tum Ten: a Story of Passion, Tragedy, Ambition and Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-536909-2. xii + 323 + 8 pp. LCCN QC174.12 .J66 2008.

Josuttis:1999:CSL

[Jos99] Nicolai M. Josuttis. The C++ Standard Library: a tutorial and reference. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-37926-0. xx + 799 pp. LCCN QA76.73.C153J69 1999. US$49.95.

Jagger:2007:ACS

[JPS07] Jon Jagger, Nigel Perry, and Peter Sestoft. Annotated C# standard. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2007. ISBN 0-12-372511-9. xxiii + 825 pp. LCCN QA76.73.C154 J35 2007.

Johnson:1989:XWA

[JR89] Eric F. Johnson and Kevin Reichard. X Window Applications Programming. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1989. ISBN 1-55828-016-2. xxii + 562 pp. LCCN QA76.76.W56 J64 1989.

Johnson:1990:AXW

[JR90] Eric F. Johnson and Kevin Reichard. Advanced X Window Applications Programming. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1990. ISBN 1-55828-029-4. xxii + 615 pp. LCCN QA76.76.W56 J63 1990.

Johnson:1991:PPL

[JR91] Eric F. Johnson and Kevin Reichard. power programming … MOTIF. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1991. ISBN 1-55828-059-6. xv + 464 pp. LCCN QA76.76.W56M68 1990. US$29.95 (book), US$59.95 (book + diskette).

Johnson:1993:PPL

[JR93] Eric F. Johnson and Kevin Reichard. power programming … MOTIF. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, second edition, 1993. ISBN 1-55828-322-6. xxxvii + 991 pp. LCCN QA76.76.W56 J636 1993. US$59.95.

Jeffers:2013:IXP

[JR13] Jim Jeffers and James Reinders. Intel Xeon Phi coprocessor high-performance programming. Morgan Kaufmann/Elsevier, Waltham, MA, USA, 2013. ISBN 0-12-410414-2 (paperback), 0-12-410494-0 (e-book). xx + 409 pp. LCCN QA76.5 .J44 2013. URL https://

www.sciencedirect.com/book/
9780124104143/intel-xeon-
phi-coprocessor-high-performance-
programming.

Jenks:1992:ASC

[JS92] Richard D. Jenks and Robert S. Sutor. AXIOM: The Scientific Computation System. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97855-0 (New York), 3-540-97855-0 (Berlin). xxiv + 742 pp. LCCN QA76.95.J46

1992.  URL http://link.
springer.com/10.1007/978-1-
4612-2940-7.

Jaffe:2018:PE

[JT18] Robert L. Jaffe and Washington Taylor, IV. The Physics of Energy. Cambridge University Press, Cambridge, UK, 2018. ISBN 1-107-01665-7 (hardcover). xxi + 874 pp. LCCN QC28 .J34 2018.

Juley:1947:BC

[Jul47] J. Juley. The ballistic computer.

Bell Laboratories Record, 24(??): 5–9, 1947. CODEN BLRCAB. ISSN 0005-8564. Reprinted in [Ran82, §6.3].

Jensen:1974:PUM

[JW74] Kathleen Jensen and Niklaus Wirth. Pascal User Manual and Report. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1974. ISBN 0-387-90144-2, 3-540-90144-2, 3-540-07167-9 (print), 3-540-37500-7 (e-book). viii + 167 pp. LCCN QA76.73.P35 J461

1975.  URL http://link.
springer.com/10.1007/978-3-

540-37500-5. See also [JW85, Ame83b].

Jenkins:1976:FO

[JW76] Francis A. Jenkins and Harvey E.

White. Fundamentals of Optics. McGraw-Hill, New York, NY, USA, fourth edition, 1976. ISBN 0-07-032330-5. xx + 746 pp. LCCN QC355.2.J46 1976.

Jensen:1985:PUM

[JW85] Kathleen Jensen and Niklaus Wirth. Pascal User Manual and Report: ISO Pascal Standard. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1985. ISBN 0-387-96048-1, 3-540-96048-1. xvi + 266 pp. LCCN QA76.73.P2 J46 1985. Revised by Andrew B. Mickel and James F. Miner. See [Ame83b, JW74].

Kornrumpf:1998:LET

[KA98] H.-J. Kornrumpf and Resuhi Abdekmen. Langenscheidt English– Turkish, Turkish–English Universal Dictionary. Altin Kıtaplar, Istanbul, Turkey, 1998. ISBN 975-405-683-8. 444 pp. LCCN PL191.L36 1998.

Kahn:1974:C

[Kah74] David Kahn. The Codebreakers. Weidenfeld and Nicolson, London, UK, abridged edition, 1974. ISBN 0-02-560460-0, 0-297-76785-2. xvi + 576 pp. LCCN Z103 .K28 1974.

Kahn:c

[Kah87] Philippe Kahn, 1987. Borland International, 1800 Green Hills Road, P. O. Box 660005, Scotts Valley, CA 95066-0005. The quote “C is a write-only language” was attributed to Kahn by a trade journal columnist.

Kahan:SIGPLAN-27-1-61

[Kah92] W. Kahan. Analysis and refutation of the LCAS. ACM SIG-PLAN Notices, 27(1):61–74, January 1992.

Kahn:1996:CSS

[Kah96] David Kahn. The Codebreakers: the Story of Secret Writing. Scribner, New York, NY, USA, revised edition, 1996. ISBN 0-684-83130-9. xviii + 1181 pp. LCCN Z103 .K28 1996.

Kane:1987:MRR

[Kan87] Gerry Kane. MIPS R2000 RISC Architecture. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-584749-4. LCCN QA76.8.M52 K36 1987.

Kanigel:1991:MWK

[Kan91]  Robert Kanigel.  The Man

Who Knew Infinity — a Life of the Genius Ramanujan. Collier Macmillan Canada, Toronto, Ontario, Canada, 1991. ISBN 0-684-19259-4. ix + 438 pp. LCCN QA29.R3 K36 1991. US$27.95.

Kane:1996:PRA

[Kan96]  Gerry Kane.  PA-RISC 2.0

Architecture. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-182734-0. various pp. LCCN QA76.8.H48K36 1996. US$34.40.

URL http://devresource.hp.
com/devresource/Docs/Refs/
PA2_0/index.html;   http://
devresource.hp.com/devresource/
Docs/Refs/PA2_0/updates/index.
html.

Kaplan:1999:NNH

[Kap99] Robert Kaplan. The Nothing That Is: a Natural History of Zero. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-512842-7 (hardcover), 0-19-514237-3 (paperback), 1-280-53045-6, 1-4294-0474-4 (e-book). xii + 225 pp. LCCN QA141 .K36 1999. US$22.00.

Kaplan:2020:BPG

[Kap20] Fred M. Kaplan. The Bomb: Presidents, Generals, and the Secret History of Nuclear War. Simon and Schuster Paperbacks, New York, NY, USA, 2020. ISBN 1-982107-29-4, 1-982107-30-8 (paperback), 1-982107-31-6 (e-book). 372 + 8 pp. LCCN U264.3 .K37 2021.

Karow:1987:DFT

[Kar87] Peter Karow. Digital Formats for Typefaces. URW-Verlag, Hamburg, Germany, 1987. ISBN 3-926515-01-5. 400 pp. LCCN Z253.3 .K371 1987.

Karski:2001:SSS

[Kar01] Jan Karski. Story of a Secret State. Houghton-Mifflin, Boston, MA, USA, 2001. ISBN 1-931541-39-6 (hardcover). vi + 391 pp. LCCN D802.P6 K3 2001.

Katzan:1978:F

[Kat78] Harry Katzan, Jr. Fortran 77. Computer Science Series. Van Nostrand Reinhold, New York, NY, USA, 1978. ISBN 0-442-24278-6. xvi + 207 pp. LCCN QA76.73 .F25 K373 1978.

Kauzmann:1957:QCI

[Kau57] Walter Kauzmann. Quantum Chemistry: An Introduction. Academic Press, New York, NY, USA, 1957. xii + 744 pp. LCCN QD453 .K3 1957.

Kaufman:1978:FCB

[Kau78] Roger Emanuel Kaufman. A Fortran Coloring Book. MIT Press, Cambridge, MA, USA, 1978. ISBN 0-262-61026-4. 285 pp. LCCN QA76.73 F25 K38 1978. US$6.95.

Kyker:1976:IMA

[KCM^{+}76] Granvil C. Kyker, Jr., D. Rae Carpenter, Jr., Richard B. Minnix, William R. Riley, and Stanley Williams. Instructor’s Manual to Accompany Physics, by Paul A. Tipler. Worth Publishers, Inc., New York, NY, USA, 1976. v + 189 pp. LCCN ????

Kopka:1993:GLD

[KD93] Helmut Kopka and Patrick W.

Daly. A Guide to LaTEX: Document Preparation for Beginners and Advanced Users. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-56889-6. xvi + 436 pp. LCCN Z253.4.L38 K66 1993. US$34.50.

Kopka:1995:GDP

[KD95] Helmut Kopka and Patrick W.

Daly. A Guide to LaTEX2e: Document Preparation for Beginners and Advanced Users. Addison-Wesley, Reading, MA, USA, second edition, 1995. ISBN 0-201-42777-X. x + 554 pp. LCCN Z253.4.L38 K66 1995. US$37.61.

Kopka:1999:GLD

[KD99] Helmut Kopka and Patrick W. Daly. A Guide to LaTEX2ε: Document Preparation for Beginners and Advanced Users. Addison-Wesley, Reading, MA, USA, third edition, 1999. ISBN 0-201-39825-7. xv + 600 pp. LCCN Z253.4.L38 K66 1999. US$44.95.

Kopka:2004:GLT

[KD04] Helmut Kopka and Patrick W. Daly. Guide to LaTEX: Tools and Techniques for Computer Typesetting. Addison-Wesley, Reading, MA, USA, fourth edition, 2004. ISBN 0-321-17385-6 (paperback). xii + 597 pp. LCCN Z253.4.L38 K66 2004. US$49.99, CAN$75.99. Includes CD-ROM.

Kelly:2017:EYS

[KD17] Scott Kelly and Margaret Lazarus Dean. Endurance: a Year in Space, a Lifetime of Discovery. Alfred A. Knopf, New York, NY, USA, 2017. ISBN 1-5247-3159-5 (hardcover). 387 + 24 pp. LCCN TL789.85.K45 A3 2017.

Keay:2000:GAD

[Kea00] John Keay. The Great Arc: the Dramatic Tale of How India Was Mapped and Everest Was Named. HarperCollins College Publishers, New York, NY, USA, 2000. ISBN 0-06-019518-5 (paperback), 0-00-257062-9 (hardcover). xxi + 182 pp. LCCN QB296.I5 K43 2000. US$13.00.

Kean:2010:DSO

[Kea10] Sam Kean. The disappearing spoon: and other true tales of madness, love, and the history of the world from the periodic table of the elements. Little, Brown and Company, Boston, Toronto, London, 2010. ISBN 0-316-05164-0. vi + 391 pp. LCCN QD466 .K37 2010.

Kean:2012:VTO

[Kea12] Sam Kean. The Violinist’s Thumb: and Other Lost Tales of Love, War, and Genius, as Written by Our Genetic Code. Little, Brown and Company, Boston, Toronto, London, 2012. ISBN 0-316-18231-1 (hardcover). ix + 403 pp. LCCN QH431 .K24 2012.

Keating:2018:LNP

[Kea18] Brian (Brian Gregory) Keating.

Losing the Nobel Prize: a story of cosmology, ambition, and the perils of science’s highest honor. W. W. Norton & Co., New York, NY, USA, 2018. ISBN 1-324-00091-0 (hardcover). xxi + 326 pp. LCCN QB991.B54 K43 2018.

Kehoe:1993:ZAI

[Keh93] Brendan P. Kehoe. Zen and the Art of the Internet: a Beginner’s Guide. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1993. ISBN 0-13-010778-6. xv + 112 pp. LCCN TK5105.875.I57 K44 1993. The first edition of this book was and is available only online, from various anonymous FTP servers.

This second edition is somewhat enlarged and updated, yet is still the shortest introductory Internet book.

Kelly:2009:MPB

[Kel09] Cynthia C. Kelly, editor. The Manhattan Project: the birth of the atomic bomb in the words of its creators, eyewitnesses, and historians. Black Dog and Leventhal, New York, NY, USA, 2009. ISBN 1-57912-808-4 (paperback), 1-57912-747-9, 1-60376-206-X (e-book). xiv + 495 pp. LCCN QC773.3.U6 M27 2009. US$10.95.

Kennedy:2012:SGE

[Ken12] Robert E. Kennedy. A student’s guide to Einstein’s major papers. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-969403-6 (hardcover). xxi + 303 pp. LCCN QC16.E5 K466 2012.

Kernighan:ditroff

[Ker81a]  Brian W. Kernighan.  A

typesetter-independent TROFF. Technical Report Computer Science Report 91, AT&T Bell Laboratories, Murray Hill, NJ 07974, USA, 1981.

Kernighan:1981:WPM

[Ker81b] Brian W. Kernighan. Why Pascal is not my favorite programming language. Computing Science Technical Report 100, AT&T Bell Laboratories, Murray Hill, NJ 07974, USA, April

2, 1981. URL https://github.
com/arnoldrobbins/cstr100.

Kernighan:pic

[Ker82] Brian W. Kernighan. PIC: A language for typesetting graphics. Software—Practice and Experience, 12(1):1–22, January 1982.

Kerrigan:1993:MF

[Ker93] James F. Kerrigan. Migrating to Fortran 90. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 1993. ISBN 1-56592-049-X. xxvi + 361 pp. LCCN QA76.73.F28 K47 1993. US$24.95. URL

http://www.loc.gov/catdir/
enhancements/fy0715/94174237-
d.html.

Kershaw:2005:MFH

[Ker05] Ian Kershaw, editor. Making Friends with Hitler: Lord Londonderry, the Nazis, and the Road to World War II. Penguin, London, UK and New York, NY, USA, 2005. ISBN 1-59420-030-0, 0-14-303607-6. xix + 488 pp. LCCN D750 .K49 2004.

Kerrisk:2010:LPI

[Ker10] Michael Kerrisk. The Linux programming interface: a Linux and UNIX system programming handbook. No Starch Press, San Francisco, CA, USA, 2010. ISBN 1-59327-220-0 (hardcover). xli + 1506 pp. LCCN QA76.76.O63 K496 2010.

Kernighan:2018:MBZ

[Ker18]  Brian W. Kernighan.  Mil-

lions, Billions, Zillions: Defending Yourself in a World of Too Many Numbers. Princeton University Press, Princeton, NJ, USA, 2018. ISBN 0-691-18277-9 (hardcover). xiv + 160 pp. LCCN QA93 .K44 2018.

Kernighan:2019:UHM

[Ker19] Brian W. Kernighan. UNIX: A History and a Memoir. Kindle Direct Publishing, ????, 2019. ISBN 1-69597-855-2. xiii + 183 pp. LCCN ????

Kessler:1988:LOS

[Kes88] Robert R. Kessler. LISP, Objects, and Symbolic Programming. Scott, Foresman and Company, Glenview, IL, USA, 1988. ISBN 0-673-39773-4 (paperback), 0-673-39773-1. xii + 644 pp. LCCN QA76.73.L23 K47 1988.

Kevles:1979:PHS

[Kev79] Daniel J. Kevles. The Physicists: the History of a Scientific Community in Modern America. Vintage Books, New York, NY, USA, 1979. ISBN 0-394-72669-3 (paperback). xi + 489 pp. LCCN QC9.U5 K48 1979.

Keyes:IBMJRD-32-1-24

[Key88] Robert W. Keyes. Miniaturization of electronics and its limits. IBM J. Res. Develop., 32(1):24– 28, January 1988.

Kane:1992:MRA

[KH92] Gerry Kane and Joe Heinrich. MIPS RISC Architecture. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-590472-2. LCCN QA76.8.M52 K37 1992.

Khayat:arabic-dot-matrix

[Kha86] M. G. Khayat. Printing Arabic text on dot matrix printers. Software—Practice and Experience, 16(2):165–172, February 1986.

Khoshnevisan:2002:MPI

[Kho02] Davar Khoshnevisan. Multiparameter Processes: An Introduction to Random Fields. Springer Monographs in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95459-7. xix + 584 pp. LCCN QA274.45 .K58 2002.

Khoshnevisan:2007:P

[Kho07] Davar Khoshnevisan. Probability, volume 80 of Graduate studies in mathematics. American Mathematical Society, Providence, RI, USA, 2007. ISBN 0-8218-4215-3. ISSN 1065-7339. xvi + 224 pp. LCCN QA273 .K488 2007.

Kiernan:2013:GAC

[Kie13] Denise Kiernan. The girls of Atomic City: the untold story of the women who helped win World War II. Simon and Schuster, New York, NY, USA, 2013. ISBN 1-4516-1752-6 (hardcover), 1-4516-1754-2 (e-book). xvii + 373 + 16 pp. LCCN F444.O3 K54 2013.

Kirk:1992:GGI

[Kir92] David Kirk, editor. Graphics Gems III, volume 3 of Graphics Gems. Harcourt Brace Jovanovich, Boston, MA, USA, 1992. ISBN 0-12-409670-0 (with IBM disk), 0-12-409671-9 (with Macintosh disk). xxxi + 631 pp. LCCN T385 .G6973 1992.

Kirschenbaum:2005:GGH

[Kir05] Valerie Kirschenbaum. Goodbye Gutenberg: hello to a new generation of readers and writers. Designer Writers series. Global Renaissance Society, New York, NY, USA, 2005. ISBN 0-9745750-3-8. 416 pp. LCCN Z116.A3 K57 2005. URL http:/

/www.goodbyegutenberg.com/.

Katz:2007:MEM

[KIR^{+}07] Victor J. Katz, Annette Imhausen, Eleanor Robson, Joseph W. Dauben, Kim Plofker, and J. Lennart Berggren, editors. The mathematics of Egypt, Mesopotamia, China, India, and Islam: a sourcebook. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-11485-4 (hardcover). xiv + 685 pp. LCCN QA22 .M3735

2007.  URL http://press.
princeton.edu/titles/8583.
html;   http://www.loc.gov/
catdir/enhancements/fy0726/
2006030851-b.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0726/2006030851-d.html;
http://www.loc.gov/catdir/
toc/ecip071/2006030851.html.

Kissinger:2011:C

[Kis11] Henry Kissinger. On China.

Penguin, London, UK and New York, NY, USA, 2011. ISBN 1-59420-271-0 (hardcover). xviii + 586 + 16 pp. LCCN DS775.8 .K47 2011.

Kittsley:1963:PCE

[Kit63] Scott L. Kittsley. Physical Chemistry: Essentials of the Introductory Course. Review Questions. Problems with Answers, volume 97 of College Outline Series. Barnes and Noble, New York, NY, USA, second edition, 1963. xvii + 217 pp. LCCN QD453 .K47 1963.

Klerer:1967:DCU

[KK67] Melvin Klerer and Granino A.

Korn, editors. Digital Computer User’s Handbook. Mc-Graw-Hill, New York, NY, USA, 1967. LCCN QA76.5 .K524.

Kylander:1999:GUM

[KK99] Olof Kylander and Karin Kylander. The GIMP User’s Manual. Coriolis Group Books, Scottsdale, AZ, USA, 1999. ISBN 1-57610-520-2 (paperback). xxiv + 895 pp. LCCN T385 .K866 1999. US$49.99, CAN$73.99, UK£34.99.

Kay:1992:GFF

[KL92] David C. Kay and John R.

Levine. Graphics File Formats. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1992. ISBN 0-8306-3060-0 (hardcover), 0-8306-3059-7 (paperback). xviii + 278 pp. LCCN T385 .K376 1992. US$24.95.

Knuth:1993:CSS

[KL93] Donald E. Knuth and Silvio Levy. The CWEB System of Structured Documentation, Version 3.0. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-57569-8. 226 pp. LCCN QA76.9.D3 K6 1993.

Kay:1995:GFF

[KL95] David C. Kay and John R. Levine. Graphics File Formats. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, second edition, 1995. ISBN 0-07-034025-0. xx + 476 pp. LCCN T385.K376 1995. US$29.95.

Kleene:1956:REN

[Kle56] S. C. Kleene. Representation of events in nerve nets and finte automata. In Claude E. Shannon and John McCarthy, editors, Automata studies, volume 34 of Annals of mathematics studies, pages 3–40. Princeton University Press, Princeton, NJ, USA, 1956. LCCN QA1.A626 no.34; TJ211 .S47. This is the original paper that introduced the concept of regular expressions, later widely implemented in software tools in the Bell Laboratories Unix operating system.

Klemens:2006:MYC

[Kle06] Ben Klemens. Ma+ h you can’t use: patents, copyright, and software. Brookings Institution Press, Washington, DC, USA, 2006. ISBN 0-8157-4942-2 (cloth). ix + 181 pp. LCCN K1519.C6 K54 2006. URL http:

//www.loc.gov/catdir/toc/
ecip0519/2005027332.html.

Klotz:1986:DDF

[Klo86] Irving M. (Irving Myron) Klotz.

Diamond Dealers and Feather Merchants: Tales from the Sciences. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1986. ISBN 0-8176-3303-0. x + 120 pp. LCCN Q173 .K5735 1986.

Klotz:1995:NMS

[Klo95]  Irving M. Klotz.  Number

mysticism in scientific thinking. The Mathematical Intelligencer, 17(1):43–51, 1995. CODEN MAINDC. ISSN 0343-6993.

Knuth:math-writing

[KLR89] Donald E. Knuth, Tracy Larrabee, and Paul M. Roberts. Mathematical Writing. Mathematical Association of America Notes Number 14, 1989.

Kahrs:ditroff

[KM84] Mark Kahrs and Lee Moore.

Adventures with typesetterindependent TROFF. USENIX Summer Conference Proceedings, pages 258–269, June 12–15 1984.

Kahaner:1989:NMS

[KMN89] David Kahaner, Cleve Moler, and Stephen Nash. Numerical Methods and Software. Prentice-Hall series in computational mathematics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-627258-4. xii + 495 pp. LCCN TA345 .K341 1989. US$50.

Knuth:string-search

[KMP77] Donald E. Knuth, J. H. Morris, and V. R. Pratt. Fast pattern matching in strings. SIAM Journal on Computing, 6(2):323– 350, June 1977. See also [BM77, Sun90, BYG92].

Kirk:2010:PMP

[KmWH10] David B. Kirk and Wen mei W. Hwu. Programming Massively Parallel Processors: a Hands-on Approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2010. ISBN 0-12-381472-3. xviii + 258 pp. LCCN QA76.642 .K57 2010. Chapter 7 (pages 125–140) discusses GPU floating-point considerations.

Kreyszig:1994:MCM

[KN94] E. Kreyszig and E. J. Normington. Maple Computer Manual for Seventh Edition Advanced Engineering Mathematics: Erwin Kreyszig. Wiley, New York, NY, USA, 1994. ISBN 0-471-31126-X. xii + 506 pp. LCCN QA401 K74 1993. There is a companion Instructor’s Maple Manual, but I have not yet found an exact citation for it.

Klabnik:2017:RPL

[KN17] Steve Klabnik and Carol Nichols. The Rust Programming Language. No Starch Press, San Francisco, CA, USA, 2017. ISBN 1-59327-828-4 (paperback), 1-59327-851-9 (e-pub). xxvii + 519 pp. LCCN QA76.73.R87 K53 2018.

Kim:2019:NPE

[KN19a] Y. S. Kim and Marilyn E. Noz.

New perspectives on Einstein’s E = mc^{2}. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2019. ISBN 981-323-770-8 (hardcover), 981-323-771-6 (ebook). xi + 192 pp. LCCN QC173.59.S65 K56 2018.

Klabnik:2019:RPL

[KN19b] Steve Klabnik and Carol Nichols.

The Rust programming language. No Starch Press, San Francisco, CA, USA, second edition, 2019. ISBN 1-09-812253-4, 1-71850-044-0 (paperback). xxix + 526 pp. LCCN QA76.73.R87. URL http://

proquest.safaribooksonline.
com/?fpi=9781098122539;
https://nostarch.com/download/
samples/RustProgrammingLanguage2018_
Sample_ToC.pdf; https://nostarch.
com/Rust2018.

Knuth:ACP68-1

[Knu68] Donald E. Knuth. Fundamental Algorithms, volume 1 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, 1968. ISBN 0-201-03803-X. xxi + 634 pp. LCCN QA76.5 .K74. US$19.75.

Knuth:ACP69-2

[Knu69] Donald E. Knuth. Seminumerical Algorithms, volume 2 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, 1969. ISBN 0-201-03802-1. xi + 624 pp. LCCN QA76.5 .K57. US$19.75.

Knuth:1970:VNF

[Knu70] Donald E. Knuth. Von Neumann’s first computer program. ACM Computing Surveys, 2(4): 247–260, December 1970. CO-DEN CMSVAN. ISSN 0010-4892. Reprinted in [AB87].

Knuth:ACP73-1

[Knu73a] Donald E. Knuth. Fundamental Algorithms, volume 1 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, second edition, 1973. ISBN 0-201-03809-9. xxi + 634 pp. LCCN QA76.6 .K641 1973.

Knuth:ACP73-3

[Knu73b] Donald E. Knuth. Sorting and Searching, volume 3 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, 1973. ISBN 0-201-03803-X. xi + 723 pp. LCCN QA76.5 .K74.

Knuth:goto

[Knu74] D. E. Knuth. Structured programming with go to statements. Computing Surveys, 6:261–301, 1974. This paper is a response to [Dij68].

Knuth:1979:TMN

[Knu79] Donald E. Knuth. TEX and : New Directions in Typesetting. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1979. ISBN 0-932376-02-9. xi + 201 + 105 pp. LCCN Z253.3 .K58 1979.

Knuth:1984:TB

[Knu84a] Donald E. Knuth. The TEXbook, volume A of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1984. ISBN 0-201-13447-0 (hardcover), 0-201-13448-9 (paperback). ix + 483 pp. LCCN Z253.4.T47 K58 1986. US$15.95 (paperback), US$32.95 (hardcover). Second printing, revised, October 1984. Sixth printing, revised, January 1986; also published as Computers & Typesetting, Vol. A. Twenty-fourth printing, summer 1993, contains final revisions. Twenty-first printing, June 1992.

Knuth:1984:LP

[Knu84b] Donald E. Knuth. Literate programming. The Computer Journal, 27(2):97–111, May 1984.

Knuth:ct-b

[Knu86a] Donald E. Knuth. TEX: The Program, volume B of Computers and Typesetting. Addison-Wes-ley, Reading, MA, USA, 1986. ISBN 0-201-13437-3. xv + 594 pp. LCCN Z253.4.T47 K578

Knuth:ct-c

[Knu86b] Donald E. Knuth. The - book, volume C of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13445-4 (hardcover), 0-201-13444-6 (paperback). xi + 361 pp. LCCN Z250.8.M46 K58 1986.

Knuth:ct-d

[Knu86c] Donald E. Knuth. : The Program, volume D of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13438-1. xv + 560 pp. LCCN Z250.8.M46 K578 1986.

Knuth:ct-e

[Knu86d] Donald E. Knuth. Computer Modern Typefaces, volume E of Computers and Typesetting. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-13446-2. xv + 588 pp. LCCN Z250.8.M46 K574 1986.

Knuth:halftone

[Knu87]  Donald E. Knuth.  Digital

halftones by dot diffusion. ACM Transactions on Graphics, 6(4): 245–273, October 1987.

Knuth:tex-errors

[Knu88] Donald E. Knuth. The errors of TEX. Technical Report STAN-CS-88-1223, Stanford University, Department of Computer Science, Stanford, CA, USA, September 1988. See [Knu89b].

Knuth:tex-3.0

[Knu89a] Donald E. Knuth. Calling all grand wizards. TEXhax, 89(98), November 1989.

Knuth:tex-errors-2

[Knu89b] Donald E. Knuth. The errors of TEX. Software—Practice and Experience, 19(7):607–681, July 1989. This is an updated version of [Knu88].

Knuth:1990:SPW

[Knu90a] Donald E. Knuth. A simple program whose proof isn’t. In Feijen et al. [FvGGM90], chapter 27, pages 233–242. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL http://

www.zentralblatt-math.org/
zmath/en/search/?an=0718.

68004. This paper discusses the algorithm used in TEX for converting between decimal and scaled fixed-point binary values, and for guaranteeing a minimum number of digits in the decimal representation. See also [Cli90] for decimal to binary conversion, [SW90] for binary to decimal conversion, and [Gri90] for an alternate proof of Knuth’s algorithm.

Knuth:virtual-fonts

[Knu90b] Donald E. Knuth. Virtual fonts:

More fun for grand wizards. TEXhax, 90(11 and 12), January

Knuth:1992:LP

[Knu92] Donald E. Knuth. Literate Programming, volume 27 of CSLI Lecture Notes. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1992. ISBN 0-937073-80-6 (paperback), 0-937073-81-4 (hardcover). xv + 368 pp. LCCN QA76.6.K644. URL

http://csli-www.stanford.
edu/publications/literate.
html.

Knuth:1996:SPC

[Knu96] Donald E. Knuth. Selected Papers on Computer Science, volume 59 of CSLI lecture notes. Cambridge University Press, Cambridge, UK, 1996. ISBN 1-881526-91-7 (paperback), 1-881526-92-5 (hardcover). xii + 274 pp. LCCN QA76.K537 1996. Dedicated “to George Forsythe, whose inspiring leadership shaped the field.”.

Knuth:1997:FA

[Knu97] Donald E. Knuth. Fundamental Algorithms, volume 1 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, third edition, 1997. ISBN 0-201-89683-4. xix + 650 pp. LCCN QA76.6 .K64 1997. US$49.50.

Knuth:1998:SA

[Knu98a] Donald E. Knuth. Seminumerical Algorithms, volume 2 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, third edition, 1998. ISBN 0-201-89684-2. xiii + 762 pp. LCCN QA76.6 .K64 1997. US$52.75.

Knuth:1998:SS

[Knu98b] Donald E. Knuth. Sorting and Searching, volume 3 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, second edition, 1998. ISBN 0-201-89685-0. xiv + 780 pp. LCCN QA76.6.K64 1997. US$49.95.

Knuth:1999:DT

[Knu99a] Donald E. Knuth. Digital Typography, volume 78 of CSLI Lecture Notes. CSLI Publications, Stanford, CA, USA, 1999. ISBN 1-57586-010-4 (paperback), 1-57586-011-2 (hardcover). xvi + 685 pp. LCCN Z249.3.K59 1998. US$90.00 (cloth), US$39.95 (paperback).

Knuth:1999:MRC

[Knu99b] Donald E. Knuth. MMIXware:

a RISC computer for the third millennium, volume 1750 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-66938-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). viii + 550 pp. LCCN QA76.9.A73 K62 1999. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t1750.htm;   http://www.springerlink.
com/content/978-3-540-66938-

                                        8;   http://www.springerlink.
                                        com/openurl.asp?genre=issue&
                                        issn=0302-9743&volume=1750.

Knuth:2000:SPA

[Knu00] Donald E. Knuth. Selected Papers on Analysis of Algorithms, volume 102 of CSLI Lecture Notes. CSLI Publications, Stanford, CA, USA, 2000. ISBN 1-57586-212-3 (paperback), 1-57586-211-5 (hardcover). xvi + 621 pp. LCCN QA9.58 .K65 2000.

Knuth:2003:SPC

[Knu03a] Donald E. Knuth. Selected Papers on Computer Languages, volume 139 of CSLI Lecture Notes. CSLI Publications, Stanford, CA, USA, 2003. ISBN 1-57586-381-2 (hardcover), 1-57586-382-0 (paperback). xviii + 594 pp. LCCN QA76.7 .K63 2002. US$75.00 (hardcover), US$35.00 (paperback).

Knuth:2003:SPD

[Knu03b] Donald E. Knuth. Selected Papers on Discrete Mathematics, volume 106 of CSLI Lecture Notes. CSLI Publications, Stanford, CA, USA, 2003. ISBN 1-57586-248-4 (paperback), 1-57586-249-2 (hardcover). xvi + 812 pp. LCCN QA39.3 .K59 2001. US$40.00.

Knuth:2005:ACPa

[Knu05a] Donald E. Knuth. The Art of Computer Programming: Volume 1, Fascicle 1. MMIX, a RISC Computer for the New Millennium. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-201-85392-2. v + 134 pp. LCCN QA76.6 .K64 2005.

Knuth:2005:ACPb

[Knu05b] Donald E. Knuth. The Art of Computer Programming: Volume 4, Fascicle 3. Generating All Combinations and Partitions. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-201-85394-9. iv + 150 pp. LCCN QA76.6 .K64 2005.

Knuth:2008:ICA

[Knu08] Donald E. Knuth. The Art of Computer Programming: Volume 4, Fascicle 0. Introduction to Combinatorial Algorithms and Boolean Functions. The art of computer programming. Addison-Wesley, Reading, MA, USA, 2008. ISBN 0-321-53496-4. ix + 216 pp. LCCN QA76.6 .K64 2005.

Knuth:2009:ACP

[Knu09] Donald E. Knuth. The Art of Computer Programming: Volume 4, Fascicle 1. Bitwise Tricks and Techniques. Binary Decision Diagrams. The art of computer programming. Addison-Wesley, Reading, MA, USA, 2009. ISBN 0-321-58050-8. vii + 260 pp. LCCN QA76.6 .K64 2009.

Knuth:2011:ACP

[Knu11] Donald E. Knuth. The Art of Computer Programming: Volume 4, Combinatorial algorithms. Part 1, volume 4A of The art of computer programming. Addison-Wesley, Reading, MA, USA, 2011. ISBN 0-201-03804-8. xv + 883 pp. LCCN QA76.6 2005.

Kogge:1981:APC

[Kog81] Peter M. Kogge. The Architecture of Pipelined Computers. Mc-Graw-Hill, New York, NY, USA, 1981. ISBN 0-07-035237-2. xii + 334 pp. LCCN QA76.5 .K587.

Kolbert:2014:SEU

[Kol14] Elizabeth Kolbert. The sixth extinction: an unnatural history. Henry Holt and Company, New York, NY, USA, 2014. ISBN 0-8050-9299-4 (hardcover), 0-8050-9979-4 (e-book), 1-4088-5123-7 (e-book), 1-4088-5122-9 (paperback), 1-4088-5121-0 (hardcover), 1-4088-5711-1 (paperback). 319 pp. LCCN QE721.2.E97.

Kolodin:2019:HMR

[Kol19] Denis Kolodin. Hands-on Microservices with Rust: Build, Test, and Deploy Scalable and Reactive Microservices with Rust 2018. Packt Publishing, Birmingham, UK, 2019. ISBN 1-78934-198-1, 1-78934-275-9. 511 (est.) pp. LCCN QA76.73.R87. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9781789342758.

Kopka:1988:LE

[Kop88] Helmut Kopka. LaTEX: Eine Einführung. Addison-Wesley Verlag, Bonn, Germany, second edition, 1988. ISBN 3-89319-136-4, 3-89319-199-2. xvi + 340 pp.

Koren:2002:CAA

[Kor02] Israel Koren. Computer Arithmetic Algorithms. A. K. Peters, Ltd., Wellesley, MA, USA, second edition, 2002. ISBN 1-56881-160-8 (hardcover), 1-4398-6371-7 (e-book). xv + 281 pp. LCCN QA76.9.C62 K67.

Korpela:2006:UE

[Kor06] Jukka Korpela. Unicode Explained. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-10121-X. xvii + 658 pp. LCCN QA268 .K67 2006. US$59.99. URL http:

//www.oreilly.com/catalog/
unicode/index.html.

Kernighan:1974:EPS

[KP74] Brian W. Kernighan and P. J.

Plauger. The Elements of Programming Style. McGraw-Hill, New York, NY, USA, 1974. ISBN 0-07-034199-0. 147 pp. LCCN QA76.6 .K47.

Kernighan:1976:ST

[KP76] Brian W. Kernighan and P. J.

Plauger. Software Tools. Addison-Wesley, Reading, MA, USA, 1976. ISBN 0-201-03669-X. 338 pp. LCCN QA76.6 .K42 1976.

Kernighan:1978:EPS

[KP78] Brian W. Kernighan and P. J.

Plauger. The Elements of Programming Style. McGraw-Hill, New York, NY, USA, second edition, 1978. ISBN 0-07-034207-5. xii + 168 pp. LCCN QA76.6 .K39 1978.

Kernighan:1981:STP

[KP81] Brian W. Kernighan and P. J. Plauger. Software Tools in Pascal. Addison-Wesley, Reading, MA, USA, 1981. ISBN 0-201-10342-7. ix + 366 pp. LCCN QA76.6 .K493. US$13.95.

Kernighan:1984:UPE

[KP84] Brian W. Kernighan and Rob Pike. The UNIX Programming Environment. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-937699-2 (hardcover), 0-13-937681-X (paperback). x + 357 pp. LCCN QA76.76.O63 K48 1984. URL

http://wiki.tuhs.org/doku.
php?id=publications:unix_
prog_environment.

Kernighan:1999:PP

[KP99] Brian W. Kernighan and Rob Pike. The Practice of Programming. Addison-Wesley, Reading, MA, USA, 1999. ISBN 0-201-61586-X. xii + 267 pp. LCCN QA76.6 .K48 1999. US$24.95, CAN$37.50. URL http://cm.

bell-labs.com/cm/cs/tpop/
code.html;   http://cseng.
aw.com/bookdetail.qry?ISBN=
0-201-61586-X&ptype=0;   http:
//tpop.awl.com.

Kessener:1986:DSR

[KPv86] L. R. A. Kessener, F. J. Peters, and M. L. P. van Lierop. Data Structures for Raster Graphics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-16310-7. vi + 201 pp. LCCN QA76.9.D35 D381 1986.

Kernighan:1978:CPL

[KR78] Brian W. Kernighan and Dennis M. Ritchie. The C Programming Language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1978. ISBN 0-13-110163-3. x + 228 pp. LCCN QA76.73 .C15 K47 1978.

Karp:TR-31-81

[KR81] R. M. Karp and M. O. Rabin.

Efficient randomized patternmatching algorithms. Technical Report TR-31-81, Harvard University, Cambridge, MA, USA, 1981. See [BYG92].

Kernighan:1988:CPL

[KR88] Brian W. Kernighan and Dennis M. Ritchie. The C Programming Language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1988. ISBN 0-13-110362-8 (paperback), 0-13-110370-9 (hardcover), 0-13-115817-1, 0-13-308621-6 (e-book), 0-13-308624-0. xii + 272 pp. LCCN QA76.73.C15 K47 1988. URL

http://wiki.tuhs.org/doku.
php?id=publications:c_prog_

language. This book is 200% a must for any C programmers . . .. The answers to the exercises can be found in [TG89].

Krauss:2011:QMR

[Kra11] Lawrence Maxwell Krauss. Quantum Man: Richard Feynman’s Life in Science. Great discoveries. W. W. Norton & Co., New York, NY, USA, 2011. ISBN 0-393-06471-

9. xvii + 350 pp. LCCN QC16.F49 K73 2011. URL http:

                                        //www.scientificcomputing.
                                        com/news-DS-Illuminating-
                                        the-Life-and-Legacy-of-Richard-
                                        Feynman-032211.aspx.

Krause:2014:CHW

[Kra14] Michael Krause. CERN: how we found the Higgs boson. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2014. ISBN 981-4623-55-5 (hardcover), 981-4623-46-6 (paperback), 981-4623-48-2 (e-book). xiii + 243 pp. LCCN QC793 .K73 2014.

Kreyszig:1993:AEM

[Kre93] Erwin Kreyszig. Advanced Engineering Mathematics. Wiley, New York, NY, USA, seventh edition, 1993. ISBN 0-471-55380-8 (hardcover), 0-471-59989-1 (paperback). xviii + 1272 + 112 + 18 pp. LCCN QA401 .K7 1993. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0707/92006989-
                                        d.html;   http://www.loc.
                                        gov/catdir/toc/onix05/92006989.
                                        html.

Kronrod:1965:NWQ

[Kro65] Aleksandr Semenovich Kronrod. Nodes and weights of quadrature formulas: sixteen-place tables. Consultants Bureau, New York, NY, USA, 1965. vii + 143 pp. LCCN QA310 .K7613.

Kronsjo:1985:CCS

[Kro85] L. Kronsjö. Computational Complexity of Sequential and Parallel Algorithms. Wiley-Intersci-ence, New York, NY, USA, 1985. ISBN 0-471-90814-2. x + 224 pp. LCCN QA76.6 .K7631 1985. US$50.00.

Krommes:fortran-web

[Kro90] John Krommes. FWEB (Krommes) vs. FWEB (Avenarius and Oppermann). TEXhax, 90(19), February 1990.

Krol:1993:WIU

[Kro93] Ed Krol. The Whole Internet:

User’s Guide and Catalog. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-025-2. xxiv + 376 pp. LCCN TK5105.875.I57 K86 1992. US$24.95.

Kruse:1984:DSP

[Kru84] Robert L. Kruse. Data Structures and Program Design. Prentice-Hall Software Series, Editor: Brian W. Kernighan. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-196253-1. xxi + 486 pp. LCCN QA76.6 .K77 1984. US$28.95.

Kruglinski:1994:IVC

[Kru94] David J. Kruglinski. Inside Visual C++: Version 1.5. Microsoft Press, Bellevue, WA, USA, second edition, 1994. ISBN 1-55616-661-8. xxx + 732 pp. LCCN QA76.73.C153K78 1994.

Kolthoff:1952:TQI

[KS52] I. M. Kolthoff and E. B. Sandell. Textbook of Quantitative Inorganic Analysis. Macmillan Publishing Company, New York, NY, USA, third edition, 1952. xv + 759 pp. LCCN QD101 .K6 1952.

Karin:1987:SE

[KS87] Sidney Karin and Norris Parker Smith. The Supercomputer Era. Harcourt Brace Jovanovich, Boston, MA, USA, 1987. ISBN 0-15-186787-9. x + 313 pp. LCCN QA76.5 .K356 1987.

Krieger:1990:IT

[KS90] Jost Krieger and Norbert Schwarz. Introduction to TEX. Addison-Wesley Europe, Amsterdam, The Netherlands, 1990. ISBN 0-201-51141-X. 278 pp. LCCN Z253.4.T47 K75 1990.

Kajiya:grey-scale

[KU81] J. Kajiya and M. Ullner. Filtering high quality text for display on raster scan devices. ACM SIGGRAPH—Computer Graphics, 15(3):7–15, August 1981. See more recent work in [NF87].

Kumar:2010:QEB

[Kum10] Manjit Kumar. Quantum: Einstein, Bohr and the great debate about the nature of reality. W. W. Norton & Co., New York, NY, USA, 2010. ISBN 0-393-07829-9 (hardcover). xvi + 448 + 16 pp. LCCN QC173.98 .K86 2010; QC173.98.K86.

Kunii:1983:CGT

[Kun83] Tosiyasu L. Kunii, editor. Computer Graphics: Theory and Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 0-387-70001-3. ix + 530 pp. LCCN T385 .I49 1983. Proceedings of Intergraphics ’83.

Kunii:1985:CGV

[Kun85a] Tosiyasu L. Kunii, editor. Computer Graphics: Visual Technology and Art. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-70009-9, 3-540-70009-9, 4-431-70009-9. ix + 382 pp. LCCN T385.C5995 1985. US$45.00. Proceedings of Computer Graphics Tokyo ’85.

Kunii:1985:FCG

[Kun85b] Tosiyasu L. Kunii, editor. Frontiers in Computer Graphics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-70004-8, 3-540-70004-8, 4-431-70004-8. ix + 443 pp. LCCN T385 .C593 1984. Proceedings of Computer Graphics Tokyo ’84.

Kunii:1986:ACG

[Kun86] Tosiyasu L. Kunii, editor.

Advanced Computer Graphics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-70011-0. ix + 504 pp. LCCN T385 .C5931 1986. US$45.00. Proceedings of Computer Graphics Tokyo ’86.

Kurani:1994:AUP

[Kur94] Bharat Kurani. Applied Unix programming, volume 2. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-304338-X (vol. 1), 0-13-304346-0 (vol. 2). xviii + 1250 pp. LCCN QA76.76.O63 K86 1994. Two volumes.

Kurlansky:1998:CBF

[Kur98] Mark Kurlansky. Cod: a biography of the fish that changed the world. Penguin, London, UK and New York, NY, USA, 1998. ISBN 0-8027-1326-2 (hardcover), 0-14-027501-0 (paperback), 1-4406-7287-3 (e-book). viii + 294 pp. LCCN SH351.C5 K87 1998.

Kurlansky:1999:BHW

[Kur99] Mark Kurlansky. The Basque history of the world. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 1999. ISBN 0-8027-1349-1. xii + 387 pp. LCCN DP302.B46 K87 1999.

URL http://lcweb.loc.gov/
catdir/toc/99026808.html;
http://www.loc.gov/catdir/
enhancements/fy0601/99026808-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0601/
99026808-d.html.

Kurlansky:2003:SWH

[Kur03]  Mark Kurlansky.  Salt: a

world history. Penguin, London, UK and New York, NY, USA, 2003. ISBN 0-14-200161-9 (paperback). x + 484 pp. LCCN TN900 .K865 2003. US$16.

Kuran:2006:HPA

[Kur06] Peter Kuran. How to Photograph an Atomic Bomb. VCE, Santa Clarita, CA, USA, 2006. ISBN 1-889054-19-4 (paperback), 1-889054-11-9. 141 pp. LCCN UG476 .K87 2006. US$24.95.

Kurzweil:2012:HCM

[Kur12] Ray Kurzweil. How to Create a Mind: the Secret of Human Thought Revealed. Penguin, London, UK and New York, NY, USA, 2012. ISBN 0-14-312404-8. xi + 336 pp. LCCN QP385 .K87 2013.

Kurlansky:2017:PPT

[Kur17] Mark Kurlansky. Paper: paging through history. W. W. Norton & Co., New York, NY, USA, 2017. ISBN 0-393-23961-6 (hardcover), 0-393-35370-2 (paperback). xx + 389 pp. LCCN TS1090 .K87 2017.

Kochan:1989:UN

[KW89] Stephen G. Kochan and Patrick H. Wood, editors. UNIX Networking. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1989. ISBN 0-672-48440-4. viii + 400 pp. LCCN QA76.76.O63 U546 1989.

Knuth:new-tex-mf

[KW90] Donald E. Knuth and Joe Weening. New TEX/ sources available on Stanford’s master archive. TEXhax, 90(13), January 1990.

Kyker:1976:SGA

[Kyk76] Granvil C. Kyker, Jr. Study Guide to Accompany Physics, by Paul A. Tipler. Worth Publishers, Inc., New York, NY, USA, 1976. ISBN 0-87901-055-X. x + 345 pp. LCCN QC21.2 .T5482.

LaPaglia:1971:IQC

[La 71] S. (Salvatore) R. La Paglia. Introductory Quantum Chemistry. Harper & Row, New York, NY, USA, 1971. xiii + 402 pp. LCCN QD462 .L35.

Lafore:1984:ALP

[Laf84] Robert Lafore. Assembly Language Primer for the IBM PC and XT. Plume/Waite, New York, NY, USA, 1984. ISBN 0-452-25497-3. viii + 501 pp. LCCN QA76.8.I2594 L34 1984. A companion book for [Mor84], written for novice assemblylanguage programmers, with considerable emphasis on the IBM PC. More elementary than [Bra84]. See [Sca84].

Laidler:1965:CK

[Lai65] Keith J. (Keith James) Laidler. Chemical Kinetics. McGraw-Hill, New York, NY, USA, second edition, 1965. ix + 566 pp. LCCN QD501 .L3 1965.

LALR:lalr

[LAL87] LALR Research. LALR 3.0, 1987. P. O. Box 50755, Knoxville, TN 37950.

Lamport:1986:LDP

[Lam86]  Leslie Lamport.  LaTEX: a

Document Preparation System: User’s Guide and Reference Manual. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-15790-X. xiv + 242 pp. LCCN Z253.4.L38 L35 1986.

Lamport:1994:LDP

[Lam94]  Leslie Lamport.  LaTEX: a

Document Preparation System: User’s Guide and Reference Manual. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-52983-1. xvi + 272 pp. LCCN Z253.4.L38L35 1994.

Langenscheidt:2001:LPD

[Lan01] Langenscheidt Editorial. Langenscheidt’s Pocket Dictionary Chinese/English, English/Chinese. Asian Language Pocket Dictionaries Series. Langenscheidt Publishers, New York, NY, USA, 2001. ISBN 1-58573-057-2. 672 pp. LCCN PL1455 .L35 2001.

Lapin:1987:PCU

[Lap87] J. E. Lapin. Portable C and UNIX Programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-686494-5. xiv + 249 pp. LCCN QA76.73.C15 L36 1987.

Lapidus:2008:SRZ

[Lap08] Michel L. (Michel Laurent) Lapidus. In Search of the Riemann Zeros: Strings, Fractal Membranes and Noncommutative Spacetimes. American Mathematical Society, Providence, RI, USA, 2008. ISBN 0-8218-4222-6. xxix + 558 pp. LCCN QA333 .L37 2008.

Larsen:1989:LPD

[Lar89] Steen Larsen. LaTEX på dansk. UNI•C, Danmarks EDB-Center for Forskning og Uddannelse, Copenhagen, Denmark, 1989. ISBN 87-7252-089-2. viii + 111 pp.

Larousse:2006:PLI

[Lar06]  Larousse, editor.  Le Petit

Larousse Illustré. Larousse, 21, rue du Montparnasse, 75283 Paris Cedex 06, France, 2006. ISBN 2-03-582492-3. 1918 pp. LCCN AG25 N75 2006. US$78.95.

Latimer:1952:OSE

[Lat52] Wendell M. Latimer. The oxidation states of the elements and their potentials in aqueous solutions. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1952. xvi + 392 pp. LCCN QD561 .L35 1952.

Lawson:sftran

[Law78] C. L. Lawson. SFTRAN language constructs supported by a portable preprocessor. Technical Report JPL Section 366 Internal Computing Memorandum 437, Jet Propulsion Laboratory, Pasadena, CA, USA, January 1978.

Lancaster:2014:QFT

[LB14] Tom Lancaster and Stephen Blundell. Quantum Field Theory for the Gifted Amateur. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-969933-X (paperback), 0-19-969932-1 (hardcover). xvii + 485 pp. LCCN QC174.45 .L36 2014.

Lewis:2017:FUL

[LBS17] Geraint F. Lewis, Luke A.

Barnes, and Brian Schmidt. A fortunate universe: life in a finely-tuned cosmos. Cambridge University Press, Cambridge, UK, 2017. ISBN 1-107-15661-0 (hardcover), 1-316-66141-5 (ebook). xvii + 373 pp. LCCN Q172.5.C45 L4845 2017.

Levy:1980:CPA

[LE80] Henry M. Levy and Richard H.

Eckhouse, Jr. Computer Programming and Architecture — the VAX-11. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1980. ISBN 0-932376-07-X. xxi + 407 pp. LCCN QA76.8 .V37 L48 1980.

Lebedev:1972:SFT

[Leb72] N. N. (Nikolai Nikolaevich) Lebedev. Special functions and their applications. Dover, New York, NY, USA, 1972. ISBN 0-486-60624-4 (paperback). xii + 308 pp. LCCN QA351 .L3613 1972. URL http://www.

                                        loc.gov/catdir/description/
                                        dover031/72086228.html. Trans-

lated to English and edited by Richard A. Silverman.

Leech:1965:CM

[Lee65] J. W. Leech. Classical Mechanics, volume 8 of Methuen’s Monographs on Physical Subjects. Science Paperbacks and Methuen & Co., Ltd., London, England, second edition, 1965. ix + 153 pp. LCCN QA807 .L4 1965.

Lehto:1998:MBH

[Leh98] Olli Lehto. Mathematics Without Borders: a History of the International Mathematical Union. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98358-9. xvi + 399 pp. LCCN QA1 .L42 1998. US$35.00.

Lehey:2003:CFD

[Leh03] Greg Lehey. The Complete FreeBSD: Documentation from the Source. O’Reilly Community Press, Sebastopol, CA, USA, fourth edition, 2003. ISBN 0-596-00516-4 (paperback), 0-596-80213-7 (e-book). xxxiii + 679 pp. LCCN QA76.76.O63 L43 2003. US$44.09, CAN$69.95,

                                        UK£31.95. URL http://press.
                                        oreilly.com/commpress.html;
                                         http://www.oreilly.com/catalog/
                                        cfreebsd/desc.html.

Leighton:2000:TBR

[Lei00] Ralph Leighton. Tuva Or Bust!: Richard Feynman’s Last Journey. W. W. Norton & Co., New York, NY, USA, 2000. ISBN 0-393-32069-3 (paperback), 0-393-02953-0. 260 pp. LCCN QC16.F49 L45 2000.

Lorentz:1952:PRC

[LEMW52] H. A. (Hendrik Antoon) Lorentz, A. (Albert) Einstein, H. (Hermann) Minkowski, and H. (Hermann) Weyl. The Principle of Relativity: a Collection of Original Memoirs on the Special and General Theory of Relativity. Dover, New York, NY, USA, 1952. ISBN 0-486-60081-5, 0-486-31840-0 (e-book). viii + 216 pp. LCCN QC 173.55 L67 1952. With notes by A. Sommerfeld. Translated By W. Perrett and G. B. Jeffery.

Lesk:1997:PDL

[Les97] Michael Lesk. Practical Digital Libraries: Books, Bytes, and Bucks. The Morgan Kaufmann series in multimedia information and systems. Morgan Kaufmann Publishers, 2929 Campus Drive, Suite 260, San Mateo, CA 94403, USA, 1997. ISBN 1-55860-459-6 (paperback). xxii + 297 + 11 pp. LCCN Z692.C65 L47 1997. US$49.95. URL http://www.

loc.gov/catdir/description/
els032/97022069.html;   http:
//www.loc.gov/catdir/toc/
els032/97022069.html.

Levin:2016:BHB

[Lev16] Janna Levin. Black hole blues:

and other songs from outer space. Alfred A. Knopf, New York, NY, USA, 2016. ISBN 0-307-95819-1 (hardcover), 0-307-94848-X (paperback). 241 pp. LCCN QC179 .L48 2016. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy1605/2015046692-
                                        b.html;   http://www.loc.
                                        gov/catdir/enhancements/fy1605/
                                        2015046692-d.html.

Levy:2019:LIT

[Lev19]  Buddy Levy.  Labyrinth of

Ice: the Triumphant and Tragic Greely Polar Expedition. St. Martin’s Press, New York, NY, USA, 2019. ISBN 1-250-18219-0 (hardcover), 1-250-18220-4 (ebook). xviii + 376 + 16 pp. LCCN G670 1881 .L48 2019.

Lewine:1991:PPG

[Lew91] Donald A. Lewine. POSIX programmer’s guide: writing portable UNIX programs with the POSIX.1 standard. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-73-0. xxvii + 607 pp. LCCN QA76.76.O63 L487 1991b. US$34.95. March 1994 printing with corrections, updates, and December 1991 Appendix G.

Lawson:sf3

[LF78] C. L. Lawson and J. A. Flynn. SFTRAN3 programmer’s reference manual. Technical Report 1846-98, Jet Propulsion Laboratory, Pasadena, CA, USA, December 1978.

Liu:1984:MSF

[LG84] Yu-Cheng Liu and Glenn A. Gibson. Microcomputer Systems: The 8086/8088 Family. Architecture, Programming, and Design. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-580944-4. ix + 550 pp. LCCN QA76.8.I292 L58 1984. A broad treatment of the Intel 8086 and 8088, with shorter surveys of the 8087, 80186, and 80286. Nothing specific to the IBM PC.

Lawson:1974:SLS

[LH74] Charles L. Lawson and Richard J. Hanson. Solving Least Squares Problems. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1974. ISBN 0-13-822585-0. xii + 340 pp. LCCN QA275 .L425 1974.

Ledgard:1979:PS

[LHN79] Henry F. Ledgard, John F.

Hueras, and Paul A. Nagin. Pascal with Style. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1979. ISBN 0-8104-5124-7. 210 pp. LCCN QA76.73.P2 L4.

Libes:1994:EET

[Lib94] Don Libes. Exploring Expect:

A Tcl-based Toolkit for Automating Interactive Programs. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, December 1994. ISBN 1-56592-090-2. xxxiii + 566 pp. LCCN QA76.755 .L52 1995. US$29.95. URL http:

//www.ora.com/gnn/bus/ora/
item/expect.html;   http:/
/www.oreilly.com/catalog/
9781565920903;   http://www.
oreilly.com/catalog/expect.

Lindley:2004:DKT

[Lin04] David Lindley. Degrees Kelvin: a Tale of Genius, Invention, and Tragedy. Joseph Henry Press, Washington, DC, USA, 2004. ISBN 0-309-09073-3 (hardcover), 0-309-53095-4 (e-book). viii + 366 pp. LCCN QC16.K3 L56 2004. US$27.59. URL ftp://

uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http:
//www.loc.gov/catdir/toc/
ecip0410/2003022885.html.

Lindley:2008:UEH

[Lin08] David Lindley. Uncertainty: Einstein, Heisenberg, Bohr, and the struggle for the soul of science. Anchor Books, New York, NY, USA, 2008. ISBN 1-4000-7996-9. viii + 257 pp. LCCN QC174.17.H4 L56 2007. URL

http://www.loc.gov/catdir/
toc/ecip0614/2006017029.html.

Lions:1996:LCU

[Lio96]  John Lions.  Lions’ Com-

mentary on UNIX 6th Edition, with Source Code. Computer classics revisited. Peerto-Peer Communications, San Jose, CA 95164-0218, USA, 1996. ISBN 1-57398-013-7. 254 pp. LCCN QA76.76.O63 L562 1996. US$29.96; CAN$41.95. URL

http://wiki.tuhs.org/doku.

php?id=publications:lions_
commentary;   http://www.
lemis.com/grog/Documentation/
Lions/;   http://www.peer-
to-peer.com/catalog/opsrc/
lions.html;   https://github.
com/kanner/lions-book. With

forewords by Dennis M. Ritchie and Ken Thompson. Prefatory notes by Peter H. Salus and Michael Tilson; a Historical Note by Peter H. Salus; and Appreciations by Greg Rose, Mike O’Dell, Berny Goodheart, Peter Collinson, and Peter Reintjes. Originally circulated as two restricted-release volumes: “UNIX Operating System Source Code Level Six”, and “A Commentary on the UNIX Operating System”.

Lippman:1991:CP

[Lip91] Stanley B. Lippman. C++ Primer. Addison-Wesley, Reading, MA, USA, second edition, 1991. ISBN 0-201-54848-8 (paperback). xvi + 614 pp. LCCN QA76.73.C15 L57 1991.

Lipkin:1999:LLV

[Lip99] Bernice Sacks Lipkin. LaTEX for Linux: a Vade Mecum. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98708-8 (paperback). xxxi + 568 pp. LCCN Z253.4 L38 L56 1999. US$49.95.

Livio:2002:GRS

[Liv02] Mario Livio. The Golden Ratio: the Story of Phi, the World’s Most Astonishing Number. Broadway Books, New York, NY, USA, 2002. ISBN 0-7679-0815-5. viii + 294 pp. LCCN QA466 .L58 2002. US$24.95, CAN$37.95. Benford’s Law is treated on pages 231–237.

Livio:2005:ECS

[Liv05] Mario Livio. The Equation That Couldn’t Be Solved: How Mathematical Genius Discovered the Language of Symmetry. Simon and Schuster, New York, NY, USA, 2005. ISBN 0-7432-5820-7 (hardcover), 0-7432-5821-5 (ebook), 0-7432-7462-8 (e-book). x + 353 pp. LCCN QA174.2 .L58 2005.

Lucas:2015:FM

[LJ15] Michael W. Lucas and Allan Jude. FreeBSD Mastery: ZFS. Tilted Windmill Press, ????, 2015. ISBN 0-692-45235-4 (paperback). xx + 212 pp. LCCN ????

Lomuto:1983:UP

[LL83] Ann Nicols Lomuto and Nico Lomuto. A UNIX Primer. Prentice-Hall Software Series, Editor: Brian W. Kernighan. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1983. ISBN 0-13-938886-9. xvi + 239 pp. LCCN QA76.8.U65 L65 1983.

Laurie:1997:ADG

[LL97] Ben Laurie and Peter Laurie. Apache: The Definitive Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997.

ISBN 1-56592-250-6. xvi + 255 pp. LCCN QA76.76.A7 L3 1997. US$34.95. URL http://

www.ora.com/catalog/apache/
;   http://www.oreilly.com/
catalog/apache.

Lloyd:1987:FLP

[Llo87] J. W. (John Wylie) Lloyd. Foundations of Logic Programming. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1987. ISBN 0-387-18199-7, 3-540-18199-7. xii + 212 pp. LCCN QA76.6.L583 1987.

Lieb:1966:MPO

[LM66] Elliott H. Lieb and Daniel C.

Mattis, editors. Mathematical Physics in One Dimension: exactly soluble models of interacting particles. Academic Press, New York, NY, USA, 1966. xiv + 565 pp. LCCN QC174.5 .L47.

Langville:2006:GPB

[LM06] Amy N. Langville and Carl D.

(Carl Dean) Meyer. Google’s PageRank and beyond: the science of search engine rankings. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-12202-4 (hardcover). x + 224 pp. LCCN TK5105.885.G66 L36 2006. URL

http://www.loc.gov/catdir/
enhancements/fy0654/2005938841-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0654/
2005938841-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0668/2005938841-t.html.

Levine:1992:LY

[LMB92] John R. Levine, Tony Mason, and Doug Brown. lex & yacc. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1992. ISBN 1-56592-000-7. xxii + 366 pp. LCCN QA76.76.U84M37 1992. US$29.95. URL http:

                                        //www.oreilly.com/catalog/
                                        9781565920002;   http://www.
                                        oreilly.com/catalog/lex.

Leffler:1989:DIU

[LMKQ89] Samuel J. Leffler, Marshall Kirk McKusick, Michael J. Karels, and John S. Quarterman. The Design and Implementation of the 4.3BSD UNIX Operating System. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-06196-1. xxii + 471 pp. LCCN QA76.76.O63 D4741 1989.

Level:1995:PTF

[LNN95] Jeff Level, Bruce Newman, and Brenda Newman. Precision Type Font Reference Guide, Version 5.0. Precision Type, Commack, NY, USA, 1995. ISBN 0-9646252-0-2. xxv + 653 pp. LCCN Z250.7 .P74 1995.

Level:2000:PTF

[LNN00] Jeff Level, Bruce Newman, and Brenda Newman. Precision Type Font Reference Guide, Version 5.0. Hartley & Marks Publishers, Point Roberts, WA, USA and Vancouver, BC, Canada, 2000. ISBN 0-88179-182-2. xxv + 653 pp. LCCN Z250.7 .P74 2000. US$39.95, CDN$59.95. Republication of [LNN95].

Le-Ngoc:1989:IPF

[LNV89] Tho Le-Ngoc and Minh Tue Vo. Implementation and performance of the Fast Hartley Transform. IEEE Micro, 9(5):20–27, September/October 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Linderberg:1973:PQC

[LÖ73] Jan Linderberg and Yngve Öhrn.

Propagators in Quantum Chemistry, volume 3 of Monographs in Theoretical Chemistry. Academic Press, New York, NY, USA, 1973. ISBN 0-12-450350-0. v + 151 pp. LCCN QD462.L56.

Linderberg:2004:PQC

[LÖ04] Jan Linderberg and Yngve Öhrn.

Propagators in Quantum Chemistry. Theoretical chemistry. Wiley, New York, NY, USA, second edition, 2004. ISBN 0-471-66257-

7. vii + 267 pp. LCCN QD462

L56 2004. URL http://catdir.
loc.gov/catdir/description/
wiley041/2003069509.html;
http://catdir.loc.gov/catdir/
toc/wiley041/2003069509.html;
 https://onlinelibrary.wiley.
com/doi/book/10.1002/0471721549.

Lockhart:2017:A

[Loc17] Paul Lockhart. Arithmetic. The Belknap Press of Harvard University Press, Cambridge, MA, USA, 2017. ISBN 0-674-97223-6. 223 pp. LCCN QA115 .L713 2017.

Lomb:2011:TVP

[Lom11] Nick Lomb. Transit of Venus: 1631 to the present. Experiment, New York, NY, USA, 2011. ISBN 1-61519-055-4. 228 pp. LCCN QB509 .L66 2012.

Longair:2016:MEL

[Lon16] Malcolm Longair. Maxwell’s Enduring Legacy. Cambridge University Press, Cambridge, UK, 2016. ISBN 1-107-08369-

9. xxi + 664 pp. LCCN

                                        ????  URL http://www.
                                        cambridge.org/us/academic/
                                        subjects/physics/general-
                                        and-classical-physics/maxwells-
                                        enduring-legacy-scientific-
                                        history-cavendish-laboratory.

Lorin:1975:SSS

[Lor75] Harold Lorin. Sorting and Sort Systems. The Systems Programming Series. Addison-Wes-ley, Reading, MA, USA, 1975. ISBN 0-201-14453-0. xv + 373 pp. LCCN QA76.5 .L577 1975.

Lorenzo:2018:ASB

[Lor18] Mark Jones Lorenzo. Adventures of a Statistician: The Biography of John W. Tukey. SE Books, 5307 West Tyson Street, Philadelphia, PA 19107, USA, 2018. ISBN 1-72201-358-3 (paperback). 340 pp. LCCN QA276.157 .T854 2018.

Lorenzo:2019:AMH

[Lor19] Mark Jones Lorenzo. Abstracting Away the Machine: The History of the Fortran Programming Language (FORmula TRANslation). SE Books, 5307 West Tyson Street, Philadelphia, PA 19107, USA, 2019. ISBN 1-08-239594-3. 325 pp. LCCN QA76.73.F25 BT22 2019.

Low:1960:PRS

[Low60] William Low. Paramagnetic Resonance in Solids, volume 2 of Solid state physics. Supplement. Academic Press, New York, NY, USA, 1960. viii + 212 pp. LCCN QC176 .L6 1960.

Lowdin:1964:MOC

[LP64] Per-Olov Löwdin and Bernard Pullman, editors. Molecular Orbitals in Chemistry, Physics, and Biology: A Tribute to R. S. Mulliken. Academic Press, New York, NY, USA, 1964. xiii + 578 pp. LCCN QD461 .L872 1964.

Lewis:1961:T

[LR61] Gilbert Newton Lewis and Merle Randall. Thermodynamics. Mc-Graw-Hill, New York, NY, USA, second edition, 1961. xii + 723 pp. LCCN QC311 .L4 1961. Revised by Kenneth S. Pitzer and Leo Brewer.

Libes:1989:LU

[LR89] Don Libes and Sandy Ressler.

Life with UNIX. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-536657-7. xx + 346 pp. LCCN QA76.76.O63 L52 1989. URL http:/

/minnie.tuhs.org/Archive/
Documentation/Books/.

Lawson:sftran-conversion

[LS78] C. L. Lawson and W. V. Snyder. Conversion of SFTRAN programs to SFTRAN 3. Technical Report 1846-99, Jet Propulsion Laboratory, Pasadena, CA, USA, December 1978.

Lesk:lex

[LS79] Michael E. Lesk and Eric Schmidt. Lex: A lexical analyzer generator. In UNIX Programmer’s Manual, volume 2, pages 388–400. Holt, Rinehart, and Winston, New York, NY, USA, 1979. AT&T Bell Laboratories Technical Report in 1975.

Lyche:1989:MMC

[LS89] Tom Lyche and Larry L. Schumaker, editors. Mathematical Methods in Computer Aided Geometric Design. Academic Press, New York, NY, USA, 1989. ISBN 0-12-460515-X. LCCN QA448.D38 M381 1989.

Ludgate:1909:PAM

[Lud09] P. E. Ludgate. On a proposed analytical machine. Scientific proceedings of the Royal Dublin Society, 12(9):77–91, 1909. Reprinted in [Ran82, §2.4].

Luke:1969:SFTa

[Luk69a] Yudell L. Luke. The Special Functions and Their Approximations, volume I of Mathematics in Science and Engineering, Volume 53-I, Editor: Richard Bellman. Academic Press, New York, NY, USA, 1969. ISBN 0-12-459901-X, 0-08-095560-6 (e-book). xx + 349 pp. LCCN QA351 .L94 1969. URL http://www.

sciencedirect.com/science/
book/9780124599017.

Luke:1969:SFTb

[Luk69b] Yudell L. Luke. The Special Functions and Their Approximations, volume II of Mathematics in Science and Engineering, Volume 53-II, Editor: Richard Bellman. Academic Press, New York, NY, USA, 1969. ISBN 0-12-459902-8. xx + 485 pp. LCCN QA351

.L797.  URL http://www.
sciencedirect.com/science/
book/9780124599024.

Luke:1975:MFT

[Luk75] Yudell L. Luke. Mathematical Functions and Their Approximations. Academic Press, New York, NY, USA, 1975. ISBN 0-12-459950-8, 1-4832-6245-6 (ebook). xvii + 568 pp. LCCN QA55 .L96 1975. URL https:

//shop.elsevier.com/books/
mathematical-functions-and-
their-approximations/luke/
978-0-12-459950-5.

Luke:1977:ACM

[Luk77] Yudell L. Luke. Algorithms for the Computation of Mathematical Functions. Academic Press, New York, NY, USA, 1977. ISBN 0-12-459940-0. xiii + 284 pp. LCCN QA351 .L7961.

Lunde:1993:UJI

[Lun93]  Ken Lunde.  Understanding

Japanese Information Processing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-043-0. xxxii + 435 pp. LCCN PL524.5.L86 1993. US$29.95. See [Rei88b].

Lunde:1999:CIP

[Lun99] Ken Lunde. CJKV Information Processing: Chinese, Japanese, Korean & Vietnamese Computing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-224-7. xxiv + 1101 pp. LCCN PL1074.5 .L85 1999. US$64.95. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0715/00700462-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/fy1008/
                                        00700462-b.html;   http:/
                                        /www.oreilly.com/catalog/
                                        cjkvinfo/.

Lord:1984:MDS

[LW84] E. A. Lord and C. B. Wilson. The Mathematical Description of Shape and Form. Ellis Horwood Series in Mathematics and its Applications, Editor: G. M. Bell. Ellis Horwood, New York, NY, USA, 1984. ISBN 0-85312-722-0, 0-85312-726-3, 0-470-20043-X (paperback). 260 pp. LCCN QA501 .L86 1984.

Lindholm:1997:JVM

[LY97] Tim Lindholm and Frank Yellin. The Java Virtual Machine Specification. The Java Series. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-63452-X. xvi + 475 pp. LCCN QA76.73.J38L56 1997. US$36.53. URL http:

//www.aw.com/cp/javaseries.
html.

Lindholm:1999:JVM

[LY99] Tim Lindholm and Frank Yellin.

The Java Virtual Machine Specification. Addison-Wesley, Reading, MA, USA, second edition, 1999. ISBN 0-201-43294-3 (paperback). xv + 473 pp. LCCN QA76.73.J38L56 1999. US$42.95.

Morse:80286

[MA86] Stephen P. Morse and Douglas J.

Albert. The 80286 Architecture. Wiley, New York, NY, USA, 1986. ISBN 0-471-83185-9. xiii + 279 pp. LCCN QA76.8.I2927 M67 1986. Morse is the chief architect of the Intel 8086. See also [Int85a].

Morris:1991:RID

[MA91] Robert A. Morris and Jacques André, editors. Raster Imaging and Digital Topography II: Papers from the second RIDT meeting, held in Boston, Oct. 14–16, 1991. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-41764-3. LCCN Z253.3 .R39 1991.

Ma:2009:FPL

[Ma09] M. Y. (Matthew Y.) Ma. Fundamentals of Patenting and Licensing for Scientists and Engineers. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2009. ISBN 981-283-420-6. xxiii + 265 pp. LCCN T339

                                        .M3 2009. URL http://www.
                                        spectrum.ieee.org/at-work/
                                        innovation/fundamentals-of-
                                        patenting-and-licensing.

MacInnes:1961:PE

[Mac61]  Duncan A. MacInnes.  The

Principles of Electrochemistry. Dover, New York, NY, USA, 1961. 478 pp. LCCN QD553 .M25 1961. Reprint of the original 1939 edition.

Macomber:1976:DST

[Mac76] James D. Macomber. The Dynamics of Spectroscopic Transitions: Illustrated by Magnetic Resonance and Laser Effects. Wiley, New York, NY, USA, 1976. ISBN 0-471-56300-5. xxiv + 332 pp. LCCN QC454.M8M3.

Mackenzie:1980:CCS

[Mac80] Charles E. Mackenzie. Coded Character Sets: History and Development. The Systems Programming Series. Addison-Wes-ley, Reading, MA, USA, 1980. ISBN 0-201-14460-3. xxi + 513 pp. LCCN QA268 .M27 1980. US$24.95.

Macrae:1992:JNSa

[Mac92] Norman Macrae. John von Neumann: The Scientific Genius Who Pioneered the Modern Computer, Game Theory, Nuclear Deterrence, and Much More. Pantheon Books, New York, NY, USA, 1992. ISBN 0-679-41308-1. x + 405 pp. LCCN QA29.V66 M33 1992. US$25.00 (US$31.50 Can.).

MacMillan:2003:PSM

[Mac03]  Margaret MacMillan.  Paris

1919: six months that changed the world. Random House, New York, NY, USA, 2003. ISBN 0-375-50826-0, 0-375-76052-0. xxxi + 570 pp. LCCN D644 .M32 2002.

Macdougall:2011:WGM

[Mac11] Doug Macdougall. Why Geology Matters: Decoding the Past, Anticipating the Future. University of California Press, Berkeley, CA, USA, 2011. ISBN 0-520-26642-0 (hardcover), 0-520-27271-4 (paperback), 0-520-94892-0 (e-book). xv + 285 pp. LCCN QE28.3 M334 2011.

Macintyre:2018:STG

[Mac18] Ben Macintyre. The Spy and the Traitor: the Greatest Espionage Story of the Cold War. Crown Publishing Group, New York, NY, USA, 2018. ISBN 1-101-90419-4 (hardcover), 1-101-90420-8 (e-book). LCCN UB271.R92 G675 2018.

Magueijo:2009:BDE

[Mag09] João Magueijo. A brilliant darkness: the extraordinary life and disappearance of Ettore Majorana, the troubled genius of the nuclear age. Basic Books, New York, NY, USA, 2009. ISBN 0-465-00903-4. xxi + 280 pp. LCCN QC774.M34 M35 2009.

Mahaffey:2009:AAN

[Mah09] James A. Mahaffey. Atomic awakening: a new look at the history and future of nuclear power. Pegasus Books, New York, NY, USA, 2009. ISBN 1-60598-040-4. xxiii + 344 pp. LCCN TK9145 .M34 2009; TK9145.M34.

Maiman:2018:LIM

[Mai18] Theodore H. Maiman. The Laser Inventor: Memoirs of Theodore H. Maiman. Springer Biographies. Springer International Publishing, Cham, Switzerland, 2018. ISBN 3-319-61939-X (hardcover)m 3-319-61940-3 (ebook). ISSN 2365-0613 (print), 2365-0621 (electronic). xxxix + 312 + 89 pp. LCCN QC16 .M356 2018.

Malcolm:j-CACM-15-11-949

[Mal72] Michael A. Malcolm. Algorithms to reveal properties of floatingpoint arithmetic. Communications of the Association for Computing Machinery, 15(11):949– 951, November 1972.

Malone:1995:MB

[Mal95] Michael S. Malone. The Microprocessor: a Biography. TE-LOS division of Springer-Verlag, Santa Clara, CA, USA and New York, NY, USA, 1995. ISBN 0-387-94145-2, 0-387-94342-0. xiv + 333 pp. LCCN TK7895.M5 M35 1995. US$49.95.

Malley:2011:RHM

[Mal11] Marjorie Caroline Malley. Radioactivity: a history of a mysterious science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-976641-X (hardcover). xxi + 267 pp. LCCN QC794.6.R3 M35 2011.

Mallory:2013:OI

[Mal13] J. P. Mallory. Origins of the Irish. Thames and Hudson Ltd., London, UK, 2013. ISBN 0-500-29330-9 (paperback), 0-500-77137-5 (e-book). 328 pp. LCCN DA910 .M25 2013.

Mandelbrot:1983:FGN

[Man83] Benoît B. Mandelbrot. The Fractal Geometry of Nature. W. H. Freeman and Company, New York, NY, USA, 1983. ISBN 0-7167-1186-9. 468 + 1 pp. LCCN QA447 .M271 1983. Updated and augment edition of [Man83], but with same ISBN.

Manber:1994:FSF

[Man94] Udi Manber. Finding similar files in a large file system. In USENIX-WINTER’94 [USE94], pages 1–10. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

Maor:1991:IBC

[Mao91] Eli Maor. To Infinity and Beyond: a Cultural History of the Infinite. Princeton paperbacks. Princeton University Press, Princeton, NJ, USA, 1991. ISBN 0-691-02511-8 (paperback). xvi + 284 pp. LCCN QA9 .M316 1991. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;

http://www.loc.gov/catdir/
description/prin021/91004396.
html;   http://www.loc.gov/
catdir/samples/prin031/91004396.
html;   http://www.loc.gov/
catdir/toc/prin051/91004396.
html.

Maor:1994:SN

[Mao94] Eli Maor. e: The Story of a Number. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03390-0. xiv + 223 pp. LCCN QA247.5.M33

1994. URL http://www-gap.
dcs.st-and.ac.uk/~history/
HistTopics/e.html. This book

chronicles the story of the ultimate version number of - .

Mao:2004:MCT

[Mao04] Wenbo Mao. Modern Cryptography: Theory and Practice. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-066943-1 (hardcover). xxxviii + 707 pp. LCCN QA76.9.A25 M36 2003. US$54.99.

Maor:2007:PTY

[Mao07] Eli Maor. The Pythagorean Theorem: a 4,000-year History. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-12526-

0. xvi + 259 pp. LCCN QA460.P8 M36 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0704/2006050969-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0726/
2006050969-t.html;   http:/

/www.loc.gov/catdir/enhancements/
fy0734/2006050969-b.html.

MAPLETECH

[MAP] Tony Scott, editor. The Maple Technical Newsletter. ISSN 1061-5733. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland. Published twice annually.

Margulis:1990:IMA

[Mar90] Neal Margulis. i860 Microprocessor Architecture. McGraw-Hill, New York, NY, USA, 1990. ISBN 0-07-881645-9. xxiii + 631 pp. LCCN QA76.5.M37 1990.

Markstein:2000:IEF

[Mar00] Peter Markstein. IA-64 and Elementary Functions: Speed and Precision. Hewlett–Packard professional books. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-018348-

2. xix + 298 pp. LCCN QA76.9.A73 M365 2000. URL

http://www.markstein.org/.

Marciano:2014:WHM

[Mar14a] John Bemelmans Marciano.

Whatever happened to the metric system?: how America kept its feet. Bloomsbury, New York, NY, USA, paperback edition, 2014. ISBN 1-60819-475-2 (hardcover), 1-60819-940-1 (paperback), 1-60819-941-X (ePub). 310 + 16 pp. LCCN QC92.U54 M37 2015; QC92.U6 M37 2015.

Martel:2014:MCW

[Mar14b] Gordon Martel. The month that changed the world: July 1914. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-966538-9 (hardcover). xxii + 484 pp. LCCN D511 .M268 2014. URL http://www.

h-net.org/reviews/showrev.
php?id=42971;   http://www.
loc.gov/catdir/enhancements/
fy1604/2013953482-b.html;
http://www.loc.gov/catdir/
enhancements/fy1604/2013953482-
d.html;   http://www.loc.
gov/catdir/enhancements/fy1604/
2013953482-t.html.

Martin:2015:DBD

[Mar15] Anthony J. Martin. Dinosaurs Without Bones: Dinosaur Lives Revealed by Their Trace Fossils. Pegasus Books, New York, NY, USA, 2015. ISBN 1-60598-703-4 (paperback). 460 + 24 pp. LCCN QE861.4 .M367 2015.

Matzinger:2019:HDS

[Mat19a] Claus Matzinger. Hands-on Data Structures and Algorithms with Rust: Learn Programming Techniques to Build Effective, Maintainable, and Readable Code in Rust 2018. Packt Publishing, Birmingham, UK, 2019. ISBN 1-78899-149-4. vii + 298 pp. LCCN Q76.73.R87.

Matzinger:2019:RPC

[Mat19b] Claus Matzinger. Rust Programming Cookbook: Explore the Latest Features of Rust 2018 for Building Fast and Secure Apps. Packt Publishing, Birmingham, UK, 2019. ISBN 1-78953-066-0, 1-78953-174-8 (PDF e-book). 434 pp. LCCN QA76.73.R87.

Mauchly:1942:UHS

[Mau42] J. W. Mauchly. The use of high speed vacuum tube devices for calculating. Privately circulated memorandum, Moore School of Electrical Engineering, University of Pennsylvania, Philadelphia, PA, USA, August 1942. Reprinted in [Ran82, §7.6].

Mauchly:1947:PPE

[Mau47] J. W. Mauchly. Preparation of problems for EDVAC-type machines. In Proceedings of a Symposium on Large Scale Digital Calculating Machinery, 7– 10 January 1947. Harvard University Press, Cambridge, MA, USA, 1947. Reprinted in Annals of the Computation Laboratory of Harvard University, 16, 203– 207 (1948). Reprinted in [Ran82, §8.2].

Mayoh:1982:PSA

[May82] Brian Mayoh. Problem Solving with Ada. Wiley, New York, NY, USA, 1982. ISBN 0-471-10025-0. viii + 233 pp. LCCN QA76.8.A15 M38 1982. US$59.00.

Mazza:1997:OPI

[Maz97] Debora Mazza, editor. Oxford paperback Italian dictionary. The Oxford Italian dictionary: Italian–English, English–Italian = italiano-inglese, inglese-italiano. Berkley Books, New York, NY, USA, 1997. ISBN 0-425-16012-2 (paperback). viii + 264 pp. LCCN PC1640 .O94 1997. US$5.99.

Mazur:2010:WLG

[Maz10] Joseph Mazur. What’s luck got to do with it?: the history, mathematics, and psychology behind the gambler’s illusion. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-13890-7 (hardcover). xvii + 277 pp. LCCN QA271 .M39 2010.

Myers:1988:MA

[MB88] Glenford J. Myers and David L. Budde. The 80960 Microprocessor Architecture. Wiley-Intersci-ence, New York, NY, USA, 1988. ISBN 0-471-61857-8. xiii + 255 pp. LCCN QA76.8.I29284 M941 1988.

Mason:1990:LY

[MB90] Tony Mason and Doug Brown. lex & yacc. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1990. ISBN 0-937175-49-8. xviii + 216 pp. LCCN QA76.76.O63 M37 1990.

Miller:2000:ASP

[MB00] Russ Miller and Laurence Boxer. Algorithms Sequential and Parallel: a Unified Approach. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-086373-4. xii + 330 pp. LCCN QA76.9.A43 M55 2000. URL http://

www.prenhall.com/books/esm_
0130863734.html.

Muller:2010:HFP

[MBdD^{+}10] Jean-Michel Muller, Nicolas Brisebarre, Florent de Dinechin, Claude-Pierre Jeannerod, Vincent Lefèvre, Guillaume Melquiond, Nathalie Revol, Damien Stehlé, and Serge Torres. Handbook of Floating-Point Arithmetic. Birkhäuser Boston Inc., Cambridge, MA, USA, 2010. ISBN 0-8176-4704-X. xxiii + 572 pp. LCCN QA76.9.C62 H36 2010. US$90 (est.).

McKusick:1996:DIO

[MBKQ96] Marshall Kirk McKusick, Keith Bostic, Michael J. Karels, and John S. Quarterman. The Design and Implementation of the 4.4BSD Operating System. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-54979-4. xxvi + 580 pp. LCCN QA76.76.O63D4743 1996. US$44.95. URL http:/

/wiki.tuhs.org/doku.php?id=
publications:4.4bsd_design.

McCarthy:1998:IS

[McC98] Linda McCarthy. Intranet Security. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, 1998. ISBN 0-13-894759-0. xxv + 260 pp. LCCN TK5105.59 .M352 1998. US$29.95, CDN$41.95.

McCartney:2001:ETT

[McC01] Scott McCartney. ENIAC: the Triumphs and Tragedies of the World’s First Computer. Berkley Books, New York, NY, USA, 2001. ISBN 0-425-17644-4. vii + 262 pp. LCCN QA76.5.

McGrath:1998:PSS

[McG98] Sean McGrath. ParseMe.1st: SGML for Software Developers. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, January 1998. ISBN 0-13-488967-3. xxiii + 341 pp. LCCN QA76.76.H94M388 1998. US$33.75.

McGrayne:2011:TWH

[McG11] Sharon Bertsch McGrayne. The theory that would not die: how Bayes’ rule cracked the Enigma code, hunted down Russian submarines, and emerged triumphant from two centuries of controversy. Yale University Press, New Haven, CT, USA, 2011. ISBN 0-300-16969-8. xiii + 320 pp. LCCN QA279.5 .M415 2011.

McKechnie:1979:WNT

[McK79] Jean L. McKechnie, editor. Webster’s new twentieth century dictionary of the English language, unabridged: Based upon the broad foundations laid down by Noah Webster. W. Collins, Cleveland, OH, USA, 1979. ISBN 0-529-04852-3. 2129 + xiv + 160 pp. LCCN PE1625 W375 1979.

McLeish:1991:SN

[McL91] John McLeish. The Story of Numbers: How mathematics has shaped civilization. Fawcett Columbine, New York, NY, USA, 1991. ISBN 0-449-90938-7 (paperback). 266 + 8 pp. LCCN QA21.M38 1991. US$14.95.

McManus:color-printing

[McM88] Paul A. McMannis. Color printing. In Robert C. Durbeck and Sol Sherr, editors, Output Hardcopy Devices, chapter 17, pages 441–462. Academic Press, New York, NY, USA, 1988.

McPhee:1974:CBE

[McP74] John A. McPhee. The Curve of Binding Energy. Farrar, Strauss, and Giroux, New York, NY, USA, 1974. ISBN 0-374-13373-5, 0-374-51598-0 (paperback). 232 pp. LCCN UF767 .M215 1974. URL

http://www.loc.gov/catdir/
bios/hol059/74001226.html;
http://www.loc.gov/catdir/
description/hol054/74001226.
html.

McPhee:1998:AFW

[McP98] John A. McPhee. Annals of the Former World. Farrar, Straus and Giroux, New York, NY, USA, 1998. ISBN 0-374-10520-0, 0-374-51873-4 (paperback). 696 pp. LCCN QE77 .M38 1998. US$35.00.

McWhorter:2008:OMB

[McW08] John McWhorter. Our magnificent bastard tongue: the untold history of English. Gotham Books, New York, NY, USA, 2008. ISBN 1-59240-395-6. xxiii + 230 pp. LCCN PE1075 .M597 2008.

Mecklenburg:2005:MPG

[Mec05] Robert Mecklenburg. Managing Projects with GNU Make. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2005. ISBN 0-596-00610-1 (paperback). xviii + 280 pp. LCCN QA76.76.U84 O73 2005. US$29.95.

Merz:1991:TBP

[Mer91] Thomas Merz. Terminal Buch, PostScript Fonts und Programmiertechnik. R. Oldenbourg, München, Germany, 1991. ISBN 3-486-21674-0. 213 pp. LCCN ???? DM 78,00.

Merz:1997:PAP

[Mer97a] Thomas Merz. PostScript and Acrobat/PDF: applications, troubleshooting, and cross-platform publishing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540w 0854-0. xiii + 418 pp. LCCN QA76.73.P67M4713 1997. US$69.50.

Merz:1997:AIW

[Mer97b] Thomas Merz. Mit Acrobat ins World Wide Web: Effiziente Erstellung von PDF-Dateien und ihre Einbindung ins Web. dpunkt-Verlag, Heidelberg, Germany, 1997. ISBN 3-9804943-1-4. x + 225 pp. LCCN ???? DM 69,00; ATS 504,00; CHF 61,00. Includes CD-ROM. English translation available as [Mer98].

Merz:1998:WPA

[Mer98] Thomas Merz. Web Publishing with Acrobat/PDF. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 3-540-63762-1. xi + 234 pp. LCCN TK5105.888.M47 1998.

URL http://www.pdflib.com/
pdfmark/index.html. Includes

CD-ROM. Revised and extended English translation of the original German edition, [Mer97b].

Messiah:1999:QM

[Mes99] Albert Messiah. Quantum Mechanics. Dover, New York, NY, USA, 1999. ISBN 0-486-40924-4 (paperback). xxi + 1136 pp. LCCN QC174.12 .M4813 1999. URL http://www.

loc.gov/catdir/description/
dover032/99055362.html;
http://www.loc.gov/catdir/
toc/dover031/99055362.html.

Metcalf:1982:FO

[Met82] Michael Metcalf. Fortran Optimization, volume 17 of A. P. I. C. Studies in Data Processing. Academic Press, New York, NY, USA, 1982. ISBN 0-12-492480-8. xii + 242 pp. LCCN QA76.73.F25 M48 1982.

Mittelbach:2004:LC

[MGB^{+}04] Frank Mittelbach, Michel Goossens, Johannes Braams, David Carlisle, Chris Rowley, Christine Detig, and Joachim Schrod. The LaTEX Companion. Tools and Techniques for Computer Typesetting. Addison-Wesley, Reading, MA, USA, second edition, 2004. ISBN 0-201-36299-6 (paperback), 0-321-51443-2 (hardcover). xxvii + 1090 pp. LCCN Z253.4.L38 G66 2004. US$59.99, CAN$86.99.

More:minpack

[MGH80] Jorge J. Moré, Burton S. Garbow, and Kenneth E. Hillstrom. User guide for MINPACK-1. Technical Report ANL-80-74, Argonne National Laboratory, Argonne, IL, USA, August 1980.

Monagan:1996:MVP

[MGH^{+}96] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, and S. Vorkoetter. Maple V Programming Guide. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94537-7, 0-387-94576-8. xii + 379 pp. LCCN QA76.95.M363 199. US$34.00. With the editorial assistance of J. S. Devitt, M. L. Hansen, D. Redfern, and K. M. Rickard.

Monagan:2003:MAP

[MGH^{+}03a] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M. Vorkoetter, J. McCarron, and P. DeMarco. Maple 9 Advanced Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2003. ISBN 1-894511-44-1. x + 443 pp. LCCN QA76.95 M364 2003.

Monagan:2003:MIP

[MGH^{+}03b] M. B. Monagan, K. O. Geddes, K. M. Heal, G. Labahn, S. M.

Vorkoetter, J. McCarron, and P. DeMarco. Maple 9 Introductory Programming Guide. Maplesoft, Waterloo, ON, Canada N2L 6C2, 2003. ISBN 1-894511-43-3. x + 388 pp. LCCN QA76.95 .M363 2003.

Mersini-Houghton:2022:BBBc

[MH22] Laura Mersini-Houghton. Before the Big Bang: the Origin of the Universe and What Lies Beyond. Mariner Books, Boston, MA, USA, 2022. ISBN 1-328-55711-1 (hardcover). xxi + 216 pp. LCCN QB981 .M477 2022.

Miano:1999:CIF

[Mia99] John Miano. The Programmer’s Guide to Compressed Image Files: JPEG, PNG, GIF, XBM, BMP. SIGGRAPH series. Addison-Wesley Longman, Harlow, Essex CM20 2JE, England, 1999. ISBN 0-201-60443-4 (paperback), 0-201-61657-2 (CD-ROM). xi + 264 pp. LCCN TA1637.M53 1999. US$44.95.

Michelin:1989:MAG

[Mic89] Michelin. Michelin Allemagne Germany (Michelin Maps). Michelin Travel Publications, Paris, France, 1989. ISBN 2-06-700984-2. 1 pp. LCCN G6081.P2 1989 .P6.

Michelin:2005:MET

[Mic05]  Michelin.  Michelin Europe

Tourist & Motoring Atlas. Atlas Seriews. Michelin Travel Publications, Paris, France, 2005. ISBN 2-06-711224-4 (paperback). 228 pp. LCCN G1797.21.P2 E876 2005.

Microsoft:2012:MMS

[Mic12] Microsoft. Microsoft Manual of Style: Your Everyday Guide to Usage, Terminology, and Style for Professional Technical Communications. Microsoft Press, Bellevue, WA, USA, fourth edition, 2012. ISBN 0-7356-4871-9 (paperback). xxiv + 438 pp. LCCN T11 .M467 2012.

Miller:1984:ENS

[Mil84] Webb Miller. The Engineering of Numerical Software. Prentice-Hall Series in Computational Mathematics, Cleve Moler, Advisor. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-279043-2. viii + 167 pp. LCCN QA297.M527 1984.

Milenkovic:1987:OS

[Mil87] Milan Milenković. Operating Systems. McGraw-Hill, New York, NY, USA, 1987. ISBN 0-07-041920-5. xix + 568 pp. LCCN QA76.76.O63 M53 1987. US$38.95.

Miller:1990:OLU

[Mil90] John David Miller. An OPEN LOOK at UNIX. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1990. ISBN 1-55851-057-5. 482 pp. LCCN QA76.76.U84 M55 1990.

Miller:1992:TTI

[Mil92] Mark A. Miller. Troubleshooting TCP/IP: Analyzing the Protocols of the Internet. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1992. ISBN 1-55851-268-3, 3-88229-025-0, 0-13-953167-X. 588 pp. LCCN TK5105.5 .M52 1992. US$44.95.

Milton:2002:TC

[Mil02] Graeme W. Milton. The Theory of Composites, volume 6 of Cambridge Monographs on Applied and Computational Mathematics. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-78125-6. xxviii + 719 pp. LCCN TA418.9.C6 M58 2001.

US$80. URL https://www.
math.utah.edu/books/tcbook.

Series editors: P. G. Ciarlet, A. Iserles, Robert V. Kohn, and M. H. Wright.

Miller:2010:JPP

[Mil10] Arthur I. Miller. 137: Jung, Pauli, and the Pursuit of a Scientific Obsession. W. W. Norton & Co., New York, NY, USA, 2010. ISBN 0-393-33864-9. xxiii + 336 pp. LCCN QC16.P37 M55 2010.

Miller:2015:BLT

[Mil15] Steven J. Miller, editor. Benford’s Law: theory and applications. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-14761-2 (hardcover), 1-4008-6659-6 (e-book). xxvi + 438 pp. LCCN QA273.6

.B46 2015. URL http://press.
princeton.edu/titles/10527.
html.

Milner:2016:PHG

[Mil16a] Greg Milner. Pinpoint: how GPS is changing technology, culture, and our minds. W. W. Norton & Co., New York, NY, USA, 2016. ISBN 0-393-08912-6 (hardcover). xx + 316 pp. LCCN G109.5 .M55 2016.

Milton:2016:ETC

[Mil16b] Graeme W. Milton, editor. Extending the Theory of Composites to Other Areas of Science. Milton–Patton Publishers, P.O. Box 581077, Salt Lake City, UT 85148, USA, 2016. ISBN 1-4835-6919-5 (print), 1-4835-6920-9 (ebook). xx + 422 pp. LCCN

                                        ???? URL https://www.math.
                                        utah.edu/books/milton.

IBM:IRS90

[Mis90] Mamata Misra, editor. IBM RISC System/6000 Technology, publication SA23-2619-00. IBM Corporation, San Jose, CA, USA, 1990.

Mitchell:32B86

[Mit86] H. J. Mitchell. 32-Bit Microprocessors. McGraw-Hill, New York, NY, USA, 1986. ISBN 0-07-042585-X. 248 pp. LCCN QA76.5 .A135 1986. US$39.95.

Mithen:2006:SNO

[Mit06] Steven J. Mithen. The singing Neanderthals: the origins of music, language, mind, and body. Harvard University Press, Cambridge, MA, USA, 2006. ISBN 0-674-02192-4. ix + 374 pp. LCCN ML3800 .M73 2006. URL http://www.

loc.gov/catdir/toc/ecip062/
2005030187.html.

Mitchell:2011:CGT

[Mit11] Melanie Mitchell. Complexity:

a Guided Tour. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-979810-9 (paperback). xvi + 349 pp. LCCN Q175.32 C65 M58 2011.

Mitchell:2018:SME

[Mit18] Alanna Mitchell. The Spinning Magnet: the Electromagnetic Force That Created the Modern World — and Could Destroy It. Dutton, New York, NY, USa, 2018. ISBN 1-101-98516-X (hardcover), 1-101-98518-6 (ebook). ix + 323 pp. LCCN QC760 .M5425 2018.

Muller:2006:SVP

[MJW^{+}06] Al Muller, Andy Jones, David E.

Williams, Stephen Beaver, David A. Payne, Jeremy Pries, and David E. Hart. Scripting VMware Power Tools: Automating Virtual Infrastructure Administration. Syngress Publishing, Inc., Rockland, MA, USA, 2006. ISBN 1-59749-059-8. xviii + 398 pp. LCCN QA76.76.O63 S385 2006eb.

Musciano:1996:HDG

[MK96] Chuck Musciano and Bill Kennedy.

HTML: The Definitive Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1996. ISBN 1-56592-175-5. xx + 385 pp.

LCCN QA76.76.H94 M87 199. US$29.95, CDN$42.95. URL

                                        http://www.oreilly.com/catalog/
                                        9781565921757.

Mlodinow:2001:EWS

[Mlo01] Leonard Mlodinow. Euclid’s Window: the Story of Geometry from Parallel Lines to Hyperspace. Free Press, New York, NY, USA, 2001. ISBN 0-684-86523-8, 0-684-86524-6 (paperback), 0-7139-9634-X. xii + 306 pp. LCCN QA443.5.M56 2001.

Margenau:1956:MPC

[MM56] Henry Margenau and George Moseley Murphy. The Mathematics of Physics and Chemistry, volume 1. D. Van Nostrand, Princeton, NJ, USA, second edition, 1956. xiv + 604 pp. LCCN QA37 .M3818 1956.

Margenau:1964:MPC

[MM64] Henry Margenau and George Moseley Murphy, editors. The Mathematics of Physics and Chemistry, volume 2. D. Van Nostrand, Princeton, NJ, USA, 1964. v + 786 pp. LCCN QA37 .M3818 1956.

McGilton:1983:IUS

[MM83] Henry McGilton and Rachel Morgan. Introducing the UNIX System. McGraw-Hill, New York, NY, USA, 1983. ISBN 0-07-045001-3. xix + 556 pp. LCCN QA76.8.U65 M38 1983. US$18.95.

McGilton:1990:TTU

[MM90] Henry McGilton and Mary Mc-Nabb. Typesetting Tables on the UNIX System. Trilithon Press, Los Altos, CA, USA, 1990. ISBN 0-9626289-0-5. xxii + 282 pp. LCCN Z253.4.U53 M33 1990. US$22.00.

Mauro:2001:SIC

[MM01] Jim Mauro and Richard Mc-Dougall. Solaris Internals: Core Kernel Architecture. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-022496-0. xli + 657 pp. LCCN QA76.76.O63 M37195 2001. URL http:/

/www.sun.com/books/catalog/
mauro/index.html.

McDougall:2007:SIS

[MM07] Richard McDougall and Jim Mauro. Solaris internals: Solaris 10 and OpenSolaris kernel architecture. Sun Microsystems Press/Prentice Hall, Upper Saddle River, NJ, USA, second edition, 2007. ISBN 0-13-148209-2 (hardcover). xlvi + 1020 pp. LCCN QA76.76.O63 M37195 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0613/2006015114.html.

McDougall:2007:SPT

[MMG07] Richard McDougall, Jim Mauro, and Brendan Gregg. Solaris performance and tools: DTrace and MDB techniques for Solaris 10 and OpenSolaris. Sun Microsystems Press/Prentice Hall, Upper Saddle River, NJ, 2007. ISBN 0-13-156819-1 (hardcover). xl + 444 pp. LCCN QA76.76.O63 M3957 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0615/2006020138.html.

McKusick:2005:DIF

[MNN05] Marshall Kirk McKusick and George V. Neville-Neil. The Design and Implementation of the FreeBSD Operating System. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-201-70245-2 (hardcover). xxviii + 683 pp. LCCN QA76.76.O63 M398745

2005.  URL http://www.
mckusick.com/FreeBSDbook.
html.

McKusick:2015:DIF

[MNNW15] Marshall Kirk McKusick, George V. Neville-Neil, and Robert N. M. Watson. The Design and Implementation of the FreeBSD Operating System. Addison-Wesley, Reading, MA, USA, second edition, 2015. ISBN 0-201-70245-2 (hardcover), 0-321-96897-2 (hardcover), 0-321-68005-7, 0-13-376180-0, 0-13-376183-5. xxx + 886 pp. LCCN QA76.774.F74 M35 2015.

MLA:2009:MHW

[Mod09] Modern Language Association. MLA Handbook for Writers of Research Papers. Modern Language Association of America, New York, NY, USA, seventh edition, 2009. ISBN 1-60329-024-9 (paperback) 1-60329-025-7 (large print). xxi + 292 pp. LCCN LB2369 .G53 2009; LB2369 .M25 2009.

Momjian:2001:PIC

[Mom01] Bruce Momjian. PostgreSQL: Introduction and Concepts. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-70331-9 (paperback). xxviii + 462 pp. LCCN QA76.9.D3 M647 2001. URL

http://www.addison-wesley.
de/projector/projector.asp?
page=bookdetails&isbn=3827318599;
http://www.ca.postgresql.
org/docs/awbook.html;   http:
//www.pearsoned.co.jp/washo/
db/wa_db27-j.html.

Monk:2013:ROL

[Mon13] Ray Monk. Robert Oppenheimer:

a Life Inside the Center. Doubleday, New York, NY, USA, 2013. ISBN 0-385-50407-1 (hardback). xvi + 825 + 32 pp. LCCN QC16.O62 M66 2013.

Moore:2016:RGT

[Moo16] Kate Moore. The Radium Girls:

They Paid with Their Lives, Their Final Fight Was for Justice. Simon and Schuster, New York, NY, USA, 2016. ISBN 1-4711-4757-6 (paperback), 1-4711-5387-8 (hardcover), 1-4711-5389-4 (e-book), 1-4926-4937-6 (e-book). xiii + 465 pp. LCCN HD6067.2.U6 M66 2016.

Morris:1983:CME

[Mor83] John Ll. Morris. Computational Methods in Elementary Numerical Analysis. Wiley, New York, NY, USA, 1983. ISBN 0-471-10419-1 (hardcover), 0-471-10420-5 (paperback). xii + 410 pp. LCCN 555 M67 1983. US$56.00 (hardcover), US$30.95 (paperback).

Morgan:1984:BAR

[Mor84] Christopher L. Morgan. Bluebook of Assembly Routines for the IBM PC and XT. Plume/ Waite, New York, NY, USA, 1984. ISBN 0-452-25498-1. x + 244 pp. LCCN QA76.8.I2594 M64 1984. A handbook collection of many assembly language routines for the IBM PC. The graphics algorithms, particular line-drawing, could be substantially speeded up. See [Laf84].

Mortenson:1985:GM

[Mor85] Michael E. Mortenson. Geometric Modeling. Wiley-Interscience, New York, NY, USA, 1985. ISBN 0-471-88279-8. xvi + 763 pp. LCCN QA447 .M62 1985.

MKS:yacc

[Mor87a] Mortice Kern Systems, Inc. MKS LEX & YACC, 1987. 35 King Street North, Waterloo, Ontario, Canada, Tel: (519) 884-2251.

MKS:awk

[Mor87b] Mortice Kern Systems, Inc. MKSAWK, 1987. 35 King Street North, Waterloo, Ontario, Canada, Tel: (519) 884-2251. See also [AKW88].

Morris:2001:TR

[Mor01] Edmund Morris. Theodore Rex. Modern Library, New York, NY, USA, 2001. ISBN 0-394-55509-0, 0-8129-6600-7 (paperback), 0-9654069-7-X, 1-58836-093-8 (ebook). x + 772 pp. LCCN E757 .M885 2001. URL http://

catdir.loc.gov/catdir/bios/
random053/2001019366.html;
 http://catdir.loc.gov/catdir/
description/random0411/2001019366.
html;   http://catdir.loc.
gov/catdir/samples/random044/
2001019366.html.

Morgan:2007:LHE

[Mor07] Michael Hamilton Morgan. Lost history: the enduring legacy of Muslim scientists, thinkers, and artists. National Geographic, Washington, DC, USA, 2007. ISBN 1-4262-0092-7 (hardcover). xviii + 301 pp. LCCN DS36.85 .M75 2007.

Moshier:1989:MPM

[Mos89] Stephen L. B. Moshier. Methods and Programs for Mathematical Functions. Ellis Horwood, New York, NY, USA, 1989. ISBN 0-7458-0289-3. vii + 415 pp. LCCN QA331 .M84 1989. US£48.00. URL http://

www.moshier.net/;   http://
www.netlib.org/cephes.

MC:68000

[Mot84]  Motorola.  MC68000 16/32-

Bit Microprocessor Programmer’s Reference Manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 1984. ISBN 0-13-541400-8. xiii + 218 pp. LCCN QA76.8.M6895 M25 1984.

MC:68020

[Mot85a] Motorola. MC68020 32-Bit Microprocessor User’s Manual. Motorola Corporation, Phoenix, AZ, USA, second edition, 1985. ISBN 0-13-566878-6. LCCN QA76.8.M6897 M37 1985.

Motorola:1985:MBM

[Mot85b] Motorola. MC68020 32-Bit Microprocessor User’s Manual. Motorola Corporation, Phoenix, AZ, USA, second edition, 1985. ISBN 0-13-566878-6. LCCN QA76.8.M6897 M37 1985.

MC:68881

[Mot85c] Motorola. MC68881 Floating-Point Coprocessor User’s Manual. Motorola Corporation, Phoenix, AZ, USA, second edition, 1985.

Motorola:1985:MFP

[Mot85d] Motorola. MC68881 Floating-Point Coprocessor User’s Manual. Motorola Corporation, Phoenix, AZ, USA, second edition, 1985.

Motorola:1989:MRM

[Mot89] Motorola. MC88100 RISC Microprocessor User’s Manual. Motorola Corporation, Phoenix, AZ, USA, second edition, 1989. ISBN 0-13-567090-X. LCCN QA76.8.M75 M3 1990.

Mui:1992:XWS

[MP92] Linda Mui and Eric Pearce. X Window System Administrator’s Guide for X11 Release 4 and Release 5. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 0-937175-83-8 (paperback), 1-56592-052-X (with CD ROM). xxiv + 346 pp. LCCN QA76.76.W56 D44 v.8 1992. US$29.95 (without CD ROM), US$59.95 (with CD ROM). URL http:

//www.oreilly.com/catalog/
9780937175835;   http://www.
oreilly.com/catalog/v8. Con-

tains CD ROM with X11R4 and X11R5 source code, plus compiled versions for Sun 3 SunOS 4.1.1, Sun 4 SunOS 4.1.1, DECstation ULTRIX 4.2, and IBM RS/6000 AIX 3.2.

Mueller:2000:CAC

[MP00] Silvia M. Mueller and Wolfgang J. Paul. Computer Architecture: Complexity and Correctness. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67481-0. xiii + 553 pp. LCCN QA76.9.A73 M845

2000. URL http://www-wjp.
cs.uni-sb.de/info/papers/#
books.

Metcalf:F8E87

[MR87] Michael Metcalf and John Reid.

Fortran 8x Explained. Clarendon Press, Oxford, UK, 1987. ISBN 0-19-853751-4 (hardcover), 0-19-853731-X (paperback). xiv + 262 pp. LCCN QA76.73.F26 M48 1987. See also [Ame87a, Ame90].

Metcalf:1990:FE

[MR90] Michael Metcalf and John Ker Reid. Fortran 90 Explained. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-853772-7 (paperback). xiv + 294 pp. LCCN QA76.73.F28 M48 1990. US$22.45. See [Ame92].

McKeown:1982:MC

[MRS82] G. P. McKeown and V. J. Rayward-Smith. Mathematics for Computing. Wiley, New York, NY, USA, 1982. ISBN 0-470-27268-6 (paperback). 428 pp. LCCN QA39.2 .M42 1982.

McCracken:1987:SCC

[MS87] Daniel D. McCracken and William I. Salmon. A Second Course in Computer Science with Modula-2. Wiley, New York, NY, USA, 1987. ISBN 0-471-63111-6. xiii + 464 + 10 pp. LCCN QA76.73.M63 M43 1987.

McCracken:1988:CES

[MS88] Daniel D. McCracken and William I. Salmon. Computing for Engineers and Scientists With Fortran 77. Wiley, New York, NY, USA, second edition, 1988. ISBN 0-471-62552-3 (paperback). xiii + 730 pp. LCCN TA345 .M3951 1988. US$59.75; US$65.95. URL http:

                                        //www.cbooks.com/sqlnut/SP/
                                        search/gtsumt?source=&isbn=
                                        0471625523.

Mazur:2015:PNR

[MS15] Barry Mazur and William A. Stein. Prime numbers and the Riemann hypothesis. Cambridge University Press, Cambridge, UK, 2015. ISBN 1-107-10192-1 (hardcover), 1-107-49943-7 (paperback). xi + 142 pp. LCCN QA246 .M49 2015.

May:1994:PAS

[MSSW94] Cathy May, Ed Silha, Rick Simpson, and Hank Warren, editors. The PowerPC Architecture: a Specification for a New Family of RISC Processors. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 1994. ISBN 1-55860-316-6. xxxi + 518 pp. LCCN QA76.8.P67 P68 1994. US$49.95.

Metropolis:1986:FS

[MSWA86] N. Metropolis, D. H. Sharp, W. J. Worlton, and K. R. Ames. Frontiers of Supercomputing. University of California Press, Berkeley, CA, USA, 1986. ISBN 0-520-05190-4. xxiv + 388 pp. LCCN QA76.5 .F76 1983.

Matsen:1986:SAA

[MT86] F. A. Matsen and T. Tajima, editors. Supercomputers: Algorithms, Architectures, and Scientific Computation. University of Texas Press, Austin, TX, USA, 1986. ISBN 0-292-70388-0. vi + 480 pp. LCCN QA76.5.S8945 1986.

Midtdal:1968:SSPa

[MTW68a] John Midtdal, Knut Thalberg, and Harald Wergeland, editors. Selected Scientific Papers of Egil A. Hylleraas, volume 1. NTH-Trykk, Trondheim, Norway, 1968. viii + 445 pp. LCCN Q143.H98 A25 1968.

Midtdal:1968:SSPb

[MTW68b] John Midtdal, Knut Thalberg, and Harald Wergeland, editors. Selected Scientific Papers of Egil A. Hylleraas, volume 2. NTH-Trykk, Trondheim, Norway, 1968. 526 pp. LCCN Q143.H98 A25 1968.

Muchnick:1997:ACD

[Muc97] Steven S. Muchnick. Advanced Compiler Design and Implementation. Morgan Kaufmann Publishers, 2929 Campus Drive, Suite 260, San Mateo, CA 94403, USA, 1997. ISBN 1-55860-320-4. xxix + 856 pp. LCCN QA76.76.C65M8 1997. US$89.95. URL http://www.

loc.gov/catdir/description/
els032/97013063.html;   http:
//www.loc.gov/catdir/toc/
els032/97013063.html.

Mueller:2000:URP

[Mue00] Scott Mueller. Upgrading and Repairing PCs. Que Corporation, Indianapolis, IN, USA, twelfth edition, 2000. ISBN 0-7897-2303-4. xvi + 1628 pp. LCCN TK7887 .M84 2001. US$49.99. Includes CD-ROM.

Muir:1960:TTD

[Mui60] Sir Thomas Muir. A Treatise on the Theory of Determinants. Dover, New York, NY, USA, 1960. 766 pp. LCCN QA191 .M85 1960. Corrected printing of the 1933 edition, revised and enlarged by William H. Metzler.

Muller:1997:EFA

[Mul97] Jean-Michel Muller. Elementary Functions: Algorithms and Implementation. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1997. ISBN 0-8176-3990-X. xv + 204 pp. LCCN QA331.M866

1997.  US$59.95.  URL

http://www.birkhauser.com/
cgi-win/ISBN/0-8176-3990-X;
http://www.ens-lyon.fr/~
jmmuller/book_functions.html.

Mundy:2017:CGU

[Mun17] Liza Mundy. Code Girls: the Untold Story of the American Women Code Breakers of World War II. Hachette Books, New York, NY, USA, 2017. ISBN 0-316-35253-5 (hardcover), 0-316-43989-4 (large print), 1-4789-2270-2 (audio book), 1-4789-2271-0 (audio download), 0-316-35255-1 (e-book). xiv + 416 pp. LCCN D810.C88 M86 2017.

Murray:1977:CWW

[Mur77] K. M. Elisabeth (Katherine Maud Elisabeth) Murray. Caught in the Web of Words: James A. H. Murray and the Oxford English Dictionary. Yale University Press, New Haven, CT, USA, 1977. ISBN 0-300-02131-3, 0-300-06310-5 (paperback). 386 pp. LCCN PE64.M8 M78. US$15.00.

Morgan:808682

[MW82] Christopher L. Morgan and Mitchell Waite. 8086/8088 16-Bit Microprocessor Primer. Mc-Graw-Hill, New York, NY, USA, 1982. ISBN 0-07-043109-4. ix + 355 pp. LCCN QA76.8.I292 M66 1982. A general book on the 8086 and 8088 with shorter treatments of the 8087, 8089, 80186, and 80286, and support chips. Nothing specific to the IBM PC.

Manber:1994:GTS

[MW94] Udi Manber and Sun Wu. GLIMPSE: a tool to search through entire file systems. In USENIX-WINTER’94 [USE94], pages 23–32. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

Myers:1979:AST

[Mye79] Glenford J. Myers. The Art of Software Testing. Business Data Processing, Editors: Richard G. Canning and J. Daniel Cougar. Wiley, New York, NY, USA, 1979. ISBN 0-471-04328-1. xi + 177 pp. LCCN QA76.6 .M888 1979.

Marshall:1997:WAS

[MZP97] I. N. Marshall, Danah Zohar, and F. David Peat. Who’s afraid of Schrödinger’s cat?: all the new science ideas you need to keep up with the new thinking. Morrow, New York, NY, USA, 1997. ISBN 0-688-11865-8. xxx + 402 pp. LCCN Q158.5 .M356 1997.

Nagar:1997:WNF

[Nag97] Rajeev Nagar. Windows NT File System Internals: a Developer’s Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997.

ISBN 1-56592-249-2. xviii + 774 pp. LCCN QA76.76.O63 N34 1997. US$49.95. URL

http://www.ora.com/catalog/
wininternals/;   http://www.
oreilly.com/catalog/9781565922495;
http://www.oreilly.com/catalog/
wininternals.

Nahin:1998:ITS

[Nah98] Paul J. Nahin. An Imaginary Tale: the Story of \sqrt{−1} [the Square Root of Minus One]. Princeton University Press, Princeton, NJ, USA, 1998. ISBN 0-691-02795-1. xvi + 257 pp. LCCN QA255 .N34 1998.

Nahin:2006:DEF

[Nah06] Paul J. Nahin. Dr. Euler’s fabulous formula: cures many mathematical ills. Princeton University Press, Princeton, NJ, USA, 2006. ISBN 0-691-11822-1 (hardcover). xx + 380 pp. LCCN QA255 .N339 2006. URL

http://www.loc.gov/catdir/
enhancements/fy0654/2005056550-
d.html.

Nakanishi:1962:IAS

[Nak62] Koji Nakanishi. Infrared Absorption Spectroscopy: Practical. Holden-Day, Inc. and Nankodo Company Limited, San Francisco, CA, USA and Tokyo, Japan, 1962. ix + 233 pp. LCCN QD95 .N36 1962.

Nash:1990:HSC

[Nas90] Stephen G. Nash, editor. A History of Scientific Computing. ACM Press history series. Addison-Wesley and ACM Press, Addison-Wesley and New York, NY 10036, USA, 1990. ISBN 0-201-50814-1. xix + 359 pp. LCCN QA76.17 .H59 1990.

NAE:1993:MT

[Nat93] National Academy of Engineers. Memorial tributes, volume 6. National Academy Press, Washington, DC, USA, 1993. ISBN 0-309-04847-8 (hardcover), 0-585-14955-0 (e-book). LCCN TA139

.N34 1993. URL https://nap.
nationalacademies.org/.

Neider:1993:OPG

[NDW93] Jackie Neider, Tom Davis, and Mason Woo. OpenGL Programming Guide: The Official Guide to Learning OpenGL, Release 1. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-63274-8. xiii + 516 pp. LCCN T385.N435 1993. US$34.95. The source code examples are

available at ftp://sgigate.
sgi.com/pub/opengl/opengl.

tar.Z. An HTML version of the book is available at http:

//arctic.eng.iastate.edu:
88/SGI_Developer/OpenGL_PG/
and http://fly.cc.fer.hr/~
unreal/index.html.

Nelson:1991:SPM

[Nel91]  Greg Nelson, editor.  Sys-

tems Programming with Modula-3. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-590464-1 (paperback). ix + 267 pp. LCCN QA76.66 .S87 1991. A description of the Modula 3 programming language by the committee that designed it, with an entertaining appendix on how various design decisions were made. Section 3.4 describes three floating-point interfaces that provide parameters of the underlying floating-point system, access primitives, and exception handling.

Ness:tv-guide

[Nes87] David Ness. The use of TEX in a commercial environment. TEXniques, 5:115–123, August 1987. Proceedings of the Eighth Annual Meeting of the TEX Users Group.

Newton:2007:CCH

[New07] Roger G. Newton. From clockwork to crapshoot: a history of physics. Belknap Press of Harvard University Press, Cambridge, MA and London, UK, 2007. ISBN 0-674-02337-4 (paperback). viii + 340 pp. LCCN QC7 .N398 2007.

Newitz:2014:SAR

[New14] Annalee Newitz. Scatter, Adapt, and Remember: How Humans Will Survive a Mass Extinction. Anchor Books, New York, NY, USA, 2014. ISBN 0-307-94942-7. 305 pp. LCCN GF86 .N485 2014.

Naiman:grey-scale

[NF87] Avi Naiman and Alain Fournier.

Rectangular convolution for fast filtering of characters. ACM SIGGRAPH—Computer Graphics, 21(4):233–242, July 1987. See [KU81].

Nguyen:2008:GG

[Ngu08] Hubert Nguyen, editor. GPU gems 3. Addison-Wesley, Reading, MA, USA, 2008. ISBN 0-321-51526-9. l + 942 pp. LCCN T385 .G6882 2008. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0720/2007023985.html.

Nielson:1997:SVO

[NHM97] Gregory M. Nielson, Hans Hagen, and Heinrich Müller. Scientific Visualization: Overviews, Methodologies, and Techniques. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7777-5. xiii + 577 pp. LCCN Q175.N438 1997. US$79.00. IEEE catalog number BP07777.

Nichols:1987:RTP

[Nic87] Major General Kenneth D. (Kenneth David) Nichols. The road to Trinity: a personal account of how America’s nuclear policies were made. Morrow, New York, NY, USA, 1987. ISBN 0-688-06910-X. 401 pp. LCCN QC774.N45 A3 1987.

Nichols:2003:ECD

[Nic03] Peter Nichols. Evolution’s Captain: the Dark Fate of the Man Who Sailed Charles Darwin Around the World. Harper-Collins College Publishers, New York, NY, USA, 2003. ISBN 0-06-008877-X (hardcover), 0-06-008878-8 (paperback). 336 pp. LCCN F2936 .N53 2003.

Nye:1990:XTI

[NO90a] Adrian Nye and Tim O’Reilly. X Toolkit Intrinsics Programming Manual, volume 4. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1990. ISBN 0-937175-34-X. xxxi + 543 pp. LCCN QA76.9.W56 N94 1990.

Nye:1990:XTIb

[NO90b] Adrian Nye and Tim O’Reilly.

X Toolkit Intrinsics Reference Manual: OSF/Motif 1.1 Edition for X11, Release 4, volume 4. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, December 1990. ISBN 0-937175-62-5 (vol. 4), 0-937175-66-8 (set). xxxi + 632 pp. LCCN QA76.9.W56N94

1990.  URL http://www.
oreilly.com/catalog/9780937175620.

Nakajima:1992:CDT

[NO92a] Heitaro Nakajima and Hiroshi Ogawa. Compact Disc Technology. Ohmsha, Ltd. and IOS Press, 3-1 Kanda Nishiki-cho, Chiyoda-ku, Tokyo 101, Japan and Postal Drawer 10558, Burke, VA 2209-0558, USA, 1992. ISBN 4-274-03347-3 (Ohmsha, Ltd), 90-5199-066-9 (IOS Press). x + 232 pp. LCCN TK7882.C56 N3613 1992. Translated by Charles Aschmann.

Nye:1992:XTI

[NO92b] Adrian Nye and Tim O’Reilly. X Toolkit Intrinsics Programming Manual: Motif Edition, volume 4M. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, August 1992. ISBN 1-56592-013-9. xxxviii + 674 pp. LCCN QA76.76.W56 N928 1992.

Nye:1993:XTI

[NO93] Adrian Nye and Tim O’Reilly. X Toolkit Intrinsics Programming Manual, volume 4. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, April 1993. ISBN 1-56592-003-1. xxxvi + 567 pp. LCCN QA76.76.W56 N945 1993. URL http:

                                        //www.oreilly.com/catalog/
                                        9781565920033.

Norton:1982:IIP

[Nor82] Peter Norton. Inside the IBM PC. Access to Advanced Features and Programming. Robert J. Brady Co., Bowie, MD 20715, USA, 1982. ISBN 0-89303-556-4. xi + 262 pp. LCCN QA76.8.I2594 N67 1983. An excellent treatment of the IBM PC.

Norton:1985:PGI

[Nor85] Peter Norton. Programmer’s Guide to the IBM PC. Microsoft Press, Bellevue, WA, USA, 1985. ISBN 0-914845-46-2. xii + 426 pp. LCCN QA76.8.I2594 N68 1985. US$19.95. An excellent treatment of the programming IBM PC. Strongly recommended for all serious users of the PC.

Norton:1994:PNC

[Nor94] Peter Norton. Peter Norton’s complete guide to DOS 6.22. SAMS Publishing, Indianapolis, IN, USA, 1994. ISBN 0-672-30614-X. xlvi + 1150 pp. LCCN QA76.76.O63N6779 1994. US$29.99 (CAN $39.99).

Newman:1973:PIC

[NS73] William M. Newman and Robert F. Sproull. Principles of Interactive Computer Graphics. McGraw-Hill Computer Science Series, Editors: Richard W. Hamming and Edward A. Feigenbaum. McGraw-Hill, New York, NY, USA, 1973. ISBN 0-07-046337-9. xxviii + 607 pp. LCCN T385 .N48.

Newman:1979:PIC

[NS79] William M. Newman and Robert F. Sproull. Principles of Interactive Computer Graphics. McGraw-Hill, New York, NY, USA, second edition, 1979. ISBN 0-07-046338-7. xvi + 541 pp. LCCN T385 .N48 1979.

Nisan:2021:ECS

[NS21] Noam Nisan and Shimon Schocken.

The Elements of Computing Systems Building a Modern Computer from First Principles. MIT Press, Cambridge, MA, USA, second edition, 2021. ISBN 0-262-53980-2 (paperback), 0-262-36100-0. xv + 325 pp. LCCN TK7888.3. URL https:

//mitpress.mit.edu/search-

                                        result-list/?keyword=978-0-
                                        262-53980-7.

Nemeth:2002:LAH

[NSH02] Evi Nemeth, Garth Snyder, and Trent Hein. Linux Administration Handbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-008466-2 (paperback). xxxi + 890 pp. LCCN QA76.76.O63 N448 2002. US$49.99.

Nemeth:2010:ULS

[NSHW10] Evi Nemeth, Garth Snyder, Trent R. Hein, and Ben Whaley. UNIX and Linux System Administration Handbook. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2010. ISBN 0-13-148005-7. xlvii + 1279 pp. LCCN QA76.76.O63 N45 2010.

Nemeth:1989:USA

[NSS89] Evi Nemeth, Garth Snyder, and Scott Seebass. UNIX System Administration Handbook. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-933441-6. xxx + 593 pp. LCCN QA76.76.O63 N45 1989.

Nemeth:2001:USA

[NSS^{+}01] Evi Nemeth, Garth Snyder, Scott Seebass, Trent R. Hein, Adam Boggs, Rob Braun, Ned McClain, Dan Crawl, Lynda McGinley, and Todd Miller. UNIX System Administration Handbook. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 2001. ISBN 0-13-020601-6. xxxv + 853 pp. LCCN QA76.76.O63 N45 2001. US$68.00. URL http://

www.phptr.com/ptrbooks/ptr_
0130206016.html.

Nocedal:2006:NO

[NW06] Jorge Nocedal and Stephen J.

Wright. Numerical Optimization. Springer series in operations research. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2006. ISBN 0-387-30303-0 (hardcover). xxii + 664 pp. LCCN QA402.5 .N62 2006. URL

http://www.loc.gov/catdir/
enhancements/fy0818/2006923897-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0818/
2006923897-t.html.

Nye:1988:XPM

[Nye88a] Adrian Nye. Xlib Programming Manual for Version 11, volume 1. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1988. ISBN 0-937175-26-9, 0-937175-89-7. xxxiii + 615 pp. LCCN QA76.76.W56 D44 v.1 1988. URL http:

//www.oreilly.com/catalog/
9780937175262.

Nye:1988:XRM

[Nye88b] Adrian Nye. Xlib Reference Manual for Version 11, volume 2. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1988. ISBN 0-937175-27-7. xiv + 701 pp. LCCN QA76.76.W56 D44 v.2

Nye:1990:XPR

[Nye90] Adrian Nye. X Protocol Reference Manual, volume 0. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1990. ISBN 0-937175-50-1. xv + 483 pp. LCCN QA76.76.W56 X215 1990.

Nye:1992:XPR

[Nye92]  Adrian Nye.  X Protocol

Reference Manual, volume 0. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 1992. ISBN 1-56592-008-2. LCCN QA76.76.W56 X215 1990. URL http:

                                        //www.oreilly.com/catalog/
                                        9781565920088.

Nystrom:2021:CI

[Nys21] Robert Nystrom. Crafting Interpreters. Genever Benning, ????, 2021. ISBN 0-9905829-3-0. vii + 626 pp. LCCN ????

Oakley:1949:AG

[Oak49] C. O. Oakley. Analytic Geometry. College Outline Series. Barnes and Noble, New York, NY, USA, 1949.

Oakley:1958:AGP

[Oak58] C. O. Oakley. Analytic Geometry Problems with Solutions. College Outline Series. Barnes and Noble, New York, NY, USA, 1958.

Othmer:1997:CSM

[OALD97] H. G. Othmer, F. R. Adler, M. A. Lewis, and J. C. Dallon, editors. Case Studies in Mathematical Modeling—Ecology, Physiology, and Cell Biology. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-574039-8. viii + 411 pp. LCCN QH541.15.M3C37 1997.

URL https://www.math.utah.
edu/books/csmmepcb/.

Ogletree:1999:URN

[Ogl99] Terry William Ogletree. Upgrading and Repairing Networks. Que Corporation, Indianapolis, IN, USA, 1999. ISBN 0-7897-2034-5, 0-585-34490-6 (e-book). xxi + 979 pp. LCCN TK5105.5 .O344 1999. US$49.99. Includes CD-ROM.

Ohanian:2008:EMH

[Oha08] Hans C. Ohanian. Einstein’s Mistakes: the Human Failings of Genius. W. W. Norton & Co., New York, NY, USA, 2008. ISBN 0-393-06293-7. xix + 394 pp. LCCN QC7 .O33 2008.

Orchin:1971:SOS

[OJ71] Milton Orchin and H. H. Jaffé.

Symmetry, Orbitals, and Spectra (S.O.S.). Wiley-Interscience, New York, NY, USA, 1971. ISBN 0-471-65550-3. xiii + 396 pp. LCCN QD461.O73.

Olver:2010:NHM

[OLBC10] Frank W. J. Olver, Daniel W.

Lozier, Ronald F. Boisvert, and Charles W. Clark, editors. NIST Handbook of Mathematical Functions. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-19225-0. xv + 951 pp. LCCN QA331 .N57 2010. US$99.00.

                                        URL http://dlmf.nist.gov/;
                                         http://www.cambridge.org/
                                        9780521140638.

Olver:1974:ASF

[Olv74] F. W. J. Olver. Asymptotics and Special Functions. Academic Press, New York, NY, USA, 1974. ISBN 0-12-525850-X. xvi + 572 pp. LCCN QA351 .O481 1974.

Omondi:1994:CAS

[Omo94] Amos R. Omondi. Computer Arithmetic Systems: Algorithms, Architecture, and Implementation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-334301-4. xvi + 520 pp. LCCN QA76.9.C62 O46 1994. US$40.00.

ONeil:2016:WMD

[O’N16] Cathy O’Neil. Weapons of math destruction: how big data increases inequality and threatens democracy. Crown, New York, NY, USA, 2016. ISBN 0-553-41881-5 (hardcover), 0-553-41883-1 (softcover), 0-553-41882-3 (e-book), 0-451-49733-3 (international edition). x + 259 pp. LCCN QA76.9.B45 O64 2016.

OSF:1991:AES

[Ope91a] Open Software Foundation. Application Environment Specification (AES) User Environment Volume, Revision B. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-043530-9. xii + 1084 + 12 pp. LCCN QA76.76.I57A665 1991.

OSF:1991:OMPa

[Ope91b] Open Software Foundation.

OSF/Motif Programmer’s Guide, Revision 1.1. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-640673-4. LCCN QA76.76.W56 O69 1991.

OSF:1991:OMPb

[Ope91c] Open Software Foundation.

OSF/Motif Programmer’s Reference, Revision 1.1. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-640681-5. xii + 1212 + 15 pp. LCCN QA76.76.W56 O7 1991.

OSF:1991:OMS

[Ope91d] Open Software Foundation.

OSF/Motif Style Guide. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-640616-5. LCCN QA76.76.W56 O833 1991.

ONW:awk

[Ope91e] OpenNetwork. The Berkeley Utilities, 1991. 215 Berkeley Place, Brooklyn, NY 11217, USA, Tel: (718) 398-3838.

OQuinn:1999:PND

[O’Q99] Donnie O’Quinn. Photoshop in a Nutshell: a Desktop Quick Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1999. ISBN 1-56592-565-3 (paperback). xxvi + 632 pp. LCCN T385 .O68 1999. US$24.95. URL http:

//www.oreilly.com/catalog/

photonut2/. Covers Adobe Photoshop version 5.0.

OReilly:1988:XWS

[OQL88] Tim O’Reilly, Valerie Quercia, and Linda Lamb. X Window System User’s Guide for Version 11, volume 3. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1988. ISBN 0-937175-29-3. xviii + 344 pp. LCCN QA76.76.W56 D44 v.3 1988.

OReilly:1990:XTI

[O’R90] Tim O’Reilly. X Toolkit Intrinsics Reference Manual, volume 5. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1990. ISBN 0-937175-35-8. xii + 543 pp. LCCN QA76.76.W56 D44 v.5 1990. URL http://

proquest.safaribooksonline.
com/9780937175354.

OReilly:2008:HI

[O’R08] Terence O’Reilly. Hitler’s Irishmen. Mercier Press, Cork, Ireland, 2008. ISBN 1-85635-589-6. 320 pp. LCCN D763.I73 O74 2008.

Organick:1972:MSE

[Org72] Elliott I. Organick. The Multics System: An Examination of Its Structure. MIT Press, Cambridge, MA, USA, 1972. ISBN 0-262-15012-3. xvii + 392 pp. LCCN QA76.5 O73.

Organick:1973:CSO

[Org73]  Elliott I. Organick.  Com-

puter System Organization: The B5700/B6700 Series. Academic Press, New York, NY, USA, 1973. ISBN 0-12-528250-8. ix + 132 pp.

Organick:1983:PVI

[Org83] Elliott I. Organick. A Programmer’s View of the Intel 432 System. McGraw-Hill, New York, NY, USA, 1983. ISBN 0-07-047719-1. xiii + 418 pp. LCCN QA76.8.I267 O73 1983. US$29.95.

NISO:1991:EMP

[Org91] National Information Standards Organization. Electronic Manuscript Preparation and Markup: American National Standard for Electronic Manuscript Preparation and Markup. Transaction Publishers, Rutgers University, New Brunswick, NJ 08903, USA, 1991. ISBN 0-88738-945-7. xv + 167 pp. LCCN Z286.E43 E428 1991. ANSI/NISO Z39.59-1988. Approved December 1, 1988, by American National Standards Institute. Developed by The National Information Standards Organization.

OShea:2007:PCS

[O’S07] Donal O’Shea. The Poincaré conjecture: in search of the shape of the universe. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 2007. ISBN 0-8027-1532-X. ix + 293 pp. LCCN QA612 .O83 2007. US$26.95.

Osborne:1979:RWN

[Osb79] Adam Osborne. Running Wild: The Next Industrial Revolution. Osborne/McGraw-Hill, Berkeley, CA, USA, 1979. ISBN 0-931988-28-4. x + 181 pp. LCCN QA76.9.C66 .O8.

Osgood:1965:M

[Osg65] William Fogg Osgood. Mechanics. Dover, New York, NY, USA, 1965. xiii + 495 pp. LCCN QA807.O8 1965. Reprint of the original 1937 edition.

OReilly:1988:MUU

[OT88] Tim O’Reilly and Grace Todino. Managing UUCP and Usenet. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, March 1988. ISBN 0-937175-09-9. xvi + 256 pp. LCCN QA76.8.U65 O64 1988. URL http:

                                        //www.oreilly.com/catalog/
                                        9780937175095.

Oram:1991:MPM

[OT91] Andrew Oram and Steve Talbott. Managing Projects with Make. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1991. ISBN 0-937175-90-0. xiv + 136 pp. LCCN QA76.76.O63 T35 199. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780937175903;   http://www.
                                        oreilly.com/catalog/make2.

Overton:2001:NCI

[Ove01] Michael L. Overton. Numerical Computing with IEEE Floating Point Arithmetic, Including One Theorem, One Rule of Thumb, and One Hundred and One Exercises. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2001. ISBN 0-89871-482-6 (hardcover), 0-89871-571-7 (paperback), 0-89871-807-4 (ebook). xiv + 104 pp. LCCN QA76.9.M35 O94 2001. US$40.00. URL http:/

/www.cs.nyu.edu/cs/faculty/
overton/book/;   http://www.
siam.org/catalog/mcc07/ot76.
htm.

Pitman:2005:MSP

[PA05] Jim Pitman and David Aldous. The Mathematics Survey Project. Notices of the American Mathematical Society, 52(11):1357–1360, December 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/200511/
comm-pitman.pdf.

Primack:2006:VCU

[PA06] J. R. (Joel R.) Primack and Nancy Ellen Abrams. The view from the center of the universe: discovering our extraordinary place in the cosmos. Riverhead Books, New York, NY, USA, 2006. ISBN 1-59448-914-9. 386 pp. LCCN QB981 .P85 2006.

Paabo:2014:NMS

[Pää14]  Svante Pääbo.  Neanderthal

man: in search of lost genomes. Basic Books, New York, NY, USA, 2014. ISBN 0-465-02083-6 (hardcover), 0-465-05495-1, 0-465-08068-5 (e-book). ix + 275 pp. LCCN GN285 .P33 2014.

Padfield:1990:HFS

[Pad90] Peter Padfield. Himmler: a Full-scale Biography of One of Hitler’s Most Ruthless Executioners. MJF Books, New York, NY, USA, 1990. ISBN 1-56731-118-0. xiv + 656 + 16 pp. LCCN D247.H46 P33 1996.

Paeth:1995:GGV

[Pae95] Alan W. Paeth, editor. Graphics Gems V, volume 5 of Graphics Gems. AP Professional, Boston, MA, USA, 1995. ISBN 0-12-543455-3, 0-12-543457-X. xx + 438 pp. LCCN T385 .G6935 1995b.

Pagels:1982:CCQ

[Pag82] Heinz R. Pagels. The Cosmic Code: Quantum Physics as the Language of Nature. Simon and Schuster, New York, NY, USA, 1982. ISBN 0-671-24802-2. 370 + 8 pp. LCCN QC174.13 .P33 1982.

Pagels:1983:CCQ

[Pag83] Heinz R. Pagels. The Cosmic Code: Quantum Physics as the Language of Nature. A Bantam new age book. Bantam Books, Toronto, ON, Canada, 1983. ISBN 0-553-23128-6, 0-553-24625-9. xiv + 333 pp. LCCN QC174.13 .P33 1983.

Pagels:1988:DRC

[Pag88] Heinz R. Pagels. The Dreams of Reason: The Computer and the Rise of the Sciences of Complexity. Simon and Schuster, New York, NY, USA, 1988. ISBN 0-671-62708-2. 352 pp. LCCN QA76.9.C66 P34 1988.

Page:2007:DHP

[Pag07] Scott E. Page. The difference:

how the power of diversity creates better groups, firms, schools, and societies. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-13854-0, 1-4008-3028-1 (e-book). xxiv + 424 pp. LCCN HF5549.5.M5 P34 2007eb.

Pagels:2011:CCQ

[Pag11] Heinz R. Pagels. The cosmic code: quantum physics as the language of nature. Dover, New York, NY, USA, Dover edition, 2011. ISBN 0-486-48506-4 (paperback). 370 pp. LCCN QC174.13 .P33 2011.

Pais:1991:NBT

[Pai91] Abraham Pais. Niels Bohr’s times: in physics, philosophy, and polity. Clarendon Press, Oxford, UK, 1991. ISBN 0-19-852048-4 (paperback), 0-19-852049-2. xvii + 565 pp. LCCN QC773 .P35 1991. US$35.00. URL

http://www.loc.gov/catdir/

                                        enhancements/fy0635/90027248-
                                        d.html.

Pais:1994:ELH

[Pai94] Abraham Pais. Einstein lived here: essays for the layman. Clarendon Press, Oxford, UK, 1994. ISBN 0-19-853994-0. xvi + 282 pp. LCCN QC16.E5 P25 1994. US$22.42.

Parkinson:1980:PL

[Par80a] Cyril Northcote Parkinson. Parkinson, the Law. Houghton-Mifflin, Boston, MA, USA, 1980. ISBN 0-395-29131-3. 207 pp. LCCN PN6231.M2 P297 1980. US$8.95.

Parlett:1980:SEP

[Par80b] Beresford N. Parlett. The Symmetric Eigenvalue Problem. Series in Computational Mathematics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1980. ISBN 0-13-880047-2. xix + 348 pp. LCCN QA188 .P3. Prentice-Hall Series in Computational Mathematics.

Parter:1984:LSS

[Par84] Seymour V. Parter, editor. Large-scale scientific computation: proceedings of a conference conducted by the Mathematics Research Center the University of Wisconsin, Madison, May 17– 19, 1983. Academic Press, New York, NY, USA, 1984. ISBN 0-12-546080-5. LCCN QA911 .L32

Parker:printer-fonts

[Par88] Mike Parker. Printer fonts.

In Robert C. Durbeck and Sol Sherr, editors, Output Hardcopy Devices, chapter 19, pages 497– 516. Academic Press, New York, NY, USA, 1988.

Parlett:1998:SEP

[Par98]  Beresford N. Parlett.  The

Symmetric Eigenvalue Problem, volume 20 of Classics in Applied Mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1998. ISBN 0-89871-402-8 (paperback), 1-61197-116-0 (e-book). xxiv + 398 pp. LCCN QA188 .P37 1998. URL

http://www.loc.gov/catdir/
enhancements/fy0708/97040623-

d.html. Corrected reprint [Par80b].

Parhami:2000:CAA

[Par00] Behrooz Parhami. Computer Arithmetic: Algorithms and Hardware Designs. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-512583-5 (hardcover), 3-540-04907-X (print), 3-540-36246-0 (e-book). xx + 490 pp. LCCN QA76.9.C62P37 1999. US$85.00.

Parker:2003:BEH

[Par03] Andrew Parker. In the blink of an eye: how vision sparked the big bang of evolution. Basic Books, New York, NY, USA, 2003. ISBN 0-465-05438-2 (paperback). xvi + 316 pp. LCCN QE770 .P37 2004.

Pauncz:1979:SEC

[Pau79] Ruben Pauncz. Spin Eigenfunctions: Construction and Use. Plenum Press, New York, NY, USA; London, UK, 1979. ISBN 0-306-40141-X. xv + 370 pp. LCCN QD462.P38.

Paul:1989:CMC

[Pau89] Kevin Paul. Chairing a meeting with confidence: an easy guide to rules and procedure. Self-Counsel reference series. Self-Counsel Press, North Vancouver, BC, Canada, 1989. ISBN 0-88908-858-6. xi + 80 pp. LCCN AS6 .P36 1989.

Paulos:1991:BNR

[Pau91] John Allen Paulos. Beyond Numeracy: Ruminations of a Numbers Man. Alfred A. Knopf, New York, NY, USA, 1991. ISBN 0-394-58640-9, 0-685-48163-8. xiii + 285 pp. LCCN QA5 .P38 1991.

Pavlidis:1982:AGI

[Pav82] Theo Pavlidis. Algorithms for Graphics and Image Processing. Computer Science Press, 11 Taft Court, Rockville, MD 20850, USA, 1982. ISBN 0-914894-65-X. xv + 416 pp. LCCN T385 .P38 1982.

Pavlidis:1999:FXP

[Pav99] Theodosios Pavlidis. Fundamentals of X Programming: Graphical User Interfaces and Beyond. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-306-46065-3 (hardcover), 0-306-46968-5 (e-book). xvii + 374 pp. LCCN QA76.76.W56P387

1999.  US$89.85.  URL

http://www.loc.gov/catdir/
enhancements/fy0902/98056154-
d.html.

Paxson:flex

[Pax88] Vern Paxson. flex: fast lexical analyzer generator. Free Software Foundation, Inc. 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1988. Electronic mail: vern@lbl-

csam.arpa or vern@lbl-rtsg.

arpa. Software also available via ANONYMOUS FTP to lbl-

csam.arpa,  lbl-rtsg.arpa,
or prep.ai.mit.edu. See also

[DS88a].

Page:1998:PCR

[PBMW98] Lawrence Page, Sergey Brin, Rajeev Motwani, and Terry Winograd. The PageRank citation ranking: Bringing order to the web. Technical report, Stanford Digital Library Technologies Project, Stanford University, Stanford, CA, USA, November 11, 1998. 17 pp. URL

http://dbpubs.stanford.edu/
pub/1999-66.

Patterson:raid

[PCGK89] D. A. Patterson, P. Chen, G. Gibson, and R. H. Katz. Introduction to redundant arrays of inexpensive disks. Digest of Papers of COMPCON Spring ’89, 1989. IEEE Computer Society Press.

pc-sales88

[PCS88] 1988. Trade journal advertisement sales volume estimates.

Pepple:2004:MSO

[PDL04] Ken Pepple, Brian Down, and David Levy. Migrating to the Solaris Operating System: The Discipline of UNIX-to-UNIX Migrations. Sun BluePrints Program. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-150263-8. xix + 251 pp. LCCN QA76.76.O63 P445 2004.

Pennington:1965:ICM

[Pen65] Ralph H. Pennington. Introductory Computer Methods and Numerical Analysis. Macmillan Publishing Company, New York, NY, USA, 1965. xi + 452 pp. LCCN QA76.5 .P38.

Penrose:2011:CTE

[Pen11] Roger Penrose. Cycles of Time: an Extraordinary New View of the Universe. Alfred A. Knopf, New York, NY, USA, 2011. ISBN 0-224-08036-9 (hardcover), 0-307-26590-0 (hardcover). xii + 288 pp. LCCN QB991.C92 P46 2011.

Peterson:1998:JRM

[Pet98] Ivars Peterson. The Jungles of Randomness: a Mathematical Safari. Wiley, New York, NY, USA, 1998. ISBN 0-471-16449-6. viii + 239 pp. LCCN QA273.15.P48 1997. US$24.95.

PeytonJones:1987:IFP

[Pey87] Simon L. Peyton Jones. The Implementation of Functional Programming Languages. Prentice-Hall International, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-453333-X. xviii + 445 pp. LCCN QA76.7.P495 1987. With chapters by Philip Wadler, Peter Hancock, David Turner.

Pharr:2005:GGP

[PF05] Matt Pharr and Randima Fernando, editors. GPU Gems 2: Programming Techniques for High-performance Graphics and General-purpose Computation. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-321-33559-7 (hardcover). xlix + 814 pp. LCCN T385 .G688 2005. URL http://www.

loc.gov/catdir/toc/ecip055/
2004030181.html.

Pasachoff:2019:CAN

[PF19] Jay M. Pasachoff and Alexei V.

Filippenko. The Cosmos: Astronomy in the New Millennium. Cambridge University Press, Cambridge, UK, fifth edition, 2019. ISBN 1-108-43138-0. ???? pp. LCCN QB43.3 .P37 2019. URL http:

//thecosmos5.com.

Press:1986:NRA

[PFTV86] William H. Press, Brian P. Flannery, Saul A. Teukolsky, and William T. Vetterling. Numerical Recipes: The Art of Scientific Computing (FORTRAN Version). Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-38330-7. xx + 702 pp. LCCN QA297 .N866 1989.

Patterson:1990:CAQ

[PH90] David A. Patterson and John L. Hennessy. Computer Architecture: a Quantitative Approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1990. ISBN 1-55860-069-8, 1-55880-169-8. xxviii + 594 pp. LCCN QA76.9.A73 P377 1990.

Patterson:1994:COD

[PH94] David A. Patterson and John L. Hennessy. Computer Organization and Design: The Hardware/Software Interface. Morgan Kaufmann Publishers, 2929 Campus Drive, Suite 260, San Mateo, CA 94403, USA, 1994. ISBN 1-55860-281-X (paperback), 1-4832-2118-0 (e-book). xxiv + 648 pp. LCCN QA76.9 .C643 P37 1994. US$74.75.

Patterson:1996:CAQ

[PH96] David A. Patterson and John L. Hennessy. Computer Architecture — a Quantitative Approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 1996. ISBN 1-55860-329-8. xxiii + 760 + A-77 + B-47 + C-26 + D-26 + E-13 + R-16 + I-14 pp. LCCN QA76.9.A73P377 1995. US$69.95.

Patterson:1997:COH

[PH97] David A. Patterson and John L.

Hennessy. Computer Organization: The Hardware/Software Interface. Morgan Kaufmann Publishers, 2929 Campus Drive, Suite 260, San Mateo, CA 94403, USA, second edition, 1997. ISBN 1-55860-428-6 (hardcover), 1-55860-491-X (softcover). xxix + 759 + 205 pp. LCCN QA76.9.C643H46 1997. US$78.95.

Patterson:2003:CAQ

[PH03] David A. Patterson and John L.

Hennessy. Computer Architecture — a Quantitative Approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, third edition, 2003. ISBN 1-55860-596-7. xxi + 883 + A-87 + B-42 + C-1 + D-1 + E-1 + F-1 + G-1 + H-1 + I-1 + R-22 + I-44 pp. LCCN QA76.9.A73 P377 2003. US$89.95. URL

http://www.mkp.com/books_
catalog/catalog.asp?ISBN=1-
55860-596-7;   http://www.
mkp.com/CA3.

Patterson:2004:COH

[PH04] David A. Patterson and John L.

Hennessy. Computer Organization: The Hardware/Software Interface. Morgan Kaufmann Publishers, 2929 Campus Drive, Suite 260, San Mateo, CA 94403, USA, third edition, 2004. ISBN 1-55860-604-1. xvii + 621 pp. LCCN QA76.9.C643 H46 2004. US$64.95.

Patterson:2008:COD

[PH08] David A. Patterson and John L. Hennessy. Computer Organization and Design: the Hardware/ Software Interface. Elsevier/ Morgan Kaufmann, San Francisco, CA, USA, fourth edition, 2008. ISBN 0-12-374493-8. xxv + 703 + A-77 + B-83 + I-26 pp. LCCN QA76.9.C643.

Patterson:2012:COD

[PH12] David A. Patterson and John L. Hennessy. Computer organization and design: the hardware/ software interface. The Morgan Kaufmann series in computer architecture and design. Morgan Kaufmann/Elsevier, Waltham, MA, USA, fourth (revised) edition, 2012. ISBN 0-12-374750-3 (paperback). xxv + 703 pp. LCCN QA76.9.C643 H46 2012. With contributions by Perry Alexander, Peter J. Ashenden, Javier Bruguera, Jichuan Chang, Matthew Farrens, David Kaeli, Nicole Kaiyan, David Kirk, James R. Larus, Jacob Leverich, Kevin Lim, John Nickolls, John Oliver, Milos Prvulovic, and Parta Ranganthan.

Patterson:2018:COD

[PH18] David A. Patterson and John L. Hennessy. Computer Organization and Design: the Hardware– Software Interface. Morgan Kaufmann/Elsevier, Waltham, MA, USA, RISC-V edition, 2018. ISBN 0-12-812275-7, 0-12-812276-5 (e-book). xxiv + 565 + 86 + i + 22 pp. LCCN QA76.9.C643 P37 2018.

Phillips:1936:BC

[Phi36] E. W. Phillips. Binary calculation. Journal of the Institute of Actuaries, 67(??):187–221, 1936. ISSN 0020-2681. Reprinted in [Ran82, §7.1].

Phillips:1986:NL

[Phi86] Jen Phillips. The NAG Library.

Clarendon Press, Oxford, UK, 1986. ISBN 0-19-853263-6. viii + 245 pp. LCCN QA297.P53 1986. US$14.50.

Pinker:1994:LI

[Pin94] Steven Pinker. The Language Instinct. William Morrow and Company, Inc., New York, NY, USA, 1994. ISBN 0-688-12141-1. 494 pp. LCCN P106.P476 1994. US$23.00.

Pinker:1997:HMW

[Pin97] Steven Pinker. How the Mind Works. W. W. Norton & Co., New York, NY, USA, 1997. ISBN 0-393-04535-8. xii + 660 pp. LCCN QP360.5.P56 1997. US$29.95.

Pinker:1999:WRI

[Pin99]  Steven Pinker.  Words and

Rules: the Ingredients of Language. Basic Books, New York, NY, USA, 1999. ISBN 0-465-07269-0 (hardcover), 0-06-095840-5 (paperback), 0-465-04971-0 (e-book). xi + 348 pp. LCCN P106 .P477 1999.

Pinker:2002:BSM

[Pin02] Steven Pinker. The blank slate: the modern denial of human nature. Viking, New York, NY, USA, 2002. ISBN 0-670-03151-8 (hardcover). xvi + 509 pp. LCCN BF341 .P47 2002.

Pinker:2003:BSM

[Pin03] Steven Pinker. The blank slate: the modern denial of human nature. Penguin, London, UK and New York, NY, USA, 2003. ISBN 0-14-200334-4 (paperback), 0-14-027605-X (paperback). xvi + 509 pp. LCCN BF341 .P47 2002.

Pinker:2011:BAO

[Pin11] Steven Pinker. The better angels of our nature: why violence has declined. Viking, New York, NY, USA, 2011. ISBN 0-670-02295-0 (hardcover), 0-670-02315-9 (paperback). xxviii + 802 pp. LCCN HM1116 .P57 2011.

Pinker:2014:SST

[Pin14] Steven Pinker. The Sense of Style: the Thinking Person’s Guide to Writing in the 21st Century! Viking, New York, NY, USA, 2014. ISBN 0-670-02585-2 (hardcover), 0-525-42792-9 (Export Edition). viii + 359 pp. LCCN PE1421 .P56 2014.

Pissanetsky:1984:SMT

[Pis84] Sergio Pissanetsky. Sparse Matrix Technology. Academic Press, New York, NY, USA, 1984. ISBN 0-12-557580-7. xiii + 321 pp. LCCN QA188 .P57 1984.

Pitzer:1953:QC

[Pit53] Kenneth S. (Kenneth Sanborn)

Pitzer. Quantum Chemistry. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1953. x + 529 pp. LCCN QC174.1 .P49.

Pitici:2019:BWM

[Pit19] Mircea Pitici, editor. The Best Writing On Mathematics: 2019, volume 2019. Princeton University Press, Princeton, NJ, USA, 2019. ISBN 0-691-19835-7, 0-691-19867-5. xvi + 272 + 16 pp. LCCN QA8.6 .B337 2019.

Peitgen:1991:FC

[PJS91] Heinz-Otto Peitgen, Hartmut Jürgens, and Dietmar Saupe. Fractals for the Classroom. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97346-X (vol. 1, New York), 3-540-97346-X (vol. 1, Berlin), 0-387-97722-8 (vol. 2, New York), 3-540-97722-8 (vol. 2, Berlin). xii + 128 (vol. 1) pp. LCCN QA614.86 .P45 1991. US$24.95 (vol. 1), US$39.95 (vol. 2). Issued in 1991–1992, with corrected second printing in 1993– 1994.

Posamentier:2007:FFN

[PL07] Alfred S. Posamentier and Ingmar Lehmann. The Fabulous Fibonacci Numbers. Prometheus Books, Amherst, NY, USA, 2007. ISBN 1-59102-475-7. 385 pp. LCCN QA241 .P665 2007. URL http://www.

                                        loc.gov/catdir/toc/ecip073/
                                        2006035406.html.

Plauger:1992:SCL

[Pla92] P. J. Plauger. The Standard C Library. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-838012-0. xiv + 498 pp. LCCN QA76.73.C15 P563 1991.

Plauger:1993:PPEa

[Pla93a] P. J. Plauger. Programming on Purpose: Essays on Software Design, volume 1. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-721374-3. viii + 236 pp. LCCN QA76.76.D47 P55 1993. US$19.95.

Plauger:1993:PPIb

[Pla93b] P. J. Plauger. Programming on Purpose II: Essays on Software People, volume 2. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-328105-1. viii + 204 pp. LCCN QA76.76.D47 P56 1993. US$19.95.

Plauger:1993:PPIc

[Pla93c] P. J. Plauger. Programming on Purpose III: Essays on Software Technology, volume 3. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-328113-2. viii + 224 pp. LCCN QA76.76.D47 P53 1994. US$19.95.

Plotnik:1982:EEM

[Plo82] Arthur Plotnik. The Elements of Editing: a Modern Guide for Editors and Journalists. Collier Books, Macmillan Publishing Company, New York, NY, USA, 1982. ISBN 0-02-597700-8. xiii + 156 pp. LCCN PN4778 .P59 1982.

Plum:1987:NDC

[Plu87] T. Plum. Notes on the Draft C Standard. Plum Hall, 1 Spruce Ave, Cardiff, NJ 08232, USA, 1987. ISBN 0-911537-06-6. 92 pp. LCCN QA76.73.C15 P585 1987. US$10.00.

Palmer:8087

[PM84] John F. Palmer and Stephen P.

Morse. The 8087 Primer. Wiley, New York, NY, USA, 1984. ISBN 0-471-87569-4. viii + 182 pp. LCCN QA76.8.I2923 P34 1984. Excellent coverage of the 8087 numeric coprocessor by the chief architects of the Intel 8087 (Palmer) and 8086 (Morse). Contains many candid statements about design decisions in these processors. A must for serious assembly language coding of the 8087 and 80287 chips. See also [Int85a].

Pennebaker:1992:JSI

[PM92] William B. Pennebaker and Joan L. Mitchell. JPEG Still Image Data Compression Standard. Van Nostrand Reinhold, New York, NY, USA, 1992. ISBN 0-442-01272-1. xviii + 638 pp. LCCN TA1632 .P45 1993.

Papagianni:2013:NRH

[PM13] Dimitra Papagianni and Michael A.

Morse. The Neanderthals rediscovered: how modern science is rewriting their story. Thames and Hudson Oxbow Books, London, UK, 2013. ISBN 0-500-05177-1 (hardcover). 208 + 16 pp. LCCN ????

Palmer:2004:SAO

[PN04] Brandon Palmer and Jose Nazario. Secure Architectures with OpenBSD. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-19366-0 (paperback). xix + 519 pp. LCCN QA76.76.O63 P335 2004. URL ftp://

                                        uiarchive.cso.uiuc.edu/pub/
                                        etext/gutenberg/;   http:
                                        //www.loc.gov/catdir/toc/
                                        ecip0413/2004001163.html.

Pogue:2005:MXT

[Pog05] David Pogue. Mac OS X, Tiger Edition: the Missing Manual. Missing manual. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00941-0 (paperback), 1-4493-7907-9 (e-book), 1-4493-9210-5 (ebook). xvi + 847 pp. LCCN QA76.76.O63 P634 2005.

Pollack:2003:USL

[Pol03] Henry N. Pollack. Uncertain science …: uncertain world. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-511-07776-9, 0-511-07465-4, 0-511-54137-6, 0-521-78188-4, 0-511-07619-3. xii + 243 pp. LCCN Q175 .P835 2003.

Pollack:2010:WI

[Pol10] H. N. Pollack. A world without ice. Avery, New York, NY, USA, 2010. ISBN 1-58333-407-6. xii + 290 pp. LCCN GB2405 .P55 2010.

Adobe:docstruct-spec

[Pos89] PostScript Developer Support Group. Document Structuring Conventions specification, version 2.1. Technical Report PN LPS5001, Adobe Systems Incorporated, 1585 Charleston Road, P. O. Box 7900, Mountain View, CA 94039-7900, USA, Tel: (415) 961-4400, January 16 1989. Available electronically

from ps-file-server@adobe.

com in response to an e-mail re-

quest send Documents struct.
ps.Zba. The request send Index

will return a complete index for the server.

Penna:1986:PGA

[PP86] Michael A. Penna and Richard R. Patterson. Projective Geometry and its Applications to Computer Graphics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-730649-0. xi + 403 pp. LCCN QA471 .P3951 1986. US$37.50.

Palais:2009:DEM

[PP09] Richard S. Palais and Robert A.

Palais. Differential Equations, Mechanics, and Computation, volume 51 of Student mathematical library; IAS/Park City mathematical subseries. American Mathematical Society, Providence, RI, USA, 2009. ISBN 0-8218-2138-5. xiii + 313 pp. LCCN QA371.P34 2009.

Perelman:1998:MHT

[PPB98] Leslie C. Perelman, James Paradis, and Edward Barrett. The Mayfield Handbook of Technical and Scientific Writing. Mayfield Publishing Company, Mountain View, CA, USA, 1998. ISBN 1-55934-647-7. xxvii + 508 pp. LCCN T11 .P394 1998. US$31.95.

Peitgen:1986:BF

[PR86] Heinz-Otto Peitgen and Peter H. Richter. The Beauty of Fractals. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-15851-0. xii + 199 pp. LCCN QA447 .P45 1986.

Prior:1955:FL

[Pri55] A. N. Prior. Formal Logic. Clarendon Press, Oxford, UK, 1955. ix + 329 pp. LCCN BC108.P8.

Price:1989:BT

[Pri89] Walter J. Price. A benchmark tutorial. IEEE Micro, 9(5):28–43, September/October 1989. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Price:2021:GWB

[Pri21] David A. (David Andrew) Price. Geniuses at War: Bletchley Park, Colossus, and the Dawn of the Digital Age. Alfred A. Knopf, New York, NY, USA, 2021. ISBN 0-525-52154-2 (hardcover), 0-525-52155-0 (e-book). 243 pp. LCCN D810.C88 P75 2021.

Prueitt:1984:AC

[Pru84] Melvin L. Prueitt. Art and the Computer. McGraw-Hill, New York, NY, USA, 1984. ISBN 0-07-050894-1 (hardcover), 0-07-050899-2 (paperback). ix + 246 pp. LCCN N7433.8.P7 1984. US$39.95 (hardcover), US$29.95 (paperback).

Chambers:CSF58-1

[Pry58a] James Pryde, editor. Chambers’s Seven-Figure Mathematical Tables, volume 1. W. and R. Chambers Ltd., London, UK, 1958. 392 pp. LCCN QA47 .P7 1958.

Chambers:CSF58-2

[Pry58b] James Pryde, editor. Chambers’s Seven-Figure Mathematical Tables, volume 2. W. and R. Chambers Ltd., London, UK, 1958. LCCN QA47 .P7 1958.

Pryor:1966:IFR

[Pry66] William A. Pryor. Introduction to Free Radical Chemistry. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1966. xvii + 110 pp. LCCN QD255 .P77.

Preparata:1985:CGI

[PS85] Franco P. Preparata and Michael Ian Shamos. Computational Geometry: an Introduction. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-96131-3, 1-4612-1098-4 (e-book), 1-4612-7010-3 (print). ISSN 0172-603X. xii + 390 pp. LCCN QA447 .P735 1985. US$32.00.

Peitgen:1988:SFI

[PS88] Heinz-Otto Peitgen and Dietmar Saupe, editors. The Science of Fractal Images. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-96608-0 (New York), 1-4612-8349-3, 3-540-96608-0 (Berlin). xiii + 312 pp. LCCN QA614.86 .S35 1988.

Plauger:2000:CST

[PSLM00] P. J. Plauger, Alexander A. Stepanov, Meng Lee, and David R. Musser. The C++ Standard Template Library. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-437633-1 (paperback). xii + 485 pp. LCCN QA76.73.C153 C17 2000. US$40.00.

Piegl:1997:NB

[PT97] Les Piegl and Wayne Tiller. The NURBS Book. Monographs in visual communication. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1997. ISBN 3-540-61545-8. xiv + 646 pp. LCCN QA224 .P54 1997. US$59.95.

Press:2007:NRA

[PTVF07] William H. Press, Saul A. Teukolsky, William T. Vetterling, and Brian P. Flannery. Numerical Recipes — The Art of Scientific Computing. Cambridge University Press, Cambridge, UK, third edition, 2007. ISBN 0-521-88068-8 (hardcover), 0-521-88407-1 (with source code CD ROM), 0-521-70685-8 (source code CD ROM). xxi + 1235 pp. LCCN QA297 .N866 2007.

URL http://www.cambridge.
org/numericalrecipes.

Pauling:1935:IQM

[PW35] Linus Pauling and E. Bright Wilson, Jr. Introduction to Quantum Mechanics: With Applications to Chemistry. McGraw-Hill, New York, NY, USA, 1935. xi + 468 pp. LCCN QC174.1 .P34. Reprinted by Dover, 1985.

Packel:1994:ACM

[PW94] Ed Packel and Stan Wagon. Animating Calculus: Mathematica Notebooks for the Laboratory. W. H. Freeman and Company, New York, NY, USA, 1994. ISBN 0-7167-2428-6. xii + 289 pp. LCCN QA303.5.D37P33 1994.

Patterson:2017:RVR

[PW17] David Patterson and Andrew Waterman. The RISC-V Reader: An Open Architecture Atlas. Strawberry Canyon, San Francisco, CA, USA, 2017. ISBN 0-9992491-1-8. xiv + 180 pp. LCCN QA76.9.A73 P388 2017.

Pyle:1981:APL

[Pyl81] I. C. Pyle. The Ada Programming Language. Prentice-Hall International, Upper Saddle River, NJ 07458, USA, 1981. ISBN 0-13-003921-7. x + 293 pp. LCCN QA76.73.A35 P94. US$14.95.

Pyne:2010:VSN

[Pyn10] Stephen J. Pyne. Voyager: Seeking Newer Worlds in the Third Great Age of Discovery. Viking, New York, NY, USA, 2010. ISBN 0-670-02183-0. xix + 444 + 8 pp. LCCN TL789.8.U6 V5275 2010.

Pyster:1980:CDC

[Pys80] Arthur B. Pyster. Compiler Design and Construction. Van Nostrand Reinhold, New York, NY, USA, 1980. ISBN 0-442-24394-

4. xvi + 357 pp. LCCN QA76.6

.P9.

Quinn:2008:MMA

[QN08] Helen R. Quinn and Yossi Nir. The Mystery of the Missing Antimatter. Science essentials. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-13309-

3. xii + 278 pp. LCCN QC173.3 .Q856 2008. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0835/2007934402-
                                        b.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0835/
                                        2007934402-d.html;   http:/
                                        /www.loc.gov/catdir/enhancements/
                                        fy0835/2007934402-t.html.

Quercia:1990:XWS

[QO90] Valerie Quercia and Tim O’Reilly. X Window System User’s Guide, volume 3. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 1990. ISBN 0-937175-14-5. xxvi + 723 pp. LCCN QA76.76.W56 Q83 1990.

URL http://proquestcombo.
safaribooksonline.com/9780937175149;
 http://www.oreilly.com/catalog/
9780937175149.

Quercia:1993:XWSb

[QO93a] Valerie Quercia and Tim O’Reilly.

X Window System User’s Guide: OSF/Motif 1.2 Edition, volume 3M. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, May 1993. ISBN 1-56592-015-5. xxx + 835 pp. LCCN QA76.76 W56Q4 1993.

Quercia:1993:XWSa

[QO93b] Valerie Quercia and Tim O’Reilly.

X Window System User’s Guide: Standard Edition, volume 3. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, May 1993. ISBN 1-56592-014-7. xxx + 835 pp. LCCN QA76.76.W56D43 1993. US$34.95. URL http:

//www.oreilly.com/catalog/
9781565920149.

Quarterman:1989:M

[Qua89] John S. Quarterman. The Matrix. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1989. ISBN 1-55558-033-5. xxvii + 719 pp. LCCN TK5105.5 .Q37 1990.

Quammen:2018:TTR

[Qua18] David Quammen. The Tangled Tree: a Radical New History of Life. Simon and Schuster, New York, NY, USA, 2018. ISBN 1-4767-7662-8. xvi + 461 + 8 pp. LCCN QH367.5 .Q36 2018.

Quigley:2002:USE

[Qui02] Ellie Quigley. UNIX Shells by Example. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, third edition, 2002. ISBN 0-13-066538-X (paperback). xix + 1015 pp. LCCN QA76.76.O63 Q54 2002.

Rogers:1976:MEC

[RA76] David F. Rogers and J. Alan Adams. Mathematical Elements for Computer Graphics. Mc-Graw-Hill, New York, NY, USA, 1976. ISBN 0-07-053527-2 (paperback), 0-07-053530-2. xv + 239 pp. LCCN T385 .R6 1976.

Rago:1993:USV

[Rag93] Steven A. Rago. UNIX System V network programming. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-56318-5. xv + 784 pp. LCCN 92-45276, QA76.76.O63 R34 1993. US$45.50.

Ramondino:1969:NWS

[Ram69] Salvatore Ramondino, editor. The new world Spanish–English and English–Spanish dictionary. Signet, New York, NY, USA, 1969. ISBN 0-451-15994-2. xviii + 1226 pp. LCCN PC4640 .N4 1968.

Ramsey:spiderweb

[Ram89] Norman Ramsey. Weaving a language-independent WEB. Communications of the Association for Computing Machinery, 32(9): 1051–1055, September 1989.

Raman:1997:AUI

[Ram97] T. V. Raman. Auditory User Interfaces: Toward the Speaking Computer. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-9984-6. xxi + 142 pp. LCCN QA76.9.U83 R36 1997.

Randell:1977:CGC

[Ran77] B. Randell. Colossus: Godfather of the computer. New Scientist, 73(1038):346–348, February 10, 1977. CODEN NWSCAL. ISSN 0262-4079, 0028-6664. Reprinted in [Ran82, §7.5].

Randell:1982:ODC

[Ran82] Brian Randell, editor. The Origins of Digital Computers: Selected Papers. Texts and monographs in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 1982. ISBN 0-387-11319-3, 3-540-11319-3. xvi + 580 pp. LCCN TK7885.A5 O741 1982.

Rankin:1989:CGS

[Ran89] John R. Rankin. Computer Graphics Software Construction. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-162793-7 (hardcover). xvi + 544 pp. LCCN T385 .R364

Rand-McNally:1991:RMR

[Ran91]  Rand McNally.  Rand Mc-

Nally road atlas: United States, Canada, Mexico. Rand McNally, Skokie, IL, USA, 1991. ISBN 0-528-80500-2. 264 pp. LCCN G1201.P2 R35 1991.

Rand-McNally:2001:RAL

[Ran01] Rand McNally. The road atlas, large scale: United States. Rand McNally, Skokie, IL, USA, 2001. ISBN 0-528-84310-9. 264 pp. LCCN G1201.P2 R349 2001.

Rayna:1987:RSA

[Ray87]  Gerhard Rayna.  REDUCE:

Software for Algebraic Computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96598-X (New York), 3-540-96598-X (Berlin). ix + 329 pp. LCCN QA155.7.E4 R39 1987.

Raymond:2004:AUP

[Ray04] Eric Steven Raymond. The Art of UNIX Programming. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-13-124085-4, 0-13-142901-9. xxxii + 525 pp. LCCN QA76.76.O63 R395 2003. US$39.99, CAN$60.99.

Roberts:1992:UDG

[RB92] Ralph Roberts and Mark Boyd. UNIX Desktop Guide to Emacs. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1992. ISBN 0-672-30171-7. xxiii + 504 pp.

LCCN QA76.76.T49 R62 1992. US$27.95.

Robbins:2005:CSS

[RB05a] Arnold Robbins and Nelson H. F.

Beebe. Classic Shell Scripting. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00595-4. xxii + 534 pp. LCCN QA76.76.O63 R633 2005. URL http:

//www.oreilly.com/catalog/

shellsrptg/. Also available in Chinese [RB08], French [RB05b], German [RB06a], Japanese [RB06c], and Polish [RB06b] translations.

Robbins:2005:ISS

[RB05b] Arnold Robbins and Nelson H. F.

Beebe. Introduction aux scripts shell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 2-84177-375-2. xxii + 558 pp. LCCN ???? URL http:/

/www.silicon.fr/getarticle.

asp?id=14015. French translation of [RB05a] by Eric Jacoboni.

Robbins:2006:KSP

[RB06a] Arnold Robbins and Nelson H. F. Beebe. Klassische Shell-Programmierung: [automatisieren Sie Ihre Unix/ Linux-Tasks]. (German) [Classic Shell Scripting: [automating your Unix/Linux jobs]]. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 3-89721-441-5. xxiii + 572 pp. LCCN QA76.76.O63 R563 2005. EUR 44.00; EUR 45.65

                                        (AT). URL http://www.gbv.
                                        de/dms/hebis-darmstadt/toc/
                                        17645067X.pdf. German trans-

lation of [RB05a] by Kathrin

Lichtenberg.

Robbins:2006:PSP

[RB06b] Arnold Robbins and Nelson H. F. Beebe. Programowanie skryptów powłoki. Helion, Gliwice, Poland, 2006. ISBN 83-246-0131-7. 557 + 2 pp. LCCN ???? URL http:

                                        //www.empik.com/b/o/19/f1/
                                        19f16b85e0d75ae1d3a1e7062569fbb0.
                                        jpg;   http://www.empik.com/
                                        programowanie-skryptow-powloki-
                                        ksiazka,360529,p.  Polish

translation of [RB05a] by Prze-

mysław Szeremiota.

Robbins:2006:SSS

[RB06c] Arnold Robbins and Nelson H. F. Beebe. Shōkai shieru sukuriputo. Orairī Japan, Tōkyō, Japan, 2006. ISBN 4-87311-267-2. 345 pp. LCCN ???? Japanese translation of [RB05a] by Aoi Hyūga.

Robbins:2008:SJB

[RB08] Arnold Robbins and Nelson H. F. Beebe. Shell Jiao Ben Xue Xi Zhi Nan = Shell Script Study Guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 7-111-25504-

6. vi + 494 pp. LCCN ???? 79.00 renminbi. Simplified Chinese translation of [RB05a].

Roberts:1964:BPO

[RC64a] John D. Roberts and Marjorie C.

Caserio. Basic Principles of Organic Chemistry. W. A. Benjamin, Inc., New York, NY, USA, 1964. xxv + 1315 pp. LCCN QD251 .R58. Second printing, with corrections, March 5, 1965.

Roberts:1964:SBP

[RC64b] John D. Roberts and Marjorie C.

Caserio. Supplement for Basic Principles of Organic Chemistry. W. A. Benjamin, Inc., New York, NY, USA, 1964. 435 pp. LCCN QD 251 R54b. Second printing, with corrections, September 6, 1965.

Reingold:2002:CCM

[RD02a] Edward M. Reingold and Nachum Dershowitz. Calendrical Calculations: The Millenium Edition. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-77752-6 (paperback), 0-521-77167-6 (hardcover). xxxii + 422 pp. LCCN CE12 .R45 2001. US$38.00 (paperback), US$100.00 (hard-

cover).  URL http://emr.
cs.iit.edu/home/reingold/
calendar-book/second-edition/

.

Reingold:2002:CT

[RD02b] Edward M. Reingold and Nachum Dershowitz. Calendrical Tabulations: 1900–2200. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-78253-8 (hardcover). xxx + 606 pp. LCCN CE12 .R46 2002. US$120.

Rogers:1987:TCG

[RE87] David F. Rogers and Rae A. Earnshaw, editors. Techniques for Computer Graphics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96492-4, 3-540-96492-4. viii + 512 pp. LCCN T385 .T43 1986. US$50.00.

Redfern:1993:MH

[Red93] Darren Redfern. The Maple Handbook. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-94054-5 (New York), 3-540-94054-5 (Berlin). 497 pp. LCCN QA76.95 .R43 1993. FF 181,00.

Redfern:1996:MH

[Red96] Darren Redfern. The Maple Handbook: Maple V Release 4. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 0-387-94538-5, 0-387-94622-5. 495 pp. LCCN QA76.95.R43 1996. US$29.00.

Rees:2000:JSN

[Ree00] Martin J. Rees. Just six numbers: the deep forces that shape the universe. Basic Books, New York, NY, USA, 2000. ISBN 0-465-03673-2. xi + 195 pp. LCCN QB 981 R434 2000.

Reeves:2008:FNF

[Ree08] Richard Reeves. A Force of Nature: the Frontier Genius of Ernest Rutherford. Great discoveries. W. W. Norton & Co., New York, NY, USA, 2008. ISBN 0-393-05750-X (hardcover), 0-393-33369-8 (paperback). 207 pp. LCCN QC16.R8 R44 2008.

Reed:2014:HSM

[Ree14a] Bruce Cameron Reed. The history and science of the Manhattan Project. Undergraduate Lecture Notes in Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-642-40296-8 (hardcover), 3-642-40297-6 (e-book). ISSN 2192-4791. xvi + 451 pp. LCCN QC702.7.H42; QC770-798; QC773.A1; QC793.5.H32-793.5.H329.

Reed:2014:PMP

[Ree14b] Bruce Cameron Reed. The physics of the Manhattan Project. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2014. ISBN 3-662-43532-2 (hardcover), 3-662-43533-0 (e-book). xvii + 222 pp. LCCN QC790 .R44 2015.

Rees:2018:FPH

[Ree18] Martin Rees. On the Future Prospects for Humanity. Princeton University Press, Princeton, NJ, USA, 2018. ISBN 0-691-17944-1. viii + 256 pp. LCCN ????

Reed:2020:MPS

[Ree20] Bruce Cameron Reed. Manhattan Project: The Story of the Century. Springer International Publishing, Cham, Switzerland, 2020. ISBN 3-030-45733-84b (hardcover), 3-030-45734-6 (e-book). xiv + 553 pp. LCCN QC773.3.U5 R4436 2020. URL http://www.

springerlink.com/content/
978-3-030-45734-1.

Reid:1988:PLP

[Rei88a] Glenn C. Reid. PostScript Language Program Design. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-14396-8. xii + 224 pp. LCCN QA76.73.P67 R45 1988. US$22.95. URL

http://www.rightbrain.com/
rightbrain.shtml.

Reischauer:1988:JTC

[Rei88b]  Edwin O. Reischauer.  The

Japanese Today: Change and Continuity. Belknap Press of Harvard University Press, Cambridge, MA and London, UK, 1988. ISBN 0-674-47181-4. viii + 426 pp. LCCN DS806.R35 1988. US$39.95.

Reid:1990:TP

[Rei90] Glenn C. Reid. Thinking in PostScript. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-52372-8. xiii + 221 pp. LCCN QA76.73.P67 R46 1990. US$22.95, CDN$29.95. URL

http://www.rightbrain.com/
download/books/ThinkingInPostScript.
pdf;   http://www.rightbrain.
com/pages/books.html;   http:
//www.rightbrain.com/rightbrain.
shtml.

Reid:2003:SSE

[Rei03] Jason Reid. Secure Shell in the Enterprise. Sun blueprints. Sun Microsystems Press, Palo Alto, CA, USA, 2003. ISBN 0-13-142900-0 (paperback). xxiii + 198 pp. LCCN QA76.76.O63 R448 2003. US$39.00.

Reid:2006:ZIW

[Rei06] Constance Reid. From Zero to Infinity: What Makes Numbers Interesting. A. K. Peters, Ltd., Wellesley, MA, USA, fifth edition, 2006. ISBN 1-56881-273-6. xvii + 188 pp. LCCN QA93 .R42 2006.

Reich:2009:PFH

[Rei09] Eugenie Samuel Reich. Plastic Fantastic: How the Biggest Fraud in Physics Shook the Scientific World. Palgrave Macmillan, New York, NY, USA, 2009. ISBN 0-230-22467-9. 266 pp. LCCN QC16.S26456 R45 2009. URL

http://www.loc.gov/catdir/
enhancements/fy0903/2008051801-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0903/
2008051801-d.html.

Renner:textyl

[Ren87] John S. Renner. TEXtyl: a linedrawing interface for TEX. Technical Report OSU-CISRC-4/87-TR9, Department of Computer Science, Ohio State University, March 1987.

Resag:2018:FHP

[Res18] Jörg Resag. Feynman and His Physics: The Life and Science of an Extraordinary Man. Springer Biographies. Springer International Publishing, Cham, Switzerland, 2018. ISBN 3-319-96835-1 (print), 3-319-96836-X (e-book), 3-319-96837-8. ISSN 2365-0613 (print), 2365-0621 (electronic). xii + 319 pp. LCCN QC16.F49 R4713 2018.

Ramabadran:crc-tutorial

[RG88] Tenkasi V. Ramabadran and Sunil S. Gaitonde. A tutorial on CRC computations. IEEE Micro, 8(4):62–75, July/August 1988.

Resnick:1977:PPO

[RH77] Robert Resnick and David Halliday. Physics, Part One. Wiley, New York, NY, USA, third edition, 1977. ISBN 0-471-71716-9. xi + 564 + A44 pp. LCCN QC21.2.R47 1977.

Rifkin:1988:UES

[RH88] Glenn Rifkin and George Harrar. The ultimate entrepreneur: the story of Ken Olsen and Digital Equipment Corporation. Contemporary Books, Chicago, IL, USA, 1988. ISBN 0-8092-4559-0. xii + 332 pp. LCCN HD9696.C62 O487 1988.

Rifkin:1990:UES

[RH90] Glenn Rifkin and George Harrar. The ultimate entrepreneur: the story of Ken Olsen and Digital Equipment Corporation. Prima Publishing, Rocklin, CA, USA, updated edition, 1990. ISBN 1-55958-022-4. xii + 336 pp.

LCCN HD9696.C62 O487 1990. US$10.95.

Riordan:1997:CFB

[RH97] Michael Riordan and Lillian Hoddeson. Crystal Fire: the Birth of the Information Age. W. W. Norton & Co., New York, NY, USA, 1997. ISBN 0-393-04124-7 (hardcover). x + 352 pp. LCCN TK7809 .R56 1997.

Riehm:2011:TTM

[RH11] Elaine McKinnon Riehm and Frances Hoffman. Turbulent Times in Mathematics: the Life of J. C. Fields and the History of the Fields Medal. American Mathematical Society, Providence, RI, USA, 2011. ISBN 0-8218-6914-0. xii + 257 + 16 pp. LCCN QA29.F54 R53 2011.

Rhodes:1986:MAB

[Rho86] Richard Rhodes. The Making of the Atomic Bomb. Simon and Schuster, New York, NY, USA, 1986. ISBN 0-671-44133-7 (paperback), 0-671-65719-4 (hardcover). 886 + 42 pp. LCCN QC773 .R46 1986.

Rhodes:1995:DSM

[Rho95] Richard Rhodes. Dark Sun: The Making of the Hydrogen Bomb. Sloan technology series. Simon and Schuster, New York, NY, USA, 1995. ISBN 0-684-80400-X, 0-684-82414-0 (paperback). 731 + 32 pp. LCCN UG1282.A8 R46

Rhodes:2015:HGC

[Rho15] Richard Rhodes. Hell and Good Company: the Spanish Civil War and the World It Made. Simon and Schuster, New York, NY, USA, 2015. ISBN 1-4516-9621-3 (hardcover), 1-4516-9623-X (ebook). xviii + 302 + 16 pp. LCCN DP269 .R457 2015.

Rhodes:2018:EHH

[Rho18] Richard Rhodes. Energy: a human history. Simon and Schuster, New York, NY, USA, 2018. ISBN 1-5011-0535-3 (hardcover), 1-5011-0537-X (e-book). xiv + 464 pp. LCCN TJ163.2 .R56 2018; TJ163.2 .R498 2018.

Rice:1981:MCM

[Ric81] John R. Rice. Matrix Computations and Mathematical Software. McGraw-Hill Computer Science Series. McGraw-Hill, New York, NY, USA, 1981. ISBN 0-07-052145-X. xii + 248 pp. LCCN QA188 .R52.

Rice:1983:NMS

[Ric83] John R. Rice. Numerical Methods, Software, and Analysis. Mc-Graw-Hill, New York, NY, USA, 1983. ISBN 0-07-052208-1. xii + 483 pp. LCCN QA297 .R49 1983. US$19.95.

Richelson:2006:SBA

[Ric06] Jeffrey Richelson. Spying on the bomb: American nuclear intelligence from Nazi Germany to Iran and North Korea. W. W. Norton & Co., New York, NY, USA, 2006. ISBN 0-393-05383-0 (hardcover). 702 pp. LCCN UB271.U5 R53 2006.

Rigden:2005:ESG

[Rig05] John S. Rigden. Einstein 1905:

the standard of greatness. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-02104-5 (paperback), 0-674-04275-1 (e-book). ix + 173 pp. LCCN QC16.E5 R54 2005.

Rimmer:1993:BMG

[Rim93]  Steve Rimmer.  Bit-Mapped

Graphics. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, second edition, 1993. ISBN 0-8306-4209-9 (hardcover), 0-8306-4208-0 (paperback). ix + 484 pp. LCCN T385 .R55 1993. US$26.95 (paperback), US$38.95 (hardcover). Explains the TIFF (Tag Image File Format) and includes sample code in C.

Ringholz:2002:UFS

[Rin02] Raye Carleson Ringholz. Uranium Frenzy: Saga of the Nuclear West. Utah State University Press, Logan, UT, USA, revised and expanded edition, 2002. ISBN 0-87421-432-7 (paperback). xiii + 344 pp. LCCN HD9539.U72 U5366 2002. URL

http://www.loc.gov/catdir/
enhancements/fy0809/2002009421-
d.html;   http://www.loc.
gov/catdir/toc/fy033/2002009421.
html.

Rowen:IEEE-MICRO-8-3-53

[RJR88] Chris Rowen, Mark Johnson, and Paul Ries. The MIPS R3010 floating-point coprocessor. IEEE Micro, 8(3):53–62, May/June 1988.

Rosenblum:2011:QEP

[RK11] Bruce Rosenblum and Fred Kuttner. Quantum Enigma: Physics Encounters Consciousness. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2011. ISBN 0-19-975381-4 (paperback). x + 287 pp. LCCN QC174.13 .R67 2011.

Rajchman:1942:REP

[RMV42] J. A. Rajchman, G. A. Morton, and A. W. Vance. Report on electronic predictors for antiaircraft fire control. Technical report, Research Laboratories, R. C. A. Manufacturing Company, Inc., Camden, NJ, USA, April 1942. Reprinted in [Ran82, §7.4].

Robertson:1968:RC

[RN68] H. P. Robertson and Thomas W. Noonan. Relativity and Cosmology. W. B. Saunders Company, Philadelphia, PA, USA, 1968. xxxiii + 456 pp. LCCN QC6 .R635.

Roberts:1959:NMR

[Rob59] John D. Roberts. Nuclear Magnetic Resonance: Applications to Organic Chemistry. McGraw-Hill, New York, NY, USA, 1959. viii + 118 pp. LCCN QD591 .R6 1959.

Robert:1989:RRO

[Rob89] Major Henry M. Robert. Robert’s Rules of Order. Berkley Books, New York, NY, USA, 1989. ISBN 0-425-11690-5. LCCN JF515 .R692 1989.

Robbins:1996:EAP

[Rob96] Arnold Robbins. Effective AWK Programming: a User’s Guide for GNU AWK. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1996. ISBN 0-916151-88-3. viii + 322 pp. LCCN QA76.73.A95 R63 1996. US$27.00. URL http://

www.ssc.com/ssc/eap/.

Robbins:1997:EAP

[Rob97] Arnold Robbins. Effective AWK Programming. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, second edition, 1997. ISBN 1-57831-000-8. x + 322 pp. LCCN QA76.73.A95 R63 1997. US$27.00. URL http://www.

ssc.com/ssc/eap/.

Robbins:1999:UND

[Rob99]  Arnold Robbins.  UNIX in

a Nutshell: a Desktop Quick Reference for SVR4 and Solaris 7. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 1999. ISBN 1-56592-427-4 (paperback). xvi + 598 pp. LCCN QA76.76.O63 R623 1999. US$24.95. URL http:

//www.oreilly.com/catalog/
unixnut3.

Robbins:2001:EAP

[Rob01] Arnold Robbins. Effective AWK Programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2001. ISBN 0-596-00070-7 (paperback), 0-596-80537-3 (e-book). xxiv + 421 pp. LCCN QA76.73.A95 R63 2001. US$39.95. URL http:

//www.oreilly.com/catalog/
awkprog3/.

Robbins:2004:LPE

[Rob04] Arnold Robbins. Linux Programming by Example: The Fundamentals. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-142964-7. xxxii + 687 pp. LCCN QA76.76.O63 R568 2004. US$39.99, CAN$57.99.

Robbins:2005:GPG

[Rob05a] Arnold Robbins. GDB Pocket Guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-10027-2 (paperback), 0-596-52866-3 (ebook), 0-596-55335-8 (e-book). v + 69 pp. LCCN QA76.9.D43 R63 2005. US$9.95, CAN$13.95.

Robbins:2005:UN

[Rob05b] Arnold Robbins. Unix in a Nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2005. ISBN 0-596-10029-9 (paperback), 0-596-52948-1 (e-book), 1-4493-9114-1 (e-book). xviii + 885 pp. LCCN QA76.76.O63 R63 2005. US$34.95, CAN$48.95, UK£24.95. URL http:/

/www.oreilly.com/catalog/
unixnut4/index.html.

Robinson:2005:EHY

[Rob05c] Andrew Robinson. Einstein:

a hundred years of Relativity. Harry N. Abrams, New York, NY, USA, 2005. ISBN 0-8109-5923-2 (hardcover). 256 pp. LCCN QC16.E5 R63 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip059/2005006593.html.

Rochkind:1985:AUP

[Roc85] Marc J. Rochkind. Advanced UNIX Programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-011818-4 (hardcover), 0-13-011800-1 (paperback). xv + 265 pp. LCCN QA76.76.O63 R63 1985. US$32.95 (hardcover), US$24.95 (paperback).

Rochkind:2004:AUP

[Roc04] Marc J. Rochkind. Advanced UNIX Programming. Addison-Wesley, Reading, MA, USA, second edition, 2004. ISBN 0-13-141154-3. xiii + 719 pp. LCCN QA76.76.O63 R63 2004. US$44.99. URL http://www.

aw-bc.com/catalog/academic/
product/0,1144,0131411543,
00.html.

Rockmore:2006:SRH

[Roc06] Daniel N. (Daniel Nahum) Rockmore. Stalking the Riemann hypothesis: the quest to find the hidden law of prime numbers.

Vintage Books, New York, NY, USA, 2006. ISBN 0-375-72772-8 (paperback). x + 292 pp. LCCN QA246 .R63 2006.

Rodgers:1990:UDM

[Rod90] Ulka Rodgers. UNIX Database Management Systems. Yourdon Press, Englewood Cliffs, NJ 07632, USA, 1990. ISBN 0-13-945593-0. xiv + 338 pp. LCCN QA76.9.D3 R65 1990.

Rodgers:2019:TAS

[Rod19] Glen E. Rodgers. Travelling with the Atom: a Scientific Guide to Europe and Beyond. Royal Society of Chemistry, Cambridge, UK, 2019. ISBN 1-78801-528-2 (paperback), 1-78801-702-1 (ebook). xxxii + 551 pp. LCCN QC171.2 .R63 2020.

Rogers:1985:PEC

[Rog85] David F. Rogers. Procedural Elements for Computer Graphics. McGraw-Hill, New York, NY, USA, 1985. ISBN 0-07-053534-5. xiii + 433 pp. LCCN T385 .R631 1985. US$19.95.

Rogers:2001:INH

[Rog01] David F. Rogers. An Introduction to NURBS: with Historical Perspective. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2001. ISBN 1-55860-669-6 (hardcover), 0-08-050920-7 (ebook), 1-281-03529-7. xvii + 324 pp. LCCN QA224.R64 2001. US$49.95.

Rogers:2011:EE

[Rog11] Alan R. Rogers. The Evidence for Evolution. University of Chicago Press, Chicago, IL, USA and London, UK, 2011. ISBN 0-226-72380-1 (hardcover), 0-226-72382-8 (paperback). 120 pp. LCCN QH361 .R64 2011.

Rohl:1984:RP

[Roh84] J. S. (Jeffrey Soden) Rohl. Recursion via Pascal, volume 19 of Cambridge Computer Science Texts, Editors: E. S. Page and C. M. Reeves and D. E. Conway. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-26329-8 (hardcover), 0-521-26934-2 (paperback). x + 191 pp. LCCN QA76.73.P2 R634 1984.

Rohl:2014:KWI

[Röh14] John C. G. Röhl. Kaiser Wilhelm II, 1859–1941: a concise life. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-07225-5 (hardcover), 1-107-42077-6 (paperback), 1-139-68039-0 (e-book). xxii + 240 pp. LCCN DD229 .R6412413 2014. Translation from the German original by Sheila De Bellaigue.

Rost:pex

[Ros88] Randi J. Rost. PEX introduction and overview. Technical Report Version 3.20, Digital Equipment Corporation, Workstation Systems Engineering, April 1988. This document is present in the X Window System Version 11 Release 3 in the file X11/

                                        X11/doc/extensions/pex/doc/
                                        intro/doc.ms.

Rose:1991:SB

[Ros91] Marshall T. Rose. The Simple Book. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-812611-9. xxix + 347 pp. LCCN TK5105.5.R68 1991.

Rosenblatt:1993:LKS

[Ros93] Bill Rosenblatt. Learning the Korn Shell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, June 1993. ISBN 1-56592-054-6. xxii + 338 pp. LCCN QA76.73.K67 R68 1993. US$27.95.

Rosenberg:2005:STW

[Ros05] Barry J. Rosenberg. Spring Into Technical Writing: for Engineers and Scientists. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-13-149863-0 (paperback). xxi + 318 pp. LCCN T11 .R663 2005.

Rosen:2010:MPI

[Ros10] William Rosen. The most powerful idea in the world: a story of steam, industry, and invention. Random House, New York, NY, USA, 2010. ISBN 1-4000-6705-7. xxv + 370 pp. LCCN TJ461 .R67 2010. US$28.00.

Rosbottom:2014:WPW

[Ros14] Ronald C. Rosbottom. When Paris went dark: the City of Light under German Occupation, 1940–1944. Little, Brown and Company, New York, NY, USA, 2014. ISBN 0-316-21743-3 (paperback), 0-316-21744-1 (hardcover). xxxii + 447 pp. LCCN D762.P3 R67 2014.

Rosbottom:2019:SCY

[Ros19]  Ronald Rosbottom.  Sudden

Courage: Youth in France Confront the Germans, 1940–1945. HarperCollins Publishers, New York, NY, USA, 2019. ISBN 0-06-247002-7 (hardcover), 0-06-247005-1 (e-book). xiv + 320 + 16 pp. LCCN D802.F8 R593 2019.

Roth:1988:RWP

[Rot88] Stephen E. Roth, editor. Real World PostScript. Addison-Wes-ley, Reading, MA, USA, 1988. ISBN 0-201-06663-7. xiv + 383 pp. LCCN Z286.D47 R4 1988.

Rovelli:2021:HMS

[Rov21] Carlo Rovelli. Helgoland: Making Sense of the Quantum Revolution. Riverhead Books, New York, NY, USA, 2021. ISBN 0-593-32888-4 (hardcover), 0-593-32890-6 (e-book). xviii + 235 pp. LCCN QC173.96 .R6813 2021. Translation to English by Erica Segre and Simon Carnell.

Roget:1990:RTE

[RR90] Peter Mark Roget and Samuel Romilly Roget. Roget’s thesaurus of English words and phrases. Portland House, New York, NY, USA, 1990. ISBN 0-517-03552-9. xxv + 800 pp. LCCN PE1591 .R7 1990. US$9.99.

Robbins:1996:PUP

[RR96] Kay A. Robbins and Steven Robbins. Practical UNIX Programming: a Guide to Concurrency, Communication, and Multithreading. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-443706-3. xiv + 658 pp. LCCN QA76.76.O63 R615 1996.

Rosenblatt:2002:LKS

[RR02] Bill Rosenblatt and Arnold Robbins. Learning the Korn Shell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00195-9 (paperback), 1-4493-7127-2 (e-book), 1-4493-7128-0 (e-book). xviii + 412 pp. LCCN QA76.73.K67 R68 2002. US$34.95.

Robbins:2003:USP

[RR03] Kay A. Robbins and Steven Robbins. UNIX Systems Programming: Communication, Concurrency, and Threads. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 2003. ISBN 0-13-042411-0 (hardcover). xvii + 893 pp. LCCN QA76.76.O63 R6215 2003.

Roads:1987:FCM

[RS87] Curtis Roads and John Strawn, editors. Foundations of Computer Music. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-18114-2. xiii + 712 pp. LCCN ML1092 .F7 1985.

Rabinowitz:1990:PC

[RS90] Henry Rabinowitz and Chaim Schaap. Portable C. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-685967-4. xi + 269 pp. LCCN QA 76.73 C15 R33 1990.

Robinson:1991:NDC

[RS91] Sinclair Robinson and Donald Smith. NTC’s Dictionary of Canadian French (NTC’s Language Definition). NTC/Contemporary Publishing Company, Lincolnwood, IL, USA, 1991. ISBN 0-8442-1486-

8. 300 pp. LCCN PC3637 .R58 1991.

Reed:2009:NEP

[RS09a] Thomas C. Reed and Danny B.

Stillman. The Nuclear Express: a Political History of the Bomb and Its Proliferation. Zenith Press, Minneapolis, MN, 2009. ISBN 0-7603-3502-8. viii + 392 pp. LCCN U264 .R44 2009.

Robson:2009:OHH

[RS09b] Eleanor Robson and Jacqueline A. Stedall, editors. The Oxford Handbook of the History of Mathematics. Oxford handbooks. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-921312-7. vii + 918 pp. LCCN QA21 .O94 2009. URL

http://www.loc.gov/catdir/
enhancements/fy0907/2008031793-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0907/
2008031793-d.html;   http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0823/2008031793.html.

Ritchie:time-sharing74

[RT74] Dennis W. Ritchie and Ken Thompson. The UNIX timesharing system. Communications of the Association for Computing Machinery, 17(7):365–375, July 1974.

Ritchie:time-sharing79

[RT79] Dennis W. Ritchie and Ken Thompson. The UNIX timesharing system. In UNIX Programmer’s Manual, volume 2, pages 20–35. Holt, Rinehart, and Winston, New York, NY, USA, 1979.

Rubinstein:1988:DTI

[Rub88] Richard Rubinstein. Digital Typography: An Introduction to Type and Composition for Computer System Design. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-17633-5. xi + 340 pp. LCCN Z253.3 .R8 1988.

Ruckert:2015:MSS

[Ruc15] Martin Ruckert. The MMIX supplement: supplement to The Art of Computer Programming, volumes 1, 2, 3 by Donald E. Knuth. Addison-Wesley, Reading, MA, USA, 2015. ISBN 0-13-399231-4 (paperback), 0-13-399289-6. xxi + 193 pp. LCCN QA76.6 .K64 2005 Suppl. 1. URL http://

                                        mmix.cs.hm.edu/.

Rudman:2007:HMH

[Rud07] Peter Strom Rudman. How mathematics happened: the first 50,000 years. Prometheus Books, Amherst, NY, USA, 2007. ISBN 1-59102-477-3. 314 pp. LCCN QA22 .R86 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0615/2006020255.html.

Ruelle:2007:MB

[Rue07] David Ruelle. The Mathematician’s Brain: a Personal Tour Through the Essentials of Mathematics and Some of the Great Minds Behind Them. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-12982-7. ix + 160 pp. LCCN QA8.4 .R84 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0726/2006049700-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0731/
2006049700-t.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0734/2006049700-b.html.

Rodale:1978:SF

[RUL78] J. I. Rodale, Laurence Urdang, and Nancy LaRoche. The Synonym Finder. Rodale Press, Emmaus, PA, 1978. ISBN 0-87857-236-8. 1361 pp. LCCN PE1591.R64 1978. Completely revised by Laurence Urdang and Nancy LaRoche.

Rabinowitz:2009:MSS

[RV09] Harold Rabinowitz and Suzanne Vogel, editors. The manual of scientific style: a guide for authors, editors, and researchers. Academic Press/Elsevier, Amsterdam, The Netherlands, 2009. ISBN 0-12-373980-2. xii + 968 pp. LCCN T11 .M36 2009. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0912/2009278258-
                                        d.html.

Raade:2004:MHS

[RW04] Lennart Råde and Bertil Westergren. Mathematics Handbook for Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fifth edition, 2004. ISBN 3-540-21141-1 (hardcover). 562 pp. LCCN QA41 .R34 2004. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0818/2006286513-
                                        d.html;   http://www.loc.
                                        gov/catdir/toc/fy0704/2006286513.
                                        html.

Rao:2001:TDC

[RY01] K. Ramamohan (Kamisetty Ramamohan) Rao and P. C. (Pat C.) Yip, editors. The Transform and Data Compression Handbook. Electrical engineering and signal processing series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2001. ISBN 0-8493-3692-9, 1-315-22052-0, 1-351-82772-3, 1-4200-3738-2 (ebook). xix + 388 pp. LCCN TK5105 .T72 2001. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0646/00057149-
                                        d.html.

Ryder:pfort

[Ryd74] Barbara G. Ryder. The PFORT verifier. Software—Practice and Experience, 4:359–377, 1974.

Sobel:1998:IL

[SA98] Dava Sobel and William J. H.

Andrewes. The Illustrated Longitude. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 1998. ISBN 0-8027-1344-0. 216 pp. LCCN QB 225 S67 1998. US$33.00.

Saari:2001:CEM

[Saa01] Donald G. Saari. Chaotic Elections!: a Mathematician Looks at Voting. American Mathematical Society, Providence, RI, USA, 2001. ISBN 0-8218-2847-9 (paperback). xiii + 159 pp. LCCN JF1001 .S227 2001.

Sabbagh:2003:RHG

[Sab03] Karl Sabbagh. The Riemann Hypothesis: the Greatest Unsolved Problem in Mathematics. Farrar, Straus and Giroux, New York, NY, USA, 2003. ISBN 0-374-25007-3, 0-374-52935-3. viii + 340 pp. LCCN QA241 .S23 2003. US$25.00. Originally published in 2002 as Dr. Riemann’s zeroes by Grove Atlantic, London, UK.

Salus:1994:QCU

[Sal94] Peter H. Salus. A Quarter Century of UNIX. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-54777-5. xii + 256 pp. LCCN QA76.76.O63 S342 1994.

URL http://wiki.tuhs.org/
doku.php?id=publications:
quarter_century_of_unix.

Salomon:1995:AT

[Sal95]  David Salomon.  The Ad-

vanced TEXbook. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94556-3. xx + 490 pp. LCCN Z253.4.T47 S25 1995. US$39.95. URL http://www.

booksbydavidsalomon.com/;
http://www.ecs.csun.edu/~
dxs/tatb.advertis/tatbAd.
html.

Salomon:1998:DCC

[Sal98] David Salomon. Data Compression: The Complete Reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98280-9. xx + 427 pp. LCCN QA76.9.D33S25 1997. US$39.95. URL http://www.

booksbydavidsalomon.com/;
http://www.ecs.csun.edu/~
dxs/DCadvertis/DcompAd.html.

Salomon:1999:CGG

[Sal99]  David Salomon.  Computer

Graphics and Geometric Modeling. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98682-0. xviii + 851 pp. LCCN T385 S243 1999. URL http://www.

booksbydavidsalomon.com/;
http://www.ecs.csun.edu/~
dxs/CGGMadvertis/CGad.html.

Salomon:2000:DCC

[Sal00] David Salomon. Data Compression: The Complete Reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 0-387-95045-1. xvi + 823 pp. LCCN QA76.9.D33 S25 2000. US$39.95. URL http://www.

booksbydavidsalomon.com/;
http://www.ecs.csun.edu/
~dxs/DCadvertis/Dcomp2Ad.
html.

Salomon:2002:GDC

[Sal02] David Salomon. A Guide to Data Compression Methods. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95260-8 (paperback). xii + 295 pp. LCCN QA76.9.D33

S28 2001. URL http://www.
booksbydavidsalomon.com/.

Salomon:2003:DPS

[Sal03]  David Salomon.  Data Pri-

vacy and Security. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 0-387-00311-8. xiv + 465 pp. LCCN QA76.9.A25 S265 2003. US$59.95. URL http://www.

booksbydavidsalomon.com/.

Salomon:2004:DCC

[Sal04] David Salomon. Data Compression: The Complete Reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2004. ISBN 0-387-40697-2. xx + 898 pp. LCCN QA76.9.D33

S25 2004. URL http://www.
booksbydavidsalomon.com/;
http://www.ecs.csun.edu/~
dsalomon/DC3advertis/DComp3Ad.
html.

Salomon:2005:CDC

[Sal05a] David Salomon. Coding for Data and Computer Communications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-21245-0. xv + 548 pp. LCCN TK5102.94 .S35 2005. URL http://www.

DavidSalomon.name/Codes/Codes.
html;   http://www.ecs.csun.
edu/~dsalomon/.

Salomon:2005:FCS

[Sal05b] David Salomon. Foundations of Computer Security. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 1-84628-193-8 (hardcover), 1-84628-341-8. xxi + 368 pp. LCCN QA76.9.A25 S2656 2005.

URL http://www.springer.
com/sgw/cda/frontpage/0,11855,
4-40007-22-65173048-0,00.
html;   http://www.springer.
com/sgw/cda/pageitems/document/
cda_downloaddocument/0,11855,
0-0-45-166687-p65173048,00.
pdf;   http://www.springer.
com/sgw/cda/pageitems/document/
cda_downloaddocument/0,11855,
0-0-45-166688-p65173048,00.
pdf;   http://www.springer.
com/sgw/cda/pageitems/document/
cda_downloaddocument/0,11855,
0-0-45-166689-p65173048,00.
pdf.

Salomon:2006:CSC

[Sal06a] David Salomon. Curves and Surfaces for Computer Graphics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-24196-5, 0-387-28452-4 (ebook). xvi + 460 pp. LCCN T385 .S2434 2005.

Salomon:2006:TPC

[Sal06b] David Salomon. Transformations and Projections in Computer Graphics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 1-84628-392-2. xiv + 288 pp. LCCN T385 .S2439 2006. US$59.95.

Salomon:2007:DCC

[Sal07a] David Salomon. Data Compression: The Complete Reference. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2007. ISBN 1-84628-602-6. xxv + 1092 pp. LCCN QA76.9.D33 S25 2007. With contributions by Giovanni Motta and David Bryant.

Salomon:2007:VLC

[Sal07b] David Salomon. Variable-length Codes for Data Compression. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-84628-958-0. xii + 191 pp. LCCN QA76.9.D33 S25 2007.

Salomon:2008:CID

[Sal08] David Salomon. A Concise Introduction to Data Compression. Undergraduate topics in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 1-84800-071-5 (paperback). xiii + 310 pp. LCCN QA76.9.D33 S34 2008. EUR 32.

Salomon:2011:CGM

[Sal11] David Salomon. The Computer Graphics Manual. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 0-85729-885-2 (print), 0-85729-886-0 (e-book). ISSN 1868-0941 (print), 1868-095X (electronic). xxiii + 1–727 + 5 + xxiii + 1–6 + 731–1503 + 5 pp. LCCN T385 .S25 2011.

Sandorfy:1964:ESQ

[San64] C. (Camille) Sandorfy. Electronic Spectra and Quantum Chemistry. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1964. xiii + 385 pp. LCCN QD95 .S218 1959. Originally published in French as Les Spectres Electroniques en Chimie Theorique, 1959.

Sandifer:2007:EML

[San07] Charles Edward Sandifer. The Early Mathematics of Leonhard Euler, volume 1 of Spectrum series; MAA tercentenary Euler celebration. Mathematical Association of America, Washington, DC, USA, 2007. ISBN 0-88385-559-3. xix + 391 pp. LCCN QA29.E8 S26 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0743/2006933948-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0743/
2006933948-t.html.

Sayood:1996:IDC

[Say96] Khalid Sayood. Introduction to Data Compression. Morgan Kaufmann Publishers, 2929 Campus Drive, Suite 260, San Mateo, CA 94403, USA, 1996. ISBN 1-55860-346-8. xviii + 475 pp. LCCN TK5102.92 .S39 1996.

Sayood:2000:IDC

[Say00] Khalid Sayood. Introduction to Data Compression. Morgan Kaufmann Publishers, 2929 Campus Drive, Suite 260, San Mateo, CA 94403, USA, second edition, 2000. ISBN 1-55860-558-4. xx + 636 pp. LCCN TK5102.92 .S39 2000. US$69.95. URL http:/

/www.mkp.com/books_catalog/
catalog.asp?ISBN=1-55860-
558-4.

Sayood:2005:IDC

[Say05] Khalid Sayood. Introduction to Data Compression. Elsevier, Amsterdam, The Netherlands, third edition, 2005. ISBN 0-12-620862-X (hardcover), 0-08-050925-8 (e-book). xxii + 680 pp. LCCN TK5102.92 .S39 2005.

Suri:2007:CAM

[SB07] Gaurav Suri and Hartosh Singh Bal. A Certain Ambiguity: a Mathematical Novel. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-12709-3 (hardcover). viii + 281 pp. LCCN QA99 .S87 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0726/2007009495-
b.html;   http://www.loc.

gov/catdir/enhancements/fy0726/
2007009495-d.html.

Schoonover:1992:GEU

[SBA92] Michael A. Schoonover, John S. Bowie, and William R. Arnold. GNU Emacs: UNIX Text Editing and Programming. Hewlett-Packard Press series. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56345-2. xxvii + 609 pp. LCCN QA76.76.T49S36.

Smith:1976:MER

[SBD^{+}76] B. T. Smith, J. M. Boyle, J. J. Dongarra, B. S. Garbow, Y. Ikebe, V. C. Klema, and C. B. Moler. Matrix Eigensystem Routines: EISPACK Guide, volume 6 of Lecture Notes in Computer Science, Editors: G. Goos and J. Hartmanis. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. ISBN 0-387-06710-8. vii + 387 pp. LCCN QA193 .M37, QA267.A1

L43 no.6. URL http://link.
springer.com/10.1007/3-540-
07546-1.

Syropoulos:2004:TXD

[SBH^{+}04] Apostolos Syropoulos, Karl Berry, Yannis Haralambous, Baden Hughes, Steven Peter, and John Plaice, editors. TEX, XML, and Digital Typography: International Conference on TEX, XML, and Digital Typography, held jointly with the 25th Annual Meeting of the TEX Users Group, TUG 2004, Xanthi, Greece, August 30– September 3, 2004: Proceedings, volume 3130 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNCSD9. ISBN 3-540-22801-2 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN Z253.3 I58 2004. URL http:/

/link.springer-ny.com/link/
service/series/0558/tocs/
t3130.htm;   http://www.springerlink.
com/openurl.asp?genre=issue&
issn=0302-9743&volume=3130;
 http://www.springerlink.
com/openurl.asp?genre=volume&
id=doi:10.1007/b99374.

Siewiorek:1982:CSP

[SBN82] Daniel P. Siewiorek, C. Gordon Bell, and Allen Newell. Computer Structures: Principles and Examples. McGraw-Hill computer science series. McGraw-Hill, New York, NY, USA, 1982. ISBN 0-07-057302-

6. xvi + 926 pp. LCCN QA76.9.A73 C65. URL https:

//archive.computerhistory.
org/resources/text/bell_gordon/
bell.computer_structures_
principles_and_examples.1982.
102630397.pdf.

Scanlon:1984:ALP

[Sca84] Leo J. Scanlon. 8086/88 Assembly Language Programming. Robert J. Brady Co., Bowie, MD 20715, USA, 1984. ISBN 0-89303-424-X. ix + 213 pp. LCCN QA76.8.I292 S29 1984. A rather short treatment of assembly language programming for the 8088 and 8086, with a short chapter on the 8087. Nothing specific to the IBM PC, and not detailed enough for a beginner. Two related books [Bra84, Laf84] contain much more detail.

Schilpp:1949:AEPa

[Sch49a] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 1 of The Library of Living Philosophers. Cambridge University Press, Cambridge, UK, 1949. ISBN 0-87548-286-4. ISSN 0075-9139. xviii + 781 pp. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Schilpp:1949:AEPb

[Sch49b] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 2 of The Library of Living Philosophers. Cambridge University Press, Cambridge, UK, 1949. ISBN 0-87548-286-4. ISSN 0075-9139. xviii + 781 pp. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Schachter:1983:CIG

[Sch83] Bruce J. Schachter, editor. Computer Image Generation. Wiley, New York, NY, USA, 1983. ISBN 0-471-87287-3. xx + 236 pp. LCCN T385 .C5934 1983. US$29.95.

Schendel:1984:INM

[Sch84] U. Schendel. Introduction to Numerical Methods for Parallel Computers. Ellis Horwood Series in Mathematics and its Applications, Editor: G. M. Bell. Ellis Horwood, New York, NY, USA, 1984. ISBN 0-470-20091-X, 0-85312-597-X. 151 pp. LCCN QA297 .S3813 1984. UK£15.00.

Schwarz:1988:ET

[Sch88] Norbert Schwarz. Einführung in TEX. Addison-Wesley, Reading, MA, USA, second edition, 1988. ISBN 3-925118-97-7. 272 pp. DM 68.

Scheifler:1989:XPR

[Sch89a] Robert W. Scheifler. X Protocol Reference Manual, volume 0. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1989. ISBN 0-937175-40-4. xv + 398 pp. LCCN QA76.76.W56 X5 1989.

Schofield:1989:OFPa

[Sch89b] Clive F. (Clive F.) Schofield. Optimising Fortran programs. Ellis Horwood series in computers and their applications. Ellis Horwood and Halsted Press, New York, NY, USA and New York, USA, July 1989. ISBN 0-7458-0654-6, 0-470-21533-X (Halsted Press). x + 309 pp. LCCN QA76.73.F25 S334 1989 WITHDRAWN. US$40.95, US$74.95. URL http://

www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=

047021533X. Cover title: Optimizing Fortran programs. Includes index. Bibliography: p. [293]-296.

Schildt:1990:AAC

[Sch90] Herbert Schildt. The Annotated ANSI C Standard: American National Standard for Programming Languages: C. Osborne/McGraw-Hill, Berkeley, CA, USA, 1990. ISBN 0-07-881952-0. xvi + 219 pp. LCCN QA76.73.C15S356 1990. US$39.95.

Schroeder:1991:FCP

[Sch91] Manfred Schroeder. Fractals, Chaos, Power Laws. W. H. Freeman and Company, New York, NY, USA, 1991. ISBN 0-671-74217-5, 0-7167-2136-8, 0-7167-2357-3. xviii + 429 pp. LCCN QD921 .S3 1990.

Schneier:1994:ACP

[Sch94] Bruce Schneier. Applied Cryptography: Protocols, Algorithms, and Source Code in C. Wiley, New York, NY, USA, 1994. ISBN 0-471-59756-2 (paperback). xviii + 618 pp. LCCN QA76.9.A25S35 1993. US$44.95.

Schneier:1996:ACP

[Sch96] Bruce Schneier. Applied Cryptography: Protocols, Algorithms, and Source Code in C. Wiley, New York, NY, USA, second edition, 1996. ISBN 0-471-12845-7 (hardcover), 0-471-11709-9 (paperback). xxiii + 758 pp. LCCN QA76.9.A25 S35 1996.

Schechter:1998:MBO

[Sch98] Bruce Schechter. My Brain is Open: the Mathematical Journeys of Paul Erdős. Simon and Schuster, New York, NY, USA, 1998. ISBN 0-684-84635-7. 224 pp. LCCN QA29.E86 S34 1998. US$13.00.

Schneier:2000:SLD

[Sch00] Bruce Schneier. Secrets and Lies:

Digital Security in a Networked World. Wiley, New York, NY, USA, 2000. ISBN 0-471-25311-1. xv + 412 pp. LCCN QA76.9.A25 S352 2000. US$29.99.

Schein:2003:DDL

[Sch03a] Edgar H. Schein, editor. DEC is dead, long live DEC: the lasting legacy of Digital Equipment Corporation. Berrett-Koehler, San Francisco, CA, USA, 2003. ISBN 1-57675-225-9. xv + 319 pp. LCCN HD9696.2.U64 D543 2003. URL

http://www.loc.gov/catdir/
enhancements/fy0715/2003041788-
d.html.

Schneier:2003:BFT

[Sch03b] Bruce Schneier. Beyond Fear:

Thinking Sensibly About Security in an Uncertain World. Copernicus Books, New York, NY, USA, 2003. ISBN 0-387-02620-7 (hardcover), 0-387-21712-6 (e-book). viii + 295 pp. LCCN HV6432 .S36 2003.

Schonfeld:2003:JH

[Sch03c] Roger C. Schonfeld. JSTOR:

a history. Princeton University Press, Princeton, NJ, USA, 2003. ISBN 0-691-11531-1. xxxiv + 412 pp. LCCN PN4836 .S36 2003. URL http://www.

loc.gov/catdir/description/

                                        prin031/2002035907.html;
                                        http://www.loc.gov/catdir/
                                        enhancements/fy0654/2002035907-
                                        b.html;   http://www.loc.
                                        gov/catdir/samples/prin031/
                                        2002035907.html.

Schweber:2008:EOM

[Sch08] S. S. (Silvan S.) Schweber. Einstein and Oppenheimer: the meaning of genius. Harvard University Press, Cambridge, MA, USA, 2008. ISBN 0-674-02828-7. xiv + 412 pp. LCCN QC16.E5 S3285 2008. URL http://www.

                                        loc.gov/catdir/toc/ecip082/
                                        2007043108.html.

Scheinerman:2011:MN

[Sch11] Edward R. Scheinerman. Mathematical notation. CreateSpace, USA, 2011. ISBN 1-4662-3052-5. viii + 84 pp. LCCN QA41 .S34 2011.

Schneier:2012:LOE

[Sch12a] Bruce Schneier. Liars and outliers: enabling the trust that society needs to thrive. Wiley, New York, NY, USA, 2012. ISBN 1-118-14330-2 (hardcover). xiv + 366 pp. LCCN BJ1500.T78 S35 2012.

Schweber:2012:NFM

[Sch12b] S. S. (Silvan S.) Schweber. Nuclear forces: the making of the physicist Hans Bethe. Harvard University Press, Cambridge, MA, USA, 2012. ISBN 0-674-06587-5 (hardcover). viii + 579 + 12 pp. LCCN QC774.B4 S39 2012. £25.95.

Schewe:2013:MGP

[Sch13] Phillip F. Schewe. Maverick Genius: the Pioneering Odyssey of Freeman Dyson. Thomas Dunn Books, St. Martin’s Press, New York, NY, USA, 2013. ISBN 0-312-64235-0 (hardcover), 1-250-02101-4 (e-book). x + 339 pp. LCCN QC16.D95 S34 2013.

Schwichtenberg:2015:PS

[Sch15a] Jakob Schwichtenberg. Physics from Symmetry. Undergraduate lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-19200-0, 3-319-66631-2 (ebook). xix + 279 pp. LCCN QC174.17.S9 S4 2015eb.

Schwinger:2015:AM

[Sch15b]  Julian Schwinger.  On An-

gular Momentum. Dover, New York, NY, USA, 2015. ISBN 0-486-78810-5 (paperback). 51 pp. LCCN QC793.3.A5 S39 2015. URL

http://www.loc.gov/catdir/
enhancements/fy1503/2014036399-
b.html;   http://www.loc.
gov/catdir/enhancements/fy1503/
2014036399-d.html.

Scott:1982:IIC

[Sco82] Joan E. Scott. Introduction to Interactive Computer Graphics. Wiley, New York, NY, USA, 1982. ISBN 0-471-05773-8 (hardcover), 0-471-86623-7 (paperback). xi + 255 pp. LCCN T385.S36.

Scott:2000:PLP

[Sco00] Michael Lee Scott. Programming Language Pragmatics. Morgan Kaufmann Publishers, San Francisco, CA, USA, 2000. ISBN 1-55860-578-9 (paperback), 1-55860-442-1 (hardcover). xxi + 856 pp. LCCN QA76.7 .S38 2000. URL http://www.

                                        loc.gov/catdir/description/
                                        els033/99047125.html;   http:
                                        //www.loc.gov/catdir/toc/
                                        els033/99047125.html.

Scott:2006:PLP

[Sco06] Michael Lee Scott. Programming Language Pragmatics. Morgan Kaufmann Publishers, San Francisco, CA, USA, second edition, 2006. ISBN 0-12-633951-1. xxxi + 875 pp. LCCN QA76.7 .S38 2006. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0632/2006272697-
                                        d.html.

Searfoss:1994:JKC

[Sea94] Glenn Searfoss. JIS-Kanji Character Recognition: Featuring the Gaiji method. Van Nostrand Reinhold, New York, NY, USA, 1994. ISBN 0-442-01813-4. xxii + 419 pp. LCCN TA1640 .S43 1994. US$49.95.

Seacord:2005:SCC

[Sea05]  Robert C. Seacord.  Secure

coding in C and C++. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-321-33572-4 (paperback). xxiv + 341 pp. LCCN QA76.9.A25 S368

2005. URL http://www.cert.
org/books/secure-coding/;
http://www.loc.gov/catdir/
toc/ecip0513/2005015012.html.

Seacord:2020:EC

[Sea20] Robert Seacord. Effective C.

No Starch Press, San Francisco, CA, USA, 2020. ISBN 1-71850-104-8 (paperback), 1-0981-2567-3 (e-book). xxviii + 240 pp. LCCN QA76.73.C15 .S433 2020.

URL https://nostarch.com/
download/samples/EffectiveC_
ToC_sample.pdf;   https://
nostarch.com/Effective_C.

Sedgewick:1983:A

[Sed83] Robert Sedgewick. Algorithms.

Addison-Wesley, Reading, MA, USA, 1983. ISBN 0-201-06672-6. viii + 551 pp. LCCN QA76.6 .S435 1983.

Sedgewick:1988:A

[Sed88] Robert Sedgewick. Algorithms.

Addison-Wesley, Reading, MA, USA, second edition, 1988. ISBN 0-201-06673-4. xii + 657 pp. LCCN QA76.6 .S435 1988. US$34.95.

Sedgewick:1990:AC

[Sed90] Robert Sedgewick. Algorithms in C. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-51425-7. xii + 657 pp. LCCN QA76.73.C15 S43 1990.

Sedgewick:1992:AC

[Sed92] Robert Sedgewick. Algorithms in C++. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-36118-3, 0-201-51059-6. xiv + 656 pp. LCCN QA76.73.C153 S38 1992.

Sedgewick:1993:AM

[Sed93] Robert Sedgewick. Algorithms in Modula-3. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-53351-0. xiv + 656 pp. LCCN QA76.73.M63 S43 1993.

Sedgewick:1998:ACP

[Sed98] Robert Sedgewick. Algorithms in C: Parts 1–4: Fundamentals, data structures, sorting, searching. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-31452-5 (paperback). xvii + 702 pp. LCCN QA76.73.C15 S43 1998. US$33.50.

Sedgewick:2002:ACP

[Sed02] Robert Sedgewick. Algorithms in C: Part 5: Graph algorithms. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-31663-3 (paperback). xiii + 482 pp. LCCN QA76.73.C15 S43 2002. US$34.99.

Segre:2007:FCS

[Seg07] Gino Segrè. Faust in Copenhagen: a struggle for the soul of physics. Viking, New York, NY, USA, 2007. ISBN 0-670-03858-X. x + 310 pp. LCCN QC15 .S427 2007. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0743/2006052807-
                                        b.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0743/
                                        2006052807-d.html.

Segre:2011:OGM

[Seg11] Gino Segrè. Ordinary geniuses:

Max Delbrück, George Gamow, and the origins of genomics and Big Bang cosmology. Viking, New York, NY, USA, 2011. ISBN 0-670-02276-4. xxi + 330 pp. LCCN QH31.D434 S44 2011.

Seifer:1996:WLT

[Sei96] Marc J. Seifer. Wizard: The Life and Times of Nikola Tesla: Biography of a Genius. Birch Lane Press, Carol Publishing Group, 120 Enterprise Avenue, Secaucus, NJ, USA, 1996. ISBN 1-55972-329-7. xiv + 542 pp. LCCN TK140.T4 S65 1996. US$32.00, CDN$41.95.

Seifert:1998:GET

[Sei98] Rich Seifert. Gigabit Ethernet:

Technology and Applications for High-Speed LANs. Addison-Wes-ley, Reading, MA, USA, 1998. ISBN 0-201-18553-9. xv + 411 pp. LCCN TK5105.8.E83S45 1998. US$49.95.

Seife:2008:SBS

[Sei08] Charles Seife. Sun in a bottle:

the strange history of fusion and the science of wishful thinking. Viking, New York, NY, USA, 2008. ISBN 0-670-02033-8. vi + 294 pp. LCCN QC791 .S45 2008.

Sen:2021:EFH

[Sen21] Paul Sen. Einstein’s fridge: how the difference between hot and cold explains the universe. Scribner, New York, NY, USA, 2021.

ISBN 1-5011-8130-0 (hardcover), 1-5011-8132-7 (e-book). xii + 305 pp. LCCN QC311 .S5118 2021.

Seroul:1989:PLT

[Ser89] Raymond Seroul. Le petit Livre de TEX. InterEditions, Paris, France, 1989. ISBN 2-7296-0233-X. 317 pp. LCCN Z253.4 T47 S47 1989. English translation available as [SL91].

Service:2018:LTN

[Ser18] Robert Service. The Last of the Tsars: Nicholas II and the Russia Revolution. Pegasus Books, New York, NY, USA, 2018. ISBN 1-68177-883-1 (paperback). xviii + 382 pp. LCCN ????

Sewell:1989:WPL

[Sew89] E. Wayne Sewell. Weaving a Program: Literate Programming in WEB. Van Nostrand Reinhold, New York, NY, USA, 1989. ISBN 0-442-31946-0. xx + 556 pp. LCCN QA76.73.W24 S491 1989.

Schreiner:1985:ICC

[SF85] Axel T. Schreiner and H. George Friedman, Jr. Introduction to Compiler Construction Under UNIX. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-474396-2. viii + 194 pp. LCCN QA76.76.C65 S37 1985. URL

                                        ftp://ftp.cs.uiuc.edu/pub/
                                        faculty/friedman/tar.

Shriner:1964:SIO

[SFC64] Ralph L. Shriner, Reynold C. Fuson, and David Y. Curtin. The Systematic Identification of Organic Compounds: A Laboratory Manual. Wiley, New York, NY, USA, fifth edition, 1964. ix + 458 pp. LCCN QD261 .S5 1964.

Stevens:2004:UNP

[SFR04] W. Richard Stevens, Bill Fenner, and Andrew M. Rudoff. UNIX Network Programming: The Sockets Networking API, volume 1. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, third edition, 2004. ISBN 0-13-141155-1 (hardcover). xxiii + 991 pp. LCCN QA76.76 .O63 S75 2004. US$69.99.

Shoemaker:1962:EPC

[SG62] David P. Shoemaker and Carl W.

Garland. Experiments in Physical Chemistry. McGraw-Hill, New York, NY, USA, 1962. ix + 471 pp. LCCN QD457 .S56.

Skillin:1974:WT

[SG74] Marjorie E. Skillin and Robert Malcolm Gay. Words into Type. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1974. ISBN 0-13-964262-5. xx + 583 pp. LCCN PN160.S52 1974. US$39.95.

Soni:2017:MPH

[SG17] Jimmy Soni and Rob Goodman.

A mind at play: how Claude Shannon invented the information age. Simon and Schuster, New York, NY, USA, 2017. ISBN 1-4767-6668-1 (hardcover), 1-4767-6669-X (paperback), 1-4767-6670-3 (e-book). xv + 366 pp. LCCN QA29.S423 S66 2017.

Scheifler:1988:XWS

[SGN88] Robert W. Scheifler, James Gettys, and Ron Newman. X Window System: C Library and Protocol Reference. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1988. ISBN 1-55558-012-2. xxix + 701 pp. LCCN QA76.76.W56 S34 1988.

Segre:2016:PPE

[SH16] Gino Segrè and Bettina Hoerlin. The Pope of Physics: Enrico Fermi and the birth of the atomic age. Henry Holt and Company, New York, NY, USA, 2016. ISBN 1-62779-005-5, 1-62779-006-3 (ebook). xi + 351 pp. LCCN QC16.F46 S44 2016.

Shachtman:1999:AZC

[Sha99a]  Tom Shachtman.  Absolute

Zero and the Conquest of Cold. Houghton-Mifflin, Boston, MA, USA, 1999. ISBN 0-395-93888-0, 0-547-52595-8, 0-618-08239-5 (paperback), 1-299-88289-7. 261 pp. LCCN QC278 .S48 1999. URL

                                        http://www.loc.gov/catdir/
                                        description/hm022/99033305.
                                        html;   http://www.loc.gov/
                                        catdir/enhancements/fy0735/
                                        99033305-b.html;   http://
                                        www.loc.gov/catdir/samples/
                                        hm051/99033305.html.

Shanley:1999:PPP

[Sha99b] Tom Shanley. Pentium Pro and Pentium II system architecture. Addison-Wesley, Reading, MA, USA, second edition, 1999. ISBN 0-201-30973-4 (paperback). xxxv + 588 pp. LCCN QA76.8.P46S48 1998. US$36.95.

Shea:1983:OHR

[She83] William R. Shea, editor. Otto Hahn and the Rise of Nuclear Physics, volume 22 of The University of Western Ontario series in philosophy of science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1983. ISBN 90-277-1584-X. 252 pp. LCCN QC773 .O87 1983. URL http:/

/link.springer.com/book/10.
1007/978-94-009-7133-2.

Shertzer:1986:EG

[She86] Margaret Shertzer. The Elements of Grammar. Collier Books, Macmillan Publishing Company, New York, NY, USA, 1986. ISBN 0-02-015440-2 (paperback). 168 pp. LCCN PE1112 .S54 1986.

Shorto:2014:AHW

[Sho14] Russell Shorto. Amsterdam: a history of the world’s most liberal city. Vintage Books, New York, NY, USA, 2014. ISBN 0-307-74375-6. 357 + 24 pp. LCCN DJ411.A55 S56 2014.

Shultis:1994:LNP

[Shu94]  J. Kenneth Shultis.  LaTEX

Notes: Practical Tips for Preparing Technical Documents. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-120973-6. xi + 180 pp. LCCN Z253.4.L38 S48 1994. US$32.00.

Shubin:2008:YIF

[Shu08a]  Neil Shubin.  Your inner

fish: a journey into the 3.5billion-year history of the human body. Pantheon Books, New York, NY, USA, 2008. ISBN 0-375-42447-4. 229 pp. LCCN QM26 .S58 2008. URL

http://www.loc.gov/catdir/
enhancements/fy0827/2007024699-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0827/
2007024699-d.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0827/2007024699-s.html;
http://www.loc.gov/catdir/
toc/ecip0720/2007024699.html.

Shulman:2008:TGC

[Shu08b] Seth Shulman. The telephone gambit: chasing Alexander Graham Bell’s secret. W. W. Norton & Co., New York, NY, USA, 2008. ISBN 0-393-33368-X (paperback), 0-393-06206-6 (hardcover). 256 pp. LCCN TK6018.B4 S58 2008.

Simpson:1959:CLD

[Sim59]  D. P. Simpson.  Cassell’s

Latin Dictionary: Latin–English, English–Latin. Macmillan Publishing Company, New York, NY, USA, fifth edition, 1959. ISBN 0-02-522570-7, 0-02-522580-4. xviii + 883 pp. LCCN PA2365.E5C3 1977. US$24.95. Reprinted 1968 with new material.

Singh:1997:FEE

[Sin97] Simon Singh. Fermat’s Enigma: The Epic Quest to Solve the World’s Greatest Mathematical Problem. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 1997. ISBN 0-8027-1331-9. xiii + 315 pp. LCCN QA244.S55 1997.

Singh:1999:CBE

[Sin99] Simon Singh. The Code Book:

the Evolution of Secrecy from Mary, Queen of Scots, to Quantum Cryptography. Doubleday, New York, NY, USA, 1999. ISBN 0-385-49531-5. xiii + 402 pp. LCCN Z103 .S56 1999. US$24.95. See also [AAG^{+}00].

Singer:2005:LSP

[Sin05] Stephanie Frank Singer. Linearity, Symmetry, and Prediction in the Hydrogen Atom. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-24637-

1. xiv + 396 pp. LCCN QC20.7.G76 S56 2005. URL

http://www.loc.gov/catdir/
enhancements/fy0662/2005042679-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0818/
2005042679-b.html;   http:/
/www.loc.gov/catdir/enhancements/
fy0818/2005042679-t.html.

Singh:2007:MXI

[Sin07] Amit Singh. Mac OS X internals: a systems approach. Addison-Wesley, Reading, MA, USA, 2007. ISBN 0-321-27854-2 (hardcover). xxxiii + 1641 pp. LCCN QA76.76.O63 S5645 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0613/2006014901.html.

Sites:1992:AAR

[Sit92] Richard L. Sites. Alpha Architecture Reference Manual. Digital Press and Prentice-Hall, 1992. ISBN 1-55558-098-X (DP), 0-13-033663-7 (PH). LCCN QA76.9.A73 A46 1992. US$35.00.

Sanders:2010:CEI

[SK10] Jason Sanders and Edward Kandrot. CUDA by Example: an Introduction to General-purpose GPU Programming. Addison-Wesley, Reading, MA, USA, 2010. ISBN 0-13-138768-5. xix + 290 pp. LCCN QA76.76.A65.

Sharma:2019:MRL

[SK19] Rahul Sharma and Vesa Kaihlavirta. Mastering Rust: Learn About Memory Safety, Type System, Concurrency, and the New Features of Rust 2018 Edition. Packt Publishing, Birmingham, UK, second edition, 2019. ISBN 1-78934-118-3, 1-78934-657-6 (paperback). 543 (est.) pp. LCCN QA76.73.R87. URL http://

                                        proquest.safaribooksonline.
                                        com/?fpi=9781789346572.

Skinner:1985:IAL

[Ski85] Thomas P. Skinner. An Introduction to 8086/8088 Assembly Language Programming. Wiley, New York, NY, USA, 1985. ISBN 0-471-80825-3. xiii + 222 pp. LCCN QA76.73.A8 S57 1985. US$17.95.

Skiena:1998:ADM

[Ski98] Steven S. Skiena. The Algorithm Design Manual. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-94860-0. xvi + 486 pp. LCCN QA76.9.A43S55 1997. US$54.95.

Skiena:2008:ADM

[Ski08] Steven S. Skiena. The Algorithm Design Manual. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2008. ISBN 1-84800-069-3 (hardcover), 1-84800-070-7 (e-book). xvi + 730 pp. LCCN QA76.9.A43 S55 2008.

Sharma:2019:CRP

[SKM19] Rahul Sharma, Vesa Kaihlavirta, and Claus Matzinger. The Complete Rust Programming Reference Guide: Design, Develop, and Deploy Effective Software Systems Using the Advanced Constructs of Rust. Packt Publishing, Birmingham, UK, 2019. ISBN 1-83882-810-9. xii + 679 pp. LCCN ????

Seroul:1991:BBT

[SL91] Raymond Seroul and Silvio Levy.

A Beginner’s Book of TEX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97562-4, 3-540-97562-4. xii + 282 pp. LCCN Z253.4.T47 S47 1991. This is a translation and adaption by Silvio Levy of [Ser89].

Shasha:1998:TML

[SL98] Dennis Shasha and Cathy Lazere, editors. Out of Their Minds: The Lives and Discoveries of 15 Great Computer Scientists. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98269-8. xi + 293 pp. LCCN QA76.2.A2 S53 1998. US$16.00.

Slater:1942:MT

[Sla42] John C. (John Clarke) Slater. Microwave Transmission. International series in physics. Mc-Graw-Hill, New York, NY, USA, 1942. x + 309 pp. LCCN QC661 .S635.

Slater:1963:QTM

[Sla63] John C. Slater. Quantum Theory of Molecules and Solids: Electronic-Structure of Molecules, volume 1. McGraw-Hill, New York, NY, USA, 1963. xv + 485 pp. LCCN QC174.1 .S551.

Sillen:1952:PPC

[SLG52] Lars Gunnar Sillén, Paul W. Lange, and Carl O. Gabrielson. Problems in Physical Chemistry. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1952. xi + 370 pp. LCCN QD456 .S514. Seventh printing, October 1965.

Sokolov:1966:QM

[SLT66] A. A. (Arsenii Aleksandrovich) Sokolov, Y. M. (IUrii Mikhailovich) Loskutov, and I. M. (Igor’ Mikhailovich) Ternov. Quantum Mechanics.

Holt, Rinehart, and Winston, New York, NY, USA, 1966. xix + 537 pp. LCCN QC174.1 .S633. Translated by Scripta Technica, Inc. from the original 1962 Russian edition, Kvantovaya Mekhanika.

Shore:1968:PAS

[SM68] Bruce W. Shore and Donald H.

Menzel. Principles of Atomic Spectra. Wiley, New York, NY, USA, 1968. xiv + 538 pp. LCCN QC454 .S47.

Simos:2007:CMS

[SM07] Theodore E. Simos and George Maroulis, editors. Computation in Modern Science and Engineering: Proceedings of the [Fifth] International Conference on Computational Methods in Science and Engineering 2007 (ICCMSE 2007), Corfu, Greece, 25–30 September 2007, volume 2A, 2B of AIP Conference Proceedings (#963). American Institute of Physics, Woodbury, NY, USA, 2007. ISBN 0-7354-0476-3 (set), 0-7354-0477-1 (vol. 1), 0-7354-0478-X (vol. 2). ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. LCCN Q183.9 .I524 2007. URL http:

//www.springer.com/physics/
atoms/book/978-0-7354-0478-
6.

Smail:1923:ETI

[Sma23] Lloyd L. Smail. Elements of the Theory of Infinite Processes. Mc-Graw-Hill, New York, NY, USA, 1923. vi + 339 pp. LCCN QA295.S52.

Smith:1987:SGMa

[Smi87a] Joan M. Smith. The Standard Generalized Markup Language and Related Issues. British National Bibliography Research Fund, London, UK, 1987. ISBN 0-7123-3082-8. 63 pp. LCCN Z286 .E43 S65 1986.

Smith:1987:SGMb

[Smi87b] Joan M. Smith. The Standard Generalized Markup Language (SGML): guidelines for authors. British Library Publications Sales Unit, Wolfeboro, NH, USA, 1987. ISBN 0-7123-3112-3. ix + 66 pp. LCCN QA76.73.S44 S65x 1987.

Smith:1990:LPV

[Smi90] Ross Smith. Learning Post-Script: A Visual Approach. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1990. ISBN 0-938151-12-6. LCCN QA76.73.P67 S55 1990.

Smith:1992:SRS

[Smi92] Joan M. Smith. SGML and Related Standards: Document Description and Processing Languages. The Ellis Horwood series in computers and their applications. Ellis Horwood, New York, NY, USA, 1992. ISBN 0-13-441056-4 (paperback), 0-13-806506-3. xviii + 151 pp. LCCN QA76.73 .S42 1992. US$50.25.

Smith:1998:SECa

[Smi98a] Crosbie Smith. The science of energy: a cultural history of energy physics in Victorian Britain. University of Chicago Press, Chicago, IL, USA and London, UK, 1998. ISBN 0-226-76420-6, 0-226-76421-4 (paperback). xi + 404 pp. LCCN QC72 .S58 1998.

URL http://catdir.loc.gov/
catdir/description/uchi051/
98024960.html;   http://catdir.
loc.gov/catdir/enhancements/
fy0608/98024960-t.html;
http://catdir.loc.gov/catdir/
enhancements/fy0609/98024960-
b.html.

Smith:1998:SXC

[Smi98b] Michael Smith. Station X: The Codebreakers of Bletchley Park. Channel 4 Books, London, UK, 1998. ISBN 0-7522-7148-2. 184 pp. LCCN D810.C88 S66 2000. UK£5.99.

Smiley:2010:MWI

[Smi10] Jane Smiley. The man who invented the computer: the biography of John Atanasoff, digital pioneer. Doubleday, New York, NY, USA, 2010. ISBN 0-385-52713-6, 0-385-53372-1 (e-book), 1-299-11995-6 (e-book). 246 + 8 pp. LCCN QA76.2.A75 S64 2010. US$25.95.

Smith:2014:SDF

[Smi14]  Gary Smith.  Standard de-

viations: flawed assumptions tortured data and other ways to lie with statistics. Duckworth Overlook, London, UK, 2014. ISBN 1-4683-1102-6, 0-7156-4974-4 (e-book). LCCN QA279 .S638 2014. URL http:

                                        //public.eblib.com/choice/
                                        publicfullrecord.aspx?p=2029045.

Schroeder:1998:VT

[SML98] William J. Schroeder, Kenneth M. Martin, and William E. Lorensen. The Visualization Toolkit. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1998. ISBN 0-13-954694-4. xx + 645 pp. LCCN QA76.64.S36 1997. With special contributors Lisa Sobierajski Avila, Rick Avila, and C. Charles Law. Includes CD-ROM with vtk-2.0. The most recent release is available on the World-Wide

                                        Web at http://www.kitware.
                                        com/vtk.html.

Strang:1988:TT

[SMO88] John Strang, Linda Mui, and Tim O’Reilly. termcap & terminfo. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, April 1988. ISBN 0-937175-22-6. xv + 253 pp. LCCN QA76.76.O63 S765 1990; QA76.8.U65 S79 1988. US$21.95. URL http:

                                        //www.oreilly.com/catalog/
                                        9780937175224;   http://www.
                                        oreilly.com/catalog/term.

Smolin:2007:TPR

[Smo07] Lee Smolin. The Trouble with Physics: the Rise of String Theory, the Fall of a Science, and What Comes Next. Mariner Books, Boston, MA, USA, 2007. ISBN 0-618-91868-X (paperback). xxiii + 392 pp.

LCCN QC6 .S6535 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0634/2006007235-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0737/
2006007235-b.html;   http:
//www.loc.gov/catdir/toc/
ecip069/2006007235.html.

Smolin:2013:TRC

[Smo14] Lee Smolin. Time Reborn: From the Crisis in Physics to the Future of the Universe. Mariner Books, Boston, MA, USA, 2014. ISBN 0-544-24559-8. xxxi + 319 pp. LCCN QB991.Q36 S66 2014.

Smith:2005:VMV

[SN05] James E. (James Edward)

Smith and Ravi Nair. Virtual Machines: Versatile Platforms for Systems and Processes. Morgan Kaufmann Publishers, San Francisco, CA, USA, 2005. ISBN 1-55860-910-5 (hardcover), 0-08-052540-7 (e-book), 1-4933-0376-7. xxii + 638 pp. LCCN QA76.9.V5

S54 2005. URL http://books.
elsevier.com/us/bookscat/
search/details.asp?country=
United+States&community=mk&
imprint=Morgan+Kaufmann&isbn=
1558609105.

Sakurai:2017:MQM

[SN17] J. J. (Jun John) Sakurai and Jim Napolitano. Modern Quantum Mechanics. Cambridge University Press, Cambridge, UK, second edition, 2017. ISBN 1-108-42241-1 (hardcover). xviii + 550 pp. LCCN QC174.12 .S25 2017.

Snow:1992:TB

[Sno92] Wynter Snow. TEX for the Beginner. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-54799-6. xii + 377 + 23 pp. LCCN Z253.4.T47 S64 1992.

Snyder:1982:MPU

[Sny82] John P. Snyder. Map Projections Used by the U.S. Geological Survey. Geological Survey Bulletin 1532. United States Government Printing Office, Washington, DC, USA, second edition, 1982. xiii + 313 pp. LCCN QE75 .B9.

Schwarz:2011:PIS

[SO11] Eric Schwarz and Vojin G. Oklobdzija, editors. Proceedings of the 20th IEEE Symposium on Computer Arithmetic, July 25– 27, 2011, Tübingen, Germany. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 0-7695-4318-9, 1-4244-9457-5. LCCN QA76.6.

Sobel:1995:LTS

[Sob95] Dava Sobel. Longitude: the True Story of a Lone Genius Who Solved the Greatest Scientific Problem of His Time. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 1995. ISBN 0-8027-1312-2. viii + 184 pp. LCCN QB225 .S64 1995. US$19.00. A biography of John Harrison, 1693–1776.

Sobel:2012:MPH

[Sob12] Dava Sobel. A more perfect heaven: how Copernicus revolutionized the cosmos. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 2012. ISBN 0-8027-7894-1. xiv + 273 pp. LCCN QB501 .S75 2012.

SoftQuad:1991:SP

[Sof91] SoftQuad, Inc. The SGML Primer. SoftQuad, Inc., Toronto, Ontario, Canada, 1991. 36 pp. LCCN QA76.73.S44 S45 1991.

Sokal:2000:SHS

[Sok00] Alan D. Sokal, editor. The Sokal Hoax: the Sham That Shook the Academy. University of Nebraska Press, Lincoln, NE, USA, 2000. ISBN 0-8032-7995-7 (paperback), 0-8032-1924-5. ix + 271 pp. LCCN Q175.37 .S65 2000.

Southwick:1999:HNI

[Sou99] Karen Southwick. High Noon:

the Inside Story of Scott McNealy and the Rise of Sun Microsystems. Wiley, New York, NY, USA, 1999. ISBN 0-471-29713-5 (hardcover). xiii + 242 pp. LCCN QA76.2.M37 S68 1999. US$25.90.

Sienko:1964:C

[SP64] Michell J. Sienko and Robert A.

Plane. Chemistry. McGraw-Hill, New York, NY, USA, Canadian edition, 1964. xiii + 623 pp. LCCN ????

SPARC:1992:SAM

[SPA92] SPARC International, Inc. The SPARC Architecture Manual: Version 8. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-825001-4 (paperback). xxix + 316 pp. LCCN QA76.9.A73 S647 1992.

URL http://www.sparc.org/
standards/V8.pdf.

Spencer:regexp

[Spe85] Henry Spencer. Regular expression pattern matching software. Usenet mod.sources and net.sources archives, November 1985.

Spivak:1986:JTG

[Spi86] Michael D. Spivak. The Joy of TEX: A Gourmet Guide to Typesetting with the AMS-TEX macro package. American Mathematical Society, Providence, RI, USA, 1986. ISBN 0-8218-2999-8. xviii + 290 pp. LCCN Z253.4.T47 S6731 1986.

Spivak:1989:LS

[Spi89] Michael D. Spivak. LAMS-TEX, The Synthesis. The TEXplorators Corporation, 3701 W. Alabama, Suite 450-273, Houston, TX 77027, USA, 1989. viii + 289 pp.

Spivak:1990:JTG

[Spi90] Michael D. Spivak. The Joy of TEX: a Gourmet Guide to Typesetting with the AMS-TEX macro package. American Mathematical Society, Providence, RI, USA, second revised edition, 1990. ISBN 0-8218-2997-1. xxii + 309 pp. LCCN Z253.4.T47 S6731 1990. US$40.00. URL https:/

/ctan.org/pkg/joy-of-tex.

Sokolnikoff:1966:MPM

[SR66] I. S. Sokolnikoff and R. M. Redheffer. Mathematics of Physics and Modern Engineering. Mc-Graw-Hill, New York, NY, USA, second edition, 1966. ISBN 0-07-059625-5, 0-07-085745-8 (student edition). xi + 752 pp. LCCN QA401 .S6 1966.

Schengili-Roberts:2000:CC

[SR00] Keith Schengili-Roberts. Core CSS. Prentice Hall PTR core series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-083456-4 (paperback). xxxvii + 635 + 4 pp. LCCN TK5105.888. S32 2000.

Stevens:2005:APU

[SR05] W. Richard Stevens and Stephen A.

Rago. Advanced Programming in the Unix Environment. Addison-Wesley, Reading, MA, USA, second edition, 2005. ISBN 0-201-43307-9 (hardcover), 0-321-52594-9. xxviii + 927 pp. LCCN QA76.76.O63 S754 2005. URL ftp://

uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/;   http://
www.aw-bc.com/catalog/academic/
product/0,1144,0201433079,
00.html;   http://www.loc.
gov/catdir/toc/ecip059/2005007943.
html.

StDenis:2006:BMH

[SR06] Tom St Denis and Greg Rose.

BigNum Math: Implementing Cryptographic Multiple Precision Arithmetic. Syngress Publishing, Inc., Rockland, MA, USA, 2006. ISBN 1-59749-112-8. xviii + 296 pp. LCCN QA402.5 2006. URL http:

//www.oreilly.com/catalog/
1597491128/index.html.

Sridhar:1996:BPC

[Sri96] M. A. Sridhar. Building Portable C++ Applications with YACL. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-83276-3. xix + 569 pp. LCCN QA76.73.C153S699 1966. Source code for Yet Another Class Library avail-

able from ftp://ftp.cs.sc.
edu/pub/yacl/yacl0150.tar.
gz.

Seltzer:1997:ASR

[SRR97] Richard Seltzer, Deborah S. Ray, and Eric J. Ray. The AltaVista Search Revolution: How to Find Anything on the Internet. Osborne/McGraw-Hill, Berkeley, CA, USA, 1997. ISBN 0-07-882235-1. xxii + 274 pp. LCCN TK5105.875 .I57 S44 1997. US$16.99.

Sargent:1984:IPC

[SS84] Murray Sargent III and Richard L. Schoemaker. The IBM Personal Computer from the Inside Out. Addison-Wesley, Reading, MA, USA, 1984. ISBN 0-201-06896-6. xii + 483 pp. LCCN QA76.8.I2594 S27 1984. A strongly hardware-oriented, but very readable, treatment of the IBM PC. Contains three chapters on the 8088 and assembly language programming, with the remainder of the book devoted to descriptions of hardware, keyboard and video display, external devices, and data communications.

Smith:1988:SUG

[SS88] Joan M. Smith and Robert Stutely. SGML: The User’s Guide to ISO 8879. Ellis Horwood series in computers and their applications. Ellis Horwood and Halsted Press, New York, NY, USA and New York, USA, 1988. ISBN 0-470-21126-1 (Halsted Press), 0-7458-0221-4. 175 pp. LCCN QA76.73.S44 S441 1988. UK£22.50.

Srikant:2002:CDH

[SS02] Y. N. Srikant and Priti Shankar, editors. The Compiler Design Handbook: Optimizations and Machine Code Generation. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2002. ISBN 0-8493-1240-X, 1-4200-4057-X (e-book). xii + 916 pp. LCCN QA76.76.C65 C66 2003. US$139.95, UK£98.00.

Senor:2009:SNS

[SS09] Dan Senor and Saul Singer.

Start-up Nation: the Story of Israel’s Economic Miracle. Twelve, New York, NY, USA, 2009. ISBN 0-446-54146-X, 0-446-54147-8 (paperback), 1-4555-0346-0 (ebook), 1-59995-857-0 (e-book). x + 304 pp. LCCN HC415.25 .S455 2009eb.

Scholze-Stubenrecht:1997:ODG

[SSStt97] W. Scholze-Stubenrecht, J. B. Sykes, the Dudenredaktion, and the German Section of the Oxford University Press Dictionary Department, editors. The Oxford–Duden German Dictionary: German–English/English– German. Clarendon Press, Oxford, UK, 1997. ISBN 0-19-860130-1, 0-19-860132-8. 1712 pp. LCCN PF3640.O94 1997. US$39.95.

Sproull:1985:DPD

[SSU85] Robert F. Sproull, W. R. Sutherland, and Michael K. Ullner. Device-Independent Graphics: with examples from IBM personal computers. McGraw-Hill, New York, NY, USA, 1985. ISBN 0-07-060504-1. xii + 546 pp. LCCN T385.S67 1985. US$29.95.

Sheldon:1952:ICP

[ST52] J. W. Sheldon and L. Tatum. The IBM card-programmed electronic calculator. In Review of Electronic Digital Computers. Joint AIEE–IRE Computer Conference. 10–12 December 1951, pages 30–36. American Institute of Electrical Engineers, New York, NY, USA, 1952. Reprinted in [Ran82, §5.4].

Sander:2012:MHA

[ST12] Richard Henry Sander and Stuart Taylor. Mismatch: How Affirmative Action Hurts Students It’s Intended to Help, and Why Universities Won’t Admit It. Basic Books, New York, NY, USA, 2012. ISBN 0-465-02996-5 (hardcover), 0-465-03001-7 (e-book). LCCN LC213.52 .S26 2012.

Stallman:gcc

[Sta88] Richard M. Stallman. Using and porting GNU CC. Technical report, Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1988. Electronic mail:

rms@prep.ai.mit.edu. Software

also available via ANONYMOUS

FTP to prep.ai.mit.edu.

Stallings:1990:RIS

[Sta90] William Stallings, editor. Reduced Instruction Set Computers. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, second edition, 1990. ISBN 0-8186-8943-9 (hardcover), 0-8186-5943-2 (microfiche). ix + 433 pp. LCCN QA76.5 .T876 1990. US$54.50.

Stansifer:SIGPLAN-27-12-61

[Sta92] Ryan Stansifer. The calculation of Easter. ACM SIGPLAN Notices, 27(12):61–65, December 1992.

Stachel:1998:EMY

[Sta98a]  John Stachel, editor.  Ein-

stein’s miraculous year: five papers that changed the face of physics. Princeton University Press, Princeton, NJ, USA, 1998. ISBN 0-691-05938-1. xv + 198 pp. LCCN QC7 .E52 1998.

US$19.95. With the assistance of Trevor Lipscombe, Alice Calaprice, and Sam Elworthy, and with a foreword by Roger Penrose.

Standage:1998:VIR

[Sta98b] Tom Standage. The Victorian Internet: the Remarkable Story of the Telegraph and the Nineteenth Century’s On-line Pioneers. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 1998. ISBN 0-8027-1342-4. ix + 227 pp. LCCN HE7631 .S677 1998. US$12.00, CAN$17.00.

Stannard:1999:NWM

[Sta99] Russell Stannard. The New World of Mr. Tompkins: George Gamow’s Classic Mr. Tompkins in Paperback. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-63009-6 (hardcover), 0-521-63992-1 (paperback), 0-511-00800-7 (e-book), 0-511-58169-6 (e-book). xii + 258 pp. LCCN QC71.S775 1999. US$24.99. URL http://www.

loc.gov/catdir/description/
cam029/98050379.html;   http:
//www.loc.gov/catdir/enhancements/
fy0732/98050379-b.html;
http://www.loc.gov/catdir/
toc/cam021/98050379.html.

Stallings:2010:COA

[Sta10] William Stallings. Computer organization and architecture: designing for performance. Prentice-Hall, Upper Saddle River, NJ 07458, USA, eighth edition, 2010. ISBN 0-13-607373-5. xviii + 774 pp. LCCN QA76.9.C643 S73 2010.

Stewart:1973:IMC

[Ste73] G. W. (Gilbert W.) Stewart. Introduction to Matrix Computations. Computer Science and Applied Mathematics, Editor: Werner Rheinboldt. Academic Press, New York, NY, USA, 1973. ISBN 0-12-670350-7. xiii + 441 pp. LCCN QA188 .S71

1973. URL http://catalog.
hathitrust.org/api/volumes/
oclc/658076.html;   http://
catdir.loc.gov/catdir/description/
els031/72082636.html;   http:
//catdir.loc.gov/catdir/toc/
els031/72082636.html;   http:
//www.gbv.de/dms/hbz/toc/
ht000832348.pdf;   http://
www.loc.gov/catdir/description/
els031/72082636.html;   http:
//www.loc.gov/catdir/toc/
els031/72082636.htmlQ.

Stevenson:COMPUTER-14-3-51

[Ste81] David Stevenson. A proposed standard for binary floatingpoint arithmetic. Computer, 14 (3):51–62, March 1981. See [IEE85b].

Steele:1984:CLL

[Ste84] Guy L. Steele Jr. Common Lisp: The Language. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1984. ISBN 0-13-152851-3, 0-932376-41-X (paperback). xii + 465 pp. LCCN QA76.73.L23 S73 1984. US$22.00. With contributions by Scott E. Fahlman and Richard P. Gabriel and David A. Moon and Daniel L. Weinreb. See also [Tat87].

Steele:1990:CLL

[Ste90a] Guy L. Steele Jr. Common Lisp: The Language. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, second edition, 1990. ISBN 1-55558-041-6 (paperback), 1-55558-042-4 (hardcover), 0-08-050226-1 (e-book), 0-13-152414-3 (Prentice-Hall). xxiii + 1029 pp. LCCN QA76.73.L23 S73 1990. With contributions by Scott E. Fahlman and others, and with contributions to the second edition by Daniel G. Bobrow and others. See also [Ste84, Tat87].

Stevens:1990:UNP

[Ste90b] W. Richard Stevens. UNIX Network Programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-949876-1. xi + 772 pp. LCCN QA76.76.O63 S755 1990.

Stern:1991:MNN

[Ste91]  Hal Stern.  Managing NFS

and NIS. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-75-7. xxiv + 410 pp. LCCN TK5105.5 .S74 1991. URL http:

                                        //www.oreilly.com/catalog/
                                        9780937175750;   http://www.
                                        oreilly.com/catalog/nfs.

Stevens:1992:APU

[Ste92] W. Richard Stevens. Advanced Programming in the UNIX Environment. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56317-7. xviii + 744 pp. LCCN QA76.76.O63 S754 1992. US$52.25. Highly recommended. A going-to-be classic on how programs work under UNIX … The source codes and errata list are obtainable by anony-

mous ftp from ftp.uu.net (in
/published/books).

Stephen:1994:SSA

[Ste94a] Graham A. Stephen. String Searching Algorithms, volume 3 of Lecture notes series on computing. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1994. ISBN 981-02-1829-X (hardcover). xii + 243 pp. LCCN QA76.9.A43 S73 1994.

Stevens:1994:TII

[Ste94b] W. Richard Stevens. TCP/IP Illustrated, Volume 1: The Protocols. Addison-Wesley Professional Computing Series. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-63346-9. xix + 576 pp. LCCN TK5105.55 .S74 1994. US$65.95.

Stevens:1996:TII

[Ste96] W. Richard Stevens. TCP/ IP Illustrated, Volume 3: TCP for Transactions, HTTP, NNTP, and the UNIX Domain Protocols. Addison-Wesley Professional Computing Series. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-63495-3. xix + 328 pp. LCCN TK5105.55.S74 1994. US$41.47.

Stevens:1998:UNP

[Ste98] W. Richard Stevens. UNIX Network Programming: Networking APIs: Sockets and XTI, volume 1. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 1998. ISBN 0-13-490012-X. xx + 1009 pp. LCCN QA76.76.O63S755 1998. US$59.00. URL http://

                                        www.phptr.com/ptrbooks/ptr_
                                        013490012X.html.

Stern:1999:EGW

[Ste99a]  Fritz Richard Stern.  Ein-

stein’s German world. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-05939-x. ix + 335 pp. LCCN DD239 .S74 1999.

                                        URL http://www.gbv.de/dms/
                                        bowker/toc/9780691059396.
                                        pdf;   http://www.loc.gov/
                                        catdir/description/prin021/
                                        99020128.htm;   http://www.
                                        loc.gov/catdir/description/
                                        prin021/99020128.html;   http:
                                        //www.loc.gov/catdir/toc/
                                        prin032/99020128.htm;   http:
                                        //www.loc.gov/catdir/toc/
                                        prin032/99020128.html.

Stevens:1999:UNP

[Ste99b] W. Richard Stevens. UNIX Network Programming, Interprocess Communications, volume 2. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1999. ISBN 0-13-081081-9 (hardcover). xvii + 558 pp. LCCN QA76.76.O63S755 1998. US$54.00. URL http://

www.phptr.com/ptrbooks/ptr_
0130810819.html.

Stern:2001:EGW

[Ste01]  Fritz Richard Stern.  Ein-

stein’s German world. Princeton University Press, Princeton, NJ, USA, 2001. ISBN 0-691-07458-5 (paperback). ix + 335 pp. LCCN DD239 .S74 2001. URL http://www.

loc.gov/catdir/description/
prin021/99020128.htm;   http:
//www.loc.gov/catdir/toc/
prin032/99020128.htm.

Stewart:2016:CCH

[Ste16] Ian Stewart. Calculating the Cosmos: How Mathematics Unveils the Universe. Basic Books, New York, NY, USA, 2016. ISBN 1-5416-1725-8 (paperback). 346 + 8 pp. LCCN QB981 .S82 2018.

Stibitz:1940:C

[Sti40] G. R. Stibitz. Computer. Unpublished memorandum, Bell Telephone Laboratories, New York, NY, USA, 1940. Reprinted in [Ran82, §6.1].

Stiddard:1975:ELS

[Sti75] M. H. B. Stiddard. The Elementary Language of Solid State Physics. Academic Press, New York, NY, USA, 1975. ISBN 0-12-671050-3. xii + 188 pp. LCCN QC176.S783.

Stoll:1989:CET

[Sto89] Clifford Stoll. The Cuckoo’s Egg: Tracking a Spy Through the Maze of Computer Espionage. Doubleday, New York, NY, USA, 1989. ISBN 0-385-24946-2, 0-307-81942-6 (ebook), 0-7434-1145-5, 0-7434-1146-3, 1-299-04734-3. vi + 326 pp. LCCN UB271.R92 H477 1989; UB271.R92 H4771 1989; UB271.R92 S47 1989. US$18.95.

URL http://vxer.org/lib/
pdf/The%20Cuckoo%27s%20Egg.
pdf.

Stoll:1999:HTH

[Sto99] Clifford Stoll. High Tech Heretic: Why Computers Don’t Belong in the Classroom and Other Reflections by a Computer Contrarian. Doubleday, New York, NY, USA, 1999. ISBN 0-385-48975-7. xv + 221 pp. LCCN LB1028.5 .S77 1999. US$24.95.

Stokes:2007:IMI

[Sto07] Jon Stokes. Inside the Machine: an Illustrated Introduction to Microprocessors and Computer Architecture. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-104-2. xx + 292 pp. LCCN TK7895.M5 S76 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0715/2005037262-
d.html;   http://www.loc.
gov/catdir/toc/ecip065/2005037262.
html.

Stone:2013:EQQ

[Sto13] A. Douglas Stone. Einstein and the quantum: the quest of the valiant Swabian. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-13968-7 (hardcover), 1-4008-4834-2 (e-book), 1-299-87423-1 (e-book). x + 332 pp. LCCN QC16.E5 S76 2013. URL http:

//www.npr.org/2013/11/01/
242356997/einsteins-real-
breakthrough-quantum-theory.

Streitweiser:1961:MOT

[Str61] Andrew Streitweiser, Jr. Molecular Orbital Theory for Organic Chemists. Wiley, New York, NY, USA, 1961. xvii + 489 pp. LCCN QD255 .S77 1961.

Stroud:1974:NQS

[Str74] A. H. Stroud. Numerical Quadrature and Solution of Ordinary Differential Equations: a Textbook for a Beginning Course in Numerical Analysis: Graduate Texts in Mathematics. Applied Mathematical Sciences Series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. ISBN 0-387-90100-0. xi + 338 pp. LCCN QA1 .A647 vol. 10.

Strang:1986:PC

[Str86a] John Strang. Programming with curses. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1986. ISBN 0-937175-02-1. 76 pp. LCCN QA76.73.C15S78 1991. US$12.95. URL http://www.

oreilly.com/catalog/curses.

Strang:1986:RWT

[Str86b]  John Strang.  Reading and

Writing TERMCAP Entries. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1986. 73 pp.

Stroustrup:1986:CPL

[Str86c] Bjarne Stroustrup. The C++ Programming Language. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-12078-X. viii + 327 pp. LCCN QA76.73.C153 S77 1986.

Stroustrup:1991:CPL

[Str91] Bjarne Stroustrup. The C++ Programming Language. Addison-Wesley, Reading, MA, USA, second edition, 1991. ISBN 0-201-53992-6. xi + 669 pp. LCCN QA76.73.C15 S79 1991.

Stroustrup:1994:DEC

[Str94] Bjarne Stroustrup. The Design and Evolution of C++. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-54330-3. x + 461 pp. LCCN QA76.73.C153S79 1994. US$26.96.

Stripp:1995:CBF

[Str95] Alan Stripp. Code Breaker in the Far East. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 1995. ISBN 0-19-280386-7 (paperback), 0-19-285316-3. xiv + 204 pp. LCCN D810.C88 S76 1989.

Stroustrup:1997:CPL

[Str97] Bjarne Stroustrup. The C++ Programming Language. Addison-Wesley, Reading, MA, USA, third edition, 1997. ISBN 0-201-70073-5, 0-201-88954-4 (paperback). x + 910 pp. LCCN QA76.73.C153 S77 1997. See also companion book [Van99].

Stringer:2013:LSH

[Str13a] Chris Stringer. Lone survivors:

how we came to be the only humans on earth. St. Martin’s Griffin, New York, NY, USA, 2013. ISBN 1-250-02330-0, 1-4299-7344-7 (e-book). xii + 320 pp. LCCN GN281 .S8746 2013.

Stroustrup:2013:CPL

[Str13b] Bjarne Stroustrup. The C++ programming language. Addison-Wesley, Reading, MA, USA, fourth edition, 2013. ISBN 0-321-56384-0 (paperback), 0-321-95832-2 (hardcover), 0-13-352283-0 (e-book). xiv + 1346 pp. LCCN QA76.73.C153 S77 2013.

Strogatz:2019:IPH

[Str19] Steven Strogatz. Infinite Powers:

How Calculus Reveals the Secrets of the Universe. Houghton Mifflin Harcourt, Boston, MA, USA, 2019. ISBN 0-358-29928-4 (paperback), 1-328-87998-4 (hardcover), 1-328-88001-X (e-book). xxiii + 360 pp. LCCN QA303.2 .S78 2019.

Syropoulos:2003:DTU

[STS03] Apostolos Syropoulos, Antonis Tsolomitis, and Nick Sofroniou. Digital typography using LaTEX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 0-387-95217-9 (paperback), 0-387-22436-X (e-book). xxix + 510 pp. LCCN Z253.4.L38 S97 2003.

Stuart:1984:WAE

[Stu84] Ann Stuart. Writing and Analyzing Effective Computer Documentation. Holt, Rinehart, and Winston, New York, NY, USA, 1984. ISBN 0-03-063892-5. x + 271 pp. LCCN QA76.9.D6 S78 1984.

Stuewer:2018:AIN

[Stu18] Roger H. Stuewer. The Age of Innocence: Nuclear Physics Between the First and Second World Wars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-186658-X, 0-19-882787-3 (hardback), 0-19-256290-8 (e-book). xv + 484 pp. LCCN QC773 .S78 2018.

Summerfield:2009:PPC

[Sum09] Mark Summerfield. Programming in Python 3: a complete introduction to the Python language. Developer’s library. Addison-Wesley, Reading, MA, USA, 2009. ISBN 0-13-712929-7 (paperback). xiv + 525 pp. LCCN QA76.73.P98 S86 2009.

Sun:SPARC

[Sun87] Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA. The SPARC Architecture Manual, part no: 800-1399-07 edition, August 8 1987.

Sunday:string-search

[Sun90] Daniel M. Sunday. A very fast substring search algorithm. Communications of the Association for Computing Machinery, 33(8):132–142, August 1990. See also [BM77, KMP77, BYG92].

Sun:1997:JEC

[Sun97] Sun Microsystems. Java Enterprise Computing: Enabling Breakaway Business Strategies. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, 1997. xvii + 202 pp.

Susskind:2008:BHW

[Sus08] Leonard Susskind. The Black Hole War: My Battle with Stephen Hawking to Make the World Safe for Quantum Mechanics. Little, Brown and Company, Boston, Toronto, London, 2008. ISBN 0-316-01640-

3. viii + 470 pp. LCCN QC174.12 .S896 2008. URL

http://www.loc.gov/catdir/
enhancements/fy0804/2007048355-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0804/
2007048355-d.html.

Shapiro:1999:IRS

[SV99] Carl Shapiro and Hal R. Varian.

Information Rules: a Strategic Guide to the Network Economy. Harvard Business School Press, Boston, MA, USA, 1999. ISBN 0-87584-863-X (paperback), 1-4221-5462-9, 0-585-16425-8 (ebook). x + 352 pp. LCCN HC79.I55 S53 1999. US$29.95.

Salthouse:1972:PGC

[SW72] J. A. Salthouse and M. J. Ware. Point group character tables and related data. Cambridge University Press, Cambridge, UK, 1972. ISBN 0-521-08139-4. v + 88 pp. LCCN QD911 .S25 1972.

Strunk:1979:ES

[SW79] William Strunk Jr. and E. B. White. The Elements of Style. Macmillan Publishing Company, New York, NY, USA, third edition, 1979. ISBN 0-02-418230-3 (hardcover), 0-02-418220-6 (paperback). xvii + 85 pp. LCCN PE1408 .S77 1979.

Skoog:1980:PIA

[SW80] Douglas A. Skoog and Donald M. West. Principles of Instrumental Analysis. Saunders College, Philadelphia, PA, USA, second edition, 1980. ISBN 0-03-021161-1. xi + 769 pp. LCCN QD79.I5S58 1980.

Steele:floating-point-output

[SW90] Guy L. Steele Jr. and Jon L. White. How to print floatingpoint numbers accurately. ACM SIGPLAN Notices, 25(6):112– 126, June 1990. See also input algorithm in [Cli90], and [Knu90a]. In electronic mail dated Wed, 27 Jun 90 11:55:36 EDT, Guy Steele reported that an intrepid pre-SIGPLAN 90 conference implementation of what is stated in the paper revealed 3 mistakes:

1. Table 5 (page 124): insert k <-- 0 after assertion, and also delete k <-- 0 from Table 6.

2. Table 9 (page 125):

for  -1:USER!("");
substitute -1:USER!("0");

and delete the comment.

3. Table 10 (page 125):

for  fill(-k, "0")
substitute fill(-k-1, "0")

Sites:1995:AAA

[SW95] Richard L. Sites and Richard L.

Witek. Alpha AXP Architecture Reference Manual. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, second edition, 1995. ISBN 1-55558-145-5. various pp. LCCN QA76.9.A73A46 1995. US$52.95.

Strunk:1999:ES

[SW99] William Strunk Jr. and E. B.

White. The Elements of Style. Allyn and Bacon, Boston, MA, USA, fourth edition, 1999. ISBN 0-205-30902-X (paperback), 0-205-31342-6 (hardcover). xviii + 105 pp. LCCN PE1408 .S772 2000. US$7.95.

Sedgewick:2011:A

[SW11] Robert Sedgewick and Kevin Daniel Wayne. Algorithms. Addison-Wesley, Reading, MA, USA, fourth edition, 2011. ISBN 0-13-276257-9, 0-321-57351-X (hardcover), 6-6130-0155-4 (e-book). xii + 955 pp. LCCN QA76.9.A43 S429 2011.

URL http://proquestcombo.
safaribooksonline.com/9780132762571.

Swartz:1965:FP

[Swa65] Clifford E. Swartz. The Fundamental Particles. Addison-Wesley, Reading, MA, USA, 1965. viii + 152 pp. LCCN QC721.S982.

Swartzlander:1976:CDD

[Swa76] Earl E. Swartzlander, Jr., editor. Computer design development: principal papers. Hayden Book Co., Rochelle Park, NJ, USA, 1976. ISBN 0-8104-5988-4. 310 pp. LCCN QA76.5 .C612565.

Swartzlander:1990:CAa

[Swa90a] Earl E. Swartzlander, Jr. Computer Arithmetic, volume 1 of IEEE Computer Society Press tutorial. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-8931-5 (hardcover), 0-8186-5931-9 (microfiche). xiii + 378 pp. LCCN QA76.6 .C633 1990.

Swartzlander:1990:CAb

[Swa90b] Earl E. Swartzlander, Jr. Computer Arithmetic, volume 2. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-8945-5. ix + 396 pp. LCCN QA76.9 .C62C66 1990. This is part of a twovolume collection of influential papers on the design of computer arithmetic. See also [Swa90a].

Swade:2001:DEC

[Swa01] Doron Swade. The Difference Engine: Charles Babbage and the Quest to Build the First Computer. Viking, New York, NY, USA, 2001. ISBN 0-670-91020-1, 0-14-200144-9 (paperback). x + 342 + 16 pp. LCCN QA75 .S954 2001.

URL http://www.gbv.de/dms/
bowker/toc/9780670910205.
pdf;   http://www.loc.gov/
catdir/toc/fy034/00068643.
html;   http://www.zentralblatt-
math.org/zmath/en/search/?
an=1017.01006.

Sweetman:1999:SMR

[Swe99] Dominic Sweetman. See MIPS Run. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 1-55860-410-3 (paperback). xviii + 488 pp. LCCN QA76.9.A73 S88 1999. US$52.95, UK£35.95.

Sweetman:2007:SMR

[Swe07] Dominic Sweetman. See MIPS Run. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 2007. ISBN 0-12-088421-6. xix + 492 pp. LCCN QA76.9.A73 S88 2007.

Swinton:2019:ATM

[Swi19] Jonathan Swinton. Alan Turing’s Manchester. Infang Publishing, Manchester, UK, 2019. ISBN 0-9931789-2-8 (paperback), 0-9931789-3-6 (e-book). 204 pp. LCCN ????

Strunk:2005:ES

[SWK05] William Strunk, Jr., E. B. (Elwyn Brooks) White, and Maira Kalman. The Elements of Style.

Penguin, London, UK and New York, NY, USA, fourth edition, 2005. ISBN 0-14-311272-4 (paperback), 1-59420-069-6 (hardcover). xvii + 147 pp. LCCN PE1408 .S772 2005. URL

http://www.penguinputnam.
com/nf/Book/BookDisplay/0,,
0_1594200696,00.html.

Sykes:1994:NOG

[Syk94] Christopher Sykes, editor. No Ordinary Genius: The Illustrated Richard Feynman. W. W. Norton & Co., New York, NY, USA, 1994. ISBN 0-297-81469-9, 0-393-31393-X (paperback), 0-393-03621-9. 272 pp. LCCN QC16.F49 A3 1994. US$19.95, CDN$25.99.

Sykes:2007:SVC

[Syk07] Bryan Sykes. Saxons, Vikings and Celts: the Genetic Roots of Britain and Ireland. W. W. Norton & Co., New York, NY, USA, 2007. ISBN 0-393-33075-3 (paperback). xvii + 306 + 16 pp. LCCN GN290.G7 S95 2007. US$16.95.

Szasz:1984:DSR

[Sza84] Ferenc Morton Szasz. The day the sun rose twice: the story of the Trinity Site nuclear explosion, July 16, 1945. University of New Mexico Press, Albuquerque, NM, USA, 1984. ISBN 0-8263-0768-X (paperback), 0-8263-0767-1. xi + 233 pp. LCCN QC773.A1 S93 1995; QC773.A1 S93 1984.

Tenenbaum:1981:DSU

[TA81] Aaron M. Tenenbaum and Moshe J. Augenstein. Data Structures Using Pascal. Prentice-Hall Software Series, Editor: Brian W. Kernighan. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1981. ISBN 0-13-196501-8. xiv + 545 pp. LCCN QA76.9.D35 T46.

Tabak:1987:RIS

[Tab87] Daniel Tabak. Reduced Instruction Set Computer: RISC: Architecture. Wiley, New York, NY, USA, 1987. ISBN 0-471-91302-2. xiv + 161 pp. LCCN QA76.9.A73 T29 1987.

Tabak:1990:RS

[Tab90] Daniel Tabak. RISC Systems.

Wiley, New York, NY, USA, 1990. ISBN 0-471-92694-9 (Wiley), 0-86380-100-5 (Research Studies Press). xii + 300 pp. LCCN QA76.9.A73.T294 1990. US$49.95.

Talbott:1986:MPM

[Tal86]  Steve Talbott.  Managing

Projects with Make. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1986.

Talbott:1988:MPM

[Tal88]  Steve Talbott.  Managing

Projects with Make. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1988. ISBN 0-937175-04-8. 77 pp. LCCN QA76.8.U65 T34 1988. URL

                                        http://www.oreilly.com/catalog/
                                        9780937175040.

Taleb:2001:FRH

[Tal01] Nassim Taleb. Fooled by Randomness: the Hidden Role of Chance in the Markets and in Life. Texere, New York, NY, USA, 2001. ISBN 1-58799-071-7 (hardcover), 1-58799-184-5 (paperback). xix + 203 pp. LCCN HG4521 .T35 2001; Hg4521 .T34 2001.

Tanenbaum:1987:OSD

[Tan87] Andrew S. Tanenbaum. Operating Systems: Design and Implementation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-637406-9. xvi + 719 pp. LCCN QA76.76.O63 T36 1987. US$34.95.

Tanenbaum:1988:CN

[Tan88] Andrew S. Tanenbaum. Computer Networks. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1988. ISBN 0-13-162959-X. xv + 658 pp. LCCN TK5105.5 .T36 1988.

Tanenbaum:1992:MOS

[Tan92] Andrew S. Tanenbaum. Modern Operating Systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-098310-1, 0-13-588187-0, 0-13-595752-4 (paperback). xx + 728 pp. LCCN QA76.76.O63 T359 1992.

Tang:1993:TSS

[Tan93] Yuanhua Tang. Theoretical Studies on Second Messenger Dynamics. Ph.D. thesis, Mathematics Department, University of Utah, Salt Lake City, UT, USA, December 1993. ix + 263 pp.

Tanenbaum:1996:CN

[Tan96] Andrew S. Tanenbaum. Computer Networks. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, third edition, 1996. ISBN 0-13-349945-6. xvii + 814 pp. LCCN TK5105.5.T36 1996. US$77.85.

Tanenbaum:2008:MOS

[Tan08] Andrew S. Tanenbaum. Modern Operating Systems. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, third edition, 2008. ISBN 0-13-600663-9. xxvii + 1076 pp. LCCN QA76.76.O63 T359 2008.

Tapscott:1996:DEP

[Tap96]  Don Tapscott.  The Digital

Economy: Promise and Peril in the Age of Networked Intelligence. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-062200-0 (hardcover), 0-07-063342-8 (paperback), 0-07-137117-6 (e-book). xviii + 342 pp. LCCN HC79.I55 T368 1996. US$24.95.

Tatar:1987:PGC

[Tat87] Deborah G. Tatar. A Programmer’s Guide to Common LISP. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1987. ISBN 0-932376-87-8. x + 327 pp. LCCN QA76.73.C28 T38 1987. See also [Ste84].

Taub:1961:JNCa

[Tau61a] A. H. Taub, editor. John von Neumann: Collected Works: Volume I: Logic, Theory of Sets and Quantum Mechanics. Pergamon Press, New York, NY, USA, 1961. x + 654 pp. LCCN ???? See also volumes II–VI [Tau61b, Tau63a, Tau62, Tau63b, Tau63c].

Taub:1961:JNCb

[Tau61b] A. H. Taub, editor. John von Neumann: Collected Works. Volume II: Operators, Ergodic Theory and Almost Periodic Functions in a Group. Pergamon Press, New York, NY, USA, 1961. x + 568 pp. LCCN ???? See also volumes I, III–VI [Tau61a, Tau63a, Tau62, Tau63b, Tau63c].

Taub:1962:JNC

[Tau62] A. H. Taub, editor. John von Neumann: Collected Works. Volume IV: Continuous Geometry and Other Topics. Pergamon Press, New York, NY, USA, 1962. x + 516 pp. LCCN ???? See also volumes I–III, V–VI [Tau61a, Tau61b, Tau63a, Tau63b, Tau63c].

Taub:1961:JNCc

[Tau63a] A. H. Taub, editor. John von Neumann: Collected Works. Volume III: Rings of Operators. Pergamon Press, New York, NY, USA, 1961–1963. ix + 574 pp. LCCN ???? See also volumes I–II, IV–VI [Tau61a, Tau61b, Tau62, Tau63b, Tau63c].

Taub:1963:JNCa

[Tau63b] A. H. Taub, editor. John von Neumann: Collected Works. Volume V: Design of Computers, Theory of Automata and Numerical Analysis. Pergamon Press, New York, NY, USA, 1963. ix + 784 pp. LCCN ???? See also volumes I–IV, VI [Tau61a, Tau61b, Tau63a, Tau62, Tau63c].

Taub:1963:JNCb

[Tau63c] A. H. Taub, editor. John von Neumann: Collected Works. Volume VI: Theory of Games, Astrophysics, Hydrodynamics and Meteorology. Pergamon Press, New York, NY, USA, 1963. x + 538 pp. LCCN ???? See also volumes I–V [Tau61a, Tau61b, Tau63a, Tau62, Tau63b].

Tausworthe:methods

[Tau76] Robert C. Tausworthe. Standardized development of computer software, part I. methods. Technical Report SP43-29, Jet Propulsion Laboratory, Pasadena, CA, USA, July 1976. 379 pages.

Tausworthe:1977:SDCa

[Tau77] Robert C. Tausworthe. Standardized Development of Computer Software. Part I. Methods. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1977. ISBN 0-13-842195-1. xiii + 379 pp. LCCN QA76.6 T39 1977.

Tausworthe:1979:SDCb

[Tau79] Robert C. Tausworthe. Standardized Development of Computer Software. Part II. Standards. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1979. ISBN 0-13-842203-6. x + 548 pp. LCCN QA76.6.T39 1976.

Taubes:1993:BSS

[Tau93] Gary Taubes. Bad science: the short life and weird times of cold fusion. Random House, New York, NY, USA, 1993. ISBN 0-394-58456-2. xxi + 503 pp. LCCN QC791.775.C64 T38 1993. US$25.00 (Canada US$32.50).

Taylor:1925:TPCb

[Tay25a] Hugh S. Taylor, editor. A Treatise on Physical Chemistry: a Co-operative Effort by a Group of Physical Chemists, volume 2. Van Nostrand Reinhold, New York, NY, USA, 1925. ix + 701– 1359 + 18 + 41 pp. LCCN QD 453 T21t 1925.

Taylor:1925:TPCa

[Tay25b] Hugh S. (Hugh Stott) Taylor, editor. A Treatise on Physical Chemistry: a Co-operative Effort by a Group of Physical Chemists, volume 1. Van Nostrand Reinhold, New York, NY, USA, 1925. xi + 603 + 18 + 41 pp. LCCN QD 453 T21t 1925.

Taylor:1963:IM

[Tay63] Edwin F. Taylor. Introductory Mechanics. Wiley, New York, NY, USA, 1963. xvii + 423 pp. LCCN QC127 .T3.

Taylor:1992:XOO

[Tay92]  Colin B. Taylor.  X/Open

and open systems, volume 1 of X/Open in action. X/Open Company, Ltd., Reading, UK, 1992. ISBN 1-872630-55-3. x + 118 pp. LCCN ????

Tamhane:2000:SDA

[TD00] Ajit C. Tamhane and Dorothy D.

Dunlop. Statistics and Data Analysis: from Elementary to Intermediate. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-744426-5 (hardcover). xiv + 722 pp. LCCN QA276 .T25 2000.

Tedlow:2003:WDF

[Ted03] Richard S. Tedlow. The Watson Dynasty: the Fiery Reign and Troubled Legacy of IBM’s Founding Father and Son. Harper-Business, New York, NY, USA, 2003. ISBN 0-06-001405-9. x + 340 pp. LCCN HD9696.2.U64 I258 2003. URL http://www.

loc.gov/catdir/description/
hc042/2003051145.html;   http:
//www.loc.gov/catdir/toc/
fy045/2003051145.html.

Tedlow:2006:AGL

[Ted06] Richard S. Tedlow. Andy Grove:

the life and times of an American. Portfolio, New York, NY, USA, 2006. ISBN 1-59184-139-9. xxii + 568 + 16 pp. LCCN HD9696.S44 T588 2006.

Temple:1991:GCY

[Tem91] Robert K. G. Temple. The Genius of China: 3,000 Years of Science, Discovery, and Invention. Prion, London, UK, 1991. ISBN 1-85375-078-6, 1-85375-292-4 (paperback). 248 pp.

LCCN M92.D00752; 02.E04877; 99.E01411.

Tent:2006:PMC

[Ten06] M. B. W. (Margaret B. W.) Tent. The Prince of Mathematics: Carl Friedrich Gauss. A. K. Peters, Ltd., Wellesley, MA, USA, 2006. ISBN 1-56881-261-2. xviii + 245 pp. LCCN QA29.G3 T46 2006.

Tent:2009:LEB

[Ten09] M. B. W. (Margaret B. W.) Tent. Leonhard Euler and the Bernoullis: mathematicians from Basel. A. K. Peters, Ltd., Wellesley, MA, USA, 2009. ISBN 1-56881-464-X. xix + 276 pp. LCCN QA28 .T46 2009.

Teska:2007:PSM

[Tes07] Kirk Teska. Patent savvy for managers: spot and protect valuable innovations in your company. Nolo Press, Berkeley, CA, USA, 2007. ISBN 1-4133-0694-2 (paperback). 278 pp. LCCN KF3114 .T37 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0714/2007013012.html.

Tondo:1989:CAB

[TG89] Clovis L. Tondo and Scott E. Gimpel. The C Answer Book: Solutions to the Exercises in The C Programming Language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1989. ISBN 0-13-109653-2. 208 pp. LCCN QA76.73 C15 T66 1989. This book contains answers to the problems in [KR88].

Thomas:2001:PRP

[TH01] David Thomas and Andrew Hunt. Programming Ruby: The Pragmatic Programmer’s Guide. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-71089-7 (paperback). xxxii + 564 pp. LCCN QA76.64 .T494 2001. US$42.95. URL http:

//cseng.aw.com/book/0,3828,
0201710897,00.html;   http:
//cseng.aw.com/catalog/isbn/
0,3837,0+0-201-71089-7,00.
html.

Unicode:1991:USW

[The91] The Unicode Consortium. The Unicode Standard: Worldwide Character Encoding. Version 1.0. Volumes 1 and 2. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-56788-1 (paperback, vol. 1), 0-201-60845-6 (paperback, vol. 2). LCCN QA268 .U55 1991, Z103 .U6 1991. US$39.95 (paperback), US$32.95 (hardcover).

Mathworks:1992:SEM

[The92] The MathWorks, Inc. The Student Edition of Matlab for MS-DOS Personal Computers. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-855974-0, 0-13-855982-1. xiv + 494 pp. LCCN QA297 .S8433 1992.

Unicode:1996:USV

[The96] The Unicode Consortium. The Unicode Standard, Version 2.0. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-48345-9 (paperback), 0-201-96623-9 (CD-ROM). various pp. LCCN QA268.U56 1996. US$54.95.

Be:1997:DGO

[The97] The Be Development Team. The Be Developer’s Guide: The Official Documentation for the BeOS. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-287-5. xi + 934 pp. LCCN QA76.76.O63 B43 1997. US$49.95. URL

http://www.ora.com/catalog/
be/;   http://www.oreilly.
com/catalog/bedev.

Be:1998:AT

[The98] The Be Development Team. Be Advanced Topics: Official documentation for the BeOS. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, July 1998. ISBN 1-56592-396-0. xi + 365 pp. LCCN QA76.76.O63 B4 1998. US$39.95. URL http:

//www.oreilly.com/catalog/

beadv. Includes BeSpecific 5 CD-ROM.

Unicode:2000:USV

[The00] The Unicode Consortium. The Unicode Standard, Version 3.0. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-61633-5. xxx + 1040 pp. LCCN QA268.U56 1996. US$50.00.

URL http://www.unicode.org/
book/;   http://www.unicode.
org/unicode/standard/versions/

. Includes CD-ROM. The principal authors and editors of The Unicode Standard, Version 3.0 are Joan Aliprand, Julie Allen, Joe Becker, Mark Davis, Michael Everson, Asmus Freytag, John Jenkins, Mike Ksar, Rick Mc-Gowan, Lisa Moore, Michel Suignard, and Ken Whistler.

Unicode:2003:USV

[The03] The Unicode Consortium. The Unicode Standard, Version 4.0. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-321-18578-

1. xxxviii + 1462 pp. LCCN QA268 .U545 2004. URL http:

//www.unicode.org/versions/
Unicode4.0.0/. Includes CD-

ROM.

Unicode:2006:USV

[The06] The Unicode Consortium. The Unicode Standard, Version 5.0. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-48091-0 (hardcover). l + 1417 pp. LCCN QA268 .U545 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0617/2006023526.html;
http://www.unicode.org/versions/
beta.html;   http://www.unicode.
org/versions/Unicode5.0.0/.

Includes CD-ROM.

Thompson:1981:MA

[Tho81] Morris M. Thompson. Maps for America. United States Government Printing Office, Washington, DC, USA, 1981. xiv + 265 pp. LCCN GA405 .T46 1981. US$11.00.

Thompson:trusting-trust

[Tho84] Ken Thompson. Reflections on trusting trust. Communications of the Association for Computing Machinery, 27(8):761–763, August 1984.

Thompson:1997:ACMb

[Tho97] William J. (William Jackson) Thompson. Atlas for computing mathematical functions: an illustrated guide for practitioners with programs in Fortran 90 and Mathematica. Wiley, New York, NY, USA, 1997. ISBN 0-471-18171-4 (cloth). xiv + 888 pp. LCCN QA331.T386 1997. Includes CD-ROM.

Thoman:2008:MMG

[Tho08a] Peter Thoman. Multigrid methods on GPUs: high performance GPGPU solvers for partial differential equations. VDM Verlag Dr. Müller, Saarbrücken, Germany, 2008. ISBN 3-8364-9970-3. viii + 61 pp. LCCN QA297 .T45 2008.

Thomson:2008:LMG

[Tho08b] Keith Stewart Thomson. The legacy of the Mastodon: the golden age of fossils in America. Yale University Press, New Haven, CT, USA, 2008. ISBN 0-300-11704-3. xvii + 386 + 16 pp. LCCN QE882.P8 T46 2008.

Thulstrup:1980:ALM

[Thu80] Erik Waaben Thulstrup. Aspects of the Linear and Magnetic Circular Dichroism of Planar Organic Molecules, volume 14 of Lecture Notes in Chemistry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 0-387-09754-6. vi + 100 pp. LCCN QD476 .T45.

Tinkham:1964:GTQ

[Tin64] Michael Tinkham. Group Theory and Quantum Mechanics. Mc-Graw-Hill, New York, NY, USA, 1964. xii + 340 pp. LCCN QC174.5 .T589 1964.

Tipler:1976:P

[Tip76] Paul A. Tipler. Physics. Worth Publishers, Inc., New York, NY, USA, 1976. ISBN 0-87901-041-X. xxvi + 1026 pp. LCCN QC21.2.T548.

Thatcher:1971:NWE

[TM71] Virginia Sarah Thatcher and Alexander McQueen, editors. The New Webster Encyclopedic Dictionary of the English Language. Consolidated Book Publishers, Chicago, IL, USA, 1971. ISBN 0-8326-0021-0. xx + 972 + 424 pp. LCCN PE1625 .N54 1974.

Todino:1987:UUU

[Tod87] Grace Todino. Using UUCP and Usenet. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, July 1987. ISBN 0-937175-10-2. xv + 194 pp. LCCN QA 76.76 O63 T63 1991.

Tolstov:1962:FS

[Tol62] Georgi P. Tolstov. Fourier Series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1962. 336 pp. LCCN QA404 .T573.

TorresyQuevedo:1915:EAS

[Tor15] L. Torres y Quevedo. Essais sur l’automatique. sa definition. Étendue théorique de ses applications (French) [Essays on automation. Its definition. Theoretical extent of its applications]. Revue Générale des Sciences Pures et Appliquées, ??(??):601– 611, November 15, 1915. ISSN 0370-7431. Reprinted in [Ran82, §2.5]. Translated by Mr. R. Basu.

Treiman:1999:OQ

[Tre99] Sam B. (Sam Bard) Treiman. The Odd Quantum. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-00926-0 (hardcover), 0-691-10300-3 (paperback), 1-4008-0810-3 (ebook), 1-4008-1368-9 (e-book), 1-4008-2309-9 (e-book). viii + 262 pp. LCCN QC 174.12.T73 1999.

Trefethen:2000:SMM

[Tre00] Lloyd N. Trefethen. Spectral Methods in MATLAB. Software, environments, tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2000. ISBN 0-89871-465-6 (paperback), 0-89871-959-3 (ebook). xvi + 165 pp. LCCN QA377 .T65 2000.

Trimmer:1998:EMS

[Tri98] Joseph F. Trimmer. The essentials of MLA style: a guide to documentation for writers of research papers, with an appendix on APA style. Houghton-Mifflin, Boston, MA, USA, fifth edition, 1998. ISBN 0-395-88316-4. 70 pp. LCCN PN147 .T75 1998.

Trimmer:1999:GMD

[Tri99] Joseph F. Trimmer. A Guide to MLA Documentation: with an Appendix on APA Style. Houghton-Mifflin, Boston, MA, USA, 1999. ISBN 0-395-93851-1 (paperback). iii + 50 pp. LCCN PN147 .T75 1999.

Trott:2006:MGN

[Tro06a] Michael Trott. The Mathematica Guidebook for Numerics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-95011-7. xxxv + 1208 pp. LCCN QA76.95 .T773 2006. Includes DVD-ROM.

Trott:2006:MGS

[Tro06b] Michael Trott. The Mathematica Guidebook for Symbolics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 0-387-95020-6. xxxviii + 1453 pp. LCCN QA76.73.M29 T765 2006. Includes CD-ROM.

Troutwine:2018:HCR

[Tro18] Brian L. Troutwine. Hands-on Concurrency with Rust: Confidently Build Memory-safe, Parallel, and Efficient Software in Rust. Packt Publishing, Birmingham, UK, 2018. ISBN 1-78839-997-8 (paperback), 1-78847-835-5. v + 449 pp. LCCN QA76.76.A65. URL http://

proquest.safaribooksonline.
com/?fpi=9781788399975.

Truss:2003:ESL

[Tru03] Lynne Truss. Eats, Shoots & Leaves: The Zero Tolerance Approach to Punctuation. Gotham Books, New York, NY, USA, 2003. ISBN 1-59240-087-6, 1-59240-203-8 (paperback). xxvii + 209 pp. LCCN PE1450 .T75 2004. US$17.50, CAN$26.00.

Trzeciak:1995:WMP

[Trz95] Jerzy Trzeciak. Writing Mathematical Papers in English: a Practical Guide. European Mathematical Society Publishing House, Seminar for Applied Mathematics, Zürich, Switzerland, revised edition, 1995. ISBN 3-03719-014-0, 83-85694-02-1. 49 pp. LCCN QA42 .T79 1995.

                                        URL http://www.ems-ph.org/
                                        books/book.php?proj_nr=34.

Todino:1993:LUO

[TSP93] Grace Todino, John Strang, and Jerry Peek. Learning the UNIX Operating System. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, August 1993. ISBN 1-56592-060-0. xv + 92 pp. LCCN QA76.76.O63 T62 1994. US$9.95. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0915/94242752-
                                        b.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0915/
                                        94242752-d.html;   http://
                                        www.loc.gov/catdir/enhancements/
                                        fy1113/94242752-t.html.

Thompson:2000:PHN

[TT00] Robert Bruce Thompson and Barbara Fritchman Thompson. PC Hardware in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 2000. ISBN 1-56592-599-8. xxi + 501 pp. LCCN TK7887.5 .T48 2000. US$29.95.

Tucker:1977:PL

[Tuc77] Allen B. Tucker, Jr. Programming Languages. McGraw-Hill Computer Science Series, Editor: Richard W. Hamming and Edward A. Feigenbaum. McGraw-Hill, New York, NY, USA, 1977. ISBN 0-07-065415-8. xv + 439 pp. LCCN QA76.7 T8 1977.

Tufte:1983:VDQ

[Tuf83] Edward R. Tufte. The Visual Display of Quantitative Information. Graphics Press, Box 430, Cheshire, CT 06410, USA, 1983. ISBN 0-9613921-0-X, 0-9613921-4-2. 197 pp. LCCN K27.S8 T84. US$34.00.

Tufte:1990:EI

[Tuf90] Edward R. Tufte. Envisioning Information. Graphics Press, Box 430, Cheshire, CT 06410, USA, 1990. 126 pp. LCCN P93.5 .T84 1990. US$48.00.

Tufte:1997:VEI

[Tuf97] Edward R. Tufte. Visual Explanations: Images and Quantities, Evidence and Narrative. Graphics Press, Box 430, Cheshire, CT 06410, USA, 1997. ISBN 0-9613921-2-6. 156 + 2 pp. LCCN P93.5.T846 1997. US$45.00.

Turnbull:1993:GM

[Tur93] Herbert Westren Turnbull. The Great Mathematicians. Barnes and Noble, New York, NY, USA, 1993. ISBN 1-56619-157-2. xi + 141 pp. LCCN QA28 .T8 1993.

Turabian:2007:MWR

[Tur07] Kate L. Turabian, editor. A manual for writers of research papers, theses, and dissertations: Chicago style for students and researchers. Chicago guides to writing, editing, and publishing. University of Chicago Press, Chicago, IL, USA and London, UK, seventh edition, 2007. ISBN 0-226-82336-9 (cloth), 0-226-82337-7 (paperback). xviii + 466 pp. LCCN LB2369 .T8 2007; LB2369 .T87 2007. URL

http://www.loc.gov/catdir/
enhancements/fy0707/2006025443-
b;   http://www.loc.gov/catdir/
enhancements/fy0707/2006025443-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0707/
2006025443-d;   http://www.
loc.gov/catdir/enhancements/
fy0707/2006025443-d.html;
http://www.loc.gov/catdir/
toc/ecip0618/2006025443;
http://www.loc.gov/catdir/
toc/ecip0618/2006025443.html.

Revised by Wayne C. Booth, Gregory G. Colomb, Joseph M. Williams, and the University of Chicago Press editorial staff.

Turing:2018:XYZ

[Tur18] Dermot Turing. X, Y and Z:

the Real Story of How Enigma Was Broken. The History Press, Gloucestershire, UK, 2018. ISBN 0-7509-8782-0 (hardcover), 0-7509-8967-X (ePub). 319 + 1 pp. LCCN D810.C88 T87 2018.

Turing:2020:CBP

[Tur20] Dermot Turing. The Codebreakers of Bletchley Park: the Secret Intelligence Station That Helped Defeat the Nazis. Arcturus Publishing Limited, London, UK, 2020. ISBN 1-78950-621-2, 1-83857-650-9 (paperback). 256 pp. LCCN ???? With an introduction by Christopher M. Andrew.

tHooft:2014:TPT

[tVW14] Gerard ’t Hooft, S. (Stefan)

Vandoren, and Steven Weinberg. Time in Powers of Ten: Natural Phenomena and Their Timescales. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2014. ISBN 981-4489-80-8 (hardcover), 981-4489-81-6 (paperback). xviii + 212 pp. LCCN QB981 .H72 2014. URL http://www.

worldscientific.com/action/
doSearch?AllField=981-4489-
80-8.

Tymms:1987:AI

[TW87] W. R. Tymms and M. D. Wyatt. The Art of Illuminating. Chartwell Books, Inc., 110 Enterprise Avenue, Secaucus, NJ 07094, USA, 1987. ISBN 1-55521-198-4. 104 pp. LCCN ND3310 .T95 1987.

Tsuji:structured-fortran

[TW88] Tatsui Tsuji and Katsumasa Watanabe. Structured Fortran preprocessors generating optimized output. Software—Practice and Experience, 18(5):427– 442, May 1988.

Taylor:1995:RTP

[TW95] Richard Taylor and Andrew Wiles. Ring-theoretic properties of certain Hecke algebras. Annals of Mathematics, 142:553– 572, 1995. This paper is a companion to [Wil95a], providing the remedy for the flaw in Wiles’ 1993 proof of Fermat’s Last Theorem. See also [Sin97].

Tanenbaum:1997:OSD

[TW97] Andrew S. Tanenbaum and Albert S. Woodhull. Operating Systems: Design and Implementation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1997. ISBN 0-13-638677-6. xvii + 939 pp. LCCN QA76.76.O63T36 1997. US$62.00. Includes CD-ROM.

Thomas:1982:UGU

[TY82] Rebecca Thomas and Jean Yates. A User Guide to the UNIX System. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, 1982. ISBN 0-931988-71-3. xi + 508 pp. LCCN QA76.8.U65 T45 1982.

Toy:1986:CHS

[TZ86] Wing N. Toy and Benjamin Zee. Computer Hardware/ Software Architecture. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-163502-6. xviii + 446 pp. LCCN QA76.9A73 T68 1986.

URL https://archive.org/
details/computerhardware0000toyw.

Ueberhuber:1997:NCM

[Ueb97] Christoph W. Ueberhuber. Numerical Computation: Methods, Software, and Analysis. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-62058-3 (vol. 1: softcover), 3-540-62057-5 (vol. 2: softcover), 3-642-59118-3 (e-book). xvi + 474 (vol. 1), xvi + 495 (vol. 2) pp. LCCN QA297 .U2413 1997. US$44.95 (vol. 1), US$49.95 (vol. 2).

Ulichney:1987:DH

[Uli87]  Robert Ulichney.  Digital

Halftoning. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-21009-6. xiv + 362 pp. LCCN T385 .U451 1987.

Unger:1987:FGF

[Ung87] J. Marshall Unger. The Fifth Generation Fallacy: Why Japan is Betting its Future on Artificial Intelligence. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1987. ISBN 0-19-504939-X. x + 230 pp. LCCN QA76.85 .U541 1987. US$19.95. See [Rei88b].

Upstill:1990:RC

[Ups90] Steve Upstill. The RenderMan Companion. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-50868-0. xix + 475 pp. LCCN T385 .U67 1990.

Urroz:2001:NSMa

[Urr01a] Gilberto E. Urroz. Numerical and Statistical Methods with SCILAB for Science and Engineering, volume 1. BookSurge Publishing, North Charleston, SC, USA, 2001. ISBN 1-58898-304-8. xv + 563 pp. LCCN TA345 .U77 2001.

Urroz:2001:NSMb

[Urr01b] Gilberto E. Urroz. Numerical and statistical methods with SCILAB for science and engineering, volume 2. Greatunpublished.com, Charleston, SC, USA, 2001. ISBN 1-58898-305-6. xv + 486 pp. LCCN TA345 .U77 2001.

USENIX:PUC94

[USE94] Proceedings of the 1994 USENIX Conference, January 17–21, 1994, San Francisco, CA, USA. USENIX, San Francisco, CA, USA, 1994. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84 1994.

Vahalia:1996:UI

[Vah96] Uresh Vahalia. UNIX Internals. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-101908-2. xxxiii + 601 pp. LCCN QA76.76.063V33

VanWyk:ideal

[Van82] Christopher J. Van Wyk. A high-level language for specifying pictures. ACM Transactions on Graphics, 1(2):163–182, April 1982.

VanWyk:awk

[Van86] Christopher J. Van Wyk. AWK as glue for programs. Software— Practice and Experience, 16(4): 369–388, April 1986.

VanWyk:1988:DSC

[Van88] Christopher J. Van Wyk. Data Structures in C. Addison-Wes-ley, Reading, MA, USA, 1988. ISBN 0-201-16116-8. x + 387 pp. LCCN QA76.73.C15 V36 1988.

VanWolferen:1989:EJP

[Van89] Karel Van Wolferen. The enigma of Japanese power: people and politics in a stateless nation. Alfred A. Knopf, New York, NY, USA, 1989. ISBN 0-333-44321-7, 0-394-57796-5. 496 pp. LCCN JQ1681 .W6521 1989. US$24.95, UK£16.95.

vanHerwijnen:1990:PS

[van90] Eric van Herwijnen. Practical SGML. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0635-X, 0-7923-9434-8. xviii + 307 pp. LCCN QA76.73.S44 V36 1990. UK£24.90, US$49.00.

VanLoan:1992:CFF

[Van92] Charles F. Van Loan. Computational Frameworks for the Fast Fourier Transform. Frontiers in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. ISBN 0-89871-285-8 (paperback), 1-61197-099-7 (ebook). xiii + 273 pp. LCCN QA403.5 .V35 1992.

vanHerwijnen:1994:PS

[van94]  Eric van Herwijnen.  Prac-

tical SGML. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, second edition, 1994. ISBN 0-7923-9434-8. xx + 288 pp. LCCN QA76.73.S44 V36 1994.

Vandevoorde:1999:CSC

[Van99]  David Vandevoorde.  C++

solutions: companion to The C++ Programming Language, Third edition. Addison-Wes-ley, Reading, MA, USA, 1999. ISBN 0-201-30965-3 (paperback). vii + 292 pp. LCCN QA76.73.C153V36 1998. US$29.95. See [Str97].

vanCalmthout:2018:SGH

[vCH18] Martijn van Calmthout and Michiel Horn, editors. Sam Goudsmit and the hunt for Hitler’s atom bomb. Prometheus Books, Amherst, NY, USA, 2018. ISBN 1-63388-450-3 (hardcover), 1-63388-451-1 (e-book). 243 + 14 pp. LCCN DS135.N6 G65613 2018.

Viswanath:2000:RFS

[Vis00] Divakar Viswanath. Random Fibonacci sequences and the number 1.13198824 . . .. Mathematics of Computation, 69(231):1131– 1155, July 2000. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://www.ams.org/journal-
getitem?pii=S0025-5718-99-
01145-X;   http://www.ams.
org/mcom/2000-69-231/S0025-
5718-99-01145-X/S0025-5718-
99-01145-X.dvi;   http://
www.ams.org/mcom/2000-69-
231/S0025-5718-99-01145-X/
S0025-5718-99-01145-X.pdf;
 http://www.ams.org/mcom/
2000-69-231/S0025-5718-99-
01145-X/S0025-5718-99-01145-
X.ps;   http://www.ams.org/
mcom/2000-69-231/S0025-5718-
99-01145-X/S0025-5718-99-
01145-X.tex.

vanLeunen:1992:HS

[vL92] Mary-Claire van Leunen. A handbook for scholars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, revised edition, 1992. ISBN 0-394-40904-3 (hardcover), 0-394-73395-9 (paperback). xi + 348 pp. LCCN PN146 .V36 1992.

vonLaue:2023:HP

[vLPO23] Max von Laue, Vesselin Petkov, and Ralph Oesper. History of Physics. Minkowski Institute Press, Montréal, Québec, Canada, 2023. ISBN 1-998902-01-3. vi + 176 pp. LCCN ???? English translation of German original Geschichte der Physik (1946, 1950, 1957).

vonNeumann:1945:FDR

[vN45] John von Neumann. First draft of a report on the EDVAC. Technical report, University of Pennsylvania, June 30, 1945. Reprinted in [Ran82, §8.2].

vonOstermann:1952:MFL

[von52] Georg F. von Ostermann. Manual of Foreign Languages for the Use of Printers and Translators. United States Government Printing Office, Washington, DC, USA, fourth edition, 1952. xii + 347 pp. LCCN Z253 .V94 1952.

Voss:2012:PL

[Voß12] Herbert Voß. Presentations with LaTEX: which package, which command, which syntax? Dante. Lehmanns Media, Berlin, Germany, 2012. ISBN 3-86541-496-6. vi + 205 pp. LCCN ????

Vulis:1992:MTA

[Vul92] Michael Vulis. Modern TEX and its Applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1992. ISBN 0-8493-4431-X. vii + 294 pp. LCCN Z253.4.T47V84 1993. US$32.95.

Walsh:1975:MO

[Wal75] G. R. Walsh. Methods of Optimization. Wiley, New York, NY, USA, 1975. ISBN 0-471-91922-5 (hardcover), 0-471-91924-1 (paperback). x + 200 pp. LCCN QA402.5 .W341.

Wallis:1982:PP

[Wal82] Peter J. L. Wallis. Portable Programming. Wiley, New York, NY, USA, 1982. ISBN 0-470-27331-3. xiv + 141 pp. LCCN QA76.6.W329 1982.

Wall:perl

[Wal87] Larry Wall. perl: Practical extraction and report language. Usenet mod.sources archives, 1987. Electronic mail: lwall@

jpl-devvax.jpl.nasa.gov.

Walsh:1994:MTW

[Wal94] Norman Walsh. Making TEX Work. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1994. ISBN 1-56592-051-1. xxxvi + 483 pp. LCCN TK5105.875.I57 M36 1994; Z253.4.T47 W34 1994. US$29.95. URL ftp:

//ftp.ora.com/pub/products/
catalogs/book.catalog;   http:
//www.oreilly.com/catalog/

9781565920514. Written specifically for database maintainers and system administrators, this Unix-based technical guide covers installing, setting up, and running Internet applications such as gopher holes, FTP and telnet sites, mailing lists, WAIS, and World Wide Web pages, and other sites, as well as keeping them secure.

Waldrop:2001:DMJ

[Wal01] M. Mitchell Waldrop. The Dream Machine: J. C. R. Licklider and the Revolution That Made Computing Personal. The Sloan technology series. Viking, New York, NY, USA, 2001. ISBN 0-670-89976-3. 502 pp. LCCN QA76.17 .W35 2001. URL

http://www.loc.gov/catdir/
toc/fy031/2001017985.html.

Walker:2003:SES

[Wal03]  Gabrielle Walker.  Snow-

ball Earth: the Story of the Great Global Catastrophe That Spawned Life as We Know It. Crown Publishers, New York, NY, USA, 2003. ISBN 0-609-60973-4 (hardcover), 1-4000-5125-8. xiii + 269 pp. LCCN QE507 .W35 2003.

Ward:1985:APT

[War85] Terry A. Ward. Applied Programming Techniques in C. Scott, Foresman and Company, Glenview, IL, USA, 1985. ISBN 0-673-18050-6. xii + 349 pp. LCCN QA76.73.C15 W37 1985. US$19.95.

Warren:2003:HD

[War03] Henry S. Warren. Hacker’s delight. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-91465-4. xiv + 306 pp. LCCN QA76.6 .W375 2003. URL

http://www.awprofessional.
com/bookstore/product.asp?
isbn=0201914654;   http:/
/www.hackersdelight.org/;
http://www.hackersdelight.
org/hackerTOC.pdf;   http:
//www.informit.com/content/
images/chap3_0201914654/elementLinks/
0201914654.pdf. While this

book does not specifically address computational aspects of floating-point arithmetic (apart from the nine-page Chapter 15), it has extensive coverage of, and clever algorithms for, integer arithmetic operations that are fundamental for implementing hardware floating-arithmetic and software multiple-precision arithmetic.

Ward:2008:UGS

[War08] Peter D. (Peter Douglas) Ward.

Under a green sky: global warming, the mass extinctions of the past, and what they can tell us about our future. Smithsonian Books/Collins, New York, NY, USA, 2008. ISBN 0-06-113792-8 (paperback). xiv + 242 pp. LCCN QE721.2.E97 W384 2008.

Warren:2013:HD

[War13] Henry S. Warren. Hacker’s Delight. Addison-Wesley, Reading, MA, USA, second edition, 2013. ISBN 0-321-84268-5 (hardcover). xvi + 494 pp. LCCN QA76.6 .W375 2013. URL http:

//www.pearsonhighered.com/
educator/product/Hackers-
Delight/9780321842688.page.

Watson:1990:FSC

[Wat90] Thomas J. Watson Jr. Father Son & Co.: My Life at IBM and Beyond. Bantam Books, New York, NY, USA, 1990. ISBN 0-553-07011-8. xi + 468 + 32 pp. LCCN HD9696.C64 I4887 1990. Memoirs of IBM President Watson, the son of the founder of IBM.

Watts:2000:OGD

[Wat00] Niki Watts. Oxford Greek Dictionary: Greek–English, English– Greek. Berkley Books, New York, NY, USA, American edition, 2000. ISBN 0-425-17600-2. x + 470 pp. LCCN PA1139.E5 O94 2000.

Watkins:2007:MEP

[Wat07] David S. Watkins. The Matrix Eigenvalue Problem: GR and Krylov Subspace Methods, volume 101 of Other titles in applied mathematics. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2007. ISBN 0-89871-641-1. x + 442 pp. LCCN QA193 .W38 2007.

Ward:2000:REW

[WB00] Peter D. (Peter Douglas) Ward and Donald Brownlee. Rare earth: why complex life is uncommon in the universe. Copernicus, New York, NY, USA, 2000. ISBN 0-387-95289-6 (paperback), 0-387-21848-3 (ebook). xxxii + 335 pp. LCCN QB54 .W336 2000. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0817/2003043437-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0817/
                                        2003043437-t.html.

Warwick:2004:HEB

[WB01] Andrew Warwick and Jed Z. Buchwald. Histories of the Electron: The Birth of Microphysics. Dibner Institute studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-02494-2 (hardcover), 0-262-52424-4 (paperback). xi + 513 pp. LCCN QC793.5.E62 H57 2004. URL

https://mitpress.mit.edu/
books/histories-electron;
https://www.scribd.com/document/
216554790/Histories-of-the-
Electron-The-Birth-of-Microphysics.

Weast:1985:CHC

[WBA85] Robert C. Weast, William H.

Beyer, and Melvin J. Astle, editors. CRC Handbook of Chemistry and Physics. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 66th edition, 1985. ISBN 0-8493-0466-0. 2384 pp. LCCN QD65 .C73 1985-86.

Wadleigh:2000:SOH

[WC00] Kevin R. Wadleigh and Isom L.

Crawford. Software Optimization for High Performance Computing: Creating Faster Applications. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-017008-9. xxxi + 377 pp. LCCN QA76.76.D47 W34 2000. US$45.00. URL http:

//vig.prenhall.com/catalog/
academic/product/1,4096,0130170089,
00.html.

Weaver:2008:OIO

[Wea08] David L. Weaver, editor. OpenSPARC Internals: OpenSPARC T1/T2 Chip Multithreaded Throughput Computing. Lulu, Inc., 860 Aviation Parkway, Suite 300, Morrisville, NC 27560, USA, 2008. ISBN 0-557-01974-5. xviii + 369 pp. LCCN ???? US$20.00. URL http://www.

                                        opensparc.net/publications/
                                        books/opensparc-internals.
                                        html.

Weart:2012:RNF

[Wea12] Spencer R. Weart. The rise of nuclear fear. Harvard University Press, Cambridge, MA, USA, 2012. ISBN 0-674-05233-1. ix + 367 pp. LCCN QC773 .W44 2012.

Wegner:1980:PAI

[Weg80] Peter Wegner. Programming with Ada: An Introduction by Means of Graduated Examples. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1980. ISBN 0-13-730697-0. xi + 211 pp. LCCN QA76.73.A35 W4 1980.

Weitzman:1980:DMM

[Wei80]  Cay Weitzman.  Distributed

Micro/Minicomputer Systems: Structure, Implementations, and Application. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1980. ISBN 0-13-216481-7. xii + 403 pp. LCCN QA76.9.D5 .W44 1980. US$22.50.

Weintraub:2011:HOU

[Wei11] David A. (David Andrew) Weintraub. How Old Is the Universe? Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14731-0 (hardcover). 370 pp. LCCN QB501 .W45 2011.

Weiner:2016:GGS

[Wei16] Eric Weiner. The geography of genius: a search for the world’s most creative places from ancient Athens to Silicon Valley. Simon and Schuster, New York, NY, USA, 2016. ISBN 1-4516-9165-3 (hardcover), 1-4516-9167-X (paperback), 1-4516-9168-8 (ebook). 353 pp. LCCN BF412 .W395 2016.

Weitz:2018:WGP

[Wei18] Eric D. Weitz. Weimar Germany: promise and tragedy. Princeton University Press, Princeton, NJ, USA, Weimar centennial edition, 2018. ISBN 0-691-18305-8 (paperback). xix + 474 pp. LCCN DD237 .W47 2018.

Wells:2005:PNM

[Wel05] D. G. (David G.) Wells. Prime Numbers: the Most Mysterious Figures in Math. Wiley, New York, NY, USA, 2005. ISBN 0-471-46234-9 (hardcover), 0-471-71892-0 (ebook). xv + 272 pp. LCCN QA246 .W35 2005. URL

http://www.loc.gov/catdir/
enhancements/fy0616/2004019974-
b.html;   http://www.loc.
gov/catdir/enhancements/fy0616/
2004019974-d.html;   http:
//www.loc.gov/catdir/toc/
ecip0422/2004019974.html.

Weaver:1994:SAM

[WG94] David L. Weaver and Tom Germond. The SPARC Architecture Manual: Version 9. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-099227-5. xxii + 357 pp. LCCN QA76.9.A73S648 1992. US$33.00. URL http:

                                        //www.sparc.org/standards/
                                        SPARCV9.pdf.

Wolcott:hershey-fonts

[WH76] Norman M. Wolcott and Joseph Hilsenrath. A contribution to computer typesetting techniques. Tables of coordinates for Hershey’s repertoire of occidental type fonts and graphics symbols. Technical Report PB-251 845, U. S. National Bureau of Standards, April 1976. NBS Special Publication 424.

Wheelan:2014:NSS

[Whe14] Charles J. Wheelan. Naked statistics: stripping the dread from the data. W. W. Norton & Co., New York, NY, USA, 2014. ISBN 0-393-34777-X (paperback), 0-393-07195-2 (hardcover). xviii + 282 pp. LCCN QA276.W458 2014.

Wheatley:2021:SWP

[Whe21] Denys N. Wheatley. Scientific writing and publishing: a comprehensive manual for authors. Cambridge University Press, Cambridge, UK, 2021. ISBN 1-108-79980-9 (paperback), 1-108-83520-1 (hardcover), 1-108-87991-8 (PDF e-book). xii + 200 pp. LCCN T11 .W4475 2021.

White:1983:MG

[Whi83] Jan V. White. Mastering Graphics. R. R. Bowker Company, 1180 Avenue of the Americas, New York, NY 10036, USA, 1983. ISBN 0-8352-1704-3. xiii + 180 pp. LCCN Z253 .W47 1983.

White:1988:GDE

[Whi88] Jan V. White. Graphic Design for the Electronic Age. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 1988. ISBN 0-8230-2121-1, 0-8230-2122-X (paperback). xi + 211 pp. LCCN Z286.D47 W5 1988.

White:1990:CEA

[Whi90a] Jan V. White. Color for the Electronic Age. Watson-Guptill Publications, 1515 Broadway, New York, NY 10036, USA, 1990. ISBN 0-8230-0732-4. 207 pp. LCCN Z 286 D47 W49 1990.

White:1990:GP

[Whi90b] Jan V. White. Great Pages. Serif Publishing, A subsidiary of Xerox Corporation, 701 South Aviation Boulevard, El Segundo, CA 90245, USA, 1990. ISBN 1-878567-01-2. 128 pp. LCCN Z286.D47 W5 1990.

Wiederhold:1977:DD

[Wie77] Gio Wiederhold. Database Design. McGraw-Hill Computer Science Series, Editor: Richard W. Hamming and Edward A. Feigenbaum. McGraw-Hill, New York, NY, USA, 1977. ISBN 0-07-070130-X, 0-07-085859-4. xiii + 658 pp. LCCN QA76.9.D3 W531.

Wiederhold:1983:DD

[Wie83a] Gio Wiederhold. Database Design. McGraw-Hill Computer Science Series, Editor: Richard W. Hamming and Edward A. Feigenbaum. McGraw-Hill, New York, NY, USA, second edition, 1983. ISBN 0-07-070132-6. xvi + 751 pp. LCCN QA76.9.D3 W531 1983. US$29.95.

Wiederhold:1983:FOD

[Wie83b] Gio Wiederhold. File Organization for Database Design. McGraw-Hill Computer Science Series, Editor: Richard W. Hamming and Edward A. Feigenbaum. McGraw-Hill, New York, NY, USA, 1983.

Wiederhold:1987:FOD

[Wie87] Gio Wiederhold. File Organization for Database Design. McGraw-Hill Computer Science Series, Editor: Richard W. Hamming and Edward A. Feigenbaum. McGraw-Hill, New York, NY, USA, second edition, 1987. ISBN 0-07-070133-4. xiv + 619 pp. LCCN QA76.9.F5 W53 1987. US$38.95.

Wilkinson:1965:AEP

[Wil65] J. H. Wilkinson. The Algebraic Eigenvalue Problem. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1965. ISBN 0-19-853403-5. xviii + 662 pp. LCCN QA218 .W686 1965.

Wilf:1986:AC

[Wil86a] Herbert S. (Saul) Wilf. Algorithms and Complexity. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-021973-8 (hardcover), 0-13-022054-X (paperback). vi + 231 pp. LCCN QA63 .W55 1986.

Will:1986:WER

[Wil86b] Clifford M. Will. Was Einstein Right?: Putting General Relativity to the Test. Basic Books, New York, NY, USA, 1986. ISBN 0-465-09088-5. xii + 274 + 8 pp. LCCN QC173.6 .W55 1986. US$18.95.

Wiles:1995:MEC

[Wil95a] Andrew Wiles. Modular elliptic curves and Fermat’s Last Theorem. Annals of Mathematics, 142:443–551, 1995. This paper contains the bulk of the author’s proof of the Taniyama-Shimura conjecture and Fermat’s Last Theorem. The companion paper [TW95] contains the solution to the flaw discovered in the proof that Wiles announced on June 23, 1993, in Cambridge, England. See also [Sin97].

Wilson:1995:IF

[Wil95b] Alistair Macintosh Wilson. The infinite in the finite. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1995. ISBN 0-19-853950-9. 524 pp. LCCN QA21 .W385 1995. US$37.50.

URL http://www.myilibrary.
com?id=90331.

Williams:1997:HCT

[Wil97] Michael R. Williams. A History of Computing Technology.

IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7739-2 (hardcover). xi + 426 pp. LCCN QA76.17.W56 1997. US$49.00. IEEE catalog number BP07739.

Wilson:2002:FCS

[Wil02] Robin J. Wilson. Four Colors Suffice: How the Map Problem Was Solved. Princeton University Press, Princeton, NJ, USA, 2002. ISBN 0-691-11533-

8. xii + 262 pp. LCCN QA612.19 .W56 2002. URL

http://www.loc.gov/catdir/
enhancements/fy0654/2002114311-
d.html;   http://www.loc.
gov/catdir/toc/prin051/2002114311.
html.

Wilson:2015:MHE

[Wil15] Edward O. Wilson. The Meaning of Human Existence. Liveright Publishing Corporation, New York, NY, USA, 2015. ISBN 1-63149-114-8 (paperback). 208 (est.) pp. LCCN BD450 .W5225 2015.

Winchester:1998:PMT

[Win98] Simon Winchester. The Professor and the Madman: a Tale of Murder, Insanity, and the Making of the Oxford English Dictionary. HarperCollins College Publishers, New York, NY, USA, 1998. ISBN 0-06-017596-6, 0-06-083978-3 (paperback), 0-06-099486-X (paperback). xi + 242 pp. LCCN PE1617.O94 W56 1998. US$23.00.

Winchester:2001:MCW

[Win01] Simon Winchester. The Map That Changed the World: William Smith and the Birth of Modern Geology. HarperCollins College Publishers, New York, NY, USA, 2001. ISBN 0-00-639422-1 (paperback), 0-06-019361-1 (hardcover), 0-14-028039-1 (paperback), 0-670-88407-3. xix + 329 pp. LCCN QE22.S6 W55 2001. US$26.00.

Winchester:2003:KDW

[Win03a] Simon Winchester. Krakatoa:

the Day the World Exploded, August 27, 1883. HarperCollins College Publishers, New York, NY, USA, 2003. ISBN 0-06-621285-5. xvi + 416 pp. LCCN QE523.K73 W56. US$25.95.

Winchester:2003:MES

[Win03b] Simon Winchester. The Meaning of Everything: The Story of the Oxford English Dictionary. HarperCollins College Publishers, New York, NY, USA, 2003. ISBN 0-19-860702-4. xxv + 260 pp. LCCN PE1617.O94 W565 2003. US$25.00.

Winchester:2005:CEW

[Win05] Simon Winchester. A Crack in the Edge of the World: America and the Great California Earthquake of 1906. Harper-Collins College Publishers, New York, NY, USA, 2005. ISBN 0-06-057199-3 (hardcover), 0-7394-6835-9 (paperback), 0-06-083628-8 (ARE), 1-59722-187-2 (large print). xiv + 462 pp. LCCN F869.S357 W56 2005.

Winchester:2008:MWL

[Win08] Simon Winchester. The Man Who Loved China: the Fantastic Story of the Eccentric Scientist Who Unlocked the Mysteries of the Middle Kingdom. Harper, New York, NY, USA, 2008. ISBN 0-06-088459-2. xii + 316 pp. LCCN Q143.N44 W56 2008.

Winchester:2011:AGS

[Win11] Simon Winchester. Atlantic: Great Sea Battles, Heroic Discoveries, Titanic Storms, and a Vast Ocean of a Million Stories. Harper Perennial, New York, NY, USA, 2011. ISBN 0-06-170258-7 (hardcover), 0-06-170262-5 (paperback). xii + 495 + 16 pp. LCCN GC481 .W56 2011.

Winchester:2014:MWU

[Win14] Simon Winchester. The Men who United the States: America’s Explorers, Inventors, Eccentrics, and Mavericks, and the Creation of One Nation, Indivisible. Harper Perennial, New York, NY, USA, 2014. ISBN 0-06-207961-1 (paperback). xxv + 463 pp. LCCN E178 W759 2014.

Winchester:2015:PSC

[Win15]  Simon Winchester.  Pacific:

Silicon Chips and Surfboards, Coral Reefs and Atom Bombs, Brutal Dictators, Fading Empires, and the Coming Collision of the World’s Superpowers.

HarperCollins College Publishers, New York, NY, USA, 2015. ISBN 0-06-231541-2 (hardcover), 0-06-231542-0 (softcover), 0-06-244073-X, 0-06-244687-8, 0-06-242163-8. xiv + 492 pp. LCCN GC771 .W56 2015.

Winchester:2018:PHP

[Win18] Simon Winchester. The Perfectionists: How Precision Engineers Created the Modern World. Harper, New York, NY, USA, 2018. ISBN 0-06-265255-9 (hardcover). xii + 395 pp. LCCN TA19 .W56 2018.

Wirth:1976:ADS

[Wir76] Niklaus Wirth. Algorithms + Data Structures = Programs. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976. ISBN 0-13-022418-9. xvii + 366 pp. LCCN QA76.6 .W561.

Wirth:1983:PM

[Wir83] Niklaus Wirth. Programming in Modula-2. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1983. ISBN 0-387-12206-0 (New York), 3-540-12206-0, 3-540-15078-1, 3-642-96757-4 (e-book), 3-642-96759-0 (print), 3-642-96880-5 (print). 176 pp. LCCN QA76.73.M63 W5713 1983.

Wirth:1986:ADS

[Wir86] Niklaus Wirth. Algorithms and Data Structures. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-021999-1, 0-13-022005-1. 288 pp. LCCN QA76.9.D35 W58 1986.

Wirth:1990:DLC

[Wir90]  Niklaus Wirth.  Drawing

lines, circles, and ellipses in a raster. In Feijen et al. [FvGGM90], chapter 51, pages 428–434. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL http:/

                                        /www.zentralblatt-math.org/
                                        zmath/en/search/?an=0718.

68004. This paper rederives the Bresenham line and circle drawing algorithms, and then extends the latter to handle ellipses.

Wirth:1992:POD

[Wir92] Niklaus Wirth. Project Oberon: The Design of an Operating System and Compiler. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-54428-8. xi + 548 pp. LCCN QA76.76.O63 W58 1992.

Williams:1948:EDC

[WK48] F. C. Williams and T. Kilburn. Electronic digital computers. Nature, 162(4117):487, September 25, 1948. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

                                        URL http://www.nature.com/
                                        nature/journal/v162/n4117/
                                        pdf/162487a0.pdf. Reprinted

in [Ran82, §8.4].

Wu:j-CACM-35-10-83

[WM92] Sun Wu and Udi Manber. Fast text searching allowing errors. Communications of the Association for Computing Machinery, 35(10):83–91, October 1992. This algorithm in this paper is implemented in the agrep program, publicly available via ANONYMOUS FTP to cs.

arizona.edu in the agrep sub-

directory. See also [BYG92].

Walsh:1999:DDG

[WM99] Norman Walsh and Leonard Muellner. DocBook: The Definitive Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-580-7. xii + 635 pp. LCCN QA76.76.H94 W35 1999. US$36.95. URL http:

//www.oreilly.com/catalog/
9781565925809;   http://www.
oreilly.com/catalog/docbook.

Includes CD-ROM.

Witten:1994:MGC

[WMB94] Ian H. Witten, Alistair Moffat, and Timothy C. Bell. Managing Gigabytes: Compressing and Indexing Documents and Images. Van Nostrand Reinhold, New York, NY, USA, 1994. ISBN 0-442-01863-0. xiv + 429 pp. LCCN TA1637 .W58 1994. US$54.95. The software for fulltext indexing described in this book, and errata for the book, are available for anonymous ftp

from munnari.oz.au in the di-

rectory /pub/mg.

Witten:1999:MGC

[WMB99] Ian H. Witten, Alistair Moffat, and Timothy C. Bell. Managing Gigabytes: Compressing and Indexing Documents and Images. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 1999. ISBN 1-55860-570-3. xxxi + 519 pp. LCCN TA1637 .W58 1994. US$54.95.

                                        URL ftp://ftp.math.utah.
                                        edu/pub/mg/;   ftp://ftp.
                                        math.utah.edu/pub/mg/mg-1.
                                        3x/bibsearch-1.02.tar.gz;
                                        ftp://munnari.oz.au:/pub/
                                        mg; http://www.cs.mu.oz.au/
                                        mg/;   http://www.cs.mu.oz.
                                        au/~alistair/arith_coder/;
                                         http://www.mkp.com/books_
                                        catalog/1-55860-570-3.asp;
                                         https://www.math.utah.edu/
                                        pub/mg/;   https://www.math.
                                        utah.edu/pub/mg/mg-1.3x/bibsearch-
                                        1.02.tar.gz.

Willard:1965:IMA

[WMD65] Hobart H. Willard, Lynne L. Merritt, Jr., and John A. Dean. Instrumental Methods of Analysis. Van Nostrand Reinhold, New York, NY, USA, fourth edition, 1965. xviii + 784 pp. LCCN QD61 .W48 1965.

Winston:1996:J

[WN96] Patrick Henry Winston and Sundar Narasimhan. On to Java. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-49826-X (paperback). xi + 329 pp. LCCN QA76.73.J38W56 1996. US$17.95. URL http://www.

                                        aw.com/.

Woit:2006:EWF

[Woi06] Peter Woit. Not Even Wrong:

the Failure of String Theory and the Search for Unity in Physical Law. Basic Books, New York, NY, USA, 2006. ISBN 0-465-09275-6. xix + 291 pp. LCCN QC794.6.S85 W65 2006. URL http:

//www.loc.gov/catdir/toc/
ecip0613/2006013933.html. See

critical review [Ber06].

Wolfram:1988:MSDa

[Wol88] Stephen Wolfram. Mathematica:

a System for Doing Mathematics by Computer. Addison-Wes-ley, Reading, MA, USA, 1988. ISBN 0-201-19334-5 (hardcover), 0-201-19330-2 (paperback), 0-201-51502-4, 0-201-51507-5 (paperback). xviii + 749 pp. LCCN QA76.95 .W65 1988. US$24.95 (paperback), US$39.95 (hardcover). This book was prepared with TEX, LaTEX, and Post-Script.

Wolfram:1991:MSD

[Wol91] Stephen Wolfram. Mathematica:

A System for Doing Mathematics by Computer. Addison-Wes-ley, Reading, MA, USA, second edition, 1991. ISBN 0-201-51507-5. xxii + 961 pp. LCCN QA76.95 .W65 1991.

Wolfram:1992:MRG

[Wol92] Stephen Wolfram. Mathematica Reference Guide. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-51502-4, 0-201-51507-5 (paperback), 0-201-19330-2. 305 pp. LCCN QA76.95 .W66 1992.

Wonneberger:1988:L

[Won88] Reinhard Wonneberger. LaTEX. Addison-Wesley Kompaktführer. Addison-Wesley Verlag, Bonn, Germany, second edition, 1988. ISBN 3-89319-152-6. xiv + 141 pp. Reprinted in 1990 and 1991.

Wood:plj

[Woo] Patrick Wood, editor. Post-ScriptLanguage Journal. Pipeline Associates, Inc., P. O. Box 5763, Parsippany, NJ 07054.

Woolf:1980:SSP

[Woo80] Harry Woolf, editor. Some strangeness in the proportion: a centennial symposium to celebrate the achievements of Albert Einstein. Addison-Wesley, Reading, MA, USA, 1980. ISBN 0-201-09924-1. LCCN QC16.E5 S63.

ACW:software

[Wor] The Austin Code Works. Purveyors of software. 11100 Leafwood Lane, Austin, TX 78750-3409.

Worsley:1950:ED

[Wor50] B. H. Worsley. The EDSAC demonstration. In Report of a Conference on High Speed Automatic Calculating Machines, 22–25 June 1949, pages 12–16. University Mathematical Laboratory, Cambridge, UK, January 1950. Reprinted in [Ran82, §8.6].

Weart:1985:HP

[WP85] Spencer R. Weart and Melba Phillips, editors. History of physics, volume 2 of Readings from Physics Today. American Institute of Physics, Woodbury, NY, USA, 1985. ISBN 0-88318-468-0 (paperback). 375 pp. LCCN QC7 .H694 1985.

Wilkes:1950:E

[WR50] M. V. Wilkes and W. Renwick.

The EDSAC. In Report of a Conference on High Speed Automatic Calculating Machines, 22–25 June 1949, pages 9–11. University Mathematical Laboratory, Cambridge, UK, January 1950. Reprinted in [Ran82, §8.5].

Wilkinson:1971:LA

[WR71] James H. Wilkinson and Christian Reinsch, editors. Linear Algebra, volume II of Handbook for Automatic Computation, Editors: F. L. Bauer, A. S. Householder, F. W. J. Olver, H. Rutishauser, K. Samelson and E. Stiefel. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 0-387-05414-6, 3-540-05414-6. viii + 439 pp. LCCN QA251 .W67.

White:1994:IRS

[WR94a] Steve White and John Reysa, editors. IBM RISC System/ 6000 Technology: Volume II. IBM Corporation, San Jose, CA, USA, 1994. A partial draft is available via anonymous

                                        ftp to ibminet.awdpa.ibm.com

in the PostScript file /pub/

                                        rs6kpapers/techbook.ps.

White:1994:PPT

[WR94b] Steve White and John Reysa, editors. PowerPC and Power2: Technical Aspects of the New IBM RISC System/6000. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, 1994. x + 237 pp. IBM order number SA23-2737-00.

Whitlam:1998:OPD

[WR98] John Whitlam and Lia Correia Raitt. The Oxford Portuguese Dictionary: Portuguese–English, English–Portuguese. Berkley Books, New York, NY, USA, 1998. ISBN 0-425-16389-X. xxii + 436 pp. LCCN PC5333 .W55 1998. US$5.99.

Wright:1939:GSW

[Wri39] Ernest Vincent Wright. Gadsby; a story of over 50,000 words without using the letter “E”. Wetzel Publishing Co., Inc., Los Angeles, CA, USA, 1939.

                                        267 pp. URL http://gadsby.
                                        hypermart.net/.

Weiss:1994:PPP

[WS94] Shlomo Weiss and James E. Smith. Power and PowerPC: Principles, Architecture, Implementation. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1994. ISBN 1-55860-279-8. xvi + 408 pp. LCCN QA76.8.P67 W45 1994. US$54.95.

Wright:1995:TII

[WS95] Gary Wright and W. Richard Stevens. TCP/IP Illustrated: Volume 2. The Implementation. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-63354-X. xxii + 1174 pp. LCCN TK5105.55 .S74 1994. US$52.75.

Wu:1999:PEP

[Wu99] Xingfu Wu. Performance Evaluation, Prediction and Visualization of Parallel Systems, volume 4 of The Kluwer international series on Asian studies in computer and information science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999. ISBN 0-7923-8462-8, 1-4613-7343-3, 1-4615-5147-1. xxii + 319 pp. LCCN QA76.58.W8 1999. US$130.

Wulf:2012:CVR

[Wul12] Andrea Wulf. Chasing Venus:

the race to measure the heavens. Alfred A. Knopf, New York, NY, USA, 2012. ISBN 0-307-70017-8 (hardcover), 0-307-95861-2 (ebook). xxvi + 304 pp. LCCN QB205.A2 W85 2012.

xopen:XPG88-5

[X/O88a] X/Open Company, Ltd. X/Open Portability Guide, Data Management, volume 5. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1988. ISBN 0-13-685876-7. xiii + 204 pp. LCCN QA76.76.O63 X423 1988.

xopen:XPG88-7

[X/O88b] X/Open Company, Ltd. X/Open Portability Guide, Networking Services, volume 7. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1988. ISBN 0-13-685892-9. xiii + 144 pp. LCCN QA76.76.O63 X423 1988.

xopen:XPG88-4

[X/O88c] X/Open Company, Ltd. X/Open Portability Guide, Programming Languages, volume 4. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1988. ISBN 0-13-685868-6. xiii + 198 pp. LCCN QA76.76.O63 X423 1988.

xopen:XPG88-6

[X/O88d] X/Open Company, Ltd. X/Open Portability Guide, Window Management, volume 6. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1988. ISBN 0-13-685884-8. xiv + 338 pp. LCCN QA76.76.O63 X423 1988.

xopen:1989:XOS

[X/O89a] X/Open Company, Ltd. X/Open security guide: programming languages. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-972142-8. x + 132 pp. LCCN QA76.9.A25X663 1989.

xopen:XPG89-3

[X/O89b] X/Open Company, Ltd. X/Open Portability Guide, Supplementary Definitions, volume 3. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-685850-3. xiv + 172 pp. LCCN QA76.76.O63 X423 1988.

xopen:XPG89-1

[X/O89c] X/Open Company, Ltd. X/ Open Portability Guide, XSI Commands and Utilities, volume 1. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-685835-X. xii + 340 pp. LCCN QA76.76.O63 X423 1988.

xopen:XPG89-2

[X/O89d] X/Open Company, Ltd. X/ Open Portability Guide, XSI System Interface and Headers, volume 2. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-685843-0. xx + 666 pp. LCCN QA76.76.O63 X423 1988.

xopen:1992:XOS

[X/O92] X/Open Company, Ltd. X/Open systems and branded products: XPG4. X/Open Company, Ltd., Reading, UK, 1992. ISBN 1-872630-52-9. various pp. LCCN QA76.753.X2 X67 1992.

xopen:1993:XOC

[X/O93a] X/Open Company, Ltd. X-Open CAE specification: message store API (XMS). X/Open Company, Ltd., Reading, UK, 1993. ISBN 1-872630-83-9. xvi + 192 pp. LCCN ????

xopen:1993:XOD

[X/O93b] X/Open Company, Ltd. X/Open DCE: remote procedure control. X/Open Company, Ltd., Reading, UK, 1993. ISBN 1-872630-95-2. xxvi + 641 pp. LCCN ????

xopen:1993:XOT

[X/O93c] X/Open Company, Ltd. X/Open Transport Interface (XTI): version 2. X/Open Company, Ltd., Reading, UK, 1993. ISBN 1-872630-97-9. xx + 244 pp. LCCN TK7887.5 .X2 1994.

xopen:1994:XOF

[X/O94a] X/Open Company, Ltd. X-Open framework and models. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-353442-1. xii + 32 pp. LCCN QA76.9.D5 X67 1994.

xopen:1994:XOG

[X/O94b] X/Open Company, Ltd. X-Open guide: internationalisation guide, version 2. X/Open Company, Ltd., Reading, UK, 1994. ISBN 1-85912-002-4. xvi + 180 pp. LCCN QA76.76.C64 X195 1994.

xopen:1995:XOCa

[X/O95a] X/Open Company, Ltd. X/Open Common Desktop Environment (XCDE) / Definitions and infrastructure. X/Open Company, Ltd., Reading, UK, 1995. ISBN 1-85912-070-9. xxii + 506 pp. LCCN ????

xopen:1995:XOCb

[X/O95b] X/Open Company, Ltd. X/Open curses: issue 4. X/Open Company, Ltd., Reading, UK, 1995. ISBN 1-85912-077-6. xviii + 294 pp. LCCN ????

Yeung:1984:ALP

[Yeu84]  Bik Chung Yeung.  8086/

8088 Assembly Language Programming. Wiley, New York, NY, USA, 1984. ISBN 0-471-90463-5. xi + 265 pp. LCCN QA 76.8 I292 Y48 1984. US$23.88.

Young:1971:ISL

[You71]  David M. Young.  Iterative

Solution of Large Linear Systems. Computer Science and Applied Mathematics, Editor: Werner Rheinboldt. Academic Press, New York, NY, USA, 1971. ISBN 0-12-773050-8. xxiv + 570 pp. LCCN QA195 .Y681 1971.

Young:1983:IA

[You83] Stephen J. Young. An Introduction to Ada. Ellis Horwood Series in Computers and their Applications, Editor: Brian Meek. Ellis Horwood, New York, NY, USA, 1983. ISBN 0-85312-535-X. 400 pp. LCCN QA76.73.A35 Y68 1983. UK£25.00.

Young:1989:XWS

[You89] Douglas A. Young. X Window Systems: Programming and Applications with Xt. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-972167-3. x + 468 pp. LCCN QA76.76 .W56 Y68 1989.

Young:1990:XWS

[You90] Douglas A. Young. X Window Systems: Programming and Applications with Xt: OSF/Motif Edition. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-497074-8. x + 533 pp. LCCN QA76.76 .W56 Y67 1990.

Young:1997:VTH

[You97] David H. Young. The Visual Tcl handbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-461674-X. xxx + 512 pp. LCCN QA76.9.U83Y68 1997. US$45.00.

Young:1992:XWS

[YP92] Douglas A. Young and John A. Pew. The X Window System: Programming and Applications with Xt: OPEN LOOK Edition. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-982992-X (paperback). vii + 589 pp. LCCN QA76.76 .W56Y67 1992.

Zachary:2014:SBR

[Zac14] G. Pascal Zachary. Showstopper!: the breakneck race to create Windows NT and the next generation at Microsoft. Open Road Integrated Media, Inc., New York, NY, USA, 2014. ISBN 1-4976-3883-6, 1-4804-9484-4 (e-book). 328 pp. LCCN QA76.765 .Z334 2014.

Zapf:2007:ASC

[Zap07a] Hermann Zapf. Alphabet Stories: a Chronicle of Technical Developments. Mergenthaler Edition, Linotype GmbH, Bad Homburg, Germany, 2007. ISBN 3-9810319-6-2. 149 pp. LCCN Z250.A2.

Zapf:2007:ACT

[Zap07b] Hermann Zapf. Alphabetgeschichten: Eine Chronik technischer Entwicklungen. (German) [Alphabet Stories: A Chronicle of technical developments]. Mergenthaler Edition, Linotype GmbH, Bad Homburg, Germany, 2007. ISBN 3-9810319-5-4. 156 pp. LCCN ???? 29.00 EUR.

Zimmermann:1995:OPU

[Zim95] Philip R. Zimmermann. The Official PGP User’s Guide. MIT Press, Cambridge, MA, USA, 1995. ISBN 0-262-74017-6. xviii + 127 pp. LCCN TK5102.85 .Z56 1995. US$14.95.

URL http://www-mitpress.
mit.edu/mitp/recent-books/
comp/pgp-user.html.

Zimmerman:2008:UMS

[Zim08] Robert Zimmerman. The universe in a mirror: the saga of the Hubble Telescope and the visionaries who built it. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-14635-7. xv + 295 + 16 pp. LCCN QB500.268. Reissued in 2010 with a new foreword.

Zhang:1996:CSF

[ZJ96] Shanjie Zhang and Jianming Jin.

Computation of Special Functions. Wiley, New York, NY, USA, 1996. ISBN 0-471-11963-6. xxvi + 717 pp. LCCN QA351.C45 1996. US$94.00.

Zlatuska:1992:EPE

[Zla92]  Jiří Zlatuška, editor.  Eu-

roTEX ’92: Proceedings of the 7th European TEX Conference, Prague, Czechoslovakia, September 14–18, 1992, Proceedings of the European TEX Conference. Masarykova Universita, Brno, Czechoslovakia, September 1992. ISBN 80-210-0480-0.

Zlotnick:1991:PSP

[Zlo91] Fred Zlotnick. The POSIX.1 Standard: a Programmer’s Guide. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, 1991. ISBN 0-8053-9605-5. xi + 379 pp. LCCN QA76.76.063 Z57 1991.

Ziliak:2008:CSS

[ZM08] Stephen Thomas Ziliak and Deirdre N. McCloskey. The Cult of Statistical Significance: How the Standard Error Costs Us Jobs, Justice, and Lives. Economics, cognition, and society. University of Michigan Press, Ann Arbor, MI, USA, 2008. ISBN 0-472-07007-X (cloth), 0-472-05007-9 (paperback). xxiii + 321 pp. LCCN HB137 .Z55 2008. URL

http://www.loc.gov/catdir/
enhancements/fy0809/2007035401-
d.html.

Zobel:1997:WCS

[Zob97] Justin Zobel. Writing for Computer Science: the Art of Effective Communication. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 981-3083-22-0 (paperback). xiii + 176 pp. LCCN T11 .Z62 199. US$21.95.

Zoellner:2009:UWEb

[Zoe09] Tom Zoellner. Uranium: War, Energy, and the Rock That Shaped the World. Penguin, London, UK and New York, NY, USA, 2009. ISBN 0-14-311672-X (paperback). xii + 354 pp. LCCN QD181.U7 Z64 2009.

Zubarev:1974:NST

[Zub74] D. N. Zubarev. Nonequilibrium Statistical Thermodynamics. Studies in Soviet Science. Consultants Bureau, New York, NY, USA, 1974. ISBN 0-306-10895-X. xix + 489 pp. LCCN QX311.5.Z813. Translated from the 1971 Russian edition by P. J. Shepherd. Edited by P. Gray and P. J. Shepherd.

Zuse:1936:VSD

[Zus36] K. Zuse. Verfahren zur selbsttätigen Durchführung von Rechnungen mit Hilfe von Rechenmaschinen. (German) [Procedure for automatic execution of calculations by calculating machines]. German patent application Z23624., April 11, 1936. Reprinted in [Ran82, §4.1].

Zuse:1962:ERE

[Zus62] K. Zuse. Entwicklungslinien einer Rechengeräte-Entwicklung von der Mechanik zur Elektronik. (German) [lines of development of computing equipment development from mechanics to electronics]. In W. Hoffman, editor, Digitale Informationswandler, pages 508–532. Vieweg & Sohn, GmbH, Braunschweig, West Germany, 1962. Reprinted in [Ran82, §4.3]. Translated by Mr. and Mrs. P. Jones.

Zwillinger:1992:HI

[Zwi92] Daniel Zwillinger. Handbook of Integration. A. K. Peters, Ltd., Wellesley, MA, USA, 1992. ISBN 0-86720-293-9, 1-4398-6584-1 (ebook). xv + 367 pp. LCCN QA299.3 .Z85 1992.